[Search for users] [Overall Top Noters] [List of all Conferences] [Download this site]

Conference asdg::cmp

Title:CMP
Notice:Welcome to the CMP notesfile
Moderator:FABSIX::B_TOWERLLE
Created:Tue Oct 31 1995
Last Modified:Fri Jun 06 1997
Last Successful Update:Fri Jun 06 1997
Number of topics:94
Total number of notes:6381

17.0. "CMP.B1 EE/PE PASSDOWNS" by SUBPAC::BJUBINVILLE (NFD Fire - Rescue) Wed Nov 29 1995 22:45

    
                                                              
        This note is designated for the Equipment and Process Engineering
    passdowns for:
    
                            CMP.B1
T.RTitleUserPersonal
Name
DateLines
17.85WAFERS NOT GETTING INTO RECEIVE #1 CASSETTESUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 05:498
    
         The wafers were getting caught slidding into RECEIVE
      ELEVATOR #1.  Darrin and I went through the calibration
      procedure and found the elevator slot height to be off.
      We went through and set a new height and cycled wafers
      through it without any more problems.
    
                                            D-shift
17.86Wafers getting stuck at receive elevatorSUBPAC::BJUBINVILLENFD Fire - RescueSat Nov 11 1995 07:249
    
    
       Yesterday, Darrin and I adjusted the Elevator #1 height 
    because the wafers weren't going in all the time.  This
    happened (4) more times again this morning but for now, 
    OPS are just giving it a hand.  They can live with it right
    now.
    
                                               D shift
17.87Vibration after polishFABSIX::R_GEESun Nov 12 1995 07:045
	There is an intermittent loud vibration sound at liftoff after
	the polish cycle. It happened about 5 times tonight no pattern
	to it. This needs to be monitored.	
			
					B shift
17.88passdownFABSIX::B_FINNMon Nov 13 1995 19:178
System is down bgn_repair, Two product wafers and 3 monitors broke on last run 
of the lot.   Carrier # 1 had a screw fall out and onto the table. Aprox.
~~ 1500 wafers on these carriers. Two carriers are being rebuilt currently. 
Found screws loose on one other carrier.  System has been cleaned and 
pad has been changed.

					A shift
17.89Severe Vibration at polish / High ParticlesFABSIX::R_GEETue Nov 14 1995 05:5940
	* Failed particles and uniformity. 

	 - Ran a particle check using the "Transfer" program, this
	   cycles the wafers thru the system without touching the
	   pad. Particle check showed 1 adder, this shows that the
	   particle problem has to be coming from the polish table 
	   area. 
	 - Removed and cleaned the conditioner.
	 - Cleaned the index table and polish table area.
	 - Re-conditioned the pad 300 sec.

	 - Ran a second particle check, no improvement.

	 - Replaced the pad. Broke in the pad with 50 wafers and ran 
	   MQC's, we had no improvement on particles. 

	  The system failed particles (overloaded the surfscan) for
	  the 2nd time. Uniformity is also very unstable.

			Head #1 = 10.2 %
			     #2 =  8.8 %
			     #3 =  6.4 %
			     #4 = 13.4 %
			     #5	= 10.6 %
	
	* There is a severe vibration sound coming from the multi-head/
	  table area during every polish cycle now, yesterday it was 
	  intermittent. It sounds and feels like it is coming mainly
	  from head #1. The program shows there is 30 more pounds
	  of downforce applied to this head than the other 4. The downforce
	  for Head #1 is over 400 pounds, this may be the reason for the 
	  vibration. The carriers have maximum downforce on them when
	  only 2/3 of the carrier is on the pad.
	  We feel the head vibration on the pad is causing the particle
	  problem, it maybe damaging the pad.

	  The system is now down to bgn_repair and will need to be 
	  evaluated by Engineering and field service this morning. 

					B shift
17.90particles, maybe head 4 relatedSUBPAC::LANDRYTue Nov 14 1995 19:4019
    The system was down this morning for particles.  The pad was replaced
    after a thorough clean.  The condition wheel was removed, cleaned and
    re-installed.  The carriers were changed out as well.  As this was
    taking place, the weekly pm was performed.
    
    Breakin and MQC's followed.  Particles were still bad and uniformity
    wasn't good.  Rust was found in a carrier.  The same thing was found
    yesterday as well.  Both times, the carrier was mounted to head #4.
    It appears that there is something leaking by (the rotary union or
    spindle seals) to cause this.  
    
    Head 4 is being taken off-line.  The carrier will remain on head 4. 
    a piece of DF200 film is being applied to it to plug up the blow off
    holes.  The pad will be replaced, the system re-cleaned and MQC's will
    be attempted again.  The hope is that the problem is definately related
    to head 4 ONLY.  
    
    						A Shift
    
17.91High particles, leak at Chiller PCWSSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 15 1995 06:3717
    
    
       Well, once again we scrubbed the entire machine, cleaned out the
     carriers and replaced the pad to try and get rid of the particle
     problems.  
    
        When we came in tonight, Leon noticed the chiller was left off
     once again.  After he turned it on, he noticed a small stream of
     mist shooting from a 1/2" PVC fitting.  It so happened that at
     different times when the pressure went up, the stream was hitting
     the heat exchanger.  I wonder if that blew out the first chiller...
     We tried PVC cementing it and letting it sit for 8 hours but 
     since it was high pressure, it started leaking again.  We ended up
     replacing the entire PVC section and stopping the leak that way.
     We used threaded fittings thus killing the downtime.
    
                                                         D shift
17.92Slurry pump hose/wafer breakage - COnd gimbal assemASDG::POIRIERFri Nov 17 1995 19:2319
    
    
    Slurry Pump Hose:
    ----------------
    A slurry hose ruptured at approx 5:00 p.m. while wafers were polishing. 
    Wafers broke on heads #'s 2, 3 & 4. The pad was removed and the broken
    wafers cleaned-up. A new pad has just been mounted on the Table. 
    Please build-up 3 more carriers and begin MQC's.
    
    
    Conditioner gimbal assembly:
    ----------------------------
    Bob Gugerty from Speedfam replaced the Conditioner Gimbal Assembly
    Bearing Seal this afternoon. While the assembly was apart, Bob also
    replaced the shims in the assembly with a stainless steel type to
    prevent any further corrosion.
    
                                                                John.
                                                                     
17.93major wafer crash, vacuum pump problemsSUBPAC::BJUBINVILLENFD Fire - RescueSat Nov 18 1995 08:4734
    
    
    
        where do I start...  the front door...
    
        The tool was running the second set of 5 wafer during the
      pad break-in when it started making a moose mating call.  A
      wafer slid out head #1 and the tool ws EMO'ed.  It was discovered
      that the slurry valves were off.  Everything was back to normal
      and was running again....  well.... for the 1st set of wafers.
      Wafer #3 was in the middle of unloading as the machine wafer
      detect tripped for a disaster.  Wafer #3 was missing 1/4 of its
      wafer.  It looks like it cracked unde the carrier and the carrier
      kept the 1/4 of the wafer.  The 1/4 wafer with the next whole 
      wafer started a chain reaction and sent little pieces of wafers
      to all of heads.  Since we are using the ID/OD pads, there isn't
      much oscillation so small chunks of wafers can slip around the
      "red eyes."  This by far was THE WAFER(S) CRASH IN HISTORY!!!!
      The inner part of the table that is exposed with no pad was
      very damaged.  There was hack marks in the steel table everywhere.
      A very fine grit stone was used to take them out but it doesn't
      look pretty.  Head #3 was the worst hit.  It now looks like it
      sends "sand"  sized wafer stuff at blow off.  C shift will be 
      looking at today.  Another problem.... the circuit breaker for the
      vacuum pump kept tripping at power-up as it was all week according
      to the vendors.  I swapped out the check valve off of the vacuum
      pump and found it stuck closed.  It was replaced with a new one.
      There are (7) known new white rings left and no plates.  There
      are (10) plates due in today.  The carriers that are on B1 we taken
      from B2 per Tim D.  
                                         We'z outta here!!!!!!!!
      
    
             
17.94Wafer Breakage Recovery / System StatusASDG::POIRIERSat Nov 18 1995 19:0223
    
    SYSTEM STATUS:
    -------------
    
    1) All Carriers, Plates & Shims left over from the wafer breakage
       event last night have been cleaned and returned to the Carrier
       Rebuild Room.
    
    2) Spindle #3, that was disabled due to contamination concerns, has
       been verified "clean" and is now back on line.
    
    3) Five good carriers from CMP.B2 have been transfered to CMP.B1.
       (There are no remaining "good" carrier plates in house for B2.
        Ten were scheduled to arrive today. However, shipping information
        was unavailable from Speedfam today). "A" shift, try checking the
        recieving dock or Package Express tomorrow. If your unable locate
        the plates, we'll have to wait until Monday morning.
    
    4)  A 20 wafer Break-in/Verification run was completed at 6:00 p.m.
        The Pad is now broken-in and the system is ready for MQC's.
    
    
                                                                 John.
17.95good start.......bad endingSUBPAC::LANDRYSun Nov 19 1995 19:2734
      The system is currently down, waiting parts.  The system ran good
      most of the day until early evening.  A wafer got loose from a
      carrier, taking out all five carriers.  Obviously, the wafer
      detection circuit didn't work properly.  There are NO pressure
      plates in-house at this time.  We expected some this weekend, but
      the shipment never arrived.

      An SBA has been generated for 29 plates to go out Monday morning.
      This is an unusually high number, but between the two systems now
      on site, the bad equipment performance since Friday night,  and the 
      lack of shipping capability from this site on weekends, we've hit
      this snag.  Shipping alternatives will need to be researched for 
      such future occurrences.

      So, the system is currently being cleaned.  B shift will have to
      pick up where we leave off.  The following items will need to be
      done:

      		-  Remove, clean and re-install the conditioner wheel
      		-  Fragments will need to be removed from the Index Table,
      		   including brushing the cups with a scrub brush
      		-  Both the tub areas will need to be thoroughly wiped
                   (the wafers were grinding away for awhile and many
                    silicon dust particles are everywhere).

      		-  The wafer detection circuitry/sensor setup will need
                   to be checked out prior to running more wafers.  This 
                   require a pad on the table (or a piece of used pad).
                   The sensor can't receive a reflection from the SS
                   table.

      		-  Maybe a good opportunity tonight to do the PM.

      						A Shift
17.96Cleaned-up wafer particles.FABSIX::R_GEEMon Nov 20 1995 07:0010
	- Dis-assembled and cleaned the conditioner.

	- Cleaned the load / unload cups.

	- Cleaned the index table, and polish table area.
	  
	- Dis-assembled and cleaned the 5 carriers, there on a cart
	  in the carrier rebuild room.

					B shift
17.97passdownSTRATA::FINNMon Nov 20 1995 19:1728
Five (5) carriers pressure plates were received today from Speedfam.  
Four were successfully built and installed on the system.  The fifth pressure 
plate is exhibiting a large range, even when built with different rings on 
different carriers. Speedfam suggests stoning the bottom of the pressure plate 
and the top of the carrier.  We have not done this yet due to down equipment in 
the area.

Twenty nine pressure plates were taken to the dock today and shipment 
confirmation has been received.

Six pressure plates are due here tomorrow from Speedfam.

The systems wafer load reference positions were adjusted by Speedfam while we 
were building carriers.  Table index, multihead, and carrier arm positions were 
reprogrammed. Wafer loss detection was also readjusted.

The system was placed up to verify at 1:30pm, but immediately failed due to 
wafers not being picked up at load by the carriers.  Table index and carrier
arm reference portions are not set correctly.  Possibly drifting?

We are currently attempting to recalibrate these positions.  Index table 
position intermittently drifts.  Index table home sensor and flag seem tight.  

					Work continues.
	
					A shift

17.98plates....carriersSUBPAC::LANDRYTue Nov 21 1995 19:0715
    Today.......
    
    A case of pressure plates (5 plates) were received in.  A decision
    was made to build them to the ring extention of 5.5 - 6.0, which
    are run exclusively on B1 (B2 currently is running ring extentions
    of 7.0 - 8.0).  The five built carriers are on the cart in the 
    carrier rebuild room.
    
    Six more plates are expected to arrive tomorrow morning.  One will be
    built for B1 to complete the desired ten carrier minimum and the other
    five will be built to B2's desire ring extention.  More plates should
    arrive on Friday.
    
    						A shift
    
17.99Elevator water level / Touch screen lockupFABSIX::R_GEEWed Nov 22 1995 05:0813
	* Cross slotting wafers in cassette #1.

		- Found the DI water level in the water track for 
		  elevator #1 was to high. Lower the overflow tubes
		  3/16". 

	* Touch screen locked up  during recipe changes.

		- Checked all the computer connections, no good.
		  
		- reset the computer, all set.
 
						Bob & Brian
17.100pad, cleaning, carrier infoSUBPAC::LANDRYWed Nov 22 1995 19:1212
    The system ran fine today up until mid-afternoon, when the pad came up
    with some poor performance.  A bubble was forming in the middle of the
    pad.  It was changed out with a new ID/OD pad.  The condition wheel was
    removed and cleaned.  The index table was drained  and cleaned.  At
    this time, a fifth carrier was installed (on head 4).  The pad wad
    broken in and, as of this typing, MQC's are underway.
    
    Carrier plates arrived today.  B1 has a full set of built carriers
    ready in the rebuild room.
    
    					Todays Crew.....
    
17.101particlesSUBPAC::BJUBINVILLENFD Fire - RescueThu Nov 23 1995 06:338
    
    
       Failed MQC's and particles.  We pulled the pad and cleaned the
     entire machine.  It passed particles but borderlined failed on
     uniformity.  After another 25 wafers, it looks like it will come
     in.
                                           D shift - Equip.+ Ops
    
17.102MQC's passed....we're running productSUBPAC::LANDRYSat Nov 25 1995 18:409
    We arrived today to find that the system hadn't passed MQC's since late
    Wednesday evening.  It was polishing center fast.  The oscillation
    distance was reduced to keep the edge of the wafer on the pad more.
    All five heads came in under the max limit for uniformity.  The system
    is currently up and flying.
    
    Special thanks to Jim Sadin for this fix.  It was his idea!
    
    						A/C shifters
17.103slurry supply blues......SUBPAC::LANDRYMon Nov 27 1995 18:4920
The slurry supply system went down early this morning.  The pump in the 
distribution unit down stairs kicked off for excessive pressure.  This
is normally caused by a "dead-head" situation.  So, we opened the by-
pass valve in the end VMB to allow flow to increase in the loop.  The
pump down stairs remained running.  As the valve was closed down to
supply pressure to the line, we were able to get the pressure up, but 
there wasn't sufficient flow.

We began taking valves and fittings apart in the VMB's.  There were many
restrictions/solidified lines in the "return" side of the loop.  The two
valves in the VMB were changed out, still not right.  Found the return
tubing to be plugged up solid.  The "slurry cement ball" was removed
from the line and the line re-installed.  All was flowing well through
the entire tool loop.

We also changed out the incoming filter during this extravaganza.

The system is currently UP and running product just fine.

      					A shifters and Tim D
17.104passdownSTRATA::FINNTue Nov 28 1995 19:2013

	System threw a wafer from carrier 5 while lifting off of the 
table at completion of polish. It did not break, one other wafers in same run 
failed to blow off carrier.  With this exception,  system has and is 
running well.

Brought 10 good Speedfam Pressure plates up from the dock. Tim is rush ordering
the white rings.


						Brendan

17.1INSTALL INFORMATIONSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:041090
                                -< 4/26 update >-
--------------------------------------------------------------------------------
    
    	On Tuesday, the rep (Mike Slepikas) and Mark Aubuchont were in to
    begin working on some of the pre-equipment startup items that can be
    addressed.  
    
    	What was accomplished Tuesday:
    
    	*  ALL the motor controllers were removed from the system and a
           jumper on the pc board was changed to a different position.
    	   This mod was a fix for any "out of control" motor performance
    	   that may/may not occur during an EMO situation.  During an EMO
    	   situation, the motors would "jerk" prior to a safe shutdown.
    	   The original fix at the factory was an inprovement but still not
    	   sufficient enough to feel comfortable working near the system.
    
    	   The new fix consisted of removing the controllers and bring them
           out of the cleanroom so that soldering could take place.  Very
    	   long and tedious work!  Lots of screws, gowning/degowning, walk
    	   up and down the stairs.
    
    	   The Speedfam people will be back in on Wednesday to continue on
    	   with the pre-startup punch list.
    
    			Have a safe an ergonomical day.............Tim
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.2                            CMP.B1                             2 of 292
SUBPAC::LANDRY                                       12 lines   6-JUN-1994 16:25
                       -< Monday installation progress >-
--------------------------------------------------------------------------------
    Monday Installation Update.........
    
    Today, the electricians ran conduit and pulled wires from the main
    breaker on the wall to the transformer.  They also ran the conduit from
    the transformer to the main system (minus the wires at this time).  The
    conduit mounted on the system was removed temperarily so that they
    could pulled the wires through it easier.  Once this was accomplished,
    the conduit was remounted to the system.
    
    There was no activity with any other hookups to this system today.
    
                                                        Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.3                            CMP.B1                             3 of 292
SUBPAC::LANDRY                                       13 lines  16-JUN-1994 15:57
                                -< 6/16 Update >-
--------------------------------------------------------------------------------
    Installation Update:
    
    All electrical work has been completed.  This includes the main system,
    the transformer and the TCU.
    
    KSI is beginning to run the GCW drains under the floor.  
    
    No exhaust or incoming air for the HEPA filters work has been performed
    at this time.
    
    No slurry, N2 or CDA installation work has started at this time.
    
                                                     Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.4                            CMP.B1                             4 of 292
SUBPAC::LANDRY                                       22 lines  24-JUN-1994 13:56
                         -< 6/24 Installation Update >-
--------------------------------------------------------------------------------
    
    ALL electrical installation work is COMPLETE!
    
    Both GCW drains have been assembled, leak checked and installed on the
    main system.  COMPLETE!
    
    Work is beginning on the TCU installation at this time.  
    
    Installation has yet to begin on the following:
    
    	CDA
    	N2
    	Slurry
    	Exhaust
    	Incoming air for the HEPA filter
    
    Started gathering labels needed per safety.  
    
    	Have:  CDA, Ethylene Glycol, DI Water
    	Need:  Slurry, N2, GCW, Exhaust
    
    						Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.5                            CMP.B1                             5 of 292
SUBPAC::TOOKER                                       35 lines   8-JUL-1994 11:02
                         -< Hook-Up Status For 7/8/94 >-
--------------------------------------------------------------------------------
                CMP.B1 Tool Hook-Up Status For 7/8/94



COMPLETED ITEMS:
----------------

    *  Bulkhead 

    *  All GCW drain connections

    *  Eelectrical connections


OPEN ITEMS:
-----------

    *  DI Water final connection and pressure gauge

    *  Soft water lines

    *  NESLAB TCU soft water, process cooling water, and drain
       connections

    *  CDA final connection

    *  N2 final connection

    *  All exhaust work

    *  HEPA filter connection

    *  Slurry lines from valve manifold box to system

    *  Slurry valve manifold box valve work

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.6                            CMP.B1                             6 of 292
SUBPAC::LANDRY                                       18 lines  12-JUL-1994 13:34
                         -< 7/12 Installation Update >-
--------------------------------------------------------------------------------
    
    The two HEPA filters for the system were installed today.  This was
    performed now due to NEVCO needing them in-place prior to AC hookup to
    the tool.  Thanks to Earl and Jim for assisting in this task!
    
    Also today, KSI was still in a state of confusion over the VMB and
    regulator hookups required on the slurry loop.  A decision has been
    made to allow KSI to hookup the regulator in the 4th VMB for containment 
    purposes.  This 4th VMB isn't needed until another planarizer is to be
    installed.  A containment vessel will be made in the future to move the
    regulator out of that VMB, thus freeing up the 4th VMB for future
    installation.
    
    Jim Tooker is currently writing the necessary paper work to allow the
    above discribed work to be performed.
    
    						Doc  
                                                                 

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.7                            CMP.B1                             7 of 292
SUBPAC::TOOKER                                       19 lines  26-JUL-1994 09:20
                             -< CMP.B1 punch list >-
--------------------------------------------------------------------------------

                CMP.B1 Tool Hook-Up Status As Of  7/25/94


PUNCH LIST ITEMS
----------------

             ITEM                                      RESPONSIBLE PERSON
   --------------------------------------------------------------------------

  1) GCW pipe union located near tool drain            D. Atkinson/KSI
     connection needs to be moved.  It is 
     interfering with the electrical cabinet 
     door.

  2) Rotate ball valve handles to inside to            D. Atkinson/KSI
     avoid damaging or guard area with uni-strut.

  3) Complete exhaust work.                            D. Atkinson/NEEVCO

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.8                            CMP.B1                             8 of 292
SUBPAC::LANDRY                                       10 lines   1-AUG-1994 13:09
                     -< 8/1 equipment installation update >-
--------------------------------------------------------------------------------
    Brian and I finished up the facilities hookups that EE personnel are
    responsible for on this system.  Those facilities are CDA & N2.  We
    still need to remove the HEPA filter access covers and remove the HEPA
    filters.  
    
    The DI in/out connections still need to be connected by the
    sub-contractors.  Also, the slurry lines / VMB's still need major work
    performed.
    
    						Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.9                            CMP.B1                             9 of 292
SUBPAC::LANDRY                                       35 lines   5-AUG-1994 16:37
                            -< Installation Update >-
--------------------------------------------------------------------------------
    Installation status is as such (as of 4:00PM on 8/5/94):
    
    DI water problem still exists, but it has improved since Thursday.  The
    pressure has increased to 35 PSI.  This may be sufficient for this
    system.  We'll need to load up the DI usage to confirm.
    
    N2 and CDA all set
    
    Slurry lines were connected today (thanks BJ) to the back of the
    system.  The work slated for the VMB hookup is still un-started.  This
    issue will be addressed Monday with the appropriate parties.  The
    hookup performed by Brian and I consists of one input line and one
    output line, tee'd off to the three slurry pumps using SS tubing.
    
    Vendor install is as such:
    
    *  The flipper gripper fingers were changed out to an improved version.
       Earl Dawson helped perform this task.  
    
    *  The slurry pumps were run today using water instead of slurry. 
       They are all set up and ready to go.
    
    *  There is a problem with ALL the motor controllers at this stage. 
       This is a major gating item at this time.  The system can not
       initialize totally until this problem is resolved.  Additional
       help is being persued to excellerate this debug.  Stay tuned!  
    
    *  Mike Slepikas, using the manual control, function tested several
       of the moving components.  All seem fine at this time
    
    *  The remote terminal was tested and used today in beginning the debug
       of the system.  This terminal is a touch screen and is working as
       it should be.
    
    						Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.10                           CMP.B1                            10 of 292
SUBPAC::LANDRY                                       20 lines   9-AUG-1994 07:25
                        -< Startup Update as of 8/8/94 >-
--------------------------------------------------------------------------------
    *  The motor controllers are all working properly now.  Five
       controllers were fixed by rewiring performed at the controller
       input.  The remaining controllers were fixed by making proper 
       connection to ground.
    
    *  Began moving components around using the manual mode.
    
    *  The system is presently experiencing a problem with the conditioner
       arm.  The arm is not lifting out of it's rinse station.
    
    *  We are still running the original software used during the source
       inspection (SPEED463).  New, upgraded software will be installed
       prior to running the system.
    
    *  Earl graciously drew out some cassettes from the stockroom.  So we
       have enough to start up with.
    
                                                     Doc
    
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.11                           CMP.B1                            11 of 292
SUBPAC::LANDRY                                       22 lines  10-AUG-1994 19:35
               -< One thing after another, but we'll get there! >-
--------------------------------------------------------------------------------
    *  A CMP milestone:  ALL glue has been removed safely from the polish
       table.  Many thanks to the Process Eng crew for their diligent work
       to make this  task a success.
    
    *  A problem exists with the flipper/gripper box.  The unload flipper
       is having problems finding home.
    
    *  Reps are having a hard time installing the new version software.
       The system is coming up with an error concerning I/O ports.  Seems
       like an easy enough fix, but it's really turning into a nightmare.
    
    *  A DI water hose began leaking severely.  Found the ferrules to be
       the problem.  It's been repaired.
    
    *  Problems redevelope around two of the motor controllers.  Found the
       wires in the connector to be the problem.  Re-did the connectors,
       all set.
    
    *  The conditioner arm still needs attention once the system software
       is correctly configured/installed.
    
    						Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.12                           CMP.B1                            12 of 292
SUBPAC::LANDRY                                       21 lines  11-AUG-1994 19:19
               -< A glimmer of light is beginning to appear!!! >-
--------------------------------------------------------------------------------
    After many hours of trouble shooting, the software has been
    successfully loaded (SPEEDFAM477).  The reason the system had problems
    accepting the new version was attributed to noisy RF coming from
    the general area around the system.  The tape drive needed to be
    down loaded on to diskettes and loaded through the A drive.  Thanks
    to Jim Martin for assisting the reps with this problem.  His insight
    with computers/software was a great benefit to the Speedfam Reps.
    
    The flipper/gripper box was re-installed and re-calibrated.  Seems to
    be working fine.  Will monitor.
    
    The condition arm still has problems at this point.  Rep's will now
    be able to address the situation now that the software problem is
    resolved.  Stay Tuned!
    
    All motor controllers are functional again.
    
    Hopefully, the startup will begin to excellerate with a few of the
    major problems out of the way.  Only time will tell.
    
                                                    Doc 

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.13                           CMP.B1                            13 of 292
SUBPAC::LANDRY                                       21 lines  12-AUG-1994 15:36
             -< The light at the end of the tunnel is brightening >-
--------------------------------------------------------------------------------
    A good/productive day!
    
    *  Software problems have been greatly reduced since the new software
       has been installed
    
    *  Both unload cassette elevators have been calibrated
    
    *  Robot alignment has been completed
    
    *  The flipper box has been taken care of.  Both flipper/grippers have
       been calibrated and are working as should
    
    *  All but one of the door interlocks are working.  One has a bent
       shaft on the Bimba cylander.  Matt will have one shipped in.
    
    *  Attempted to reduce the vibration caused from the robot vacuum pump.
       There is some slight improvement, but the problem still exists.
    
    *  Still a long way to go, but things are beginning to shape up.
    
    							Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.14                           CMP.B1                            14 of 292
SUBPAC::LANDRY                                       37 lines  16-AUG-1994 18:06
                              -< Progress Report >-
--------------------------------------------------------------------------------
    Accomplishments over the last day or so are as such:
    
    *  First off........The infamous traveling crate has arrived.  The
       crates "trek" started in Chandler AZ, then ended up somehow on
       a flight to Hong Kong (really!).  From there it returned to the
       states, hung out in customs for awhile, then showed up at the Fab 6
       dock on Monday morning.  All this in three weeks!  The crate  had
       more stickers than Herbie the Love Bug!  
    
    *  Changed out the Automation Frame door.
    
    *  All safety guards for electrical connections were installed.
    
    *  The "timeout" problem on the unload flipper to the paddle has
       been resolved.  One of the three sensors in the unload flipper
       box was out of adjustment.  Mechanically re-adjusted said flipper
       is working with positive results.
    
    *  The "timeout" problem with the carriers at the load cups still
       exists.  The problem has been identified as a problem with carrier
       # 4.  It will be addressed further. 
    
    *  Initial "run-out" was performed on the polish table.  The reading
       came out less than 1 mil tolerance.  It will be re-performed again
       for any and/or all CMP personnels viewing pleasure!!!
    
    *  Both unload cassettes were re-setup, due to the new software in-
       stall and are working the nads.
    
    *  The calibration of the touch screen has been performed.  The Input/
       Output screen still has a problem with a couple of dead zones.  
       Software guru's in Chandler are researching the problem as I type!
    
    *  The Cleanroom compatible TCU arrived late this afternoon.  It will
       be de-crated and brought up stairs within the next couple of days.
    
                                                 Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.15                           CMP.B1                            15 of 292
SUBPAC::LANDRY                                       28 lines  19-AUG-1994 16:07
                            -< another day closer! >-
--------------------------------------------------------------------------------
    
    *  Head #4 had a problem showing on the Infranor motor controller
       indicating that their was an over-temp condition.  After many
       hours of trouble shooting, the problem was found to be a short
       in the connector going to the motor controller.  It has been
       repaired.
    
    *  All 5 heads are now moving upon command from the computer.  There
       was an initial problem with heads 3 & 4 where there was no move-
       ment.  Again, many hours were spent trouble shooting the wiring /
       connectors / and motor hardware.  Head #4 was found to have
       overdriven.  This triggered the override circuit.  The head had to
       be physically moved by hand prior to any clearing of the override
       sensor (not much fun doing this).  Head #3 had the same situation
       as well as a suspected poor connection at one of the connectors.
    
    *  Slurry is available but has not been turned on to the system.
    
    *  Both flipper/gripper boxes were worked on again.  They were mechan-
       ically adjusted and home positions were set through the software.
    
    *  The Neslab TCU plumbing has a problem.  It requires a shutoff valve
       for the drain resevoir.  The system was hard plumbed without this 
       feature.  The proper people have been notified and it will be 
       addressed.
    
                                                    Doc
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.16                           CMP.B1                            16 of 292
SUBPAC::LANDRY                                       32 lines  22-AUG-1994 19:44
                    -< two steps forward....one step back! >-
--------------------------------------------------------------------------------
    
    
    
       A very productive day!
    
    *  There was a problem with a couple of the heads.  Somehow, the 
       pneumatic over-rides for those particular heads were in the 
       activation mode.  Returned them to their proper positions and
       all seems fine.
    
    *  Had a problem with some heads not picking up wafers.  Found the 
       vacuum pump off.  A wire had come loose from the connector.
       Reattached wire and all is fine.  
    
    *  Aligned the heads to the load cups and set home position.
    
    *  Aligned the heads to the unload cups and set home position.
    
    *  Cycled 10 wafers from the load station through to the unload
       elevators (without touched down on the polish table due to
       no pad).  All alignments are fine.
    
    *  Removed four wiring harnesses from the flex track.  The wire 
       originally installed has proven to be a problem on the other
       system in the semi-conductor field.  The wire doesn't stand up
       well to movement.  It is being replaced now to avoid anticipated
       problems in the future.  Rewiring to begin in the morning.  About
       a couple of hundred wires will need to be run, cut to size, spliced,
       and attached to the terminal blocks in the multihead.  Oh Joy!!!!
    
                                                      Doc
                             

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.17                           CMP.B1                            17 of 292
SUBPAC::LANDRY                                       18 lines  23-AUG-1994 20:24
                -< What a day!!!  I'm WIRED, so's the machine! >-
--------------------------------------------------------------------------------
    Lots of work done today!
    
    *  All four cables needing replacement were changed out and hooked up.
    
    *  The TCU was swapped out today.  Awaiting KSI to repair improper
       PVC plumbing hookup as well as re-do one of the connection
       configurations.  
    
    *  Ran the interface wires from the main system to the chiller.
    
    *  Repaired broken wires/connectors to the system vacuum pump.
    
    *  Upon completion of the re-wiring, the system would not power up
       properly.  Slurry pumps #1 & #2 were traced to be the problem.
       It's late and it's been a long day.  We'll address this in the
       morning.
    
    						Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.18                           CMP.B1                            18 of 292
SUBPAC::LANDRY                                       33 lines  24-AUG-1994 21:07
                     -< Lost a day+ due to mis-labeling! >-
--------------------------------------------------------------------------------
    Todays activities.......
    
    *  Fixed the slurry pumps problem.  Found that the wires we ran the day
       before to be done correct.  The problem was that the female
       connector for the four wire harnesses (which is mounted in the
       system) was labeled wrong.  Put the proper connector to the proper
       spot and this problem cleared.  UNFORTUNETLY......read on!
    
    *  Since the wire cables were in the wrong spots on power up, many
       sensors saw a bit more current than they should have.  Many sensors
       will be replaced over the next couple of days.  This will further 
       delay the scheduled mechanical acceptance date.
    
    *  Spent many hours trouble-shooting the system after the mishap.
    
    *  We did replace four sensors in the multihead for arms 1, 2, 3, & 4.
       This took 6 hours to do by three people.  Definetly not located in
       a technician friendly spot.  We had to actually remove the multihead
       top cover and change these out from the top of the system.  After 
       they were replaced, all arms needed the sensor and accompanying 
       inhibit switches adjusted.  That BJ for your assistance with this
       task.  (The pedistals definetly got a workout with the two of us
       on top of the machine!!!!)
    
    *  A piece of the the Index Table ring was damaged by one of the heads.
       It has been repaired by welding the polypropylene piece back on.  It
       was originally damaged due to the wiring problem mis-communicating
       with the sensors.  
    
    *  More sensors are due in tomorrow from Speedfam.
    
                                                      Doc
                                                                     

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.19                           CMP.B1                            19 of 292
SUBPAC::BJUBINVILLE                                  10 lines  25-AUG-1994 17:26
                             -< Passdown for 8/24 >-
--------------------------------------------------------------------------------
    
    
       Replaced the final round sensors that were found blown.
    
       There was no vacuum on arm #4.  We traced it to a blown
    fuse on the sensor.
    
       The slurry is on and flowing.  We had to leave the fab
    due to a gas detector alarm.  Mike is going in when the
    all clear is given and is going to do some more work.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.20                           CMP.B1                            20 of 292
SUBPAC::LANDRY                                       22 lines  26-AUG-1994 15:53
                         -< We're getting closer!!!!! >-
--------------------------------------------------------------------------------
    
    *  All sensors have been checked for proper functionality....all set.
    
    *  The load station vacuum pump has been turned off (due to excessive
       vibration causing wafers to wander in the cassettes).  BJ and I
       took the side panel off the system.  We then used a "Greenlee" to
       punch an access hole in it.  BJ then ran a vacuum hookup to the 
       house vacuum system.  
    
    *  The systems "Wet Mode" was tried today.  It is functional and the 
       system has been left in that mode for the weekend.  
    
    *  We began hooking up the chiller to the main system.  Found a leak
       in one of the lines.  It was substantial.  Found the o-ring to be
       missing.  Procured an o-ring and the leak is fixed.  It appears now
       that we have no process cooling water coming to the system.  All
       valves in the room have been checked and are ALL open.  I'll notify
       Dave Atkinson to have valves in the sub-fab checked.
    
                                                 Doc
    
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.21                           CMP.B1                            21 of 292
SUBPAC::LANDRY                                       31 lines  29-AUG-1994 19:46
                          -< Monday, Monday, Moanday >-
--------------------------------------------------------------------------------
    *  All five heads had their respective "wafer sense" sensors adjusted
       today.  They have all been set at 17" of mercury as the acknowledged
       setting.  Once the sensors are made, the system allows for 3" of
       mercury drop until the sensor reaches the low limit of 14".
    
    *  The scrubber section had a problem with the bottom position not
       sensing when it should.  All the I/O's checked out find but the
       computer didn't recognize it.  While trouble shooting the problem,
       it cleared up somehow.  The sensor may be on the "edge" of passing
       or failing.  We'll monitor it.
    
    *  The chiller hookup is in adequate at this time.  The problem is two
       fold.  1)  There is still no PCW coming into the room.  2)  The line
       that supplies the coolant to the main system needs to be regulated
       to 10 PSI.  There is currently NO regulator and or pressure gauge on
       this line.  Stay tuned!
    
    *  The system began cyclying wafers this afternoon.  All was working
       well until the scrubber began to get a "timeout error" (see above).
       Also, a wafer come loose from one of the polishing heads.  It's
       unclear at this time why this happened.  Will monitor future runs.
       When this happened, the system was only using one head down on the
       table at a time.  Nothing was damaged.
     
    *  We need to find a "10X Magnifying Glass" or equivalent to use now
       that wafers are being run on the system.
    
    *  Yet another couple of interlocks were tried today (unplanned of
       course!!!).
    
                                                    Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.22                           CMP.B1                            22 of 292
ASDG::POIRIER                                         4 lines  31-AUG-1994 23:26
                          -< Speedfam 1000 wafer run >-
--------------------------------------------------------------------------------
    
    * The speedfam reps ran into a few minor problems and thus decided
      to postpone the 1000 wafer run until tomorrow. They will be here
      at 7:00 a.m..

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.23                           CMP.B1                            23 of 292
SUBPAC::LANDRY                                       37 lines   6-SEP-1994 18:48
                        -< next wave of reps are here! >-
--------------------------------------------------------------------------------
    
    *  The TCU was fired up today:
    
    	-  The PCW pressure is capable of attaining at least 80 PSI.
    
    	_  The TCU had a flow valve restriction on the incoming valve to
           the system.  It required a manual reset to clear the fault.
    
    	-  The TCU also had a problem with the internal cooling fan.  This
    	   also required a manual reset to clear.  This reset was necessary
           due to an excessive pressure problem.  This may have happened
           possibly as long as two weeks ago when the initial PCW pressure
           first started.
    
    	-  The main system/TCU interface is being tested out for proper
           communication/functionality.  All appears fine but will continue
           to monitor.
    
    	-  Two new reps in here this week (Craig Howard, software & Tom
           Crosby, electro/mechanical).  Both attended site protocol/safety
           training.
    
    	-  Craig began looking into the software issues in the short time 
           available.
    
    *  Carriers:
    
    	-  I received a message from Mike Slepikas this afternoon.  He
    	   informed me that the carriers that were shipped out last Thurs.
           have arrived at Speedfam.  The down side was that they were
    	   damaged.  He reported that they were shipped in poor condition.
    	   Many pins were bent and some carriers were shipped face to face
           with no protective coatings in-place.  He says it will take some
    	   additional time to fix these, possibly delaying their return.
    
    						Doc
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.24                           CMP.B1                            24 of 292
SUBPAC::LANDRY                                       66 lines   7-SEP-1994 18:59
             -< Software changes will make the Hardware run good! >-
--------------------------------------------------------------------------------
    Todays activities were........
    
    *  Software:  Extensive work was done
    
    	-  In a "carrier up" segment, the pressure alarm limit was changed
    	   to look for an excess of 100 lbs. as opposed to the previous
           setting of 50 lbs.
    
    	-  Elevator #2 was experiencing an intermittent alarm saying that
    	   track 2 was blocked when it wasn't.  The sensor was fine.  The
    	   problem was discovered to be within the software.  It has been
           fixed.
    
    	-  The acceleration/deceleration of the polish table was changed by 
    	   a factor of 10. 
    
    	-  The prealigner was getting timeout errors.  The timeout value
    	   was changed to 15 seconds.  Previous setting was 8.
    
    	-  A check was added for the robot sensing the wafer before going 
           to the prealigner.  This was added as an equipment safety
    	   feature in case of vacuum failure.
    
    	-  Match mode bugs were fixed.  Let me explain match mode to you!
    	   Match mode is used to keep track of the wafers from the load
    	   cassettes right on through until they end up in the unload
    	   cassettes.  By the means of using the wafer mapper on the
    	   system, the system knows what slots contain wafers and what
    	   slots don't.  This feature, combined with the match mode, will
    	   put the wafer into the unload cassette in the exact orientation
    	   as they came out of the load cassettes.  For example, if the
           mapper detected say 22 wafers, with slots 3, 8, and 19 empty,
    	   the unload cassette would end up with slots 3, 8, and 19 empty.
    	   Basically, it's a nice feature for experiments/individual wafer
    	   tracking.
       
    
    *  Speedfam Fed-X shipment:
    
    	-  The parts from Speedfam for the vacuum cabinet upgrade arrived
           this morning.  Prep work is underway to perform the task.
    
    	-  Still waiting on one more part (being shipped directly from the
           Furon Company) to arrive.  It's a valve needed to perform the
    	   upgrade.
    
    *  TCU (temperature control unit):
    
    	-  Now that the CMP-V is controlling the Neslab TCU performance,
    	   including set points, I noticed that the Neslab will not con-
    	   trol temperature properly WHEN THE CMP-V SYSTEM IS OFF.  This
    	   is a non-issue but one that should be communicated.  When the 
    	   main system is turned off, the Neslab does NOT receive a set
           point and assumes you want to hit 0 degrees.  There are several
    	   ways to get around this condition.  They are:
    
    		a)  Turn the interface switch to manual and dial in a set
                    point on the front of the TCU.
    		b)  Simply power down the chiller while the main system
                    is off.  
    		c)  Do nothing and let the cooling solution temperature
   		    change.  The only draw back to this is that your fluid
                    temperature will require time to recover to an opera-
                    ting temperature thats acceptable.
    
    							Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.25                           CMP.B1                            25 of 292
SUBPAC::LANDRY                                       26 lines   9-SEP-1994 17:08
                               -< befuddled!!!! >-
--------------------------------------------------------------------------------
    Friday activity:
    
    *  Craig Howard continued to work on the software.  The software is
       nearing completion.  
    
    *  Craig attempted to backup the software but ran into two seperate
       stoppers.  The first one was due to someone bumping the EMO, hence
       killing power to the computer.  The other was caused by a "dip" in
       the incoming power to CMP, which again killed power to the computer.
       We'll attempt this again Monday.  It takes about two hours to per-
       form this task.
    
    *  Tom Crosby has made good progress with the vacuum up-grade to the
       carriers.  The tank has been modified, tubing has been layed out
       (in some cases it's in place), and all necessary fittings have been
       procured.  He was waiting on the valve from Furon to arrive to com-
       plete the job.  Said valve has arrived and the balance of the
       upgrade will be completed Monday.
    
    *  The system is close to running the 1000 wafer mechanical acceptance
       run.  There is one major holdup left.  Speedfam management will not 
       allow the run to take place until the new carriers are available.
       They should be available soon, but I do not have the due date at
       this time.
    
    						Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.26                           CMP.B1                            26 of 292
SUBPAC::LANDRY                                       26 lines  12-SEP-1994 17:38
                            -< Mondays activities >-
--------------------------------------------------------------------------------
    Todays activities........
    
    *  Vacuum Upgrade:  
    
    	1)  The vacuum hose was changed out today to 1/2" diameter.  The
            original size was 3/8".  Sounds like an easy task, but guess
            again.  The line had to be snaked through the flex track and
            the multihead, not to mention the path under the index table.
    
    	2)  The new valve has not been installed yet.  The area it needs
            to go into is pretty crowded.  We'll make it fit though.  Stay
            tuned!
    
    	3)  All wiring harness that were removed (due to them being in the
            way of the hose installation) have been re-installed. 
    
    *  Software:
    
    	1)  Continues to be a problem with backing up the system.  Last
            week's problems were due to power glitches.  The backup tape 
    	    was reformated and tried again.  It proved unsuccessful.  
            Craig Howard will bring in a new tape tomorrow and retry to
            backup the system.  He may also need to reboot the software.
            The power glitches may have corrupted a few files.
    
    							Doc 

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.27                           CMP.B1                            27 of 292
SUBPAC::LANDRY                                       29 lines  13-SEP-1994 19:30
                      -< Tuesdays activities........... >-
--------------------------------------------------------------------------------
    Tuesday's activities.......
    
    Vacuum Upgrade:
    
    	-  The new valve was installed.
    
    	-  The check valve on the vacuum tank was removed and cleaned.
           This was necessary due the tank being drilled into.  Shavings
    	   of PVC cement got into the valve, causing a reduction of 
    	   vacuum.  
    
    	-  Re-installed valve and tested the vacuum system for proper
    	   functionality.  Alls well.
    
    Software:
    
    	-  The systems still needs to be backed up.
    
    Cycling: 
    
    	-  Began cycling the system now that the vacuum upgrade is
    	   complete.
    
    	-  Confirmed proper use of match mode.
    
    	-  All hardware functions are beginning to move.
    
                                                             Doc
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.28                           CMP.B1                            28 of 292
SUBPAC::LANDRY                                       31 lines  15-SEP-1994 21:03
                    -< Polishers do it on a padded table! >-
--------------------------------------------------------------------------------
    Thursday's activities..............
    
    TCU:  Came in this morning and found the Neslab powered off.  The 
          "power on" switch was in the on position but the electrical
          breakers was NOT tripped.  After some quality CMP team work
    	  trouble shooting was performed, traced the problem back to a
    	  hidden switch that was in the wrong position.  Corrected said
          switch and presto.........power was back on.
    
    Pressure switches:  All facility pressure switches were verified for
    	  proper functionality.  Facilities consisted on N2, CDA, DI, and
          vacuum.  Done deal!
    
    Polishing:  Yes, polishing was tried late this afternoon!  Ten wafers
          (Oxided) were run performing single head polishing.  Two wafers
          were done on each head (one at a time).  They were then scrubbed
          and dried on the OnTrak.  Larry C. to measure them in the AM.
          Stay tuned for these potential milestone process results.
    
    Carriers/Pressure Plates:  After the ten wafer polish run, Mike
          Slepikas and your truely changed out the pressure plates on the
          five carriers used.  The pressure plates had come in from
          Speedfam today.  Pre-teardown and post-buildup measurements 
          were taken on each carrier and the information is logged in
          the newly started Speedfam log book.  A good experience!  They
          are pretty easy to change out as they arrived with new film on
          each plate.
    
                                                         Doc
    
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.29                           CMP.B1                            29 of 292
SUBPAC::LANDRY                                       42 lines  19-SEP-1994 18:46
                             -< Monday.....Monday >-
--------------------------------------------------------------------------------
    
    *  Began the day cycling wafers (with the carriers up at the polish
       table).  The systems was running smooth.
    
    *  Began experiencing some minor problems.  Seems to be a few bugs
       still in the software.
    
    *  Tried verifying flow rates for the slurry.  Both Pump 1 & 2 had
       insufficient flow.  Pump 1 was really bad.  Traced problem to be
       a couple of kinked lines in the flex track.  The hoses seemed to
       bunch up for some reason.  The problem has been fixed, but we'll
       definetly have to keep an eye of this.  The flex track has given
       us multiple problems, starting at the source inspection.  At the
       source insp., an air hose was pinched causing a spindle to react
       slowly to commands.
    
    *  The vacuum tank had a problem.  We were getting periodic "low 
       vacuum" alarms.  Seems the tank was extremely full of water.  It
       needed to have the "blow out" performed several times to deplete 
       the excessive water containment.
    
    	NOTE:  The tank is designed much like the Strasbaugh vacuum tank.
    	       It's has a interior wall with a valve in it.  When the 
               water needs to be drained, the internal valve closes, the
               bottom half is opened up to drain (assisted by the intro-
               duction of CDA to blow out the water), while the top half
               retains vacuum for continuous operation of the system.
    
    *  Softwater flow is insufficient at this time.  We're getting about
       25 - 30 PSI to the system.  Looks like this will need to be jacked
       up some.
    
    *  a log book is next to the system to record system activities and/or
       malfunctions as we begin to cycle more and more wafers.  Please feel
       free to use it if necessary.
    
    *  Larry C is scheduled to polish some wafers tomorrow morning (about
       10 or so).  This will be the second attempt at actual polishing.
       Not sure how reliable the initial attempt is due to the unknown
       flow rate of the slurry (was it a problem or not).
    
                                                         Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.30                           CMP.B1                            30 of 292
SUBPAC::LANDRY                                       25 lines  20-SEP-1994 19:37
                     -< interesting day to say the least >-
--------------------------------------------------------------------------------
    Todays events.............
    
    *  Attempted to run 12 wafers (actual polish), 10 survived....2 didn't.
       One wafer came loose from the carrier taking the second wafer along
       with it.  The root cause is believed to be the carrier.  The new
       carriers will help in reducing this situation, so say'ith the rep!
       By the way, I don't have the data from the surviving 10 wafers.
    
    *  There is presently a low flow of softwater going to the system.  We
       broke into the PVC plumbing/regulator/flow meter this morning.  Low
       and behold, a multitude of PVC shavings.  Cleaned affected area's,
       flowed waterthrough the plumbing for 20 - 30 minutes, re-assembled
       the regulator/flow meter and pressure improved.......temporarily.
       It's quite possible some of the shavings had already made their way
       into the internal system regulators/tubing.  The flow is fine for
       time being, but we'll have to address improving it in the near
       future.  We'll have to break into the internal plumbing.
    
    *  The flipper/gripper box was modified this evening.  The mod was to
       lengthen the flex/sealtight hosing.  The original hosing caused a
       lot of stress on the wires running through it.  This, coupled with
       the 180 degree swing in position, was only going to cause inter-
       mittant problems in the near future. 
    
    							Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.31                           CMP.B1                            31 of 292
SUBPAC::LANDRY                                       38 lines  21-SEP-1994 17:59
                -< Wednesdays activities........not so smooth! >-
--------------------------------------------------------------------------------
    Another day........another problem....or 2
    
    *  Began the day with the usual system wipe down/clean.
    
    *  Proceeded to cycle wafers with the heads up.  All was well to the
       system came up with an error {All Cups are NOT Down}.  The monitor
       said to fix the problem and hit any key to continue.  Well, all cups
       were down, so we (Jim S and myself) hit any key and away went the
       system to cycle more wafers.  About 25 wafers or so along, the
       problem re-occurred.  The cups were again down.  So, I stopped the
       system and checked out the I/O's for the cup positions.  Low and
       behold, cup # 2's sensor was not being made.  So, the sensor was
       adjusted and tested.  All is well with this problem.
    
    *  As we were in the process of fixing the sensor, the cup to the right
       of cup # 2 (this one has no number as it is a stationary cup) was
       found to be free twisting.  This needs to be locked into a safe
       position when the index table rotates.  Found the dowel pin to be
       recessed into the wall, thus not locking the cup in place.  Moved
       the dowel to it's proper position and tested.  All is well with
       this problem.
    
    *  The reps made a modification to the scrubber assembly.  It is an air
       driven mechanism that would bang when each limit was made.  They
       installed springs to help alleviate the banging situation.  They
       then re-setup the sensors and tested several strokes.  All seemed
       fine.  Again, wafers were cycled.  Ten wafers or so into running,
       the scrubber started acting up.  Timeout errors were happening,
       the travel rate wasn't consistant and sensors were sporatically
       working/not working.  As I type, the scrubber is being torn down
       to investigate what exactly is happening.  See tomorrows news reel
       for the out-come.
    
    *  Mike S. rebuilt one of the carriers with Jim S. this afternoon.  It
       will be tested out when the scrubber situation clears.
    
    							Doc
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.32                           CMP.B1                            32 of 292
SUBPAC::LANDRY                                       32 lines  26-SEP-1994 18:04
                 -< Monumental day.......the Gyro's are here! >-
--------------------------------------------------------------------------------
    Today......
    
    TCU:  Bobby Gee fixed the last remaining leak on the Neslab.  
    
    Carriers:  The new "Gyro" carriers arrived today.  Upon unpacking, it
    	was noticed that two carriers were damaged.  Both had one pin that
    	was sheared off.  It was obvious that the currier didn't handle
    	them gently at all.  There was a hole in the side of one of the
    	boxes (pretty much the same size as a pin which sheared off...hmmm)
    	So, Speedfam will rush out a couple of replacement pins.  Should be
    	here tomorrow.  Four of the five new Gyro's are presently on the
    	machine.  The fifth will be installed upon replacing the broken
    	pins.
    
    Vacuum Tank:  At startup this morning, there was no vacuum going out
        to supply the heads.  There  was vacuum at the tank in idle mode,
    	but it went to atmosphere when it performed it's "dump" of the
    	tank.  Dump?  Whats that you say?  Well, a dump is a normal occur-
    	ance of the system during elevator servicing or system initializa-
    	tion.  The internal vacuum tank valve closes off the bottom half
    	from the top half.  The top half retains vacuum while the bottom
    	half opens to drain the water accumulated in the tank.  Upon the
    	completion of the dump, the valve opens back up and the entire tank
    	again acts as a vacuum resevoir.  Well, upon completion, the tank
    	never recovered it's proper vacuum pressure.  Found the problem to
    	be PVC shavings in the check valve located at the bottom of the 
    	tank.  Removed the check valve, cleaned out all the shavings and
    	voila, 25 inches of mercury (good vacuum pressure to others).
    
    Cycling:  Countinued to cycle various configurations of wafers.
    
    							Doc
17.3COMPUTER ERRORSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:0520
SUBPAC::LANDRY                                       17 lines   6-DEC-1994 09:33
                          -< Help me.....I'm drying! >-
--------------------------------------------------------------------------------
    This morning,
    
    	I walked in to discover that the system was not in it's proper
    state, which is "wet mode" of course.  So, I turned on the screen to
    observe what state it was in........low and behold, it was in a state
    of "computer dump"!  Tried to reboot the system and it came up with a
    "low vacuum" error.  This is somewhat normal as the vacuum tank needs
    to perform a water dump.  So, I gave it ample time to dump but the
    error didn't go away.  It was discovered that the vacuum pump breaker
    had tripped.  Reset the breaker and then rebooted the computer.  Alas,
    the vacuum pump/tank performed as should.  The system was then
    re-initialized and returned to wet mode, where it is properly wetting!
    
    The pump functionality will need to be watched for awhile as it is
    unclear as to how this problem originated (was it the computer?....was
    it the pump?)
                                                    Doc
17.4SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:0527
SUBPAC::LANDRY                                       24 lines  13-DEC-1994 18:46
                         -< Upgrades performed today >-
--------------------------------------------------------------------------------
    Tom Crosby, Speedfam rep extrodinaire, arrived to day to begin some
    modifications to the system.  Heres whats been accomplished:
    
    *  A new slurry drip ring was installed at the bottom of the polishing
       table.  This mod will re-route the slurry/DI waste to drain directly
       into the tub as opposed to draining into the small drain lines
       originally designatedas the path.  These lines are feared to be too
       small in diameter, with potential clogging of slurry to happen.
    
    *  The main vacuum pump had a safety feature installed.  The mod is a
       cover over the exposed wiring going to the 240VAC connection.
    
    *  Safety covers were installed over all the exposed muffin fans in 
       the system.
    
    *  The "park" station for the conditioner arm had a cascade pipe
       installed.  This will allow the conditioner wheel to be submerged
       in water when in an idle condition.
    
    Tomorrows goal is to install the new inverter into the system.  Other
    minor projects may also happen depending on time restraints.
    
    							Doc
    
17.5VACUUM BREAKER TRIPPEDSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:069
SUBPAC::BJUBINVILLE "Quit crying and play hockey!"    6 lines  12-DEC-1994 08:29
                            -< 3rd shift passover >-
--------------------------------------------------------------------------------
    
       The main vacuum pump circuit breaker tripped when
    I was taking the left and right receive elevators off
    line to dump the DI.  I had to reset the breaker and
    power the machine back up.  
                                 BJ
17.6SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:0627
SUBPAC::LANDRY                                       24 lines  13-DEC-1994 18:46
                         -< Upgrades performed today >-
--------------------------------------------------------------------------------
    Tom Crosby, Speedfam rep extrodinaire, arrived to day to begin some
    modifications to the system.  Heres whats been accomplished:
    
    *  A new slurry drip ring was installed at the bottom of the polishing
       table.  This mod will re-route the slurry/DI waste to drain directly
       into the tub as opposed to draining into the small drain lines
       originally designatedas the path.  These lines are feared to be too
       small in diameter, with potential clogging of slurry to happen.
    
    *  The main vacuum pump had a safety feature installed.  The mod is a
       cover over the exposed wiring going to the 240VAC connection.
    
    *  Safety covers were installed over all the exposed muffin fans in 
       the system.
    
    *  The "park" station for the conditioner arm had a cascade pipe
       installed.  This will allow the conditioner wheel to be submerged
       in water when in an idle condition.
    
    Tomorrows goal is to install the new inverter into the system.  Other
    minor projects may also happen depending on time restraints.
    
    							Doc
    
17.7COMMUNICATION PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:0720
SUBPAC::LANDRY                                       17 lines  14-DEC-1994 21:10
                        -< communicate.......will ya! >-
--------------------------------------------------------------------------------
    
    Todays activities:
    
    *  The new Furon valve for the carrier/spindle vacuum was installed
       along with it's accompanying bracket.
    
    *  The new inverter was installed today.  At the present time, it is
       having difficulty communicating properly with the computer.  Much
       information was exchanged between Tom Crosby and Speedfam home base.
       Still unable to change some parameters.  A recovery plan will be
       determined tomorrow amongst the powers that be.  The system has been 
       lockedout / taggedout for the evening.  The hand sprayer has been
       "fudged" to spray water into the index table to allow the films to
       remain wet.
    
    						Doc
    
17.8SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:0711
                     -< no progress on the inverter issue >-
--------------------------------------------------------------------------------
    No work was performed today concerning the inverter problem.  A
    decision was made to wait for the software guru who will be here
    on Monday (12/19).
    
    I removed the carrier from the system.  The pressure plates were
    dismantled and an SBA has been generated to send them back to Speedfam
    to be re-filmed.
    
    						Doc
17.9SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:0738
SUBPAC::LANDRY                                       35 lines  19-DEC-1994 18:14
              -< Inverter, software and teamwork!  What a trio! >-
--------------------------------------------------------------------------------
    This morning, Craig Howard arrived to begin software changes to the
    system, as well as a couple of other items.  Her's whats transpired:
    
    *  The new version of software Craig arrived with was installed.  We
       couldn't verify it's performance until the inverter problem was
       fixed.
    
    *  The inverter problem has been solved.  After verifying that the 
       hardware was indeed installed properly, the software was checked
       out.  A problem was found within the "config.sys" file.  The error
       was fixed, alas, the inverter began to communicate properly.  We
       then initialized the system and manually ran the table.  All is
       well with this issue!
    
    *  Next, we tried to make a run (with all heads remaining up at the
       polish table).  It didn't go so well.  The system ran was performing
       well until the carriers picked up ,the wafers from the load cups.
       Low and behold....Low Vacuum Alarm.  We found the wafers to NOT be
       loaded properly inside the carrier ring.  Traced the problem back
       to be the Index Table not at home position.  Not sure why this was.
       It could have been from the new software parameters installed or
       possibly because the sensor bracket was a tad loose.  Anyway, we
       secured the bracket and performed the Index Table calibration.
       All seems fine at this time.  
    
    *  We ended the day there and will attempt to run the system in the 
       morning.
    
    *  Talked with Mike Slepikas just a few minutes ago.  He will talk 
       with the Speedfam Films group (led by the incomparable Mark 
       Aubuchont) to try and arrainge for more pressure plates to arrive
       here before the first week on January.  We will definetly need a
       back up set of plates for the only set we currently have in-house.
    
    						Doc 
17.10PEDISTAL BENT, SOFTWARE PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:0824
SUBPAC::LANDRY                                       21 lines  20-DEC-1994 17:18
                             -< Todays activities >-
--------------------------------------------------------------------------------
    Today's activities:
    
    *  Craig attempted to run the system (with his newly installed
       software) but had some problem with one of the load cups, cup 5.
       We found the cylinder that raises the cup to be bent, causing the 
       cup to have problems.  We were able to get a spare cylinder from
       stock but we were unable to come up with the necessary extender
       piece that was also mangled.  Speedfam was notified and that piece
       will be here prior to our need to run said cup.  It will probably
       be hand carried by Mike Slepikas who will be here 1/4/95.  Presently
       the system is running with head 5 disabled.  No problem due to Craig
       will be spending the balance of this week on software
       changes/upgrades.
    
    *  Some software upgrades have been checked out.  Still a couple of
       bugs but also some favorable results.  One such pleasant result was
       the ability of programming in multiple recipes per lot run, instead
       of stopping processing between cycle to change the recipes.  Other
       mile stones will be published as they come along.
    
    						Doc
17.11SOFTWARE AND RPM CHECKSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:0917
SUBPAC::LANDRY                                       14 lines  21-DEC-1994 17:10
                         -< Coming along quite nicely >-
--------------------------------------------------------------------------------
    *  The software was given a test trial today.  At this time, all looks
       good as far a strictly cycling goes.  The heads were run UP with no
       contact being made to the polish table.
    
    *  The table rotation was strobed today.  A setting of 29 RPM was pro-
       grammed in.  The tach readout (handheld tach, external of system)
       was 28.9, can't get much more precise than that.
    
    *  Head five is still in the "disabled" mode until a part we're waiting
       on arrives.
    
    *  More indepth runs to be performed tomorrow.
    
    						Doc
17.12DOWNFORCE CHECKEDSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:1040
SUBPAC::LANDRY                                       37 lines   4-JAN-1995 19:28
                   -< Very good day...lots accomplished!   >-
--------------------------------------------------------------------------------
    A very productive day on the system!
    
    *  The new load cell calibration fixture was tested out today.  It
       works real good.  We then proceeded to calibrate all five heads
       at a down pressure of 300 lbs.  Most heads were fairly close and
       needed only to be tweaked in.
    
    *  Cup 5's cylinder (which was installed last week but was awaiting
       it's lifting pedestal to arrive) was finally fixed totally.  The
       sensor gap for cup at home was re-set-up as well.  Works the nads.
       This also gave us the opportunity to check out the new cup align-
       ment tool.  It appearance is similar to a carrier, except for it's
       diameter and color.  Works good.
    
    *  The new cascade drain set up for the carrier park station was semi-
       fixed today.  It will need to be fit with the proper size o-ring to
       work at optimum performance.
    
    *  The conditioner wheel was removed and checked out.  A decision will
       be made on keeping it or replacing it tomorrow.  A new wheel is in-
       house if needed.
    
    *  Carriers were rebuilt today.  Larry C and the reps worked this
       issue.  New ring gapping is being checked out for best performance
       of the wafer as well as the hardware.  Stay tuned.
    
    *  A couple of negatives happened today.  Nothing serious though:
    
    	-  One of the polypropylene flaps that keep's slurry/water away
           from the multihead drive mechanism's was damaged.  I will
           attempt to reweld it in the morning.
    
    	-  Somehow, unbeknowst to everyone, the film on carrier five got
           gouged.  The carrier has been dis-abled for now and four head
           cycling is underway.  Carrier five will be addressed in the AM. 
    
                                                     Doc
17.13UNIFORMITY, CONDITIONERSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:1026
SUBPAC::LANDRY                                       23 lines   6-JAN-1995 16:27
              -< Still needs to improve, but it's getting better >-
--------------------------------------------------------------------------------
    Todays activities consisted of the following:
    
    *  Larry C, Earl D, and the reps ran some wafers .  Head to head
       uniformity comparisons will be made.  Data presently being
       compiled.  Wafers that were run the prior evening showed a 
       range of 4 - 10.  This needs to be improved, shooting to 
       eventually get down below the 3 range.
    
    *  The diamond conditioning wheel was replaced with a new one.
    
    *  The conditioner wheel park station cascade tube was fitted with
       an o-ring to prevent water leakage.  Functionality to be verified
       tomorrow.
    
    *  Earl D rebuilt five carriers today.  All went well.
    
    *  I gathered up all the items needing to be returned to Speedfam.
       There are 11 pressure plates going back to be re-filmed, 5 carriers
       that are of the obsolete style being returned, a diamond condition-
       ing wheel, and the old inverter that was removed last month.  An
       SBA will be generated next week to ship these out.
    
    							Doc
17.14SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:1113
SUBPAC::LANDRY                                       10 lines   9-JAN-1995 16:50
                      -< Door locking mechanism adjusted >-
--------------------------------------------------------------------------------
    
    	The unload door locking mechanism alignment was out.  There wasn't
    much adjustment space to use, in fact so little that I could not get it
    to align sufficiently.  Had to remove the metal guide that is attached
    to the door and open up the holes used to attach the guide to the door.
    This enabled the guide to have just enough adjustment to get the lock/
    guide to line up sufficiently.  
    
    						Doc
    
17.15VACUUM PUMP CHECK VALVESUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:1112
SUBPAC::LANDRY                                        9 lines  10-JAN-1995 17:51
                             -< Check that valve! >-
--------------------------------------------------------------------------------
    This afternoon.........the system failed several times for a low vacuum
    error.  After searching for an hour or so, the problem was finally
    traced to be a defective (or sticking) check valve.  The valve showed
    signs of slurry in it (hell, thats what the valves there for!).  Tried
    rebuilding the valve with skeptical results.  Therefore, the check
    valve was changed out and normal vacuum pressures were exhibited.  
    Special kudo's to Larry C for his part in the deal!
    
    						Doc
17.16SCRUBBER BELTSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:1126
SUBPAC::LANDRY                                       23 lines  23-JAN-1995 14:55
               -< Warning, warning.......we got a belt problem! >-
--------------------------------------------------------------------------------
    	The Speedfams scrubber currently is being run under a warning
    status.  The bottom set of brushes on the scrubber are not rotating.
    This past Saturday, the problem was investigated.  It showed that
    one of the spliced endings on the drive belt has broken, causing the
    belt to slip off of the gears.  With no way of repairing the belt
    properly, the belt was put back on the gears and the tension adjusted
    to try and help hold it on to the gears better.  It ran that way four
    at least four hours.  Upon arriving Monday morning, it was noted that
    the brushes were not turning again.  The top belt is still working as
    should, rotating the brushes most important as they are the ones that 
    make contact with the device side of the wafer.  Regardless of the 
    Speedfam scrubber status, the wafers ARE ALWAYS run on the OnTrak 
    scrubbers after polishing.  
    
    	Speedfam has been contacted and they will ship out two belts, one
    to install and one as a spare.  Speedfam is also in the process of try-
    ing to come up with a different belt that will last longer than what
    we're seeing now.
    
    	Speedfam will also be sending a check valve for the vacuum system.
    This was replaced once already by Larry C and your truely.
    
    						Doc
17.17SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:1213
SUBPAC::LANDRY                                       10 lines  26-JAN-1995 17:02
                -< Good things are starting to happen here!!! >-
--------------------------------------------------------------------------------
    Manufacturing ran several lots today.  The belt for the scrubber
    section has yet to arrive.  Jim T to have dialog with rep concerning
    status.  System still being run with a warning status.
    
    Ops techs currently changing the polishing pad and carrier films as I
    type.  Film life is proving to last much longer on this system than
    what is used on the Strasbaughs!  An SBA will be generated to ship
    out the next set of pressure plates for re-filming.
    
    						Doc
17.18COMMUNICATION PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:1225
SUBPAC::LANDRY                                       22 lines   8-FEB-1995 16:43
                 -< What happened?  wheres the communication? >-
--------------------------------------------------------------------------------
    Not a pleasant day......
    
    The day started with the unload flipper mis handling a wafer.  It
    didn't crunch it or anything, but it did require EMO'ing the system.
    No big deal.  Retrieved the wafer and re-powered the system.  Nice
    try, but the vacuum pump needed to be reset.  Power down, open the
    safety interlocked door, reset the breaker and rebooted.  All looked
    good.  Unload some wafers to check the flipper, it looked real good.
    Then, went to load some more wafers, lo and behold, the pre-aligner
    isn't responding to it's commands.  Rebooted, same thing.  Tried a 
    third time, same thing.  Powered down and started checking out the 
    communication, electronics, power supply, boards, cables, connectors,
    you name it...it was checked.  The only problem seemed to be in the 
    communication to drive II board. No LED's lit up.   Power was there 
    though.  What there to lose....rebooted the old software and the pre-
    aligner was back in motion.  Went back and forth with the software
    versions to confirm that it was the problem.  The problem stayed with
    one software set and cleared with the other set very consistantly.
    I am currently having dialog with the reps and hope to have a recovery
    plan soon.
    
    							Doc
17.19SCRATCHESSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:1312
SUBPAC::LANDRY                                        9 lines  13-FEB-1995 17:07
               -< You scratch my back, I'll scratch your wafer! >-
--------------------------------------------------------------------------------
    The system is currently going through a polishing pad change and
    carrier swap out.  The system was scratching wafers.  Larry C and
    Jim S working the issue.
    
    The system vacuum pump breaker needed resetting a couple of times
    today.  It was definetly overworked as it took a while for the
    over-temp protector to reset prior to restarting.
    
    						Doc
17.20SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:1319
SUBPAC::BJUBINVILLE "Brooks, Tritt and Diffie"       16 lines  20-FEB-1995 06:58
                 -< SLURRY SYSTEM SHUT DOWN, FILTER INSTALLED >-
--------------------------------------------------------------------------------
  
         An in-line slurry filter has been installed and all
     stainless steel fittings in the slurry section of the 
     machine have been replaced.  A 5 micron filter was installed
     for now.  Finished things up around 5am. 
    
         Around 6:30am, there was no slurry flow on any of the
     polishers.  The slurry alarm did NOT come on.  Ashland is
     on their way in now.  A1 was logged to BGN_EXTERNAL.  Checked
     system to see if the upgrade caused the trip but no leaks were
     found. 
    
                                                 BJ
                                                  
                                                     
 
17.21SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:1323
SUBPAC::LANDRY                                       20 lines  20-FEB-1995 16:32
                       -< Software / Hardware problems >-
--------------------------------------------------------------------------------
    A couple of problems today................................
    
    *  The wrong software was booted up by accident.  This software (486)
       is capable of running the system but should only be used as an
       emergency backup.  Reloaded the latest version (487) and tested.
       Heres where the next problem occurred.
    
    *  The Hardware addresses were mixed up.  Re-addresses the proper line
       items and tested.  Heres where problem 3 occurred.
    
    *  The load flipper was overtraveling at the pre-aligner station,
       causing the wafers to miss-load.  Actually, they weren't loading at
       all.  The grippers were not grasping the wafers properly, leaving
       them on the pre-aligner pedistal.  Re-calibrated the flipper to 
       aligner position.  Tested and all was fine.
    
    *  Ops/Process personnel changed out the pad & carriers as well!
    
    							Doc
    
17.22SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:1447
SUBPAC::LANDRY                                       44 lines  24-FEB-1995 15:24
                         -< Load Flipper isn't homing >-
--------------------------------------------------------------------------------
    The system is currently down for a "load flipper homing" problem.  
    
    It all started with a wafer mis-unloading to the scrubber (yes, this is
    the unload flipper side).  It appeared it was going to be broken, so
    the EMO was hit to prevent such a scene.  Got the wafer out of harms
    way and rebooted the system.  Multiple problems appeared.
    
    *  The vacuum pump breaker kicked off a couple of times.  Found water 
       (with some slurry content) had made it's way back to the pump dia-
       phrams.  This caused reductions in filter pores, causing the pump
       to work like hell.  Eventually mpopping the breaker.  Tore down the
       pump's filter assembly's, cleaned the filter material (thanks
       Larry) and diaphrams as well as the housing, and rebuilt the
       assembly's.  Pump, performance was greatly improved.  This will now
       become a normal check.  
    
       The only way water can make it's way back to the pump is if the
       drain time is too short or if the drain/blowout it's happening.  
       Will be investigated as time allows.
    
    *  The system then had trouble homing the load flipper and the
       scrubber.  Wiped off the sensor for the scrubber and that is now
       working properly.  
    
    *  The system still is experiencing a flipper home problem.  I've
       checked the wiring / connectors, which appear fine.  The digital
       input isn't responding or receiving a signal, which means the 
       sensor isn't working.  The sensor isn't located in a technician
       friendly spot.  It'll require the removal of the flipper box 
       assembly.  If someone gets to it, be careful not to touch the
       adjusting screws as this is a critical area upon re-assembly.
       If messed with, we're in for a bunch of hardware / software
       adjustments.
    
    *  PLEASE NOTE:  The carriers have been removed from the spindles
       and have been placed in the load/unload cups of the index table.
       DO NOT MOVE THE MULTIHEAD without first removing these carriers,
       as major damage WILL happen.  The carrier films need to stay in 
       there if possible to stay wet.  If removed, make sure you keep
       them in a wet situation.  WET MODE is unavailable due to the load
       flipper homing problem prohibits automatic modes.  MANUAL movement
       can still take place though.
    
    					Good Luck.............Doc
17.23FLIPPER BOX PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:1421
SUBPAC::LANDRY                                       18 lines  25-FEB-1995 12:15
              -< The User Friendly Flipper Box has been yanked! >-
--------------------------------------------------------------------------------
    The Flipper Box has been removed from the system.  Unfortunetly, this
    had to come out as all the easy fix possibilities had dwindled down to
    nil.  Thank the Lord that I had a couple of other people in CMP when
    this removal took place.  Definetly an ergonomical hardship.  The unit
    weighs in excess of 40 pounds and is burrowed in a cavity several feet
    inside the panels.  Thanks to Jim S and Eric B for there participaion
    in this endeavor.
    
    The positional sensor amplifier for the load flipper appears to be the 
    problem.  The proper 24 VDC is coming from the connector with it's next
    junction the amplifier (could be a broken wire too).  The amplifier 
    isn't doin' it's communication with the sensor.  By the way, the sensor 
    is fine.  Tried it on the unload flipper amplifier, checks out good.
    
    The system is currently locked out/tagged out.  The carriers have been
    left in the index table so that the films may still remain wet.
    
    				Can't wait for Monday! %^)>         Doc 
17.24SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:1418
SUBPAC::LANDRY                                       15 lines  27-FEB-1995 16:21
                    -< Flipped out......then re-installed! >-
--------------------------------------------------------------------------------
    The system is back to manufacturing........
    
    The Load Flipper's sensor amplifier was found to be biscuits.
    The amplifier was replaced, the new one tweaked in, and the flipper
    positions were retaught.  Thanks go out to Brian to get the flipper
    box amplifier installed dispite the swamping of work he had Sunday
    evening.
    
    Upon attempting to run wafers, the slurry wasn't flowing.  The nozzles
    had become clogged from the system going down on Friday.  Unable to
    put the system in wet mode over the weekend due to the sensor problem
    prohibiting this from happening.  Jim S cleaned up the nozzles and
    the system was off and running.  Good Job Jim!
    
    					Doc  /  BJ  /  Jim T
17.25WATER LEAKSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:1630
SUBPAC::LANDRY                                       27 lines   9-MAR-1995 17:55
                -< Lets hope this won't be a standard Weekly! >-
--------------------------------------------------------------------------------
    Some interesting happenings while doing the Weekly PM today!
    
    *  Did the minor wiping, cleaning, etc.
    
    *  Went to check the numerous pressure gauges in the pneumatics
       cabinet, Whoa!  I do detect liquid.  The vacuum pump was sucking
       the water right out of the vacuum resevoir tank and blowing it
       out it's (be nice) orifice vent.  Well, we powered down the system
       and yanked the pump out of the cabinet.  The diaphram area was
       tore down, cleaned and rebuilt.  Nothing unusual found here.  We
       cleaned up the cabinet and re-installed the pump.
    
       Upon start up of the system, the pump fired up with NO problem.
       Seconds later, it was blowing water again.  The resevoir tank was
       extremely full.  We performed several tank dumps to return the
       tank to proper operating levels.
    
       Now, this situation that happened shouldn't be considered normal.
       The tank should never reach this level.  So, we checked out the 
       dump time and we feel confident that is isn't being allowed enough
       time to rid itself of water.  Hence the gradual rise in level.  The
       value has been increased from 10 to 20 in the program to allow more
       dumping time.  It will be closely monitored over the next couple of
       weeks.
    
    					Doc & the Poiriers X 2
    
17.26GREASE BUILD-UP, WATER LEAKSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:1628
SUBPAC::LANDRY                                       25 lines  15-MAR-1995 16:48
                       -< No runs, one drip, one error! >-
--------------------------------------------------------------------------------
    The Weekly PM was performed today.  Found a couple of items to note.
    
    *  Some lubricant was found sitting on the anidized metal of polish
       arm #2.  This was no where near the polish area.  It was spotted
       internal to the multihead.  We could not locate where the sub-
       stance came from as there were no drips, streaks, moist areas
       other than where it was located.  We wiped up the area and will
       monitor it.
    
    *  The fitting on the bottom of the vacuum tank had developed a leak.
       The leaking substance is water (yes, there should be water in this
       section of the tank.  don't be misled by the vacuum word).   Anyway,
       we remove the fitting and re-taped the threads.  After a few hours,
       the leak began to drip again.  This time it was from the barb
       fitting/hose clamp.  Slightly tightened the clamp and the leak has
       stopped.
    
    *  While wiping up the minor amount of water that had collected in the 
       spill containment pan, we accidently had a wet wipe fall on the
       spill detector.  Nice test to assure the spill detection circuit is
       working fine, which it is.  The reason to why it didn't alarm for 
       the drip was that the water amount was so insignificant that it
       never reached the sensor.
    
    					Doc  /  JP  /  Roger
17.27CONDITIONER ARMSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:1730
SUBPAC::LANDRY                                       27 lines  21-MAR-1995 17:48
                -< the Tale of the Statuesque Conditioner Arm >-
--------------------------------------------------------------------------------
    Today, the conditioner arm movement crapped out........
    
    Well, Roger and I took a quick look at it this morning.  The computer
    has a tendency to occasionaly lock up every so often, which causes
    componants like this to stall out.  It usually just takes a reboot of
    the computer and all is fine.  No such luck this time.
    
    After the reboot, the arm movement was tried in manual mode.  It raised
    then moved out over the table where it decided to pull the statue
    routine, just suspended there.  All screen touch pads were dead and we
    couldn't exit the screen.  Rebooted.  Upon initialization, the
    multihead began to home while the conditioner arm was still over the
    table.  To avoid a massive and expensive crash, we powered down.
    
    Looked into the situation.  All I/O's we could check were functional.
    The motor was fine, gears too!  Checked the connectors associated with
    the loop, one looked suspicious.  Redid the connector and rebooted.
    After assuring the hardware program, we initialized and all was well.
    
    Some interesting findings, there are 7 sensors associated with the
    conditioner arm.  4 for the left to right movement and 3 for the up /
    down movement.  Incredible!  But all warranted as the lose of position
    of the arm could be a major financial hardship, not to mention the
    weeks of down time that would accompany it.
    
    						Doc  /  Roger
    
17.28VACUUM PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:1820
SUBPAC::LANDRY                                       17 lines  30-MAR-1995 18:10
                       -< Better today than yesterday! >-
--------------------------------------------------------------------------------
    Today...............
    
    	The system ran pretty good.  Ran all the wafers fine until the last
    cycle of wafers went through.  The system alarmed for low vacuum.  It 
    quickly recovered (as it should).  Upon restarting the cycle, the
    system interpreted this as polish again.  Jim therefore went into
    manual and finished the cycle.  I could not get the system to repeat
    this phenomena.  We'll watch closely tomorrow.
    
    Talk to Speedfam concerning the vacuum system.  They didn't tell me
    anything I didn't know.  I'll retrace some steps when the system is
    available for a few hours.  I think we need to take a good look at our
    mapping procedures/head disabling/program segments, we may be missing
    something.  Anyway, the system is UP and functional for manufacturings
    use.
    
    						Doc
17.29SCRATCHESSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:1811
SUBPAC::LANDRY                                        8 lines  22-MAR-1995 16:23
                            -< Stop Scratching!!! >-
--------------------------------------------------------------------------------
    This afternoon...........
    
    	The scrubber brushes were changed per Engineerings request. 
    Although the polishing pad had been changed out, scratches were still
    showing up, hence the brush change request.
    
    						1st Shifters
    
17.30COMPUTER BOOT-UP PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:1951
SUBPAC::LANDRY                                       48 lines   6-APR-1995 18:04
                       -< To boot or not to boot....... >-
--------------------------------------------------------------------------------
    Bizarre happening occured today........
    
    We couldn't boot up the computer.  Kept getting two faults:
    
    	*  Controller not responding
    	*  ELODEV not installed
    
    Dispite these faults, MS-DOS booted up to where we got  the proper
    prompt.  So, we attempted to boot-up the executable file.  No response
    from the keyboard what_so_ever.
    
    Well, I had no idea as to what controller this refered to as we have
    many controllers.  So, I checked each and everyone out individually.
    They were all ok with NO fault lights lit up.  Checked the inverter,
    ok too!  Now what?  Well, proceeded to check for the obvious.  Cables,
    connectors, various hardware, etc, etc.  No help!  Removed the CPU 
    module from the rack, looked good / connections clean and in tact.
    Did the same to the floppy module, same results.
    
    Talked to Craig out at Speedfam.  He didn't sound optomistic at all.
    From the errors and the way the system was responding, he suspected
    some DOS problems.  They ain't gonna get fixed from freakin Phoenix!
    
    
    Well, the situation was bleak, but low and behold, Craig suggested a
    few items to check.  Utopia didn't seem possible.  One of the ideas
    was to try another keyboard.  Why, it passed without any keyboard
    error?  Anyway, I located a keyboard that could be hooked up in the 
    computer cabinet.  Due to the cable connection, it couldn't be hooked
    up to the front panel.  So, a rebootin' we'll go.  Unbelievable!  The
    system booted up just fine.  Tried with the old keyboard, no go.  Tried
    the new keyboard, go.  System back up and humming.
    
    Should someone run the system before the new keyboard from Speedfam
    arrives........and the system is powered down in any way.......the
    temporary "Digital" keyboard will need to be hooked up in the computer
    cabinet to execute the bootup software.  It's located on the module
    that is positioned furthest to the left in the cabinet (labeled CPU).
    It's the top cable connection.  Once the windows program is booted,
    re-attach the original cable back to the CPU.  Any questions or
    concerns???  Feel free to call me at home.
    
    The keyboard will be shipped out today.  We'll have it Monday at the
    latest, possibly tomorrow.  The keyboard is not needed to run the
    system.
    
    							the Doctor
     
17.31PARTICLESSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:1925
SUBPAC::LANDRY                                       22 lines  12-APR-1995 17:48
                               -< Particles!!! >-
--------------------------------------------------------------------------------
    The system was experiencing high particles this morning.  Several
    things were done prior to running any MQC's.
    
    *  The hoses on both slurry pumps were changed out.  The old hose
       was badly worn and stretched.  Not sure if this was contributing to
       the particle issue, but it was definetly a possibility.  By the way,
       it's an arduous task to say the least.
    
    *  The slurry flow from the VMB (valve manifold box) was poor.  The
       filter was suspected to be clogging as the secondary gauge was
       fairly stagnant.  The filter was removed and inspected.  It didn't
       appear to be the problem.  Non-the-less, the filter was flushed out
       until a steady flow of clear water was exhibited.  Re-installed the
       cleaned filter and same senario with the gauges existed.  Low and
       behold, the other systems were experiencing the same results.  The
       Ashland folks were contacted to check their end.  The pump down
       stairs had stalled out.  They reset it.  Decent flow was back.
    
    *  Jim Sadin then conditioned the polish pad several times.  MQC checks
       came in fine.
    
    						1st shifters
17.32CUPS NOT GOING DOWN, PEDISTAL BROKESUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2046
SUBPAC::LANDRY                                       43 lines  13-APR-1995 19:58
                          -< We're waiting on parts! >-
--------------------------------------------------------------------------------
    What an incredible day.  Can't say I've ever had a worse one!
    
    Arrived to find that the system wasn't detecting all cups down, despite
    the fact that they were.  Traced problem to a sheared off lifting
    pedistal.  Well, I located one and attempted to install it.  Something
    wasn't quite right.  The cup height was different from before.  Turns
    out that of the six cups that lift, one was a specialty!  Yup, thats
    the one that broke and we ain't got one.  
    
    Proceeded to check into why this happened.  Found that some cups were
    barely making it to the down sensor, right at the hairy edge as they
    say.  Of course I wasn't there when it originally happened, but i'd say
    it's safe to say that the sensor was made by the top edge of the
    pedistal but the pedistal was just low enogh to catch the cups bottom
    mechanism.  
    
    Anyway, we worked the cups up and down a few times and all seemed
    workable.  Wrong!  Somehow, Load Cup four never dropped after picking
    up the wafers from the index table.  The heads began to move in and
    the wafer got wedged/broken.  Minor damage was done to the cup.  The
    carrier film ripped pretty good on that one carrier.  That got rebuilt.
    
    Began looking into why this happened.  It's still unclear.  There are
    many possibilities and none are obvious at this time.  We'll need to
    check a few things while waiting on the obvious parts:
    
    	1)  Need to check if the cup feedthrough is cracked or the weld
            broken on the index table feed through.  If the tube is shifted
            it could be causing the cup(s) to hang up.
    
    	2)  Need to check the index table feedthroughs for burrs.
    
    	3)  Need to check the stems of the cups for defects/differences 
            between suspected bad ones and know good ones.
    
    Anyway, the system is down for parts, being pedistals for the air 
    cylinders.  The above items don't require power on to the system.
    The system can't be properly initialized at this time due to a 
    missing pedistal that won't make a sensor.  Hence, half the system
    will not be properly initialized.
    
    					Doc  /  BJ  /  Roger
    
17.33LOAD/UNLOAD CUPSSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2014
PATE::GEE                                            11 lines  13-APR-1995 23:36
                            -< load & unload cups >-
--------------------------------------------------------------------------------
	De-burred and cleaned the #4 load cup so that it now slides 
        smoothly. Straightened the lift cylinder rod for load #5 the
        best we could. 

	Need to replace the load #5 sensor bracket, it's bent. The
        load and unload cups will need to be dismantled and cleaned,
        there are silicon particles between the two sections.

						2nd Shift
         
        
17.34PEDISTAL REPLACEMENTSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2129
SUBPAC::LANDRY                                       26 lines  14-APR-1995 17:43
                           -< We're 80% functional. >-
--------------------------------------------------------------------------------
    Somewhat successfull day........
    
    *  Received the two lifter cylinder pedistals from Speedfam.  They 
       were installed, alignment performed and a test run made.  We used
       the transfer program as we were just concerned with wafer handling
       issues.  The test run went as such:
    
    	-  Load cups 1, 2, 3, 5 all performed acceptably.
    	-  Load cup 4 was in bad shape.
    	-  The unload cup was raising fine but lowering was a bit iffy.
    
    *  From the above information, we determined that: 1) Load cup 4 would 
       be disabled in the software, which means the cup would stay put in 
       the bottom of the Index Table (for water retaining purposes).  We 
       will be able to run the other four heads for the time being.  2) The
       unload cup specifically (as well as the other cups) would need to be
       watched closely for the next couple of days.  This is due to the
       cups not seating as well as they should.  This is what caused the
       two days of down time.  The cup didn't seat properly, causing a
       crash  of the cup and the lifter pedistal.  3)  Cup 4 will need to 
       be examined much more closely to locate exactly what the problem is.
       Field Service may need to be called in.  Stay tuned.  Kudo's to Jim
       Sadin and Brian Connelly for their assistance in the post-fix
       activities.
    
    							Doc
17.35ARM WOULDN'T HOMESUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2129
SUBPAC::LANDRY                                       26 lines  18-APR-1995 18:01
                                  -< DTWYDK >-
--------------------------------------------------------------------------------
    Rough start today...........
    
    Came in to find the system half powered up (still sitting in DOS). 
    Upon rebooting, arm #1 would not home.  The arm had driven by the home
    sensor and tripped the inhibit switch, killing power to the arm. 
    Couldn't get the arm to respond.  This, in addition to the load cup #4
    problem, put the system into a 60% runable system.  Due to the number
    of lots backed up in the room, it was decided that we would run three
    heads for now and address the two problem areas as time permitted.
    
    Later on in the day, there was an hour window of oportunity for
    maintenance work.  Went back into DOS from where the motors could be
    forced on.  Got arm # 1 to move off of the inhibit switch.  Enabled the
    arm and rehomed it.  All went well.  A fourth head was available to the
    manufacturing folks.  Due to MQC's prior to the fix, they will not run
    this head today.  It will be used tomorrow as checks are done again. 
    The head has been disabled for the night.
    
    Moved on to the load cup 4 problem.  Through swapping out of cups, cup
    positions, utilizing different air cylinders, etc. it was determined
    that the index table cup feed-through is the problem with this area.
    It had taken a whack mid week and made have deformed the shape of the
    feed-through.  Will take a closer look tomorrow, when time permits.
    
    					Doc / with some help from Jim T
    
17.36CONDITIONER ARMSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2230
SUBPAC::LANDRY                                       27 lines  19-APR-1995 17:59
                -< Couple of problems, but we produced anyway! >-
--------------------------------------------------------------------------------
    Couple of happenings today............
    
    *  A problem occurred with the conditioner arm.  Despite it's obvious
       physical "up" position, we received a "timeout for condition arm
       up" error.  Would work sometimes, sometimes not.  Traced problem to
       be two fold, 1)  a loose wire in the top sensor molex connector, and
       2)  the top sensor position was on the edge of make or break.  Of
       course, the wire/connector was repaired and the sensor adjusted.
    
    *  The scrubber timed out several times during manufacturing.  Let them
       finish up the lot, then checked out scrubber performance.  The
       in/out movement of the unit was slower than normal.  The rails were
       lubricated to correct this.  Also wiped off both the in and out
       sensors.  Working fine now.
    
    *  Arm #1 was qualified and used today, bring the system back to four
       heads polishing.  We may give the fifth head a try tomorrow!
    
    *  Had dialog with Speedfam reps today concerning the performance of
       Load Cup #4.  The feethrough needs to be thouroughly checked out for
       integrity (no broken welds, no scratched/pitted areas, possibly try
       to check for any oval shape that may be causing the sticking
       problem).
    
    *  Brought 9 pressure plates to the dock for shipment, be re-filmed.
    
    					Doc  /  Bobby
17.37CUP PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2227
SUBPAC::LANDRY                                       24 lines  20-APR-1995 17:21
                          -< more cup research info >-
--------------------------------------------------------------------------------
    Today..........the four running heads ran well.  I took some time to
    continue investigating the load cup 4 issue.
    
    *  The seal for the cup feedthrough was removed and examined.  Found
       some minor chaffing, but nothing substancial enough to cause the 
       problem.  Tried the cup up/down movement without the seal.  The 
       movement was the same whether the seal was in or out.  
    
    *  Took the locating pin out of play.  No luck.  No luck with the seal
       in or out too.
    
    All possibilities/configurations/etc have been tried to this point,
    except for the feedthrough "roundness" intergrity.  We'll have to
    figure out a way to measure the circumference to see if the shape has
    become ovaled or not.
    
    Dialog continues with the Speedfam reps.  They are also trying to
    figure out a way for us the measure the feedthrough.  
    
    *  One last item, they have sent out the load cup adjustment fixture.
       I will be checking all cup alignments as time permits.
    
    							Doc
    
17.38Machine powered downSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2316
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              13 lines  21-APR-1995 07:40
                           -< Machine powered down >-
--------------------------------------------------------------------------------
    
    
         OPS couldn't get the display to appear at the monitor.  
    Come to find out that it appears the machine had apparently 
    had been accidentally EMO'ed and all DI out of the LOAD table 
    indicating it had been off for at least (4) hours.  The pad 
    was still damp but I told Gwen to check the MQC results carefuly.  
    The particle count was way out of spec.  Larry and Erik will 
    have to decide it the pad should be replaced.  Full MQC's have 
    to be run again.         
                                          
                                          3rd shift
    
17.39CUP PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2321
SUBPAC::LANDRY                                       18 lines  22-APR-1995 11:34
                                 -< a first! >-
--------------------------------------------------------------------------------

    
    
      I had an opportunity to work on the system today.  Continued on
      with the load cup 4 research.  All cups were removed from the Index
      Table.  Throughly cleaned and dried as well.  Upon suggestion from
      Speedfam, I checked out the table for flatness using a dial indicator.
      Seeing that the surface to be measured was made of polypropylene, it
      was expected that there would not be a strict flat tolerance.  They
      just wanted me to find out if there was a big difference in the area
      of the troubled cup as opposed to the known working cups.  Much to my
      shugrin, there didn't seem to be any major change change.

      Right now, all thats left to verify is the inner circumference of the
      feedthrough.  We'll need to come up with a method of measuring it.
      Any ideas on how to measure this are greatly welcomed.

      						Doc
17.40ARM #1 homing sensor failed SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2418
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              15 lines  24-APR-1995 08:28
                        -< ARM #1 homing sensor failed >-
--------------------------------------------------------------------------------
    
    
        Machine was being set-up for running exercise wafers from wet
    idle when ARM #1 failed to move out to the LOAD position.  After
    cycling the power, ARM #1 failed to initialize.  This appear to be
    the same problem as last week, same arm.  It appears the arm missed
    the HOME sensor and tripped the inhibit switch which killed the 
    power to the arm.  I spoke with Doc on the phone and there is a
    extensive procedure to go into DOS and manually bring the head.
    That procedure is not available to us at this time.  Doc is calling
    SpeedFam as soon as they open and will be calling in with the 
    procedure.  Mr. Zimage is requesting a Service Rep comes in and looks
    at this ASAP.....
    
                                                      BJ
17.41Pressure plate, Load/Unload Flipper Box SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2416
FABSIX::R_POIRIER                                    13 lines  24-APR-1995 23:40
                  -< Pressure plate, Load/Unload Flipper Box >-
--------------------------------------------------------------------------------


Removed carrier pressure plates from the carriers on
CMP.B1. Tim will be sending them back to vendor.


Helped Tim remove Load/Unload Flipper box from CMP.B1.
Removed cover from Load/Unload Flipper Box. No visible 
damage seen. SpeedFam rep will be coming in tomorrow 
to work on this and other problems.


				2nd Shift
17.42FLIPPER BOXSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2449
SUBPAC::LANDRY                                       46 lines  25-APR-1995 19:30
                   -< The Lord has surely graced me today! >-
--------------------------------------------------------------------------------

      Very good day for equipment recovery........

*  The Load Flipper not homing problem was identified and repaired.
   The problem was two-fold.  1st, the sensors amplifier was out of
   tolerance.  Adjusted it.  Upon test out, problem 2 was discovered.
   Seems that there was a mechanical problem with the motor shaft/cam.
   Couldn't get the LED's on the amplifier to jive with it's perform-
   ance.  Found the problem to be that the cam was binding a bit with
   the shaft housing.  Repositioned the cam and voila, success.  Re-
   setup the amplifier to match cam position.

   After the Flipper Box was reinstalled, the Load Flipper positions 
   were set and tested.  All went smoothly.  At this time, we checked
   out the Unload Flipper to assure all was still working.  All posi-
   tions remained as was.  We did, however, mechanically adjust one of
   the grippers as this gripper was not making contact with picking up 
   wafer.  Retested and all was fine.  Big kudo's to Roger for his
   assistance in the flipper box ordeal.  

*  The Scrubber timeout problem was checked out.  All sensors are re-
   sponding as should.  Was unable to duplicate past failure occurances.
   We'll have to monitor it for any re-occurances.

*  As the system was powered up today, the Arm #1 not homing occurred.
   Along with it came no rotation.  Found both of these problems to have
   stemmed from the DCX boards (1 & 2).  There are several cable con-
   nectors that were either partial dislodged or totally dislodged.  
   Reconnected the cables and tried to relieve some of the tension on
   the wiring.  The tensioned wires, coupled with the plugging in of 
   the remote terminal, seems to have caused this atrocity.

*  Jim, JP and myself checked out the feedthrough for Load Cup #4 today 
   for any inner diameter flaws.  This was done with percision tools.  
   The feedthrough doesn't seem to be out of round.  I don't know what
   Speedfam plans next, but I'm about out of ideas to troubleshoot this 
   baby.  A rep will be here tomorrow to address this issue.  Stay tuned!

*  The "double loading" problem has been identified as a software problem,
   as was suspected earlier this week.  The rep will address this too!

*  Load Cup height adjustments will be performed tomorrow to assure that
   the proper gapping still exists.

      			Doc, with guest appearances by Roger, JP & Jim

17.43CUP PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2534
SUBPAC::LANDRY                                       31 lines  26-APR-1995 19:25
                              -< good day again >-
--------------------------------------------------------------------------------
    A second good day in a row..........
    
    *  Load Cup # 4 is now functional to use.  Found the feedthrough
       mildly "out of round".  This, coupled with a less than perfect
       cup cylinder, combined to be the culprit.  The Speedfam rep
       (Mark Aubuchont) made the decision to hone the feedthrough and
       lightly smooth the cup cylinder with crocus paper.
    
    *  The Unload cup was adjusted to better sit once the air cylinder
       retracted.  Working better now.
    
    *  The scrubber performance wasn't up to par as well.  found the
       pedistal wafer rotation guides on two pedistals uneven.  Fix that
       problem by reseating.
    
    *  Had Brian Connelly build a fifth carrier for the fix load cup.
       He was shown some techiques by Mark in rebuilding these.  Then 
       installed all five rebuilt carriers.
    
    *  Installed a polishing pad.  IT HAS YET TO BE CONDITIONED!  It
       WILL NEED conditioning prior to MQC runs.
    
    *  Using the TRANSFER program, a cassette of wafers were run to 
       mechanically test all moving components.  NO wafers touch down
       on the pad when using this program.
    
    *  The system will also need to run 50 or so "break-in" wafers, as is
       standard procedure for this system. 
    
    						Doc
    
17.44Flipper box problem SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2515
FABSIX::R_POIRIER                                    12 lines  28-APR-1995 01:47
                            -< Flipper box problem >-
--------------------------------------------------------------------------------



Removed Load/Unload Flipper box from CMP.B1.
Removed cover from Load/Unload Flipper Box. 
Load Home Cam does not seem loose or out of
sync. No visible damage seen. Set up Flipper
box on the bench using jumper connectors for
Tim to work with in the morning. 


				2nd Shift
17.45FLIPPER PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2534
SUBPAC::LANDRY                                       31 lines  28-APR-1995 16:43
                      -< between a rock and a hard place >-
--------------------------------------------------------------------------------
    Some good news.........some bad
    
    *  Bad news:
    
    	-  The Load Flipper has a bad sensor amplifier.  There was one in
           stock.  Installed it and it was also bad.  The original one
           couldn't get the LED's to light when commanded.  The new
           couldn't get the LED'd to shut off!  Yes, I'm sure somethings
           not hooked up correctly.  One is being Fed-X'd for Saturday
           morning (hopefully).  Plans are to install it and test it that
           day.  If all works fine, final installation of the flipper box
           and gripper alignments will be done.
    
    *  Good news:
    
    	-  The Paddle travel problem (shaking) has been cleared up. 
           Cleaned and lubricated the guide rails.  Adjusted the sweep
           pressure.  Much better!
    
    	-  The scrubber travel (hard stopping) has been cleared up.  
           Adjusted the sweep pressure.
    
    	-  Got the carriers into Wet Mode by disabling flipper box.
    
    	-  Fix the software to allow the carrier arms to spread prior to
           servicing the carriers.  This will allow a better ergonomical 
    	   situation as well as allow the use of two hands when using the
           quick disconnects.  Please refrain from using the channel locks
           on the quick disconnects.
    
    					Doc  /  Speedfam  /  Jim T
17.46lipper Box was removedSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2634
SUBPAC::LANDRY                                       31 lines   2-MAY-1995 17:46
                                -< not again!! >-
--------------------------------------------------------------------------------
    Guess what happened today?  The Flipper Box was removed.  Surprised?
    
    Upon arriving this morning, RTC messages showed that the system was on
    warning for scrubber timeouts.  To fix this problem, one must first
    clean the travel rails by wiping them with a dry wipe.  DO NOT use any
    isopropyl alcohol to clean these.  Just the wipe will do.  At this time
    we are NOT adding any lubricant to the rails.  This is to try an use up
    some of the lubricant that is jammed in the bearings.  Speedfam will
    look into whether we have hardened rails or not.  They may be changed
    out in the near future.  Stay tuned!  Follow the above procedure for
    now, until Speedfam gets back to us.
    
    Now the Flipper Box.  Again, the Load Flipper would not home properly.
    The box had to be removed.  Upon checkout, moisture was spotted inside
    the sensor amplifier.  Previous checkouts when the amplifier failed
    showed no signs that moisture had been introduced to the amplifier.
    Moisture was found in the box on two of the three occassions it was
    removed and opened though.  The leaking sections were identified and
    hopefully were repaired enought to last until new seals arrive from
    Speedfam.  Knowing we had no more amplifiers in stock this morning, all
    looked bleak.  But, thanks to Greg Gravelle, we were able to install a 
    new one.  He happened to have one squirreled away from years gone by.
    
    The new one is working fine.  It has been adjusted.  The box has been
    sealed with RTV in the affected leaking area's and a loose bolt was
    tightened down (also evidence of moisture leakage).  The box was
    re-installed and final load cup/pre-aligner alignments were made.
    
    	Hopefully we've seen the last of the interior of the flipper box!
    
    						Doc
17.47Slurry system shut down SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:269
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"               6 lines   3-MAY-1995 08:30
                          -< Slurry system shut down >-
--------------------------------------------------------------------------------
    
        The slurry system shut down causing a few "problems."
    The slurry ran out and the (5) exercise crashed and the
    carriers and pad are being replaced.
    
                                        BJ
17.48 wafers were bit due to slurry supply outage SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2712
SUBPAC::LANDRY                                        9 lines   3-MAY-1995 17:02
                -< wafers were bit due to slurry supply outage >-
--------------------------------------------------------------------------------
    Late third shift last night, the slurry supply went off.  This caused
    the system to loose wafers, breaking them on the polish table.  The
    pad and carriers were changed out.  The system was cleaned up.  
    Standard breakin procedures were done.
    
    The system ran fine for the duration of the shift.
    
    							Doc
    
17.49scrubber timeoutsSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2717
SUBPAC::LANDRY                                       14 lines   4-MAY-1995 18:04
                       -< A decent day was had by all! >-
--------------------------------------------------------------------------------
    Finally, a productive day!
    
    *  System ran all day.  We did get a couple of scrubber timeouts.
       The rails were wiped clean.  We are going to implement a wiping
       of the rails once a shift.  This will be communicated to the 
       people who operate the system over the next day or two.
    
    *  We received in 10 wafer guide rings for the carriers.  I will be
       submitting a SID sheet to have these stocked on site.  They are
       considered a consumable, so we'll have to purchase them.  Other
       consumables will be added to the list.
    
    							Doc
    
17.50Wafer breakage incident SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2818
ASDG::POIRIER                                        15 lines   8-MAY-1995 19:09
                 -< Wafer breakage incident / Warnings logged >-
--------------------------------------------------------------------------------
    
    * Wafer breakage incident today:
    
      The system alarmed for the "vacuum system" for quite some time
      this morning but continued to function normally. The system 
      evidently lost a wafer or wafers during an auto polish cycle
      around mid-morning. Noone actually had visual contact with the 
      carriers during the incident. In addition, the wafer detection
      system, designed to detect broken wafers on the pad, failed to
      operate. The auto polish cycle continued to operate for some time 
      reducing the wafers to a powder. Two Warnings were logged against
      the tool today, One for the Vacuum system and one for the wafer
      detection system.
    
                                                             John. 
17.51Conditioner Arm problem SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2834
SUBPAC::LANDRY                                       31 lines  11-MAY-1995 18:52
                          -< Conditioner Arm problem >-
--------------------------------------------------------------------------------
    Interesting day......
    
    With the uniformity problems happening this week, we had been checking
    out the normal causes (pads, films, parameters, etc., even % solids of
    the slurry) from past experience.  None were panning out to well.  So,
    we began observing the opeation of the system.  Looks like we came up
    with a first.  The conditioner arm was experiencing some strange move-
    ment.  It would sweep inward fine, but the outward movement showed some
    erratic behavior.  It would temorarily stop rotating.  Even when it
    would rotate, it was much slower than the normal speed.
    
    We decided to tear into the conditioner arm and have a see.  Several
    problems were prevalent.  1)  The anodized housing containing two 
    pressed bearings had a problem with less than adequate rotation, 2)
    The shaft assembly had a great deal of rust around the gimbal bearing.
    The rust was so bad it had frozen the gimbal in a stationary position,
    hence defeating it's purpose.
    
    It's believed that the lack of gimballing is definately causing the pad
    to be conditioned poorly.  This will cause the uniformity to sky-rocket
    badly.  It's also possible that the poor pad condition is causing the 
    unusually high number of wafer breakage.
    
    Speedfam will be shipping out the necessary parts to repair the
    problem.  They should be here tomorrow morning.  And will be installed
    then.  Speedfam has also experienced this problem elsewhere and are
    currently working on redesigning the assembly.
    
    						1st Shifters
    
    
17.52condition arm parts arrived and installed SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2824
SUBPAC::LANDRY                                       21 lines  12-MAY-1995 17:31
                 -< condition arm parts arrived and installed >-
--------------------------------------------------------------------------------
    Well fans, the conditioner arm assembly came in this morning.  Leon and
    I installed it.  It was tested out and is working fine.  The wheel
    spins smoothly and uniformly.  In fact, the noise the arm would make
    when conditioning has been reduced to virtually nothing (no more cry of
    the Moose in heat).
    
    The system status is now in "verify".  Process Eng. wants to check out
    the rebuilt carriers before they are installed and used.  The ring
    extention may need to be adjusted prior to running the system.  This is
    due to changes made previously to try and correct the uniformity
    problem that existed.  See Larry for details.
    
    Upon working on the conditioner arm this week, I have noticed excessive
    build up of slurry within the polish area and index table.  It appears
    we will need to increase the cleaning frequency.  I will be getting
    with all those who are currently working on the system and showing them
    the areas that will need to be addressed more often.  It's obvious that
    we will have better control of particles and equipment longevity if we
    institute some cleaning measures.
    
    						Doc  /  Leon
17.53scrubber timed out SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2930
SUBPAC::LANDRY                                       27 lines  17-MAY-1995 19:18
                         -< one wafer per head please >-
--------------------------------------------------------------------------------
    Bad, bad day.............
    
    Upon arriving, I was told that the scrubber timed out a couple of
    times.  So, I did the standard wipe down of the rails.  The system
    began to run again and the timeouts persisted.  What I observed was
    that the scrubber was indeed making the sensor in plenty of time but
    the scrubber would back off the sensor once made, hence showing
    scrubber wasn't at home.  So, I was going to adjust the speed of the
    scrubber when IT happened again.  IT = the system double loading then
    heading off to polish.  Not good, people.  Crunch!  The software bit
    us again.  The system was shut down and Speedfam was called to get the
    software bug fixed.  They sent a person down from Burlington, VT.
    
    He has taken what software we have and will analyse it and remove said
    software bug.  He will be back in the morning to reload the necessary
    files and test out.
    
    Also, we are currently down to two pressure plates.  Six more are being
    Fed-X'd for tomorrow morning along with five rings.  Eight other plates
    are currently in the delivery loop.
    
    Back to the scrubber..........to prevent the scrubber from backing off
    the sensor, Booby and I took off some of the spring loops on the rails.
    These springs were initially added at startup to help the scrubber slow
    down as it approached the home position.
    
    							Doc
17.54NEW SOFTWARESUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2936
SUBPAC::LANDRY                                       33 lines  19-MAY-1995 18:02
                               -< excellent day >-
--------------------------------------------------------------------------------
    Big day............
    
    *  The new boot-up software is revision 491 (speed491).
    
    *  The new software has added some new features to the touch screen.
       I will be getting with all who choose to learn these features in
       the very near future.  The changes made will not affect the way the
       tool has been operated in the past.  The new items are more in the
       realm of convieniences.
    
    *  Started the day with a pad change due to a drying situation of the
       pad.  The scrubber air pressure had dropped, causing the scrubber
       to fail for "waiting scrubber in".  This stops the wet mode cycle
       until the problem is corrected.  It had stopped in a position such
       that the pad was NOT being supplied water.  Adjusted the air pressure 
       to the regulator and the scrubber ran flawlessly all day.
    
    *  Checked out the mis-loading problem from the day previous.  It was
       confirmed that it wasn't software related.  We'll chalk this one up
       as a poor pickup.
    
    *  With the help of Brian C, 200 additional wafers were cycled through
       the system today.  NO failures or assists what-so-ever.
    
    *  The system status is in VERIFY for the purpose of giving it an MQC.
    
    *  We received in an additional 8 pressure plates today, bringing us to
       a spare stock of 11 plates.
    
    *  Also received in 5 sets of load cup alignment fixtures.  
    
    							Doc
    
17.55WAFER BREAK, CONDITIONER RUSTSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:2945
SUBPAC::LANDRY                                       42 lines  25-MAY-1995 19:15
                            -< will it never end? >-
--------------------------------------------------------------------------------
    Many problems today..........
    
    *  System crashed and broke 5 wafers at the polish table.  Wafers were
       being polished at the time of crash.  The two layered polish pad was
       on the table at the time.
    
    *  Performed cup alignment with the new fixtures just received in this
       past week.  Found cup three to be off substantially and cup one of
       marginally.  Both were re-aligned by the rep.
    
    *  Took apart the conditioner arm gimbal assembly, as this was a major
       problem a few weeks back with wafer breakage.  Even though the
       gimbal assembly was only a couple of weeks old, it is already show-
       ing signs of rust/deterioration.  Cleaned and dried the assembly the
       best I could and re-assembled.  The gimbal effect is a tad better
       than adequate, but it's making the loud noise of conditioning days
       gone by.  Speedfam is aware of this design flaw as all of their
       customers have this problem.  Redesign is being addressed.  
    
    	-  We'll have to dismantle the gimbal assembly every week to clean
           and lubricate.  Work the lubrication in and wipe off the excess
           left around the edges.  Wipe the teflon encased spring seal
           before re-assembly.
    
    	-  The condition arm park station is filling up to high, causing
           the gimbal area to take on more water than is needed.  Due to
           the process, we can not rid the assembly of water, but we sure
           can reduce the time it see's water.  Roger (if time permits)
           will be cutting the drain tube down to allow only the diamond
    	   surface to make contact with the water on a continuous basis.
    
    *  After the adjustments were made, the system was run.  Load cup #1
       mis-loaded several times.  It needed to be re-aligned.  Showed Roger
       how to perform this task, we were able to get the spindle to cup
       tolerance fairly tight.
    
    *  The system is going down to an ENG state to research further on the
       excellerated amount of wafer breakage that has been happening of
       late.  Please do NOT attempt to run the system tonight.
    
    						Doc  /  Roger
    
17.56CUP ALIGNMENTSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3022
SUBPAC::LANDRY                                       19 lines  26-MAY-1995 17:31
                               -< 60% for know! >-
--------------------------------------------------------------------------------
    Well, this is what happening with the system.........
    
    *  Cup alignment was checked on all heads this morning.  This was done
       due to vacuum problems on heads 1 & 2.  Heads two and three were 
       adjusted, the remaining three heads were fine.
    
    *  Carrier tolerence were checked out first thing this morning.  All
       carriers were verified and re-installed.
    
    *  Upon exercising of the system, heads 1 & 2  errored for vacuum
       problems.  The problem could be one of several different items. 
       Lack of polisher equipment in the UP state swayed us to run three
       heads for the time being.  Heads 1 & 2 were disabled.
    
    *  Nylon washers were received in today.  I will bring them into the
       room Tuesday.
    
    						Doc
    
17.57VACUUM switches adjustedSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3126
SUBPAC::LANDRY                                       23 lines  30-MAY-1995 18:43
                  -< switches adjusted..........temporarily >-
--------------------------------------------------------------------------------
    The system was turned over to EE late in the afternoon for "no more
    than forty-five minutes, ok?" to look at the vacuum problem with heads 
    # 1 + 2.  
    
    Yes, there is a slight vacuum deficiency.  No, it's not severe.  In
    fact, it's marginal.  It's missing set point by 1" Hg.  Due to the 
    amount of lots in the room and the amount of systems functioning in the
    room right now, a decision was made to adjust the pressure switches to
    trip off at about 15" Hg, instead of the customary 17" +/- 2" range.
    It will remain as is until ample time can be arrainged to properly 
    address the problem.
    
    After adjusting the switches, we went to change the off-sets, but the 
    settings were long gone.  Jesus is in the process of trying to contact
    Larry C for those settings.  If he does not get through to Larry
    tonight, Jesus will continue to run heads 3, 4, & 5.  If he does get
    the off-set info, heads 1 + 2 will need to be re-qualified and heads 3, 
    4, + 5 will need removal rates checked.
    
    Also today, the scrubber timed out a couple of times.  Wiped the rails
    clean and re-lubed them.  
    
    						Doc  /  Roger
17.58vacuum restriction found SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3122
SUBPAC::LANDRY                                       19 lines  31-MAY-1995 17:50
                             -< restriction found >-
--------------------------------------------------------------------------------
    Some good results today with the vacuum problem...............
    
    We experienced vacuum problems again today with spindles 1 & 2. 
    Looking at the gauge, they were at the same reading as the previous
    day.  It was thought that maybe we set the pressure switches too close
    to the "cross-over-point".  Not the case this time.
    
    Head 5 errored as well, so Earl and myself took a look at the vacuum
    pump.  It was obvious that the pump had seen some water/slurry residue.
    We tore the pump down and found a build-up of "crust_ated" slurry on
    the internal filters and diaphrams.  We cleaned up the mess and
    re-assembled the pump.  Pressure improved to about 5" Hg.  All heads
    were reaching approximately 19-21" except for head four which was at
    25".  Found the gauge to be defective as it was reading about 4" at
    atmosphere.  This "convienience" gauge will be replaced in the future.
    As time permits, the pressure switches on heads 1 + 2 will be re-set
    for 17" Hg.
    
    					Doc  /  Earl 
17.59cup problem SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3213
SUBPAC::LANDRY                                       10 lines   1-JUN-1995 17:45
                           -< cup problem resolved  >-
--------------------------------------------------------------------------------
    The system had an equipment failure at the tail end of 3rd shift last
    night.  The system came up with the error "one or more cups not down".
    Through the use of the I/O functions, the problem was traced to a
    sensor not being made on cup #3.  Checked out the sensor functionality
    and found it to be "good".  Therefore, the physical position of the
    switch was adjusted.  Twenty five wafers were run to verify the fix.
    Got some assistance from Earl and Rich with this endevor.
    
    					Doc  /  BJ
                                                             
17.60a lack of N2 pressureSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3215
SUBPAC::LANDRY                                       12 lines   2-JUN-1995 17:45
                      -< a lack of pressure for a change >-
--------------------------------------------------------------------------------
    The system had problems with the incoming N2 pressure today.  We were
    getting @ 70 PSI this morning, causing erractic system behavior.  The
    N2 was being supplied by the backup system.  Air Products was contacted
    and about 4 hours later, we were back on line with 100 PSI incoming.
    
    The system ran after this and high uniformity was exhibited on two of
    the heads.  The carriers were removed and measured.  One was definately
    out of spec and the other was tweaked in a bit tighter.  Roger and I
    ran 25 breakin wafers after the carriers were installed.  Next step due
    is MQC's.
    
    						Doc
17.61vacuum pump down SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3217
SUBPAC::LANDRY                                       14 lines   6-JUN-1995 17:25
                      -< vacuum pump down for the count >-
--------------------------------------------------------------------------------
    The system is currently down awaiting a vacuum pump.
    
    We had trouble pulling acceptable levels of vacuum.  Found the fan
    portion on one of the pistons in disarray.  Looks like some internal
    workings have come loose.
    
    So, I have shut off the vacuum pump breaker, edited a couple of files
    to ignore the vacuum checks, and rebooted the system.  It is currently
    in wet mode.  The pump is due in tomorrow.
    
    Thanks to Earl and Brian for their assistance.
    
    							Doc
    
17.62the pump has been installed SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3331
SUBPAC::LANDRY                                       28 lines   7-JUN-1995 18:04
                        -< the pump has been installed >-
--------------------------------------------------------------------------------
    The system is currently back on line..........
    
    The vacuum pump arrived this morning.  It was wired and installed.
    We ran 50 wafers utilizing the "transfer" receipe (no contact made
    at the polish table).  The vacuum pressure at the most troublesome
    head (head #1) was reading 21" Hg, which is very good.
    
    Now, with the new pump installed, we have not rid ourselves of the 
    original vacuum problem with the system.  The new pump will not pre-
    vent the pump from sucking waste water/slurry through it.  Some ideas
    were exchanged between Speedfam and myself.  We both feel that the
    problem resides within the vacuum tank.  The feeling is that the
    internal tank valve is failing to actuate to it's designed limits.
    The vacuum tank internal working can not be checked out right now
    due to the location of the unit.  Therefore, we have made a pressure
    adjustment to the "Miscallaneous" regulator.  The pressure has been
    increased to ~70 PSI (from the original setting of 40 PSI).  This
    increase is to supply more air pressure for the valve actuation.
    
    Preliminary results look favorable at this time.  We'll have to run the
    system for a day or two before we can begin to feel that this may have
    solved the problem.  We'll continue dialog with Speedfam until the
    problem has been solved.
    
    So, Jesus Roa is beginning to run some MQC wafers through the system.
    He and Brian C just finished up running 50 breakin wafers.
    
    						Doc 
17.63Wafer breakage, machine downSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3313
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              10 lines   8-JUN-1995 01:47
                    -< Wafer breakage, machine downage.... >-
--------------------------------------------------------------------------------
    
          A wafer broke on the pad and took everything else out.  
      Carrier #2 was the one that lost it's wafer.  There are no
      more rings so we can't get it back on line.  The pad was removed
      and not reinstalled.  The scrubber brushes were removed and 
      placed in DI water.  The entire machine was cleaned up and and 
      the carriers were broken down and cleaned.  Cause of the problem 
      is still under investigation.
    
                                                      Leon, Gwen and BJ
17.64CARRIERS, PLATES, RINGS AND TABLE ISSUESSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3445
SUBPAC::LANDRY                                       42 lines   8-JUN-1995 18:22
               -< waiting on pressure plates / retaining rings >-
--------------------------------------------------------------------------------
    Depressing day............
    
    *  Came in to a disaster with broken wafers.  Only one plate was left
       to use.  No rings either.
    
    *  We decided to run a single head using the best of the old rings on
       that one carrier.  At least this head ran good all day.
    
    *  The rings have not been ordered for stock yet due to some
       communication problems with Speedfam.
    
    *  The polish table needed to have a few dings / scratches removed
       from it today.  The slurry ring also has some dings in it.  Please
       refrain from using an unacceptable implements when removing the
       pads.  
    
    *  Had dialog with several Speedfam people today concerning the vast
       problems we've exhibited over the last few weeks.
    
    	-  Carriers:  the .010" shims we've been using appear to be causing
           some problems with hardware integrity.  The inside diameter of
           these shims are smaller than the other two sizes we have (.002"
           + .003").  The tightness is causing the retaining ring receive
           more stress than is expected.  New shims are on the way.
    
    	   A minimum of 5 pressure plates (potential of 10) will be here 
           tomorrow.  We'll be back on line to run 5 heads.
    
    	   We should also be receiving 20 retaining rings.
    
    	   They have FAX'd out a new parts listing for the carriers.  I
           will be pulling together a list of various items and get them
           stocked on site in the near future.
    
    	-  Wafer Breakage:  Speedfam has asked that I perform a couple of 
           checks for them that may be affecting our breakage problems. 
    	   The conditioner arm is the focus concerning this check.  I need
           to gap the conditioner gimbal assembly when the wheel is at the
           table.  Down force pressure changes will need to be adjusted for
           this check.
    
    							Doc
17.65plates/rings and assorted things SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3433
SUBPAC::LANDRY                                       30 lines   9-JUN-1995 17:21
                     -< plates/rings and assorted things >-
--------------------------------------------------------------------------------
    Today was like Christmas, 4 boxes of consumables showed up!
    
    *  Received in 10 pressure plates.
    
    *  Received in 20 retaining rings (valued at $412.50 each).  
    
    *  No new shims showed up though.  But I have found plenty in our 
       fab to suffice for the time being.  I have weeded out several bad
       ones.  I'll weed out the remaining as time permits.
    
    *  Larry is currently constructing some carriers.
    
    *  Had more dialog with Speedfam today about wafer breakage.  The
       problem has been ellivated there.  Many people are focusing on our 
       situation.  Work between us and them forges forward.
    
    	-  We will be checking the gimbal gap of the conditioner assembly
           as the system becomes available.
    
    	_  Segments of the recipe will be thoroughly researched.
    
    	-  Carrier construction will be altered.  Received a procedure that
           Speedfam recommends using while building carriers.  
    
    	-  New .010" shims to be procured. 
    
    				Stay Tuned!
    
    						Doc
                                           
17.66gap measurement of the gimbal assemblySUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3520
SUBPAC::LANDRY                                       17 lines  10-JUN-1995 11:35
                  -< gap measurement of the gimbal assembly >-
--------------------------------------------------------------------------------
    This morning,
    
    	The gap was measured at the conditioner arm gimbal assembly. 
    Speedfam had ask me to check this measurement.  The conditioner arm /
    gimbal assembly has a lot to do with potential wafer breakage.  The gap
    was measured at .011" @ 100% down force, taking 8 measurements in a 360
    degree radius.  The gap was consistant.  The gimbal assembly seems to
    be holding up better since the conditioning disk stations drain pipe
    height was altered.  The potential exposed gimbal seal no longer sit's
    in the water in an idle state.  
    
    	Took a shot at building one carrier.  It is on the table and
    labeled.  It should be checked out by Larry prior to anyone attempting
    to use it.  He has been notified of it.
    
    						Doc
    
17.67wafer got loose without smashing! SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3536
SUBPAC::LANDRY                                       33 lines  16-JUN-1995 18:21
               -< 1st time a wafer got loose without smashing! >-
--------------------------------------------------------------------------------
    Rough start.............
    
    Upon arrival, the machine just spit out a wafer from under one of the
    carriers.  We found that another wafer was hanging over the edge of the
    extention ring of the carrier.  It was unclear as to how this happened.
    We then manually moved the wafers to the Mutihead to the Index Table to
    blow the wafers off the carriers.  We found that one of the wafers got
    crunched a little on the edge of the wafer, breaking a small chunk off
    (maybe about the size of one die).  So, we checked the pad and carriers
    for any damage that may have occurred.  Low and Behold, there was
    silicon fragments on the polish pad and some loose fragments on the
    film of carrier 3 (I think).  The carrier film wasn't damaged., so we
    did an extended rinse cycle to flush away anything left behind.  We did
    however remove the pad as this is common procedure, couple with the pad
    life situation.
    
    The conditioning ring was dismantled, cleaned, rebuilt and reinstalled.
    A new pad was installed, double conditioned and breakin wafers were
    run.  MQC's came in OK.
    
    Some other news, during manually unloading the wafers (by software), we
    ran into a problem with the unload flipper grippers at the unload cup. 
    It wasn't lining up properly to get the wafer.  BJ and I calibrated and
    set the position.  Ran fine.  Upon restarting the automatic running,
    the unload flipper position was off again, this time it was under
    traveling to the unload cup.  Re-calibrated the position and all ran
    fine for the entire day.  Don't ask me why it was off in the manual
    mode!  Seems like the original positions needed to stay put.
    
    Cleaned and lubricated all three scrubber rails.
    
    						Doc  /  BJ  /  Brian C
    
17.68uniformity problemsSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3514
SUBPAC::LANDRY                                       11 lines  22-JUN-1995 19:54
                             -< uniformity stuff >-
--------------------------------------------------------------------------------
    Uniformity problems today.  The carriers have had several weeks of wear
    on them.  So, they have been removed and are currently being rebuilt.
    
    Earlier in the day, BC broke down the conditioner arm, cleaned and then
    rebuilt it.  While that was taking place, BJ and I tore down the gimbal
    assembly, lubricated it and re-installed it.  That portion of the
    Weekly PM has been completed along with the cleaning of all contamin-
    ants within the tubs.
    
    					a bunch of us...........
    
17.69Wafer Breakage-Water Pressure Drop Suspected SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3622
ASDG::POIRIER                                        19 lines  26-JUN-1995 20:25
               -< Wafer Breakage-Water Pressure Drop Suspected >-
--------------------------------------------------------------------------------
    
    
    * Wafer Breakage:
    
      The system broke wafers this morning. A strong vibration was felt
      through out the room just prior to the breakage. We strongly
      suspect a temporary D.I. water pressure drop. A HOTLINE call
      was logged and people in the CUB building responsible for the D.I.
      system were contacted. They have had no indication of fluctuations
      in the D.I. water system.
    
      In a related matter, on Tuesday June, 13th, at 4:50 P.M., the D.I. water
      pressure in CMP went to zero for 10 sec. This seems to be an alarming
      trend. Lets keep a close watch on the water pressure in CMP until we
      can identify and solve the cause of this problem.
    
                                                                   John.
     
       
17.70Unload Flipper/Unload cUps #'s 4 & 5 - Mechanical Interference SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3624
ASDG::POIRIER                                        21 lines  27-JUN-1995 18:58
      -< Unloed Flipper/Unload cUps #'s 4 & 5 - Mechanical Interference >-
--------------------------------------------------------------------------------
    
    * Wafer Pick-up Problem:
    
      The unload flipper was contacting the top of unload cups #4 & #5
      during the wafer unload sequence. The guide pins on these two cups
      had retracted into the bodys of the assemblies causing the cups to
      free-float/rotate. This rotating motion caused the flipper tips
      to contact the top of the cups instead of lowering into the cut-out
      slots.
    
      Roger and I gently tapped the pins back into position and aligned
      them with the guide slots on both cup assemblies. This activity
      required rotating the 2 cups 180 degrees. We tested the load cup 
      up/down motion by actuating the air cylinders in manual mode. The
      motion was not very smooth or consistant in any of the load cup 
      assemblies. Speedfam has a fix in the works for this problem so
      I guess we'll just have to keep a close watch on the situation
      until the upgrade is completed. 
    
    
                                                        John & Roger.
17.71Low Vacuum at head SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3615
SUBPAC::LANDRY                                       12 lines  29-JUN-1995 19:50
                           -< Low Vacuum at head 1 >-
--------------------------------------------------------------------------------
    
    Today...........
    
    	Carrier # 1 was coming up with intermittent vacuum problems.  The
    wafer was loading fine, but the vacuum at this head was hovering around
    the make/break point of setting.  The head was soaked with water for
    awhile.  This seemed to improve it's performance.  It's possible the
    film needed to be soaked down.  Ran fine from early afternoon >>-->
    this memo. 
    
    						Doc
               
17.72scrubber motor being shipped today SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3711
SUBPAC::LANDRY                                        8 lines   7-JUL-1995 17:56
                    -< scrubber motor being shipped today >-
--------------------------------------------------------------------------------
    The system is currently logged down awaiting a scrubber motor.  It is
    being shipped out this afternoon and will hopefully arrive Saturday
    morning.  The package is being sent to "Packaging Plus", which is
    located down the street.  I'll pick it up in the AM and install it
    that day.
    
    							Doc
    
17.73Scrubber Rotation Problem / Motor Replaced SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3727
ASDG::POIRIER                                        24 lines  10-JUL-1995 20:23
                -< Scrubber Rotation Problem / Motor Replaced >-
--------------------------------------------------------------------------------
    
    * Scrubber Brush Rotation Problem:
    
      The Scrubber Motor arrived this morning and was installed into the
      Scrubber assembly. The brushes performed normally for a few cycles
      then stopped rotating after contact with a wafer on every scrub
      cycle thereafter. We called Speedfam and they recommened the
      following action:
                         
                      1) Check for a good ground on the scrubber motor.
    
                      2) Check the voltage on the motor. It is supplied by
                         by an unregulated power supply so don't be sur-
                         prized by a voltage up to 50 V.
    
                      3) Pull-up the scrubber menu and run one off the test 
                         recipe's. Check for reduced torque sysmtoms at 
                         various rpm settings.
    
                                              
       Bob has begun the above mentioned testing activivties.
    
    
                                                                    John.
17.74scrubber motor problemSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3831
SUBPAC::LANDRY                                       28 lines  12-JUL-1995 19:54
                      -< we're getting much, much closer >-
--------------------------------------------------------------------------------

      A long, busy, educational day..........

The system is still down for a scrubber motor problem.  The Servo-Amplifier
arrived this morning.  This was the recommended next step in finding what
exactly is the problem.  The amplifier was installed and an immediate problem
appeared.  No documentation arrived with the amplifier.  Found some DIP
switches internal to the amp that needed setting.  Duplicated the original
amp DIP switch configuration.  

The Amplifier still wasn't up to speed.  Had to have some more dialog with 
the reps on the loop, current, offset, and gain pots.  Got it tweaked properly
and tyested the results.  I would occasionally work / occasionally fail.  The
motor would work fine and then just cut out sometimes.  The problem then 
shifted to wiring.  With a mighty stroke of luck, the problem was narrowed to
a short section of wiring within a liquid tight flex tube.

Unfortunetly, the wire is connected to a connector that is sealed up real 
tight.  We'll have to splice into the connector for the time being with a short
section of new wire.  On the other end of the flex tube, we'll have to cut in
and use a molex connector to make the junction, as the cable is mixed in with
about 100 other cables, snaking back many, many feet to it's destination.

I've communicated with Speedfam about a replacement in the near future.  They
will take care of it asap.

      						Doc

17.75Scrubber repairedSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3815
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              12 lines  13-JUL-1995 06:25
                             -< Scrubber repaired >-
--------------------------------------------------------------------------------
    
        Well....  the (3) conductor wire was replaced and re-run 
     through all the holes and tight fitting places and into the
     smaller than needed water tight teflon fittings.  It took 
     quite a while and I don't recommend replacing "the work" with
     the new harness unless it breaks.  We ran into some problems 
     with the manual operation of the scrubber.  All and all...
     its in VERIFY and needs 25 wafers through it to verify the 
     scrubber operation and then will need full MQC's.
    
                                                   Bob and BJ
     
17.76Unload Flipper adjusted SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3922
SUBPAC::LANDRY                                       19 lines  13-JUL-1995 20:11
                  -< Unload Flipper adjusted / PM completed >-
--------------------------------------------------------------------------------
    The system bit some wafers early in the shift this morning.  No one was
    able to capture details as to how or why.  The only info was that the
    unload module had an alarm around the time of the mishap.  Of course,
    the unload module is no where near the polish table.
    
    After the mess was cleaned up and the carriers changed out, the system
    ran fine, right through the rest of the day.  MQC's came in good.
    
    There was one minor problem that occurred during the shift.  The unload
    flipper failed to properly pick up a wafer from the unload cup.  The
    unload flipper positions were adjusted at the "flipper to unload cup"
    and the "flipper to scrub".  The wafers are currently picking up
    satisfactory for now.  Mechanical adjustment may be needed in the
    future.
    
    The Weekly PM was performed today while the MQC data was being
    retreived.
    							Doc
    
17.77vacuum pump problemSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3946
SUBPAC::LANDRY                                       43 lines  14-JUL-1995 19:33
                            -< vacuum pump problem >-
--------------------------------------------------------------------------------
      The system is currently down for a vacuum problem.  Many, many items
were checked out today.  Here's the recap:

Arrived today to a "main vacuum low pressure" problem.  So, I first looked at
the vacuum tank gauge for the reading.  It was sitting at 10" Hg, totally
unacceptable.  The switch minimum is set at @ 18" Hg, but it's recommended
we run at a greater number (normally runs about 22 - 24").  From there, the
vacuum pump was checked out for possible slurry/water (as has been the case
in the past).  The pump turned out to be clean as a whistle.  From there, we
proceeded to check for loose fittings/hoses, crimped hose, etc.  All checked 
out good.  

At this point, the vacuum line from the pump was removed and a pressure gauge
was put in-line to read the pump performance.  The pump pulled 27" Hg.  So,
we then rehooked up the pump and tried again, 12" was attained at best at the
vacuum tank gauge.  The pressure gauge was then installed between the vacuum 
tank and the hose junction manifold.  We barely got any response at all, a 
couple inches at best.  So, we felt we had it isolated to the vacuum tank and
the hosing that links the tank to the pump and to the junction manifold.

Check the vacuum tank drain check valve for any debris, slurry, etc that may
have prevented the valve from seating properly.  It was removed, checked and
re-installed.  No change in tank pressure.  Checked the hosing/fittings, all
good.  Really getting rediculous now.

By the way the system responded to vacuum and the almost calibrated reading it
kept attaining, I figured that possibly the <10 PSI blow off valve may not be 
working well, letting some air leak by into the tank.  So, I turned off the
valve and checked for results.  No chnage again.

Now it's looking like the vacuum tank has a leak.  We pumped it down the best
we could, then shut off the vacuum pump breaker and observed the leakback rate.
It held vacuum remarkably well.  That was the last physical check we could 
perform and nothing checked out bad.

Jim T suggested we try hooking up the process vacuum (house) to the system.
With the process vacuum hooked up to all the existing hoses/fittings/vacuum 
tank, etc.; the system attained a pressure of 25+" Hg.  So, we have request
they send us a vacuum pump, which they will.  It could be here in the morning.
Thanks to Earl and Barry for they're assistance in performing some of the 
checks.  The bulkheaded walls got to go.

      					Doc  /  Jim T  /  Earl
17.78vacuum pump check valveSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:3918
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              15 lines  15-JUL-1995 07:41
                        -< vacuum pump checked out ok >-
--------------------------------------------------------------------------------
    
        Upon concluding that the vacuum system on A2 could handle
      B1 as a last resort, we decided to see what we could do about
      checking the pump one more time.  We located a vacuum pump
      from HLO1 to try some trouble shooting.  Upon checking the
      plumbing coming out of the pump, it was discovered that the 
      check valve was clogged with slurry.  We replaced the check 
      and everything came back to normal.  The vacuum tank was back
      to spec.  We ran a boat of wafers and after the last (5) wafers,
      it appears the vacuum went to atmosphere.  I quickly unplugged
      the vaccum pump after the fittings and there was plenty of vacuum.
      After I plugged it back in, the vacuum came back to normal.  
      doc is in now and will be checking things out....
    
                                                BJ and Brian P.
17.79Low Nitrogen Flow / Vacuum pump SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:4013
PATE::GEE                                            10 lines  17-JUL-1995 07:33
                      -< Low Nitrogen Flow / Vacuum pump >-
--------------------------------------------------------------------------------
	Ran 50 wafers using the "transfer" program, no problems. Installed
	a pad and started to break it in. After about 15 wafers an error
	came up "low Nitrogen flow", checked the guage at the system it 
	read 50 psi. this is good. This error forced use to power the system
        down and unload the wafers. 
	When we tried to re-initialize the vacuum pump kicked off, we did
	get the system into Wet mode but still no vacuum. Will pass on
    	to 1st shift. 
		          
        
17.80vacumm pump circuit breaker tripping at boot-up SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:4018
ASDG::POIRIER                                        15 lines  17-JUL-1995 19:26
              -< vacumm pump circuit breaker tripping at boot-up >-
--------------------------------------------------------------------------------
                                   
    
     Vacuum Pump Circuit Breaker Problems:
    
     The vacuum pump breaker was being tripped during every boot-up cycle
     this morning. The problem was the vacuum bleed-off valve adjustment.
     The valve was opened 3/4 turn and the system now boots-up without
     tripping circuit breaker #13.
    
    
     NOTE: The electrical conections at the vacuum pump were in bad shape.
           The copper strands in the wires were frayed and the connectors
           were loose. It appears that the cause was excessive vibration 
           generated by the vacuum pump. The wires were soldered to the
           connectors for improved vibration resistance.
17.81Brush station drive chain SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:407
PATE::GEE                                             4 lines  18-JUL-1995 06:44
                         -< Brush station drive chain >-
--------------------------------------------------------------------------------
    	The lower set of brushes on the scrub station stopped rotating.
    	We disassembled the unit and found the drive chain was broken.
    	Found a drive chain in the Speedfam storage case, we replaced
    	it and reassembled. All set.
17.82wafer breakageSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:4111
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"               8 lines  19-JUL-1995 06:36
                              -< wafer breakage >-
--------------------------------------------------------------------------------
    
    
          Wafer #4 appeared to be the culprit in the 5-head crash.
       The machine was all cleaned up, carriers rebuilt and pad replaced.
       The was only (4) plate avaible so Brian Connoley stayed late to
       get us back on track.  Pad broken-in, machine needs full MQC's.
    
                                                    Bob and BJ   
17.83computer lockupSUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:4121
SUBPAC::LANDRY                                       18 lines  20-JUL-1995 20:17
                              -< computer lockup >-
--------------------------------------------------------------------------------
    Pretty decent day of running, except for one incident.....
    
    The computer locked up.  The condition arm would not recess into it's
    home station, the MHTA didn't recognize it's position, and a poly-
    propylene cover over the MHTA rails broke off.  So, we rebooted the
    software to rtecover fromm the lockup condition.  The cover suffered
    severe damage.  I was able to reconstruct enough of it to re-install on
    the system.  It poses no safety, process, or equipment problems.  It's
    purpose in life is to shield any contaminants from reaching the rails,
    which is doing!
    
    Speedfam has been contacted to ship out the necessary parts to repair
    the problem.  They should arrive by Saturday morning.
    
    Pressure plates should be here in the AM, should we need them.
    
    							Doc
                                                                  
17.84Intermittent Electrical Problems SUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:4128
ASDG::POIRIER                                        25 lines  14-OCT-1995 19:48
                     -< Intermittent Electrical Problems >-
--------------------------------------------------------------------------------
    
    * System Performance:
    
      There have been 3 instances of intermittent electrical problems
      during "C" shift today. Twice during power-up breakers #13 & #6
      tripped. #13 is the vacuum pump and #6 connects to Power Supply
      #4 which supplies power to the INFANOR Motor Controllers for 
      (Head #4 Rotation), (Index Table Rotation), (Head #5 Rotation)
      & (Multihead Transport).  
    
      The 3rd instance was at 7:20 p.m.. The system errored for a Head
      Rotational Failure. We EMO'd the system and attempted to power-up
      and re-initialize. The same Rotational Failure occured. The system
      was EMO'd again, this time the cable harness was wiggled and the 
      power-up/intialization was sucessful. "D" shift, you may want to
      investigate the cable harness further. The Speedfam rep has indicated
      that the cables have given them intermittent problems on the other
      tools in the field.
    
      NOTE: Both Speedfam reps have left the East Coast. Bob will be
      returning Monday evening at 9:00. For support between now and Monday,
      page their replacement, Mark Aubuchont, at 1-800-980-6682. 
    
                                                                       
                                                                     John.
17.2MECHANICAL ACCEPTANCESUBPAC::BJUBINVILLENFD Fire - RescueWed Nov 29 1995 22:533921
SUBPAC::LANDRY                                       30 lines  28-SEP-1994 21:11
                         -< The Marathon Man runs!!!! >-
--------------------------------------------------------------------------------
    *  After a long and arduous installation/startup, we are finally
       running the 1000 wafer mechanical acceptance marathon run.
    
    *  We started at 9:30 this morning.
    
    *  All mechanical components are huming along just fine.
    
    *  We had to pause for several minutes a couple of times today due
       to a lack of wafers and the length of time on the OnTrak scrubber
       program.  The Speedfam was polishing at the table for only 30
       seconds (much shorter than an actual run).  This was necessary
       due to the amount of oxide on each wafer.  We needed to remove
       about 100 angstroms or less to make the wafers last.
    
    	NOTE:  by the way, the OnTrak is also huming along nicely.  It's
               like running a second marathon on this system!
    
    *  The marathon run has seen one "assist" occurrance happen several
       times.  The problem is for low vacuum.  This happens when the
       mapper feature is used.  As the mapper is used, the vacuum tank 
       performs it's dump on water that accumulates in the tank.  As the
       start button is activated to begin a run, the vacuum system isn't 
       quite within operating range.  After waiting 10 seconds or so, the 
       tank vacuum pressure improves and the alarm clears.
    
    *  Other than the above problem, the system is running like a champ!
    
    *  Wafer count as of 9:00PM is at 575 (give or take a few).
    
    						Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.34                           CMP.B1                            34 of 292
ASDG::POIRIER                                         8 lines  29-SEP-1994 00:55
                         -< 1000 wafer marathon run. >-
--------------------------------------------------------------------------------
    
    * 1000 wafer run:
     
      777 completed @ 12:20 A.M.
      Brian Jubinvile is taking over on 3rd shift.
    
                                                                    John. 
      

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.35                           CMP.B1                            35 of 292
SUBPAC::LANDRY                                       33 lines  29-SEP-1994 14:26
                         -< The Marathon is Complete! >-
--------------------------------------------------------------------------------
    
    *  THE MARATHON IS COMPLETE!!!
    
    *  The run funished about 6:00 AM
    
    *  There were a couple of problems that occurred:
    
    	1)  Low Vacuum Alarm, almost every time the wafers were mapped in.
            The tank still need additional time to recover proper pressure.
    	    The fault only required a couple extra seconds to clear.
    
    	2)  The system locked up twice, both induced by operator error.  
    	    Cassettes were trying to be swapped as the system ran, hence 
    	    confusing the computer as to cassette status.
    
    *  After the run was complete, we continued to run the system for a
       couple of hours.  Ran great.
    
    *  Late morning, Larry C and I polished 35 wafers for initial process
       results.  We found slurry pump #1 not flowing.  Found the problem to
       be a pinched hose in the flex track.  I had to unlatch the retaining
       pieces of the track and pull out some excessive hose.  This is the 
       2nd time this has happened since the system has been on sight.  
       Speedfam will be made aware of this.
    
    *  The system is currently in Wet Mode.
    
    *  I'd like to say thanks to all who help out with the marathon. 
       Especially Jesus Roa, Brian Jubinville and Earl Dawson for their
       dedicated participation with the run.  Thanks again.
    
                                                       Doc
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.36                           CMP.B1                            36 of 292
SUBPAC::LANDRY                                       16 lines   4-OCT-1994 17:46
                          -< To regulate or not???? >-
--------------------------------------------------------------------------------
    The softwater flow to the system has been inadequate since the startup
    of the system.  Many times the PVC plumbing was removed, dismantled
    and cleaned.  Each time this happened, many pieces of pvc shavings were
    found in the regulator, severely reducing the flow of water.  Well, it
    was cleaned again today and once the shavings were removed, the flow
    was still not quite acceptable.  Therefore, a decision was made to 
    eliminate the regulator (which had a small 1/4" oriface) and the flow
    gauge.  This was replaced with  1/2" PVC pipe, hence providing more
    flow.  Prior to the modification we were seeing a flow of about 1 1/2
    GPM.  After the mod, we were able to achieve about 4 GPM.  The system
    has internal regulators to adjust the flow accordingly.
    
    Also today, wafers were polished.  The average uniformity was ~6%.  So,
    we're getting better with each attempt.
    
    						Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.37                           CMP.B1                            37 of 292
SUBPAC::LANDRY                                       17 lines   7-OCT-1994 16:09
               -< Tighten your belt boy!  Clean your plate too! >-
--------------------------------------------------------------------------------
    A little late...........
    
    The polish table was experiencing a "jerking" situation upon intial
    rotation of the table, but would quickly smooth out (in about a second
    or two).  It was found that a belt had stretched and was a little
    loose.  The rep when in and made a few adjustments to improve the per-
    formance.  There was a little bit of a learning curve to adjusting this
    belt, but it got done just fine by the time it was over.
    
    The new -2 micron plates for the carriers came in and have been
    installed on the system.  Their respective serial numbers and final
    height adjustments have been recorded in the log book.
    
    Process Engineering personnel are beginning to exercise the system to
    familiarize themselves with parameter adjustments.
    
    							Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.38                           CMP.B1                            38 of 292
PATE::GEE                                             2 lines  14-OCT-1994 11:11
                             -< new carrier heads >-
--------------------------------------------------------------------------------
    The new carrier heads came in and are on the table next to the system.
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.39                           CMP.B1                            39 of 292
SUBPAC::LANDRY                                       25 lines  18-OCT-1994 17:00
                -< I knew this day of re-attaching was coming! >-
--------------------------------------------------------------------------------
    The system update is as such...............
    
    *  We began to re-attach alot of components/covers back onto the
       system.  They are:
    
    	-  All five vacuum sensors for the spindle heads were remounted
           into their respective positions and bolted down.
    
    	-  The &$#%*#$ multihead cover was re-installed.  It had taken us
           several hours to remove it (back a couple of months ago).  Well,
    	   putting it back on was just as much fun.  We had to loosen all
           the surrounding covers prior to the top cover fitting in
           properly.
    
    	-  Four of the five internal arm covers were put back on.  The
           fifth will remain off the system for now until some additional
           checks are performed.
    
    *  Tom Crosby, Speedfam rep extrodinare, is presently changing out the
       plates on the five carriers.  I'm not sure at this time what the 
       minus micron reading is for the new plates.  It will be recorded
       as usual in the on-line log book along with the respective serial
       numbers.
    
    							Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.40                           CMP.B1                            40 of 292
SUBPAC::LANDRY                                        7 lines  27-OCT-1994 13:12
                         -< Whats that on your plate? >-
--------------------------------------------------------------------------------
    We shipped out fifteen (15) carrier pressure plates to Speedfam today.
    They are being returned to have new film put on them.  Yesterday, we
    received in five new filmed plates.  They have not been installed on
    the carriers yet.  
    
    				Doc, BJ, Bob 
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.43                           CMP.B1                            43 of 292
SUBPAC::LANDRY                                       21 lines  10-NOV-1994 12:03
                  -< Here's the "condition" of the "pad"!!! >-
--------------------------------------------------------------------------------
    This morning, we (Brian Bobby, and I) took a look at the problem
    concerning the Wet Mode functionality.  It was noted that the center
    section of the pad was not getting during wet mode operation.  After 
    a quick examination of the plumbing, we decided to just swap the DI
    water line with the "optional" slurry line at the base of the rinse
    tub.  The slurry line is not used at this time.  It is an option
    that may be experimented with in the future.  Anyway, this solved the
    problem.  If it decided that slurry will be used at the conditioner
    arm, we will return the lines to their proper spots.
    
    Also, when we removed the cover from the conditioner arm, we found the
    drive belt to the conditioner head/drive motor was not hooked up.  This
    is not to be confused with slipping off!  We mean that the belt was
    removed and tucked away neatly in the arm cavity.  It is believed that
    Process Eng. and Speedfam reps performed this task due to the drive
    mechanism causing excessive wear to the pad.  It will be left in this 
    state until it is deemed needable (is that a real word?  If not, it 
    is now!).
    
    					Doc  /  BJ  /  Bobby
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.44                           CMP.B1                            44 of 292
ASDG::POIRIER                                         6 lines  10-NOV-1994 23:35
                             -< conditioning pad >-
--------------------------------------------------------------------------------
    
    
    * The pad did not dry out during WET IDLE mode tonight. Looks like 
      swapping the tubing is working out fine. Good idea 1st shift!
    
                                                               John. 

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.45                           CMP.B1                            45 of 292
ASDG::POIRIER                                         5 lines  11-NOV-1994 23:51
                    -< RTC status changed to BGN_EXTERNAL >-
--------------------------------------------------------------------------------
    
    * The status of all CMP tools in RTC was changed to 'BGN_EXTERNAL'due
      to the exhaust shutdown this weekend.
    
                                                                   John. 

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.46                           CMP.B1                            46 of 292
SUBPAC::BJUBINVILLE "Quit crying and play hockey!"    9 lines  15-NOV-1994 16:17
                            -< Passdown for 11/15 >-
--------------------------------------------------------------------------------
    
    
         Doc, Earl and myself rebuilt the carriers
    that were on the machine with the new plates
    with the white film already on them.  Doc
    reinstalled the carriers on the machine and
    the pad was conditioned by Larry.  B1 is 
    currently in MQC for Larry.
                                        BJ

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.47                           CMP.B1                            47 of 292
SUBPAC::BJUBINVILLE "Quit crying and play hockey!"    6 lines  16-NOV-1994 15:16
                            -< Passdown for 11/16 >-
--------------------------------------------------------------------------------
    
         I finished up rebuilding as set of plate and pads
    with the DF-200 film on them.  We will be checking them
    on the system tomorrow.
    
                                             BJ

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.48                           CMP.B1                            48 of 292
SUBPAC::LANDRY                                       19 lines  17-NOV-1994 16:09
                          -< Whats up with the Cup? >-
--------------------------------------------------------------------------------
    The system went down today for "one or more cups not down" fault.
    ALL cups were down put the system didn't acknowledge it.  Using
    the I/O's, it was determined that cup # 5 was the culprit.  The
    "cup down" sensor needed to be adjusted.  Brian and myself performed
    this task.  We throughly tested the results by manually moving the
    cup up/down through the use of the software.  The switch responded
    correctly to the respective positions.
    
    Also, upon arriving at the system, we found one of the process doors
    to be off it's track.  This has happened numerous times since we first
    acquired the system.  We were able to "coax" it back on it's track
    without removing any hardware.  It needs to be mentioned that if the
    doors are not sliding freely, stop moving it and call EE to correct
    the problem.  Moving the doors once they have shifted off track has
    a tendency to generate alot of particles, not to mention excessive wear
    to mechanical components.
    
    						Doc  /  BJ
                             

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.49                           CMP.B1                            49 of 292
PATE::GEE                                             3 lines  17-NOV-1994 23:39
                                    -< TCU >-
--------------------------------------------------------------------------------
    Insulated all the TCU lines and wrapped with aluminum tape.
    
    Bob              

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.50                           CMP.B1                            50 of 292
SUBPAC::BJUBINVILLE "Quit crying and play hockey!"    9 lines  18-NOV-1994 16:54
                            -< Passdown for 11/18 >-
--------------------------------------------------------------------------------
    
    
          A wafer let loose under possibly carrier #1 and
    took the rest of the (4) wafers out to lunch, free of
    charge.  Larry investigated the carriers and really
    found nothing wrong.  B1 will stay down until Monday when
    the carriers are taken care of.
    
    BJ

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.51                           CMP.B1                            51 of 292
SUBPAC::BJUBINVILLE "Quit crying and play hockey!"   11 lines  21-NOV-1994 16:46
                            -< 1st shift passdown >-
--------------------------------------------------------------------------------
    
    
       We rebuilt all (5) carriers and installed a new pad.
    After the pad was conditioned, Larry ran some wafer to 
    check things out.  After (15) wafers, LOAD CUP #5 sensor
    screwed up again.  Doc tweaked it in and we checked it
    by repeatedly sending the cup up and down.  
    
                                              BJ
    
                        

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.52                           CMP.B1                            52 of 292
SUBPAC::LANDRY                                        6 lines  28-NOV-1994 16:19
                           -< NO PCW means NO RUN! >-
--------------------------------------------------------------------------------
    The system was re-powered this morning after the holiday shutdown.  No
    problems were observed.  Late in the morning, fab 6 lost the process
    cooling water which goes to the Neslab TCU.  Therefore, the system will
    remain idle until the PCW problem has been rectified.
    
    							Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.53                           CMP.B1                            53 of 292
SUBPAC::LANDRY                                        7 lines   1-DEC-1994 17:13
                       -< around the rim and........off >-
--------------------------------------------------------------------------------
    Today.....
    
    	The system broke another wafer.  It wasn't totally annihilated,
    	just a small section of the outer rim was broken off.  Not sure
    	if the wafer was a little stressed or the system caused it.
    
    							Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.54                           CMP.B1                            54 of 292
SUBPAC::LANDRY                                       17 lines   6-DEC-1994 09:33
                          -< Help me.....I'm drying! >-
--------------------------------------------------------------------------------
    This morning,
    
    	I walked in to discover that the system was not in it's proper
    state, which is "wet mode" of course.  So, I turned on the screen to
    observe what state it was in........low and behold, it was in a state
    of "computer dump"!  Tried to reboot the system and it came up with a
    "low vacuum" error.  This is somewhat normal as the vacuum tank needs
    to perform a water dump.  So, I gave it ample time to dump but the
    error didn't go away.  It was discovered that the vacuum pump breaker
    had tripped.  Reset the breaker and then rebooted the computer.  Alas,
    the vacuum pump/tank performed as should.  The system was then
    re-initialized and returned to wet mode, where it is properly wetting!
    
    The pump functionality will need to be watched for awhile as it is
    unclear as to how this problem originated (was it the computer?....was
    it the pump?)
                                                    Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.171                          CMP.B1                           171 of 292
ASDG::POIRIER                                        24 lines  25-JUL-1995 19:28
                      -< Carrier Plates / Unload FLipper >-
--------------------------------------------------------------------------------
    
    
    
    
    * Carrier Plates:
    
      The system broke wafers for the 3rd time in 24 hrs. This has depleted
      our stock of "filmed" carrier plates. There are currently no usable
      carrier plates in-house. Seventeen were shipped back to Speedfam
      this afternoon for film application. We're expecting a shipment
      of 10 plates tomorrow morning. The system will remain down until
      the plates arrive and carriers are built. 
    
    
    
    * Unload Flipper:
    
      The Unload Flipper was mis-aligned with the wafers in the Index
      Station Load Cups. This condition was corrected by recalibrating
      Unload Flipper Pick-up location by 25 counts towards the "front"
      direction.
               
    
                                                           

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.172                          CMP.B1                           172 of 292
SUBPAC::LANDRY                                       12 lines  26-JUL-1995 19:56
                                -< 80% running >-
--------------------------------------------------------------------------------
    	We received in 5 pressure plates for a set of carriers.  The powers
    that be have decided to run only four heads today, leaving head # 4
    disabled for the time being.  
    
    	Jesus and I built four carriers from set A to use.  All new
    extention rings were used.  Uniformity was acceptable on all heads,
    although heads three and five were a bit close to the spec limit.  It's
    possible the omission of head four is the cause, as slurry flows past
    the track of least resisstance best.
    
    						Doc
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.174                          CMP.B1                           174 of 292
SUBPAC::LANDRY                                       26 lines  27-JUL-1995 19:34
                       -< a first, the re-use of film! >-
--------------------------------------------------------------------------------
    What a week we're having..............
    
    Arrived to find the system down and out for the count.  A wafer spit
    out at the table again last night.  No, it wasn't head four!  Well,
    with only one new pressure plate in-house, they could go no further
    than to clean up the debris and sit tight.
    
    After the morning meeting, a plan was made to attempt to salvage two
    of the carriers involved in the crash.  The film on them looked ok, but
    the extention rings would need to be replaced.  These two, along with
    the known good plate, were built and installed.  
    
    Prior to running the system, a couple of major alterations were made. 
    Some new pads had come in, so we reverted back to them.  Also, a new
    diamond conditioning ring was installed, incorporating the new style
    mounting plate equiped with groumets.  The new ring eliminated the
    infamous "moose" call feature associated with the old style.
    
    Break-in wafers went well.  MQC's, the first attempt were marginal. 
    One of the heads had 11.*% uniformity and the other two wavered on the
    border line.  The second pass went better with all heads improving a
    few %.  So, we're running three heads.
    
    A shipment of plates should be here tomorrow (late morning hopefully).
    
    						Doc  /  Jesus

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.175                          CMP.B1                           175 of 292
PATE::GEE                                            15 lines  31-JUL-1995 08:18
                        -< Low N2 Flow / Scrub station >-
--------------------------------------------------------------------------------
	Came in tonight and found the system wasn't in "wet mode".
	There was an error "scrub station not home", this error
	must have kicked the system out of wet mode.
	The pad was dry, no idea how long it was out of wet mode.
	Cleaned and lubricated the slide rails for the scrub station
	and put it into wet mode, replaced the pad.
        
	At 6:30 this morning the system errored for "low n2 flow"
	twice, left a message on 2 numbers for Air Product to find
	out if work is being done. George C. will bring it up in 
	with them this morning.

	Earl noticed some vibration during the polish cycle this morning
	just before the N2 Error. We need to bring this up with the feild 
	rep. He will be here this morning.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.176                          CMP.B1                           176 of 292
SUBPAC::LANDRY                                       17 lines   4-AUG-1995 19:35
                           -< whats left for the PM >-
--------------------------------------------------------------------------------
    The system continues to run well since the new conditioning wheel was
    installed last week.  Most of the month has been performed in intervals
    this week.  All that remains is the changing of the peristaltic pump
    lines.  If time permits, please perform the task.  If the system is not
    available, I'll get to it tomorrow sometime.
    
    Please try the clear tygon tubing (that we use on the Strasbaughs) if
    you change it.  Use the manual polish screen to flow slurry to see if
    the rate changes. 
    
    If the lines get changed, a check sheet can be filled out and put in
    the Notes file.  ALL other items on the monthly are complete.
    
    Don't forget to restart the tape as it's limit gets close.
    
    						Doc
                                                              

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.177                          CMP.B1                           177 of 292
PATE::GEE                                             6 lines   9-AUG-1995 07:18
                      -< Broken Wafers / Retainer rings >-
--------------------------------------------------------------------------------
	* While running verification wafers after a scheduled pad and
    	  carrier change, the system broke wafers on the pad. Two Speedfam 
    	  Rep.s were here and saw it happen, also the video was on at the
    	  time. The Rep.s will be back around 8:00 am this morning. Two
    	  carriers have been built ( by them ). 
    	  There is only one more retainer ring left in-house.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.180                          CMP.B1                           180 of 292
SUBPAC::LANDRY                                       37 lines  10-AUG-1995 16:34
                    -< Some runouts done, carrier rebuilt! >-
--------------------------------------------------------------------------------
    After the wafer breakage............
    
    It appears that the wafer breakage problem has been happening at the
    head #4 spot more than all other heads combined, but not exclusively
    there.  Anyway, we performed several checks today to try and determine
    the exact cause.  Here's what transpired:
    
    *  Spindle runout was performed an head #4.  There was some play, but
       not that much to cause a problem.  Two other heads were checked 
       (heads 1 + 3) and they had approximately the same amount of play.
    
    *  The table runout was then done.  No problem there.
    
    *  Next to look at would be the carriers used on head #4.  Four of the 
       last five head 4 breakages happened on set B (carrier 4B).  So, 
       carrier 4B was dismantled and checked out.  Nothing unusual was
       found.  The gimballing effect was smooth and sturdy.  No play was
       evident within the assembly.  So, the carrier was cleaned up and
       rebuilt.  We'll give it a try after the up-grade is complete.
    
    *  The old diamond wheel had been re-installed prior to this last wafer
       breakage problem.  The newer one was scratching wafers abit too
       much.  Speedfam will send us out a new diamond wheel replacement.
    
    *  A couple of recipes had also been changed prior to the breakage.  
       This change is not believed to be related to any of the breakage
       problems.
    
    *  This afternoon, we unpacked 5 huge boxes of up-grade materials and
       brought them to the fab.  They are currently on the cart between A2
       and B1.  Please leave them there, ON THE CART!
    
    *  The 1st wave of up-grades will be starting tomorrow at 8:00 AM.  
       Although, there is a possibillity that some items may start tonight.
       
    			Doc / Jim / Larry / Speedfam Reps
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.182                          CMP.B1                           182 of 292
SUBPAC::TOOKER                                       17 lines  11-AUG-1995 16:10
                          -< phase 1 upgrade status >-
--------------------------------------------------------------------------------

             CMP.B1 Phase 1 Upgrade Status Report for 8/11/95
             ------------------------------------------------

The Phase 1 upgrade is well under way.  The original schedule called for the
work to begin at 8:00 am on 8/11.  Due to the system being down all day on 
8/10/95 because of wafer breakage, SpeedFam was able to get a 7+ hour head
start on the work.  As of 4:00 pm, all Day 1 tasks have been accomplished, 
and they are 1 shift ahead of schedule.  This included all necessary dis-
assembly work.  The 2nd shift team has arrived and work has begun on items
listed for Day 2.

No problems have been encountered so far.  We are looking for continued smooth
sailing.


Jim T.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.184                          CMP.B1                           184 of 292
SUBPAC::TOOKER                                      106 lines  19-AUG-1995 16:08
                     -< CMP.B1 UPGRADE STATUS FOR 8/19/95 >-
--------------------------------------------------------------------------------
   Digital Equipment Corporation	  I N T E R O F F I C E   M E M O



To:  Lou Golato        	            Date:  08/19/95
     Marco DeMartino                From:  Jim Tooker
     Steve Otenti                   Dept:  FAB6 CMP Equipment Engineering
     Jim Zimage                     DTN:   225-7073
     karl Koch                      Loc/Mail Stop:  HLO1-1/T09
     George Cumming                 DEC NET: SUBPAC::TOOKER

CC:  CMP Team


SUBJECT:  CURRENT STATUS OF CMP.B1 UPGRADE AND RECOVERY PLAN


Background Information:
-----------------------

On 8/11/95 a scheduled 6 day major upgrade began on the SpeedFam CMP-V Planar-
izer (CMP.B1).  These upgrades were offered to Digital free of charge as a part
of SpeedFam's ongoing tool reliability improvement program.  This upgrade in-
cluded the following items:

     1)  Installation of a wafer loss detection system to detect the 
         presence of whole or broken wafers on the polishing pad and
         shut down the tool before those wafers can cause damage to
         other wafers and wafer carriers.  The addition of the wafer 
         loss detection system to this tool will have a major impact
         on tool availability.  Just since 5/20/95 we have experienced
         13 down events for broken wafers.  These 13 down events equate
         to 65 broken wafers and ~208 hours of lost production time.
         While the wafer loss detection system does not eliminate wafer
         breakage, it does minimize the number of broken wafers and sub-
         sequent recovery time.  If wafer loss detection had been pre-
         sent during the previously mentioned 13 down events for breakage,
         the impact could have been potentially reduced to 13 broken 
         wafers and ~52 hours of lost production time.

     2)  Relocation and reconfiguration of stepper controllers to prepare
         for eventual upgrade from software to hardware cassette elevator
         limit switches.  This upgrade was combined with #1 because the
         stepper controllers would have been removed anyway.

     3)  Replacement of the optical wafer pre-aligner with a mechanical
         pre-aligner.  Also performed at this time because of tie in to 
         #1, several problems have occurred in the field with optical pre-
         aligners at other SpeedFam customer sites.


Problems encountered to date:
-----------------------------

In the performance of this upgrade, it was necessary for SpeedFam to completely
remove from the system several major components such as the robot, wafer mapper,
13 servo controllers, wafer pre-aligner, frequency inverter, stepper motors, and
lots of cables and wires.  The component removal and re-installation phases of
the upgrade went very well and according to the planned schedule.  Since initial
power-up after component re-installation, the following problems have occurred
and have caused major slips in the planned schedule:

     1)  Three damaged servo controllers. (new controllers have been installed)

     2)  Electrical noise interference causing problems with the flipper 
         box controller. (new shielded ribbon cables have been installed)

     3)  Carrier rotation problems. (a wiring problem has been fixed)

     4)  Damaged DCX controller board modules. (some modules replaced and more
         replacement modules are due in today)

     5)  Cassette elevator operation problems. (new encoders are due in today)

Trouble shooting of the above listed problems, testing of the wafer loss
system, and re-calibration of the various components has caused a three day
slip in the planned schedule to date.  More schedule slips are possible as
not all of the problems have been solved.


Recovery plan:
--------------

To complete the upgrade work and minimize further schedule slips, the follow-
ing recovery plan has been implemented:

     1)  The original four member team from SpeedFam which was on site 
         from 8/11 thru 8/17 has been relieved by a fresh four member team.
         The new team includes SpeedFam's leading engineer on software and
         motion controllers and an electrical engineer.

     2)  SpeedFam management has committed to having their people work in 
         two person shifts around the clock until all problems are identified
         and resolved.

     3)  Once production has resumed on the tool, SpeedFam has committed to
         providing a combination of on site/on call local coverage until
         Operations has cleared away the backlog of lots now sitting in CMP.

     4)  There will be an immediate re-evaluation of the SpeedFam consignment
         parts package to ensure that all critical parts will be stocked on
         site.


If anyone has any questions regarding the upgrade work, please do not hesitate
to contact me.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.185                          CMP.B1                           185 of 292
SUBPAC::TOOKER                                        7 lines  21-AUG-1995 15:36
                   -< PLEASE READ.......CMP.A2 EMO PROBLEM >-
--------------------------------------------------------------------------------
Due to the intermittent EMO problem we are experiencing in CMP.A2, it will be
necessary to activate the side sliding door interlocks until the parts arrive
to repair this problem.  The necesary parts are on order and should arrive on
8/22.  Until that time, please to not disable any system interlocks.

Thanks,
Jim Tooker

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.186                          CMP.B1                           186 of 292
ASDG::POIRIER                                        27 lines  21-AUG-1995 20:25
      -< Invertor Communication Timeout error / Scrubber Travel Problem >-
--------------------------------------------------------------------------------
    
    * "Invertor Communication Timeout" error problem:
    
      The Speedfam rep increaced the rate of Table decceleration. This 
      change was made to ensure that the Table comes to a complete stop
      after the conditioning cycle. It appeared that the Table was still
      rotating slightly at the beginning of the Auto Polish Cycle. If the
      Table has not completly stopped when the Polish Cycle initiates, a
      Invertor Communication Error will result.
    
    
    
    * Scrubber movement problem:
    
      The Scrubber required numerous assists today. Although there appears to
      be adequate lubrication on the rails, the scrubber continually binds
      during its travel. A disassembly, cleaning & inspection of this unit
      should be done asap. A future planned upgrade of this assembly will
      take care of this problem. 
    
                                             
                                                         John & Brendan
      
      
      time allows. 
     
      

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.187                          CMP.B1                           187 of 292
PATE::GEE                                             7 lines  22-AUG-1995 07:35
                               -< Clean station >-
--------------------------------------------------------------------------------
	The clean station wasn't making it to the front position
	so the field Rep. adjusted the air solenoid for forward
	travel 2 1/2 turns. Relubricated the guide rails with
	oil. 
	We are only running with 3 heads do to do to poor uniformity,
	carriers need to be built.


          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.188                          CMP.B1                           188 of 292
ASDG::POIRIER                                        24 lines  22-AUG-1995 19:06
                  -< Pads / Carriers / Scrubber / Uniformity >-
--------------------------------------------------------------------------------
    
    * Carriers:
    
      Rebuilt 5 Carriers on A shift today. We need more plastice washers 
      for the retaining ring screws. We couldn't find them in SGMR, does
      anyone know if we have them in house?
                              
    
    * Uniformity problems:
    
      Ran with 3 spindles all day due to poor uniformity on 2 heads.
    
    
    * Pads:
    
      Still experiencing short pad life. (<375 wafers per pad)
    
    
    
    * Scrubber:
    
      The scrubber continues to limp along. It needs constant lubrication 
      to prevent binding on the rails. This causes constant "scrubber
      timeout" errors/alarms. 

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.190                          CMP.B1                           190 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              19 lines  24-AUG-1995 07:35
                            -< Wafer detect failed >-
--------------------------------------------------------------------------------
      
          Late on 2nd shift the system halted when the second
      set of (5) wafers where ready to touch down.  The error
      on the screen was "INVERTER ERROR" and the head by head
      error of "CARRIER ROTATION TIMEOUT."  At first we suspected
      that the slurry never came on.  We started things over again 
      and it happened again.  It we ran 25 wafers without the 
      wafer dectect on-line and there were no errors.  When we
      put the dectect back on line, it failed.  Tony noticed that
      the detector sensor for head #1 was out of position and may
      of cause a false reading and prevented the machine from 
      starting up.  The sensor was repositioned and we tried things
      out again.  It failed again...  It was decided upon that we
      would run the system without the wafer dectect on-line just to
      give the machine back to production.  When time permits, Tony
      is ready to go after the sensor alignment.
    
    
                                                   Bob/BJ

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.191                          CMP.B1                           191 of 292
STRATA::FINN                                         23 lines  25-AUG-1995 16:26
                            -< B1 "low vac alarm" >-
--------------------------------------------------------------------------------

		   Speedfam "LOW VACUUM ALARM"

Low vacuum alarm  occurred on B1 today. When tech tried to reboot.
Circuit breakers for carriers 4 and 5 tripped.  Reset breaker.  Homed table. 
Vacuum pump was not draining.  Tried several mappings of cassettes.  Top 
half of vacuum tank read 23"HG, when purge N2 turned on no dump occurred, but 
pressure leaked back into vacuum tank top half indicated by rise in vac 
gauge to 5" HG. Intermittently the tank would drain properly, and no rise 
in top vacuum gauge was noticed until N2 was shut off and inline pneumatic 
isolation valve opens, causing momentary increase in pressure (as expected), 
as top and bottom compartments are evacuated.   Conclusion was that 
isolation valve was leaking by, possibly caused by dried slurry or wafer chips. 

Hooked water line to input of vac tank and flushed with water 
while cycling isolation valve.  Replaced water line with N2 line and flushed 
pump.

Ran from 11:30 until 4pm. Has not failed yet for "vac low" but tank is draining 
intermittently again.  DOC has isolation valve coming in tomorrow.

We currently looking "requirements and time" to take A2 off line and hook up 
B1 to A2's pump.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.192                          CMP.B1                           192 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              19 lines  26-AUG-1995 07:49
                                -< Wafer crash >-
--------------------------------------------------------------------------------
    
    
       Good news and bads news this morning....
    
       The good news...   The mega upgrade with the introduction 
       of the wafer loss detect system works great.  An execise
       wafer slid out of head #5 and the machine stopped dead in 
       its tracks.  All of the heads stayed down and the table
       stopped as it is suppose to.  Carriers #1, #2 and #3 are
       fine.  #5 looks ok but the ring got chewed.  Head #4 didn't
       break its wafer but has a trashed ring.  We cleaned up the
       mess, installed a new pad and conditoned it.  It's now in
       wet idle awaiting C shift evaluation....
    
          The camera did pick up the crash but the really isn't a 
      direct view of head #5.  Just hit play and the crash is within
      the few seconds of viewing.
    
                                                Brian x 2

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.193                          CMP.B1                           193 of 292
PATE::GEE                                            13 lines  29-AUG-1995 06:56
                              -< Indexing table >-
--------------------------------------------------------------------------------
	The indexing table ran into all five active pedestals while
	in the up position and became jammed. Had to remove all the 
	pedestal air cylinder assembles to release the indexing table.
	Re-installed the cylinder assembles and re-aligned the pedestals
	and load/unload cups, as well as adjusted the pedestal sensors.
	Replace damaged index table feed thru seals for cups #3 and #4.
	
	There are 2 new air cylinders and pedestals arriving tonight
	along with another field service rep. The system is running 
	fine but these parts will need to be replaced due to damage
	from the crash. 
       
                                       Bob/Doc/Roger

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.194                          CMP.B1                           194 of 292
STRATA::FINN                                         13 lines  29-AUG-1995 18:03
                              -< Load cup sensor >-
--------------------------------------------------------------------------------

			Load Cup Sensor

	Errors on load cup # 5.  Sensor not seeing down position of cup. 
Bob tried to adjust sensor but found (cup actuator/pedistel) bent out of round 
due to crash. Doc and Bob repalced this pedisteland readjusted the sensor.


	Also had "Head 5 Rotation Error".  Reset circuit breaker. aok


						Bob/Doc


          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.195                          CMP.B1                           195 of 292
STRATA::FINN                                         60 lines  29-AUG-1995 18:04
                         -< Manual Tank Dump Procdure >-
--------------------------------------------------------------------------------

Scope:	The following documents the addition of a quarter turn isolation valve
        to simplify the purging of B1's Vacuum tank.  And the recommended 
	procedures and frequency for the purge.


Safety: After insuring there are no wafers on the system. Isolated the 
        electrical power by actuating the systems EMO switch.  Tagout the 
        switch.  Inform others in the area that this system is locked out.
	Attempt to operate the system to insure that power has been removed.

Procedure: 

	Remove side panel containing vacuum pump and pneumatic solenoids.

	Remove the next side panel to access vacuum tank and drain line

	Closed new 1/4 turn isolation valve so the handle is across the 3/8 
        line as compared to inline with it. This valve isolates the vacuum pump 
        from the tank and is located near "R6 Head Ball adj."
	
	Turn the red actuator on Solenoid "SV 28 Vacuum tank dump" 180' 
        clockwise until it stops. 

	Turn red actuator on solenoid "SV 19 Vac tank blow out" Counter 
        clockwise until it stops.
	

	Pressure will start to build inside the tank and can be read on
	the tanks vacuum gauge.  When the gauge reads higher than 0"HG the 
        drain will start to purge. 

	When most of the water has been removed, you may increase the flow
	to through the solenoid by the adj. needle valve under this solenoid 
        and you may read this increase pressure on the associated gauge.

	When the water has been removed and air is running through the tank 
        drain, you must close the  red actuator on solenoid 
        "SV 19 Vac tank blow out" clockwise until it stops. The dot on this 
        solenoid should now be inline with the other solenoids in the manifold.

	Readjust the pressure of the solenoid via the needle valve if it had 
        been adjusted.  Insure that it NEVER EXCEEDS 10 PSI as stated on the 
        solenoid gauge. Severe damage to the tank could resultif the pressure
        as read on  tank gauge exceeds 10 PSI

	Then turn actuator on solenoid "SV 28 Vacuum tank dump" counter 
        clockwise until it stops. The dot on this solenoid should 
        now be inline with the other solenoids in the manifold. 

	Open the green 1/4 turn isolation valve so the handle is inline with 
        the 3/8 hose.

	Replace all covers.

	Return EMO to its operational state. 

	Refer to the maintenance Spec for system power up procedure.



          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.196                          CMP.B1                           196 of 292
ASDG::POIRIER                                        33 lines  29-AUG-1995 18:27
         -< Follow-up to last night's Index Table collision passdown >-
--------------------------------------------------------------------------------
    
    * Indexer Problem - Follow-up from last night's passdown:
    
      The following is a description of the events leading up to 
      the collision between the Index Table and the Send Cup
      Cylinder Assemblies: 
    
    1) While loading wafers from the Load Station to the Index Table
       in Auto Cycle, the Unload Flipper began its sequence without
       a wafer present in the Unload Cup. (SHOULDN'T HAVE HAPPENED)
    
    2) Four of the Send Cup Cylinder Assemblies rose-up while the 
       Index Table was rotating. This caused a collision between
       the Load Cups, which are mounted in the Index Table, and
       the Cylinder Assemblies, which are mounted on the deck just
       below the Index Table. (THE CYLINDER ASSEMBLIES SHOULD
       NOT RISE WHILE THE INDEX TABLE IS ROTATING).
    
    3) Two errors repeatedly appeared on the screen, "All Send Cups
       are up" and "One or more Send Cups is up".   
    
    
    4) Since the Index table never completed its sequence, the Multihead
       and the Conditioner Arm locked-up.
    
    5) The system was then rebooted but the initialization sequence 
       would not complete due the Index Table collision. Also, the
       two error messages mentioned in item #3, would not clear from
       the initialization screen.
    
    
    
                                                  John, Roger & Brian C.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.197                          CMP.B1                           197 of 292
PATE::GEE                                            27 lines  30-AUG-1995 01:53
               -< Air Cylinder replacement / Vacuum tank valve >-
--------------------------------------------------------------------------------
	The Speedfam rep. came in tonight at 8:00 pm and checked the
	system out while it was running. He said that it was functioning
	well enough to run until morning. He brought 4 air cylinders and
	2 pedestals, these are all that were available. We have 1 air 
	cylinder in stockroom #6. These are not the new upgrade type, these
	are the same type we have in the system now.
	
	He will be back in at 8:30 am to discuss what needs to be done,
	and will be here until Friday.

     *  Air Cylinder replacement.

	To replace the 5 air cylinders and fix the water leaks under the
	air cylinder assemblies, it will required 3 to 4 hours. This includes
	removing and re-installing all 5 air cylinder brackets and assembles
	that attach them to the base of the machine.
    
     *  Vacuum Tank valve replacement.

	This has never been done and can possible be a very difficult task,
	he estimates between 4 and 6 hours.

	His name is: Bob Gugerty   Beeper # 1-800-593-5880
	    
	
	


          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.198                          CMP.B1                           198 of 292
PATE::GEE                                            17 lines  30-AUG-1995 05:50
                     -< Vacuum tank purge / scrub station >-
--------------------------------------------------------------------------------
    

      * We have been blowing the water out of the vacuum tank every 
	4 to 5 hours, using the new procedure. We found a large quantity
	of water in the tank each time. The last time the procedure was
	performed was at 6:00am. 

      * Had to clean and lubricate the scrub station guide rails twice
	tonight. The DI water cascades onto the guide rails and keeps
	washing off the lubricant.

      * The Field service rep. recommends we lubricate the load cup
	air cylinder shafts on a weekly bases with the same lubricant
	used for the scrub station guide rails. This will help the cups
	travel up and down smoothly.

						B & D shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.199                          CMP.B1                           199 of 292
SUBPAC::LANDRY                                       28 lines  30-AUG-1995 20:26
                  -< cylinders and no rotating head problems >-
--------------------------------------------------------------------------------
    The ride continues rough.........
    
    The system continues to run on three heads (1, 2, 5).  Anyway, those
    heads were running fine when an alarm appeared for one or more cups
    not down.  Found all cups to be seated, but the I/O module for cup 2
    wasn't being made.  Discovered that the sensor was positioned under the
    pedistal, hence it couldn't see the pedistal position.  Mechanically
    adjusted the sensor and tested numerous times.  All looked functionally
    well.  After several runs, the problem reappeared.  Same results, the 
    sensor under the pedistal.  
    
    We then made a decision to remove the cylinder and install a new one.  
    The new one was setup and tested.  All is well.  
    
    Upon inspection of the old cylinder, the pedistal showed signs of
    slight warpedness.  Also, we'll need to start lubricating the cylinder
    shafts occasionally.
    
    After that, the system was run again and the rotation on head 5 wasn't
    working as it touched down to polish.  This cause the wafer to crash.
    The breaker for that motor was tripped.  This has tripped several times
    since the retrofit.  Speedfam has been notified of this occurance and 
    will investigate.  The mess was cleaned up and D shift is taking over
    from here.  Some carriers will need to be rebuilt.  No word on whether
    any new rings have come in yet.  SGMR needs to be checked for this
    info.
    
                        

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.200                          CMP.B1                           200 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              21 lines  31-AUG-1995 07:04
                       -< Wafer breakage, LOAD problems >-
--------------------------------------------------------------------------------
    
    
       Came in to helping Doc clean up after a wafer crash.
    Worked with rep with some leaks from under the index
    table. Two carriers were rebuilt and a pad was installed.
    Heads #1, #2, #3 and #5 were activated.  After 40 or so 
    wafers, the wafer under #1 slid out and the machine halted.
    The wafer did not break.  I used a questionable ring since
    there none in stock.  Rather than risk a crash, #1 was taken
    off line.  MQC's failed for the first round and Earl is
    going to check it out to see what needs to be done.  The
    pad that was used came from Larry's office.  The other two
    on the Vidmar came from the same lot.
    
        While running the final break-in wafers, the LOAD flipper dropped
    the wafer at the aligner and never knew it.  The wafer sat on the 
    aligner when the next wafer was placed on top of it.  Still no error.
    I had to ABORT the run since the two wafers were now double stacked
    in the index table.
    
                                                 BJ

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.201                          CMP.B1                           201 of 292
SUBPAC::LANDRY                                        9 lines  31-AUG-1995 20:20
                  -< valve replaced........vacuum tank one. >-
--------------------------------------------------------------------------------
    The vacuum tank internal valve was replaced today.  Took about 2 and
    a half hours, first time ever performed in the field.
    
    Most of the quarterly is done.  All the weekly section is done.  All
    but the replacement of the peristaltic lines on the monthly is
    complete.  Just the quarterly section and the slurry pump lines remain.
    
    						Doc
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.202                          CMP.B1                           202 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              10 lines   1-SEP-1995 03:18
                            -< Down for the count >-
--------------------------------------------------------------------------------
    
         A 50 wafer verify was done with SpeedFam Rep, Bob, and
      the vacuum tank looks good.
    
         The Quarterly PM was finished up.
    
         B1 failed the final round of MQC's.  Management's request
      was to log the machine down until the pad show up today.  The
      pad was removed and placed in wed idle.
                                                      D shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.203                          CMP.B1                           203 of 292
SUBPAC::LANDRY                                       26 lines   1-SEP-1995 18:13
                        -< pads, hoses, breakers..... >-
--------------------------------------------------------------------------------
    Couple of things today........
    
    *  New Rodel pads showed up.  In a rush to get them on line, I was
       subject to mule work.  Pads came in at HLO1 dock.  Had to be
       received in at fab 6 dock.  I had to haul 20 boxed pads all the 
       way over to the sub fab, get them received and truck a few up 
       stairs to the fab.  A job not fit for man nor beast.  
    
       Five Speedfam pads were brought into the room and the other five
       went into stock.
    
    *  Upon breakin of the pad, the slurry flow in pump one stopped.  The
       line was pinched at the peristaltic pump head.  Repositioned the
       line and resecured same.
    
    *  Later on, the multihead never picked up the five wafers from the
       load cups.  Not sure what happened here, but it appears that the 
       computer interface went askewed.  It also had no rotation on 
       carriers 4 & 5.  The breaker had blown, again.  This seems to be
       happening daily since the retrofit took place.  The rep has been
       told of this.  He will investigate the issue.  
    
    *  Currently running MQC's.
    
    							Doc
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.204                          CMP.B1                           204 of 292
SUBPAC::LANDRY                                       45 lines   2-SEP-1995 19:33
                      -< Scrubber rotation non-existant >-
--------------------------------------------------------------------------------

      The system is currently down for NO scrubber brush rotation.  

Whats been done so far:

      *  The system was working fine until Jim noticed no brush rotation.

      *  I checked for a broken belt, but the belt(s) and coupling were intact.

      *  Checked the circuit breaker, all was fine.

      *  Checked the motor controller (#18).  Found a connector on it that
         had two wire attachments.  One wire was "holding on" by 1 strand of 
         wire from the sheathing.  The other wire was connected by two strands.
         It's obvious that the same connector was used from the original 
         installation as both wires were spliced about an inch back from
         the connector.  The strands that were loose from the connectors were
         all over the place.  Vibration from the MHTA travel could have
         caused problems.  The pins used in the connector were pretty mangled
         as well.  Checked around and couldn't locate any replacement pins.
         Therefore, I soldered a couple of pins, taped them individually for
         insulation purposes and reconnected to the proper pins.  No success.

      *  Voltage was checked at the scrub brushes motor.  The voltage isn't
         getting to it.  Don't quote me, but I believe we should be seeing
         about 4 VDC when the system is idle and the brushes rotating (not
         spinning at high speed during the clean step, which should read
         about 60 VDC when it's high speed spinning).  

      *  So, we have voltage at the motor controller and lose it before it
         reaches the motor.  The cable in the flex track has got to be bad 
         again.  I began pulling the old/running the new wire through the 
         system.  It is currently snaked through a little more than half way.  
         Not to make any excuses, but one person trying to snake the wire is 
         almost impossible.  I got as far as I could.  Hopefully, Bobby and 
         Roger won't have a tough a time as I.  Sorry to leave the job in this 
         kind of shape.

      *  Once done, this should take care of it.  If not, the Speedfam Hotline
         number is 800-276-9149.  I've have talked with them (Ian) today about
         this situation.  They feel confident that this is the solution too.

      						Doc
      
                 

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.204                          CMP.B1                           204 of 292
SUBPAC::LANDRY                                       45 lines   2-SEP-1995 19:33
                      -< Scrubber rotation non-existant >-
--------------------------------------------------------------------------------

      The system is currently down for NO scrubber brush rotation.  

Whats been done so far:

      *  The system was working fine until Jim noticed no brush rotation.

      *  I checked for a broken belt, but the belt(s) and coupling were intact.

      *  Checked the circuit breaker, all was fine.

      *  Checked the motor controller (#18).  Found a connector on it that
         had two wire attachments.  One wire was "holding on" by 1 strand of 
         wire from the sheathing.  The other wire was connected by two strands.
         It's obvious that the same connector was used from the original 
         installation as both wires were spliced about an inch back from
         the connector.  The strands that were loose from the connectors were
         all over the place.  Vibration from the MHTA travel could have
         caused problems.  The pins used in the connector were pretty mangled
         as well.  Checked around and couldn't locate any replacement pins.
         Therefore, I soldered a couple of pins, taped them individually for
         insulation purposes and reconnected to the proper pins.  No success.

      *  Voltage was checked at the scrub brushes motor.  The voltage isn't
         getting to it.  Don't quote me, but I believe we should be seeing
         about 4 VDC when the system is idle and the brushes rotating (not
         spinning at high speed during the clean step, which should read
         about 60 VDC when it's high speed spinning).  

      *  So, we have voltage at the motor controller and lose it before it
         reaches the motor.  The cable in the flex track has got to be bad 
         again.  I began pulling the old/running the new wire through the 
         system.  It is currently snaked through a little more than half way.  
         Not to make any excuses, but one person trying to snake the wire is 
         almost impossible.  I got as far as I could.  Hopefully, Bobby and 
         Roger won't have a tough a time as I.  Sorry to leave the job in this 
         kind of shape.

      *  Once done, this should take care of it.  If not, the Speedfam Hotline
         number is 800-276-9149.  I've have talked with them (Ian) today about
         this situation.  They feel confident that this is the solution too.

      						Doc
      
                 

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.205                          CMP.B1                           205 of 292
FABSIX::R_POIRIER                                     6 lines   3-SEP-1995 04:19
                          -< Scrubber brush rotation >-
--------------------------------------------------------------------------------
    We found, removed and replaced the section of cable that had the broken
    wire in it. After putting the scrubber back together we tested it and
    scrubber works fine.
    
    						B Shift
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.206                          CMP.B1                           206 of 292
PATE::GEE                                            18 lines   5-SEP-1995 07:25
                       -< Scrub station travel problems >-
--------------------------------------------------------------------------------

      * The scrub station was timing out going in and out everytime.
 	Lubricated 2 inches out from the in position with NSK#2.
       
      * We noticed that when the scrubber return to the home position
	it would bounce out away from the home position and time out,
	also the home sensor is being used as a hard stop. The travel
	speed seemed to be good. To cushion the sensor and limit the
	bounce we cut a very small piece of polishing base pad and placed 
	it were the sensor hits the steel wall. It worked fine the rest
	of the night. Hopefully this temporary fix will last until the
 	scrub station can be permanently repaired.

      * The unload flipper problems that caused a wafer to break on 
	A shift were not observed tonight on B shift. The flipper
	operated with no problems, very smooth.
	
					B shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.207                          CMP.B1                           207 of 292
SUBPAC::TOOKER                                       12 lines   5-SEP-1995 13:06
--------------------------------------------------------------------------------
SpeedFam now has a 24 hour technical support line.  The number is 800-276-9149.
This number must be called by anyone seeking technical support no matter what
time of day and for ordering parts after hours.  SpeedFam personnel will be
available from 11:00 AM to 8:00 PM to take the calls.  After hours, weekends,
and holidays the call will be picked up by an answering service and they will
notify the SpeedFam person currently on call.  Do no contact any SpeedFam
people (such as Mike Slepikas) at their direct lines for technical assistance.
Many times these people are away from desks or traveling and will not get your
call.  This could slow down the whole process and cause longer down situations.

Thanks,
Jim

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.208                          CMP.B1                           208 of 292
SUBPAC::LANDRY                                       19 lines   5-SEP-1995 19:15
                             -< pads.....carriers >-
--------------------------------------------------------------------------------
    Today.......
    
    *  The polishing pad was changed.
    
    *  5 new rebuilt carriers were installed.
    
    *  The system was caked with slurry at the polish table section.
       We are in need of a new slurry drain feedthrough to attach to the
       slurry containment ring.  A couple have been broken off lately.
       Please do not use these plastic fittings as "grips" to turn 
       the 800 pound table!
    
    *  There is currently a set of 5 carriers that need to be dismantled,
       plates wrapped up and ready to ship out.  Put the used plates  on
       the cart, not the carrier rebuild table as that is reserved for new
       plates.
    
    							Doc
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.209                          CMP.B1                           209 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              10 lines   7-SEP-1995 06:02
                      -< Wafer hooked by Unload Flipper >-
--------------------------------------------------------------------------------
    
    
        One finger on the unload flipper hooked a corner of a the
     wafer after it was lifting from dropping the wafer onto the
     cleaning station.  This caused the wafer to lift up and fall
     between the (4) supports.  When the shuttle came in, it wedge
     the wafer more.  Luckily, the shuttle was manually moved back 
     and the wafer did not break.  Further wafer runs went through
     without any problems.
                                               D shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.210                          CMP.B1                           210 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"               8 lines   8-SEP-1995 00:14
                   -< head rotation and flipper alignments >-
--------------------------------------------------------------------------------
    
    
       Found broken wire from power supply #3 to the #5 motor
     controller and #3 controller which is a likely suspect for
     the rotation and homing problems.
    
       Re-adjusted the unload flipper to pick-up and drop off
     wafers smoother.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.211                          CMP.B1                           211 of 292
PATE::GEE                                             6 lines  10-SEP-1995 06:18
                               -< slurry lines >-
--------------------------------------------------------------------------------
    Had to replace a broken slurry line at the peristaltic pump.
    We replaced both lines and cleaned the system. 
    
    Replaced the pad do to high wafer count 1066 and poor removal rate.
    
    				B Shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.212                          CMP.B1                           212 of 292
SUBPAC::LANDRY                                       13 lines  10-SEP-1995 19:47
                     -< slurry problems......minor stuff >-
--------------------------------------------------------------------------------
    A weird things happened today.  Slurry was constantly flowing and
    not shutting off.  Found the the incoming pressure fromt the loop
    was excessively high (@50 PSI).  The pumps can only handle pressures 
    of 40 PSI, hence in blowing through.  Reduced the pressure at the
    loop ball valve to about 25 PSI.  Please periodically check it out
    tonight.
    
    Alsop, slurry pump #1 has it's hose kinked.  This could have been
    caused by the excessive pressure problem.  Loosened the locking
    collar and stretched the kink out.  Re-locked the collar.
    
    						Doc  /  Brendan
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.213                          CMP.B1                           213 of 292
PATE::GEE                                             8 lines  11-SEP-1995 05:48
                          -< Unload cup air cylinder >-
--------------------------------------------------------------------------------
	The unload flipper only grabbed the wafer with 2 fingers.
	This happened 3  times tonight on unload 2,3, & 4 but 
	we never dropped a wafer. We noticed that the unload cup
	air cylinder is leaking air badly around the plunger shaft.
	We need to order another and replace it as soon as possible,
	this could be the reason for the flipper problem. 

					B shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.214                          CMP.B1                           214 of 292
STRATA::FINN                                          7 lines  11-SEP-1995 16:21
                             -< Two SBA's to dock >-
--------------------------------------------------------------------------------
    
    Brought 13 used speedfam carrier plates to the dock for
    shipment/rebuild. Also brough cart load of misc. speedfam hardware to
    doc for shipment. Spoke to the dock to insure double boxing and
    carefull handling of carrier plates.
    
    						Brendan

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.215                          CMP.B1                           215 of 292
PATE::GEE                                             8 lines  12-SEP-1995 07:24
                          -< Unload Cup air cylinder >-
--------------------------------------------------------------------------------

	We noticed that the unload cup air cylinder is leaking air 
	badly around the plunger shaft, it's still working properly.
    	We need to order a replacement, the Speedfam 800 hot line number
     	was called tonight, but we found that it was out of service, we'll
     	ask "A" shift to contact them for a new one.

					B shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.216                          CMP.B1                           216 of 292
FABSIX::R_POIRIER                                    10 lines  12-SEP-1995 07:44
                              -< Unload Flipper >-
--------------------------------------------------------------------------------

The Unload Flipper was still having a problem picking up wafers.
It would either grab the wafer but lose it when the Unload 
cup went back down or it would only grab the wafers with two
or three of the four fingers. We found that the fingers needed 
adjusting to equalize the pressure on the wafer from all four 
fingers. Made adjustments to the fingers and Flipper is working
much better.
					B Shift


          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.217                          CMP.B1                           217 of 292
SUBPAC::TOOKER                                       16 lines  12-SEP-1995 17:29
                    -< SpeedFam Field Serv. onsite tonight >-
--------------------------------------------------------------------------------
    The SpeedFam field service eng.(Bob Gugerty) will be onsite tonight.
    He should arrive between 7:00 and 8:00 PM.  Please have him check out
    the unload flipper and unload cups for proper adjustment.  Also, he
    will be looking for the scrubber upgrade parts.  These parts are lo-
    cated on the cart behind cmp.b1.  He may want to do some sub-assembly
    of these parts.  If on the outside chance that cmp.a1 and .a2 come back
    alive, Bob wants to take cmp.b1 to perform the scrubber upgrade.  The
    time duration to perform the upgrade is 10 hours, so please check with
    the Ops supervisor to see if it is OK.  Also, please let Bob know that
    the left over parts from the wafer detect upgrade were shipped out on
    Monday.  If there are any questions, please call me at home
    (508-874-0839).
    
    Thanks,
    Jim Tooker
          

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.219                          CMP.B1                           219 of 292
ASDG::POIRIER                                        12 lines  14-SEP-1995 19:15
                     -< Unload Air Cylinder replacement. >-
--------------------------------------------------------------------------------
    
    
     * Unload Position Air Cylinder:
    
       The Unload Air Cylinder, located under the Index Table, was replaced
       this afternoon. The distance from the bottom of the pedestal to the 
       top of the Air Cylinder was set to 1 inch. In addition, the Unload
       Flipper was taught a new position. It is now 25 counts lower into
       the Index Table Cup. Earl has begun the 50 wafer verification run.
    
     
                                                       John, Earl & Jim T. 

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.221                          CMP.B1                           221 of 292
SUBPAC::TOOKER                                       35 lines  18-SEP-1995 09:12
             -< PLEASE READ: SpeedFam Technical Service Bulletin >-
--------------------------------------------------------------------------------
================================================================================
SpeedFam Technical Support Bulletin:

Bulletin #: 9509TSB012

Date: 09/08/95


  Tight Load Cup Causing Wafer Misload and Loss:


  PROBLEM:

  Several wafer misloads have been detected that would have caused a wafer
  crash.  When the cups are raised (with MTHA over the polish table), the 
  outer guide ring on the cups is being pulled down by the tight fit on the
  index table.  When this happens, the wafer is not being properly guided
  into the carrier pocket.  This problem may exist even when all normal
  checks/inspections are made and are not related to arm alignments or cup
  height adjustments.


  FIELD ACTION:

  *  Check the cup position with the MHTA over the polish table and see if
     any resistance is holding the outer guide ring down.  Any cup that does
     not slip smoothly in and out of the index table may have this problem.
     Remove the load cup and check it dimensionally.

  *  The Lower Shaft Guide, P/N AD3170, is causing the problems in the field
     due to the tight fit between the shaft and the hole in the index table.
     Check your inventory and make sure all shafts are to print and the parts
     did not change  over time due to temperature fluctuations.

================================================================================

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.222                          CMP.B1                           222 of 292
SUBPAC::TOOKER                                       65 lines  18-SEP-1995 09:13
             -< PLEASE READ:  SpeedFam Technical Support Bulletin >-
--------------------------------------------------------------------------------
================================================================================
SpeedFam Technical Support Bulletin:

Bulletin #: 9509TSB010

Date: 09/12/95


  Resolver - Analog/Digital Cables Replacement


  PROBLEM:

  Recently, cable problems have been occuring with four cable failures on
  Tool #10 alone during a short interval of 1.5 weeks.

  The affected assembly part numbers are AD1022A1-1-10 and AD1023A1-1
  through AD1023A1-4.  New cables (with new part numbers) have been
  utilized on Tool #08 for four weeks with no problem.


  SYMTOMS:

  Typical symtoms of a broken cable are:

      1.  Carier will not rotate
      2.  Arm will not oscillate
      3.  Carrier time-out error

  You will not necessarily see error lamps illuminate on the Infranor
  amplifier or see DCX errors.  However, if you do see Infranor error lamps
  illuminate or you see DCX errors, this could be a sign of broken cables.


  FIELD ACTION:

  EN #CMP00407 contains the following data which SpeedFam recommends replacing
  all old cables with new cables (with new part numbers):

      1.  CABLE, RESOLVER, FLEX

          NEW PART NUMBER     OLD PART NUMBER
        ---------------------------------------
              950324             AD1022A1-1
              950325             AD1022A1-2
              950326             AD1022A1-3
              950327             AD1022A1-4
              950328             AD1022A1-5
              950329             AD1022A1-6
              950330             AD1022A1-7
              950331             AD1022A1-8
              950332             AD1022A1-9
              950333             AD1022A1-10


     2.  CABLE, A I/O, FLEX

          NEW PART NUMBER     OLD PART NUMBER
        ---------------------------------------
              950334             AD1023A1-1
              950335             AD1023A1-2
              950336             AD1023A1-3
              950337             AD1023A1-4

================================================================================

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.223                          CMP.B1                           223 of 292
STRATA::FINN                                         28 lines  19-SEP-1995 19:36
                          -< scrubber and vac errors >-
--------------------------------------------------------------------------------

Problems:     1. Scrubber brushes not working

	      2. Vacuum Errors "Low Tank Pressure"


	Found brushes would turn "on", when bottom wire (within corrugated tube)
was flexed.  Cut out bad (open) section of wire.  Rewired and reassembled 
scrubber. (Note: ohm'ed out bad section of wire harness and found open was in 
black wire at the apex of the curve). Tested scrubber, but found it to be at a 
slow rpm without much torque.  Traced to inside motor housing.. found motor 
brushes worn and touching commutator intermittently. Cleaned and realigned 
worn brushes. Tested AOK.  Brushed are worn an will not last forever. 
Speedfam is shipping a new motor to us. Thanks to Larry Cedrone for his 
astute assistance. 

	Found vacuum pump speed to be extremely slow.  Used vac gauge to 
test speed at pump. Found it aok. Then tested just before tank. Found poor 
pumping speed.  Worked back, shut off valve was aok. Tested without checkvalve 
and speed was good.  Replaced checkvalve.  Tested by mapping several 
cassettes. aok. Also adjusted needle bleed valve, which was open significantly.

        Ran 24 wafer verification lot.  aok.   Placed system up on equipment 
tracking system.



						Doc/Brendan/Larry

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.224                          CMP.B1                           224 of 292
SUBPAC::LANDRY                                       15 lines  20-SEP-1995 12:36
                             -< plates.....motors >-
--------------------------------------------------------------------------------
    13 pressure plates were received in this week.  Brendan and I have
    brought them to the room and are on the carrier rebuild table.  
    Please do not stack any items on top of these plates, as imprints
    can and will occur.  And it will affect uniformity, big time!
    
    Also received in today was a spare scrubber brush motor.  As you read
    yesterdays note, the motor currently in the system was temporarily
    fixed and is currently working.  We do not expect this motor to last
    to long though.  Should the motor quit on your shift, the spare motor
    is located in the black storage bin (Speedfam's) under the carrier
    rebuild table.  PLEASE, do not discard the defective motor as it will
    need to be returned to Speedfam as a swap.  Label it and leave it in
    the carrier rebuild table draw.  Thank you!
    
    					B & D.......the A shifters

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.225                          CMP.B1                           225 of 292
STRATA::FINN                                         12 lines  20-SEP-1995 20:13
                               -< broken wafer >-
--------------------------------------------------------------------------------

	Speedfam broke a wafer. Larry found all screws loose. Issue was raised 
that small Plastic or nylon washer are not in stock.  Doc ordered 5200 on sid 
sheet with immediate order of 800 two months ago.  Part appears on stock but 
with zero balance.   John/Mike please check on this with the stock room.  

	Pad was changed, carriers were dropped,  as of now Doc and I have 
check and verified the exposure and tightness of 4 of the five carriers.  Dan 
was trying to second source the washers. I believe he has spoken to Mike.

						a shift


          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.226                          CMP.B1                           226 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              10 lines  21-SEP-1995 06:49
                               -< Pad breakage >-
--------------------------------------------------------------------------------
    
     
       After the 2nd batch of (5) wafers of the break-in wafers
     were polishing, the wafer from head #2 slid out and blew away
     taking out #1 as well.  The pad was replaced and carriers #2
     and #1 were rebuilt.  A new pad was put on and brek-in went 
     smooth.  However..  it failed MQC's.  Another pad was installed
     and broken it and we are now waiting for the MQC results.  
    
                                             D shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.227                          CMP.B1                           227 of 292
ASDG::POIRIER                                         9 lines  21-SEP-1995 19:52
                        -< Head #2 Rotational Failure >-
--------------------------------------------------------------------------------
    
    * Head #2 Rotation Failure:
    
      Head #2 stopped rotating during an Auto Cycle today and broke
      a wafer. The system is now runing on 4 heads. Head #2 will remain 
      off line until we can find the cause of the failure.
    
    
                                                                   John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.228                          CMP.B1                           228 of 292
ASDG::POIRIER                                        17 lines  22-SEP-1995 20:05
               -< Vacuum Problems / Invertor & Rotation Errors >-
--------------------------------------------------------------------------------
    
    * Vacuum Problem:
    
      The Vacuum Pump and Check Valve were replaced on the system today.
      The vacuum is now holding at a steady 24 in.
    
    
    
    * Invertor Communication & Spindle Rotation Errors:
    
      The system is failing to initialize due to the Rotation errors
      at the present time. We have jumped-out the wafer loss detection
      system per Speedfams instructions. We've been working with the rep
      via telephone and they will have a rep here sometime tomorrow.
      "D" shift is picking-up the trouble shooting activities.
    
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.229                          CMP.B1                           229 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              20 lines  23-SEP-1995 08:56
                            -< breaker #6 tripped >-
--------------------------------------------------------------------------------
    
    
        Heads #4 and #5 were off-line due to the circuit breaker #6
     tripped at boot-up.  Reset breaker, put #4 and #5 back on line.
     a 100 wafer verify was run and there were now problems.  However...
     Head #1 lost its wafer at segment #3 which is before the polish step. 
     Head #1 was taken off line and there were now problems after that
     with any other head.  B1 flunked MQC's due to particles and bad 
     uniformity.  The pad was changed and broken in.  B1 flunked MQC's
     again for the same reasons.  The pad and brushes are being changed
     as I type.
            
    
         Randy Spears from SpeedFam will be in on Monday to re-wire
     the motor controllers.  I informed him on what was needed as far
     as suppliers (wire, labels.. exc..).  I also relayed to him that
     circuit breaker #6 trips on boot up from time to time and is likely
     tied into the wiring problems.
    
                                                     BJ

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.232                          CMP.B1                           232 of 292
PATE::GEE                                            44 lines  25-SEP-1995 07:35
                      -< Elevator #1 / Arm #3 not homing >-
--------------------------------------------------------------------------------
	* Elevator #1 wouldn't move. 

	  We found a wire on the 24 pin connector had come loose, we
	  soldered it back on. 
	  The elevator would then move down to the home sensor, but
	  would continue to try and travel down, hitting the hard stop.
         	  
		We checked the 2 bottom sensors, they are good. The input
		rack also indicated that they were working properly.
		We swapped the motor controllers with one not being used,
		the elevator would travel down 1 inch smoothly and then
		start to act like it was binding then timeout. We powered
		the system down and checked for binding, it moved very
		smooth manually. We re-installed the old controller.

	* During all of this we had a problem trying to boot the system up.
	  It would say, "Arm #3 not at home position". This wouldn't
	  allow us to initialize the system, the touch screen would lock up.
          DCX board #2 light #4 was on meaning that this was a real error.

	        We tried resetting the computer, no go. Checked the Inverters
		for spindle #3, no error lights. Checked wire connectors
		for spindle #3 and found some loose, tightening these
	 	didn't help. Called the Speedfam hot line and talked to 
		Brian, he asked use to check most everything that was 
		already done. 
		He told use how to get around the screen lockup and get
		into the system functions. We tried to move the arms out
		to the service position, they wouldn't move and it would
		come up with "elevator not homing " error, and lock up.
 		Tried to rotate them, no problems with that. He said
		that until the elevator was fixed we couldn't go any
		further.

	* Elevator status as of now. 

		If you try to home the elevator now, the first time it will
		time out. If you hit any key to continue and tell it to
		home a second time it will home. Feild service should be 
		in today.
	   

		
						B shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.233                          CMP.B1                           233 of 292
FABSIX::B_FINN                                       36 lines  25-SEP-1995 20:23
                                 -< Passdown >-
--------------------------------------------------------------------------------


	Powered system up found tb6 tripped. Reset and re init.
Received error head 3 not homing on power up.  Went into DOS and disabled head 
3. Error moved to head 4 with corresponding dcx error. Went into DOS and 
disabled all heads. Error moved to conditioning arm. REP arrived check all 
limit switches aok, moved all heads via his pc aok. Found corrupted data files 
for servo gains and error limits.  Corrected with backup files he had from last 
months upgrade.  This corrected all init errors except the original elevator 1 
problems (creep and binding).   

Swapped both cables at motor , drift and motor binding problem swapped to 
elevator 2 with cables and electronics. Motors seem ok.  

Swapped cables at DCX no change in problem

Swapped DCX and amp cables and motor problem remained on elevator 1 

Exchanged amps no change.

Lowered value of current reference resistor on pre amp to 8.2k no change in 
problem.

Swapped motor cable from elevator 1 to 2 and at amp.  motor drift problem moves 
to elevator 2.  the drift problem seems to be in this cable.

Re- verified that moving both cables from 1 to 2 causes binding problem to move 
from motor 1 to motor 2.  It does!

There does not seem to be a mechanical bind.

Swapped DCX modules no change

Need to check out "encoder" cable for motor 1 all the way from dcx to motor.  
 
						A shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.234                          CMP.B1                           234 of 292
PATE::GEE                                            52 lines  26-SEP-1995 07:25
                                -< elevator #1 >-
--------------------------------------------------------------------------------
       * Continuing on with the Elevator #1 problem:

	The field Rep. (Randy Smith) swapped the elevator #1 ribbon cable
 	at the DCX board with the one from elevator #2. Swapped both
	the motor and encoder cables at the elevator #1 with the ones
	from elevator #2. One molex connector had to be repaired.
	
	Ran elevator #1 on DCX elevator #1, no hard stop or drift.
 
	Swapped limit sensors, but didn't need to swap home sensors, 
	those go to the IO's only. The elevator homed and ran fine.

	Removed the old elevator #1 cabling, it needs to be inspected.

      * Tried to initialize the system and found carrier #1 wouldn't move.

	Found the ribbon connector pulled off of DCX board #2, re-seated.

      * Adjusted the lost wafer detection, it was to sensitive.
	
	He left at 12:00pm. He would like to be called early morning if 
        there are any problems with the systems operation. Otherwise, he'll 
	be here mid-morning.  
   

			Pager # 1-800-823-6094
			Hotel # 1-508-634-9500
     
      * After he left 5 carriers and a pad were installed on to the system.
	When we tried to start the break-in wafers it wouldn't start, and
	came up with an error "elevator #2 disabled". We enabled the elevator
	then the system started. Randy was contacted about this.

      * We completed the break-in wafers and are now running mqc's.

	Things the rep. needs to look at:
        ---------------------------------

      * Found that when the transfer paddle traveled up, the scrub station 
	hadn't moved out of the way completely and the edge of the paddle 
	would nick the bottom front of the scrub station. Turned the air
	pressure down to slow the up travel. I believe the scrub station
 	home sensor needs to be replaced. Randy was contacted about this.


      * Slight problem with wafers going into cassette #1. 	

						B shift





          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.235                          CMP.B1                           235 of 292
SUBPAC::LANDRY                                       45 lines  26-SEP-1995 19:33
                      -< a multitude of issues addressed >-
--------------------------------------------------------------------------------

Another rough ride today, but we're currently running product.....

Some of the days events:

*  The scrubber brushes were not rotating this morning.  The rotation
   motor was replaced.  We knew this motor was on it's way out as it
   was overhauled last week for intermittently stalling out.  Speedfam
   had sent a spare to last week.  We were waiting for an opportunity
   to change it.  I have the old motor in my posession and will send it
   back to Speedfam as a swap for a new one at no cost.

*  Head one alarmed today for no rotation.  It is believed that the alarm
   is actually duee to the slipping out of a wafer during polish.  The 
   detection would cause the head rotation, table rotation, etc to shut
   down, preventing all wafers/heads from getting wiped out.  The computer
   would look for the first alarm (which is head 1 rotation) and display
   it on the screen.  We will have to monitor it to assu5re this in fact
   is tyhe real reason.

*  Still need to check out all motor controller wires/connections as time
   permits.  System availability will be a factor in this issue.

*  Elevator #1 had a failure mid afternoon.  Don't panic!  It wasn't the
   pulsing problem we've been seeing.  It was just a simple re-calibration
   of the elevator home position for the fix.  The wafers were not making
   it into the cassette slots properly, hence hanging up and triggering
   an alarm.  Good training spot for the newer folks to see.

*  The original elevator situation has been moved.  Elevator #2 remains
   off line.  We are expecting the arrival of two cables associated with
   the problem.  Said cables should be here in the morning.

*  The scrubber home position sensor is currently faulty.  It doesn't 
   create a down status though.  We will have one arriving in the morning
   and will install it upon system availability.

*  Jim S and Larry Cedrone tried installing a double stacked pad today.  
   Uniformity was out to lunch, so they removed the pads and installed 
   one pre-stacked pad.  All five carriers were run for the break-in/MQC
   runs.  All five passed said tests.  Heads 4 & 5 were high, but acceptable.


      					the A shifters........
  

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.236                          CMP.B1                           236 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              21 lines  27-SEP-1995 08:41
                     -< Head #2 giving us some problems. >-
--------------------------------------------------------------------------------
    
    
         A wafer slid out of head #2 and took out #1 and #5.
      The mess was cleaned up and a new pad was installed.  
      Carriers #1, #2 and #5 were rebuilt and insatalled on the
      machine.  After 10 wafers, a wafer from #2 once again
      slid out but I made the save...  The wafer and machine
      were ok.  We swapped heads #2 with the #3 to see if it 
      is a carrier issue.  After we got everything started back
      up again, head #4 errored out because it wasn't rotating.
      The breaker was not tripped and it would not come on in
      manual mode.  After was cycled the power to the machine, 
      head #4 started to rotate again.  We started up the break-in
      again and head #2 lost another wafer but didn't break.  This
      rules out a carrier issue.  We took the panels off of head #2
      assembly and checked alot of things out but we came up empty.
      We took #2 off-line so OPS could continue along with break-in
      and to reduce the many numbers of lots in the area. 
    
                                               D shift
                                        

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.237                          CMP.B1                           237 of 292
SUBPAC::LANDRY                                       32 lines  27-SEP-1995 19:42
                             -< rotation problems >-
--------------------------------------------------------------------------------
    Today........
    
    *  Calibrated head two "pickup" position.  It was slightly off
       originally.  This may have contributed to the wafer slip-out
       problems on that head.  Time will tell.
    
    *  Check the vacuum switch set point for head two. It was kicking out
       at 13" Hg.  This is an acceptable situation.  We believe that this 
       would not contribute to the wafer slippage.
    
    *  Checked all five heads for proper wafer detection acknowledgement.
       They are all responding.  Not sure how the system could have had
       three heads damaged the night before.  See below though....
    
    *  Attempted to run all five heads.  Nice try!  Head four had NO
       rotation.  Aborted the run.  Powered down to open some cabinets and 
       then rebooted.  All heads were rotating fine now.  Re-attempted to
       run the system, crash!  Head five wafer slipped out taking head four
       along with it.  The wafer detection system should probably be a bit 
       more sensitive to halt it faster.  Research to be done.
    
    *  The rep has had a DCX board sent out and it should be here in the
       morning.  He feels this may be the problem with the non-rotating
       head issues.  This board sends a direct signal to the motor
       controllers for what currant is desired.
    
    *  Two more Speedfam people will be here in the morning to assist with
       the repairs/problems.  Stay tuned!!!
    
                                               the Modified C shift
    
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.238                          CMP.B1                           238 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              12 lines  28-SEP-1995 09:09
                        -< Elevator #2 cables replaced >-
--------------------------------------------------------------------------------
    
    
         Replaced the ELEVATOR #2 cables. The machine was powered
      up without any problems with the elevators.  Carriers #4 and
      #5 were installed on the machine and they are all now at the
      index table rotating and staying wet.  The PVA brushes for the
      scrubber are still submerged on the ELEVATOR #2 because the
      Service Rep is replacing the HOME sensor for the scrubber.
      All three Reps should be here around 9:30am or so.  Elevator #2
      is still going to be looked at by Randy to verify positioning.
    
                                                       D shift      

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.239                          CMP.B1                           239 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              20 lines  29-SEP-1995 05:42
                 -< Machine running but particles everywhere >-
--------------------------------------------------------------------------------
    
    
         The Reps were here all day and left around ~8:30.
      the found several problems with the machine which they
      fixed or are going to fix.  They removed several cups 
      and rubbed them down with VERY fine sand paper and got
      them to go up and down real smooth.  The LOAD air cylinder
      is leaking and they have a new one coming in this morning.
      They also feel that there is also alot of mis-alignments 
      with the load flippers and cups and that they are going 
      to go through the machine today.  They noticed some kind
      of vibration in HEAD #2 and they are not quite sure what 
      it is.  They took #2 off line for us to use the machine.
      We ran 50 wafers for break-in without any problems.  The
      MQC's were run and it passed on uniformity but failed
      bad on particles, all from the pad.  There were also some 
      scratches showing up on the TRANSFER program.  The PVA
      brushes are going to be replaced as well.
    
                                               D shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.240                          CMP.B1                           240 of 292
ASDG::POIRIER                                        23 lines  29-SEP-1995 20:04
                 -< Invertor Comm Error Prob. - Damaged Wire. >-
--------------------------------------------------------------------------------
    
    * Invertor Communication Timeout Errors:
    
      The Speedfam Reps changed-out the 24v power supply only to find
      it being pulled down by a short. An unlabled wire on the buss 
      was then traced back to the (Multihead "Table" Home Sensor)
      (Digital Input #61). The wire leading to the sensor was badly damaged,
      it appears that it may have got pinched by the Conditioner
      mechanisim. The wire was spliced and soldered. We are of the opinion 
      that this wire was the cause of the Invertor Communication Error problem.
    
      The system is now running a 25 wafer verification run. If sucessful
      MQC's will commence.
    
    
    
    * Load Station Air Cylinder:
    
      A replacement Air Cylinder arrived from Speedfam today. It was
      installed by the Speedfam reps this afternoon.
         
    
                                                                  John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.241                          CMP.B1                           241 of 292
ASDG::POIRIER                                        14 lines  30-SEP-1995 19:57
       -< Pad Cond Down Force Reduced / N2 Pressure Switch Replacement >-
--------------------------------------------------------------------------------
    
    
     Activities today:
    
     A replacement Carrier was built for head #1. Head #2 remained off-line
     due to oscillation errors. The conditioner down force was reduced
     proportionally to the amount that the Strasbaugh conditioners were
     reduced yesterday.(Approx 23%)  We ran 75 wafers with no fly-outs.
     At this point, The tool errored for (Main N2 LOW). We are now in the 
     process of replacing the (Nitrogen OK - DG IN 25)  Pressure Switch.
     This switch is located next to the incomming nitrogen regulator. 
    
    
                                                                     John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.242                          CMP.B1                           242 of 292
PATE::GEE                                            55 lines   1-OCT-1995 07:27
                        -< N2 switch / dropping wafers >-
--------------------------------------------------------------------------------
      * Replaced the n2 sensor and set it per specification.
	The sensor didn't trip the way the specification said it would,
	when we adjusted the pod it didn't respond. We had no
	n2 low errors all night.

      * Ran 24 test wafers and began mqc's when a wafer fell off of 
	carrier #2 before polish and the wafers on #4 and #5 were
	mis-loaded, not seated in the carrier properly. Field 	
	service was here and determined the problem was the alignment 
	at the load cups.

	We performed an alignment between the spindles and the load cups
	using the alignment fixtures. We found #2 and #5 needed to be 
	adjusted.

	We checked the multi-head repeatability, it looked good.
 
	We checked the index table repeatability, it looked good.

	We adjusted the load flipper so it seats the wafer into the 
	load cup better, and checked repeatability. This was check on 
	all load cups.

      * When the wafers were not seated properly there was no error for
	low vacuum. The vacuum switches on heads #1,3,4,5 were re-adjusted
	to trip at 20hg. Head #2 was not adjusted do to it not having
	a carrier, this will need to be done when a carrier is mounted.

	Manually mis-aligned wafers on all 4 carriers to check the 
	functionality of the switches, all 4 worked properly.

      * Ran 40 wafers using the break-in program to verify the repair.

      * At 6:15 am running mqc's using the mqc program a wafer dropped off 
	of carrier #1 and landed on the polish table. The carriers never touch
	the table do to the wafer detection, detected the wafer. This is the 
	second that wafers have fallen off the carriers in the mqc recipe.
	
        Contacted field service (Joe) and he asked use to run it again and 
	watch the vacuum at the spindle. We ran the break-in and then the mqc
	programs. Break-in ran and the mqc lost a wafer again, before polish.
	Vacuum on head #4 was on.

	He wants use to contact process engineering and have them check
	the MQC program.

		Speedfam beeper #'s
				Joe 1-800-936-0984
	 		     Randy  1-800-823-6094

						B shift





          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.243                          CMP.B1                           243 of 292
FABSIX::B_FINN                                       15 lines   1-OCT-1995 19:29
                                 -< passdown >-
--------------------------------------------------------------------------------
	
	Jim and Eric zeroed back pressure in all recipes.  Deleted segment 3 
on recipes (contained table speed at 11 rpm).  Set speed to 23 rpm for all 
recipes.  Ran 24 wafers without a problem.. Noticed scrubber brushes were not 
working when the scrubber indexed out to the scrub position. However, they did 
work at home position and low speed.  You guessed it,  found open in scrubber 
motor wire at apex of curve.  Splice new wire and repaired.

 	Verified scrubber fix and continued to cycle wafer using "test MQC" 
recipe..  At approximately 6:50 pm system through a wafer from head 5 and broke 
it.  Changing pad now

Note:  5 pressure plates are packed and will be shipped out tomorrow on sba
       box is labeled


          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.244                          CMP.B1                           244 of 292
PATE::GEE                                            29 lines   2-OCT-1995 07:01
                     -< Slurry line / Verification wafers >-
--------------------------------------------------------------------------------
	Three Speedfam field Rep's came in tonight. These are things
	they checked.

	      * They checked some alignments for drifting.

	      *	Slurry flow and PH.
	
	      * Voltage going to the load cell amplifier and load cells
		for each carrier.

	No problems were found. They wanted to run wafer but we hadn't
	finish rebuilding the carriers. We were asked to run 100 wafers 
	using the "break-in" program when the carriers were completed.
	
      * We ran 80 wafers when an error appeared " waiting for one or more
	cups down." Checked the LED's and found #1 load cup wasn't being
	detected down, but infact it was down. I adjusted the down position
	sensor slightly and the LED came on.
	After that the touch screen froze so we had to reboot. 
	
      * Found that a slurry line had burst, replaced both slurry lines.

	Still need to run 20 more wafers using the Breakin program. 
	
					B shift





          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.246                          CMP.B1                           246 of 292
PATE::GEE                                            20 lines   3-OCT-1995 07:25
                            -< 2 wafers on head #5 >-
--------------------------------------------------------------------------------
	Ran product all night with only one error "Wafer not detected
	on carrier". We found no wafer on carrier #4 when the multi-head
	picked up the wafers at the index table and two on head #5.
	
	The run was a touch up run with a monitor wafer setup for
	head #1 and #4 and product on head #3 and #5. This was in the
	middle of a 24 wafer run. We found a monitor seated on carrier #1,
	a product wafer seated on #3, no wafer on #4 and a monitor seated
 	in #5 with a product wafer stuck to the monitor. No visable
	damage to the wafer the die side was facing down.

	Field service called (Joe) and was told about the error. We
	ran some dummy runs with no problem and monitored the system
	for the rest of the night, no other errors.

					B shift





          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.247                          CMP.B1                           247 of 292
FABSIX::B_FINN                                       53 lines   3-OCT-1995 08:52
                              -< tim's passdown >-
--------------------------------------------------------------------------------
    
This is mondays a shift passdown for B1.
    
From:	SCOMAN::TDYER "TIM....CMP/ETCH ENGINEER, X5072, P997  02-Oct-1995 1937"  2-OCT-1995 19:39:53.58
To:	@CMP.DIS
CC:	TDYER
Subj:	CMP passdown

     
 
    +---+---+---+---+---+---+---+
    | d | i | g | i | t | a | l |	  
    +---+---+---+---+---+---+tm-+
        Semiconductor
					I N T E R O F F I C E   M E M O

     To: cmp.dis		               Date: October 2nd, 1995
                                               From: Timothy P. Dyer
                  	                       Dept: Engineering
                          	               Ext:  5072/4515  
     CC:Larry Camiletti                            Node: Scoman::Tdyer
	Jim Tooker		               Lock/Mail Stop: HLO1-1/p-11
         

================================================================================
 SUBJECT: CMP.A1,A2,B1   AND    WSTK.1.2 PASSDOWN
++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++

 B1 - The standard MQC recipe was determined to be corrupted today. Over the
      past several days wafers have been breaking during the MQC recipe but 
      not during the BREAKIN recipe. A new "MQC" recipe has been created.
      ~100 Wafers, including 3 lots, have run with out wafer loss.  The
      floppy drive is not working. The wafers loss sensors were calibrated
      and tested.
      
      DO NOT USE RECIPES 31 - 90 (TUP, QUICKIES,TDP_MD, etc). 
      FOR ANY TOUCH UP PROCESS USE  "T_UPA" - "T_UPL", make your time 
      adjustments to each of these as required per head.
    
      If B1 goes down please page Tony (speedfam rep) at pager 1-800-516-4197.
       or Joe at 1-800-936-0984
                                        
 A1 - Running with 1 IC1000/SUBA-IV stacked ontop of an other SUBA-IV

 A2 - Pad was at ~400 and the system broke a wafer. 

 
 WSTK.1 - Running, hand stacked pads. Press bar is due in tomorrow.

 WSTK.2 - Please plan to shut down WSTK.2 tomorrow at 8:30am, Westech
          field service will be here to put in the arm run away SAFETY upgrade
          and the new software for RPC-II.


          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.248                          CMP.B1                           248 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              12 lines   5-OCT-1995 09:04
                         -< Pad change and PM started >-
--------------------------------------------------------------------------------
    
    
        Started some of the monthly PM items.  No adjustments made to the
      the machine, just the observation items.
       
        The pad was replaced due to bad uniformity.  (24) wafers 
      were used as break-in.  B1 passed uniformity but failed bad on 
      particles.  Another pad was installed but it failed particles
      once more.  Every MQC wafer had a 1"-2" imprint on the side of 
      wafer in only one spot.  C shift is looking into a possible fix 
      as I type.
                                              D shift 

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.249                          CMP.B1                           249 of 292
SUBPAC::LANDRY                                       42 lines   5-OCT-1995 19:23
                         -< Particles, sensors, etc. >-
--------------------------------------------------------------------------------
    A bad, bad day......
    
    *  The scrubber brushes were changed out today to help address the 
       particle issue.
    
    *  The slurry filter was changed out today too.  The removed filter
       was not clogged what so ever (for what it's worth).  Upon re-
       introducing slurry through the filter/system, they was a big air
       bubble that was difficult to move.  We may be having slurry loop
       problems as the same thing happened when A2's filter was replaced.
       Please check on it occasionally.
    
    *  The system was wiped down throughly this morning, both in the polish
       area as well as the index table aea.
    
    *  The carriers were removed and checked for any potential particle
       problems.  There seemed to be evidence of contaminated water within
       the carriers.  This was removed and the carriers were re-installed.
    
    *  Ran 24 break-in wafers once ready to go.  Had a problem with an
       "inverter error".  The table wouldn't rotate.  Traced the problem
       back to the wafer loss detection board.  The sensors for heads 3 + 5
       were alarming, dispite the presence of any "beam" interuption.  The
       sensors were adjusted by Tony and off we were running.
    
    *  Initial MQC's showed high particles were still there.  This is when
       we decided to change the filter and brushes.
    
    *  We attempted to put head #2 back on line but were getting a "no
       rotation" alarm.  The motor controllers, DCX board 1 and wafer 
       detection showed no signs of failure.  Head 2 was the only head
       on-line at the time.  Retried running and got an inverter error
       again.  Yup, head #5's sensor was triggered off again for no
       apparent reason.  
    
    *  Tony has been paged (by Tim Dyer).  Have him get a procedure from 
       Speedfam on the correct way to adjust this.  He appears to be 
       "winging it" to me.  I check a couple of other sensors and they
       seem to be operating differently from one another.  The LED 
       sequences don't repeat from sensor to sensor.
    
    							Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.251                          CMP.B1                           251 of 292
SUBPAC::LANDRY                                       33 lines   6-OCT-1995 19:35
                   -< Wafer detection, bootup problems, etc >-
--------------------------------------------------------------------------------
    Not much improvement today.......
    
    *  Particles still out of spec at the beginning of the day.
    
    *  Upon attempting to run the system, head # 5's wafer detection LED
       lit up again.  The beam was missing at the table.
    
    *  The wafer detection circuitry was taken of line after this 
       reoccurrance.
    
    *  Breakins went fine.  An ILD lot was started and the very first wafer
       run had a mishap with a wafer loss.  Head 3's wafer was the culprit.
       Cleaned up the mess, put on a new pad, new carriers AND THE WAFER
       DETECTION CIRCUITRY.
    
    *  Checked the wafer loss sensors on all heads and adjusted a few (#5
       for one).
    
    *  Upon re-trying to run, the wafer centering pedistal (load station)
       became non-finctional.  The pnuematic cylinder is defective.
    
    *  We had various nuances all day with the system:  Vacuum pump breaker
       a couple of times, flipper pick up problems in the load station,
       the now usual bootup problems, the scrubber brushes not
       excellerating to high speed, scrubber water not flowing, and a
       few more that escape me.
    
    *  Many Speedfam people, many new parts and many new hopes will be
       here in the morning sometime.
    
    *  Many thanks to Earl and Samuel for there help today!
    
    		A on C shifter / Tim D / Larry C

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.252                          CMP.B1                           252 of 292
STRATA::TDYER                                        63 lines   6-OCT-1995 20:11
                                  -< cmp pd >-
--------------------------------------------------------------------------------
    +---+---+---+---+---+---+---+
    | d | i | g | i | t | a | l |	  
    +---+---+---+---+---+---+tm-+
        Semiconductor
					I N T E R O F F I C E   M E M O

     To: cmp.dis		               Date: October 6th, 1995
                                               From: Timothy P. Dyer
                                               Dept: Engineering
                          	               Ext:  5072  
     CC:Matt Vanhanehem                        Node: Scoman::Tdyer
        Jim Tooker                             Lock/Mail Stop: HLO1-1/p-11
        Larry Camilletti
        Eric Bodensiec 
        Frank Krupa

==============================================================================
 SUBJECT: FAB4/FAB6 CMP passdown
++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++

CMP PARTICLES/BREAKAGE TEST: Today we installed a FAB6 triple stack (ic/s4/s4)
    on WSTK.1 in FAB4 to evaluate particles and breakage potential.  
    Before any pads were removed from WSTK.1 we ran a particle check with
    out going to the final pad (simulating a FAB6 condition), particles were
    ~2200 with 300 large. .050" of shim has been removed from the carrier shaft
    and the triple pad installed. Particles on this new pad with NO final
    process was 900 with 55 large. Although this is only two data points, we
    are pretty confident that the PAD is NOT the particle problems in FAB6. 
    Also Frank had the particles from A2 analyzed, the majority was SiO2, 
    silicon dioxide from the slurry (slurry particles). This will be repeated
    on B1 in the morning. As far as the triple stack causing wafer breakage, 
    this tests not yet completed. WSTK.1 will cycle monitors through the night, 
    if no breakage occurs by morning, we will not focus on the pad as a source 
    of PARTICLES or the cause for BREAKAGE.

WSTK.1 - Pro_ENg, cycle monitors through out the night.
 
B.1 - The system is still having a verity of problems, currently the centering
      station is not working. The plan for the night  and tomorrow is to have
      Speedfam go through the entire system THOROUGHLY, correct any hardware
      and software problems, especially problems related to wafer detect (this
      appears to be a catalyst for everything else). There will be NO product
      run on this Polishers until Speedfam has completed there tasks, than
      the system will be hardware accepted and released by Engineering. If
      B1 is able to cycle wafers tonight, go ahead and check MQCs, we still 
      have not resolved the particle problem and we will need particle data.
      The following Speedfam personnel will be here through the weekend, please 
      make sure a representative from Equipment Engineering is with them at all
      times. I will be in in the morning. Tony will work with the system 
      through the night with D shift.
   
  PAUL HOLZAPHEL
  SCOTT RAY. 		
  Toby Jordon
  Allan Tam
  Tony
	
 A1 - Running, when a pad is needed, put down 1 SUBA-IV and than put an
      IC1000/SUBA-IV on top of it.

 A2 - Running, when a pad is needed, put down 1 SUBA-IV and than put an
      IC1000/SUBA-IV on top of it.


          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.253                          CMP.B1                           253 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              31 lines   7-OCT-1995 07:41
                            -< a bunch of bugs... >-
--------------------------------------------------------------------------------
    
    
        Well....  
    
            The water wasn't flowing on the scrubber again.
            Checked everything out only to reboot the machine
            and it works now.
    
            The pre-aligner fingers weren't closing, cycled
            power, all set.
    
            Scrubber brushes weren't turning during the scrub
            cycle.  Lowered speed from 155 to 145, now they spin.
    
            Wafer detect sensors #3 and #5 flaking out, now working.
    
            The unload cup went down but the pedistal didn't go all
            the way down... SNAP... replaced... adjusted... all set...
    
            The multihead motor controller went south at boot-up.
            Swapped motor controller for #5 rotation with the 
            multihead and got that working.  We took head #5 off-line.
            Carrier #5 was removed and placed on the back table and
            is being kept wet.  There isn't any controllers in stock
            so that is being sent out today.
    
            The SpeedFam party train showing up soon....
    
            Can someone say SOFTWARE BUGS?????????
                                                     
                                              D-shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.254                          CMP.B1                           254 of 292
FABSIX::B_FINN                                       28 lines   7-OCT-1995 19:41
                                 -< passdown >-
--------------------------------------------------------------------------------


Changed slurry pump lines
Cleaned and lubricated scrubber
Cleaned and lubricated paddle rails
Cleaned drip pan
Completed weekly and monthly

Reps;

	Replaced floppy drive

	Adjusted CDA to all cups so they don't slam on accent.
	Cup three sensor is not working and needs to be replaced, Checking 
        stockroom. If not Reps will rush order. Don't run this cup single test.
        until sensor is replaced.

	Replaced bad motor controller on head 5

	Replacing pneumatic solenoid for prealigner.  Sticking has recently
        been noticed at other customers as well. Reason for premature failure 
        is being evaluated.

	Changed Pad conditioner.

	Software upgrade is planed for tonight.

					Brad/Brendan

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.255                          CMP.B1                           255 of 292
PATE::GEE                                            18 lines   8-OCT-1995 06:54
                   -< Robot / Index Table / Vacuum Sensors >-
--------------------------------------------------------------------------------
	Four Field Service Rep.'s Activities tonight :

	* Re-taught the robot positions. Replaced a blown fuse 
	  inside the robot power supply (2.5amp 250v).
	* Adjusted the index table home position, the home 
	  sensor was loose.
	* Checked the spindle to load cup alignments.	
	* Installed carriers and pad.	
	* Checked the scrub sensors, front and back are working.
	* Tried to run wafers, but the system errored for no vacuum 
	  found on all heads. Adjusted the vacuum sensors, this was done
	  last week for not detecting lose of vacuum.
	
                        		B shift
  




          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.256                          CMP.B1                           256 of 292
FABSIX::B_FINN                                       35 lines   8-OCT-1995 19:39
                                 -< passdown >-
--------------------------------------------------------------------------------
	Replaced cup #3 sensor (wire had been torn out by cup)

	Ran test mqc, it placed 2 wafers in carrier 5 and 1 in head 3, and 2 
        wafers on table. ??? not under stood how? Could not get it to 
        repeat this feat.

	Vacuum pump tripped at least 3 times at bootup during the day.

	Ran 75 wafers run in breakin program aok

	Tried passivated wafers 5 wafer 
	results:

		non uniformity 15-22% particles were 5k avg.
	        edge fast removal.

	Changed Bin 1-3 and breakin programs to 120 sec

	Changed pad suba iv and ic 1000 (old pad adhesive was really 
        accumulated on table, needed razors iso and scotch-brite to 
        remove it all)
	
        Head four had carrier down error, adjusted sensor

	Currently running 5 wafers in bin 1, 5 wafers in bin 2, and 5 wafers 
        in bin 3, then will switch to breakin for last 10 wafers.

	There are 5 pre measured wafers, preceded by five breakin wafers in 
        cassette on table left of speedfam.  Data is on A4 and all particles 
        are under 10.

					a shift




          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.257                          CMP.B1                           257 of 292
PATE::GEE                                            18 lines   9-OCT-1995 06:52
                 -< Unload Flipper / Field service activities >-
--------------------------------------------------------------------------------
	Field Service Rep.'s Activities tonight :

	* Adjusted the unload flipper at the unload cup.

	* Ran MQC wafers two times, both times they failed for
	  high uniformity 11.0 % to 15.0 % and also overloaded
	  on particles 8000+.

	* Checked some programs for problems, they will talk to 
	  process engineering this morning.

	* Ran 300 exercise wafers using the Breakin program. 

                        		B shift
  




          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.258                          CMP.B1                           258 of 292
FABSIX::B_FINN                                       46 lines   9-OCT-1995 19:27
--------------------------------------------------------------------------------

	Changed direction of carrier clean station nozzles 90',
        at Jim Tookers request.

	Lowered clean station rpm at rinse per factory recommendation.

	Ran particle wafer for Lab analysis ( came back later in day with 
        aluminum and si02 )

	Checked vacuum sensors setting to determine if it will detect misplaced 
        wafer. It did not. Adjusted all heads for 20"hg trip point. 

	Ran MQC wafers, it fail uniformity and particles.   Removed pad and 
        replaced with ic 1000 prestacked only.

	Uniformity was in on 3 heads, Eric found screw to pressure plate loose.
        tightened and ran MQC again.

	Started to build replacement carriers.  System loaded first 5 from 
        cassette #1 and did not pick up any wafers from cassette #2. I was 
        paged back from the dock at this point and unloaded in Manual.  
        Error out for rotation head 3+4 and numerous vacuum pump circuit 
        breaker trips. Yes we did wait on power up ! Screen locked up on Eric 
        in manual. 

	Frank took slurry samples from all systems including b1. Wafer detect 
        tripped  as well as screen lock up again and later dump from windows 
        to DOS when I acknowledged "waiting for carrier 4 vac".

	Added speed291 back to autoexec.bat file.

        Ran 10 MQC wafers, ran fine waiting for readings	

	If tests at scrubber don't improve particle problems 2 jugs of slurry 
        are at systems door, please hook them in before internal slurry pump 
        to eliminate house sully line as possible cause of contamination.
        Run MQC wafers and check particle levels pre/post run.

	Rep cleaned water from wafer detect 4 sensor.

	Brought 10 pressure plates up from dock and cut sba to send 5 pressure 
        plates and 200 monitor (source inspection) wafers to speedfam.


						Brendan/Tim/Frank


          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.259                          CMP.B1                           259 of 292
PATE::GEE                                            22 lines  10-OCT-1995 06:53
                             -< Flipper Assembly >-
--------------------------------------------------------------------------------
	Field Service Rep.'s Activities tonight :

	* Removed the load / unload flipper assembly, adjusted 
	  the wafer detection sensors. Re-installed and adjusted
	  the load and unload flipper positions.
	
	We removed the pad and cleaned the adhesive off of the table.
	Also, cleaned the table area and conditioner. By the time
	the rep.'s finished with the flipper repair there was no
	time to do the slurry test.
	
	There is 2 gallons of slurry in a green container at the rear 
	of the system and 2 gallon containers filled with slurry also.
	
	The spindles are down in the rinse station rotating with
	rinse on. We built 5 speedfam carriers, there on the cart.

                        		B shift
  




          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.260                          CMP.B1                           260 of 292
SUBPAC::LANDRY                                       42 lines  10-OCT-1995 19:33
                       -< Product is flowing........... >-
--------------------------------------------------------------------------------
We're running product!!!  Zippity do-dah!

This morning, we disconnected the slurry loop from the system. 
We then connected a rung of tubing from a tub of slurry (slurry
from fab 4's loop in it) to the inlet of B1's slurry junction.
A new pad was installed and conditioned prior to running the
new slurry.  Also, we flushed the internal slurry tubing in B1
prior to the expirement.  Once ready, we loaded up 20 wafers.
first 5 ran the B_IN.1, 2nd 5 ran B_IN.3, the 3rd 5 ran BREAK_IN
and the last 5 ran MQC.  The system ran fine for the expirement.
Post readings showed that particle readings were still high.  
Our Slurry loop has been ruled out as the source of the particles.

After this, Larry Camilletti edited the MQC recipe (not sure what
he actually called it at this time).  Runs were made again, and
particle and uniformity was checked.  4 heads stayed under 8%
uniformity with the 5th at 9%.  Particles were somewhat high, but
acceptable.  Product is running.
    
During the day, a couple of occurrances happened:

      -  The Multihead timed out for some reason.  The head was in
         between the tables when it stalled.  An assist was performed
         to move it along.  No reason was found for this problem.
         No failures since.

      -  The bottom section of the scrubber had a chain break.  Non
         in stock.  Speedfam is sending one out for tomorrow.  When
         assemblying, make sure it is routed such that the brushes
         all rotate in the same direction.  Process Eng blesses the use
         of the scrubber until the new chain arrives.

      -  Five more pressure plates were sent out fr re-filming.

      -  A shortage request was sent today to Annie Hargraves for
         a multitude of shortages (shims, brushes, chains, etc).

      -  See the mail message from Frank Krupa/Robin Sacco about the
         metrology equipment thats down.

      						A shifters
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.261                          CMP.B1                           261 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              18 lines  11-OCT-1995 07:30
                          -< Wafer handling problems >-
--------------------------------------------------------------------------------
    
    
       Things started off somewhat ok at first.   The unload
    flipper started giving us some problems with handling.  It
    would intermittingly grab only two corners of the wafer and
    drop the wafer in the scrub area.  Tony checked out the flipper
    and load cups and re-seated the cups and the problem went away.
    OPS were questioning the uniformity so MQC's were started.  Head
    #2 stopped dead in its tracks and failed to ocillate.  Tony took 
    #2 off-line and MQC's were ran.  The uniformity was boarder line
    but ok to run.  About 20 minutes ago, the robot failed to pick
    wafers up at the right LOAD cassette.  After the third slot, it
    started picking up the wafers.  OPS were not comfortable with 
    thses problems since it was a hot lot on the machine so they
    logged it down.  SpeedFam was called and they will be here in 20
    minutes.
     
                                            D shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.262                          CMP.B1                           262 of 292
STRATA::TDYER                                        43 lines  11-OCT-1995 16:31
                               -< cmp pd 10/11 >-
--------------------------------------------------------------------------------
    +---+---+---+---+---+---+---+
    | d | i | g | i | t | a | l |	  
    +---+---+---+---+---+---+tm-+
        Semiconductor
					I N T E R O F F I C E   M E M O

     To: CMP.dis		               Date: October 11th, 1995
                                               From: Timothy P. Dyer
                                               Dept: Engineering
                          	               Ext:  5072  
     CC: Matt Vanhanehem                       Node: Scoman::Tdyer
         Jim Tooker                            Lock/Mail Stop: HLO1-1/p-11
         Larry Camilletti
         Eric Bodensiec 
         Frank Krupa

==============================================================================
 SUBJECT:  FAB4/FAB6 CMP PD
++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++
 
WSTK.2 - Reported RPC errors (again) to Westech. Waiting on their next move...
         Continue to contact EEG techs when RPC-II locks up, eeg please record
         exact errors on alarm page.

B.1 - System has been running MOST of the day with head 2 off line and load
      cassette 2 off line. Head 2 will not be looked at until a new infinor
      arrives tomorrow, then we will set up time with MFG. The scrubber belt
      arrived today, the reps installed and calibrated it in between lots.
      Speedfam will have 24 hour support until the WIP BUBBLE is clear and 
      when we have some confidence in system performance. If the reps are not 
      in the room when a problem occurs, page them ASAP. If Digital EEG 
      technicians are able to recover the system, do so, but page Speedfam as 
      well to review the failure.

      (6pm-6am coverage)   Tony - 800 516 4197
      (6am-6pm coverage)   Bob -  800 593 5880

A.1 - Running, use  a IC1000/SUBA-IV over another SUBA-IV.

A.2 - Running, use  a IC1000/SUBA-IV over another SUBA-IV.

* Sent request for more 28" and 32" ic1000/suba-iv and suba-iv to materials.


          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.263                          CMP.B1                           263 of 292
ASDG::POIRIER                                        29 lines  11-OCT-1995 19:49
     -< Head #2 Oscil Failure / Right Load Sta Pick-up Prob/ Scrubber Be >-
--------------------------------------------------------------------------------
                                                   
    * Head #2 Oscillation Failure:
    
      Head #2 has been taken off-line do to an oscillation failure error.
      The Speedfam reps have ordered a replacement Motor Controller to
      address this problem. We expect it to arrive tomorrow. We're
      currently running with 4 heads.
    
    
    
    * Right Load Station Handling Problems:
    
      The Robot is having problems removing wafers from slots #'s 1 & 2
      on the Right Load Station. The Left Load Station is not experiencing
      this problem. Tim Dyer has aquired an ADE corp diskette and manuals that
      will allow us to troubleshoot the Robot and ,if necessary, teach it
      new locations for slots 1 & 2 on the Right Load Station. The file has
      been installed on the hard drive in the Dell system in the CMP
      service Chase. The Manuals are in the Speedfam tool box.
    
    
    
    * Scrubber Belt:
    
      The lower brush drive belt in the scrubber has been replaced. The
      belt was routed such that the brushes now rotate in the proper 8 inch
      configuration. (per Speedfam's recomendation)
    
       

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.264                          CMP.B1                           264 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              20 lines  12-OCT-1995 07:47
                          -< more cylinder probs... >-
--------------------------------------------------------------------------------
    
    
        here we go again...
    
           Cylinder #4 failed to go down all of the way.  The
       pedistal got bent but repairable.  Actually, the cylinder
       assembly was swapped with #2 until a new one can be sent
       in.  The sensor for #2 was detecting that the cup was not
       down even though it was off line (It still looks for it for
       safety reasons). Tony tried to adjust the sensor it ended
       up breaking off.  That will be replaced today.
    
          Some software bugs...  the touch screen froze up several
       times as well as the scrubber brushes failed to rotate.
    
           Gwen did MQC's and it failed bad for scratches.  The
      system is now DOWN and the pad is being replaced as well as
      a massive cleaning to the entire system.
    
                                                  D shift 

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.265                          CMP.B1                           265 of 292
ASDG::POIRIER                                         8 lines  12-OCT-1995 19:37
                -< Neslab TCU -  Solid State Relay Burnt-out. >-
--------------------------------------------------------------------------------
    
    Neslab TCU Problems:
    
    The TCU failed to power-up at 5:00 p.m., closer inspection revealed
    a burnt-out SS relay. The TCU intended for B2 is being swapped with
    B1's unit. We expect the work to be completed on "D" shift tonight.
    
                                                                  John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.266                          CMP.B1                           266 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              12 lines  13-OCT-1995 05:15
                         -< still failing uniformity >-
--------------------------------------------------------------------------------
    
        The chiller was turned on around 3am without any
    problems.  B1 had failed MQC's for uniformity twice prior
    to this.  We are working on the 3rd pad on D shift as of
    5am.   Tony replaced the Infranor on head #2 as well as 
    the data board for #2.  He wasn't able to continue as we
    took the machine back.  The right send elevator was 
    re-taught and tested without any problems.  Please keep
    an eye on it.
                                  D shift
    
      

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.267                          CMP.B1                           267 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"               9 lines  14-OCT-1995 05:40
                        -< Vacuum pump breaker tripped >-
--------------------------------------------------------------------------------
    
    
        The vacuum pump circuit breaker tripped for an unknown
      reason.  Nothing broke but we had to run to unload cycles 
      to start things over.  The breaker has tripped in the past
      with the old vacuum pump.  A slightly larger breaker might
      be considered.
    
                                          D shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.269                          CMP.B1                           269 of 292
SUBPAC::LANDRY                                       36 lines  15-OCT-1995 19:35
                   -< carrier timeout, unload problems.... >-
--------------------------------------------------------------------------------
The system failed this morning for a "carrier timeout".  Well,
this is one both Brendan and I have never seen before.  The MHTA
was over the Index Table, the arms were spread to prepare to
unload the wafers into the cups, the cups were up.....but every-
thing just halted, with the alarm sounding.

Checked the I/O's and all were fine.  The screen was prompting us 
to hit any key to continue, which we did.  The timeout happened
again.  We could not get off the screen or abort the process, so
we had to EMO and reboot.  Upon rebooting, several problems arose.
The unload flipper and the unload cup were the main hitters.  The
flipper wouldn't pickup the wafers consistantly and the unload
cups were the same case.  Upon exam, we found that the down stroke 
of the unload pedistal was slow/sticky.  Adjusting the air only 
helped slightly.  We checked the cups/feedthroughs.  Found eight
of the ten cups out of position to their assigned feedthrough
ports.  Returned all cups to the designated spots and tried again.
A big improvement, but still an occassional sticking situation.
Checked the feedthrough seals.  Found a couple not quite sitting
in their grooved seats properly.  Assured all were properly seated.
Used crocas paper on the cup shafts to "clean up" any dings and/or
buildup on them.  All cups were working sufficiently.  Successfully
unloaded 15 wafers manually to confirm performance.  

Back to the original problem...............
We then tried to repeat the "carrier timeout" problem.  We used the
"transfer" program to exercise the carrier arms.  Ran a couple of
cycles fine.  No problems!?!?  The system ran the rest of the shift 
without this alarm re-appearing.

The ID/OD pad was changed today.  The wafers were edge-light, meaning
the outside of the wafer wasn't polishing enough.  So, a "full" sized
pad was installed.  MQC's looked real good.  Uniformity is 3% to 5%
range.  Still running with head two off line.

      						A shifters

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.270                          CMP.B1                           270 of 292
SUBPAC::LANDRY                                       22 lines  15-OCT-1995 19:47
                             -< Speedfam tidbits >-
--------------------------------------------------------------------------------
    Let me suggest a few things we all should try and do with this system.
    
    	-  When powering down the system for any reason, DO NOT repower the
           system back up for ~ 1 minute or so.  The quick pwoering up is
           believed to be causing many bootup problems (ie. breakers need-
           ing to be reset, scrubber water flow/brush rotation problems,
           rotation problems, etc.).  If given the minute plus delay, the
           system seems to boot up fine.
    
    	-  When the system is powered down for any time greater than 3
    	   minutes, shut the chiller power off.  DO NOT hit the mute button
           as that only shuts up the alarm and the chiller continues to
           dive in temperature.  This no_let contributed to the death of
           the first chiller.
    
    	-  Leave the carriers on all heads when the system is running,
           especially Wet Mode.  With  the carrier removed, the wafer de-
           tection sensors are exposed to potential water mishaps.  When
           the water hits the sensor, it thinks a wafer is detected, not
           a pretty scene.
    
    						Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.271                          CMP.B1                           271 of 292
PATE::GEE                                            29 lines  16-OCT-1995 07:09
                             -< Carrier Timeout  >-
--------------------------------------------------------------------------------
	* The system error for "timeout waiting for carriers".
	  This happened everytime before loading the second 4 wafers
	  onto the carriers using the ID_OD program. It would lock the 
	  touch screen up and we would need to re-boot.

	  The first time it happened we noticed on DCX board #2 the
	  number 3 error light was on "Carrier #2 oscillation error",
	  carrier #2 is off line so no error should occur.
	  The 3 times after that all the led's looked good. Called the
	  field rep, no response. 

	  We decided to remove carrier #2 from the multi-head, it ran 
	  fine without it last night. Changed the program from ID_OD
	  to ID_OD75. At the same time we placed a call to the speedfam
	  hot line. We received a call back from Tony within minutes.

	  The system ran fine after that using the ID_OD75 ID_OD_PD program. 
	  We had the same problem when we tried to use ID_OD240 and
	  ID_OD200 programs. Tony said that there must be something in 
	  those programs telling the system to reference head #2. We 
	  checked the programs, they looked o.k. to us.

					 Hope this makes sense

	                            		B shift
  




          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.272                          CMP.B1                           272 of 292
FABSIX::B_FINN                                       13 lines  16-OCT-1995 19:19
                              -< System hang up >-
--------------------------------------------------------------------------------


	Multihead stopping at index table with arms extended over cups. Error 
read " time out waiting for carriers".  Found no DCX errors or missing I/O's.
Checked them all.  Traced to spindle 3 out overtravel switch.   Found bad 
connection.  Jumpered connection system ran fine.   Removed connector and pins and
hard wired switch,  tested aok.   Returned to Manufacturing.  Thanks to 
Larry Cedrone (operations) for his assistance.


					Brendan



          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.273                          CMP.B1                           273 of 292
PATE::GEE                                             9 lines  17-OCT-1995 08:07
                                  -< Head #5 >-
--------------------------------------------------------------------------------
	* Errored for "Waiting for carriers" over the index table.

 		Had to adjust Head #5 limit switch.

	                            		B shift
  




          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.274                          CMP.B1                           274 of 292
STRATA::TDYER                                        44 lines  17-OCT-1995 18:53
                                    -< pd >-
--------------------------------------------------------------------------------
    +---+---+---+---+---+---+---+
    | d | i | g | i | t | a | l |	  
    +---+---+---+---+---+---+tm-+
        Semiconductor
					I N T E R O F F I C E   M E M O

     To: CMP.dis		               Date: October 17th, 1995
                                               From: Timothy P. Dyer
 					       Dept: Engineering
                          	               Ext:  5072  
     		                               Node: Scoman::Tdyer
                        		       Lock/Mail Stop: HLO1-1/p-11
      Matt Vanhenehem    
      Larry Camiletti
      Eric Bodensieck
      Frank Krupa
                                        
==============================================================================
 SUBJECT:  CMP
++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++

Speedfam coverage - If any problems arise call Bob Guberti  800 -593 5880

WSTK.2 - Unif was poor, after alignment uniformity was still center slow.
       * CHECK TO MAKE SURE THE CAM/ROLLER IS NOT RESTRICTING THE ARM DOWN
         MOVEMENT. (When the arm is down the ROLLER should NOT be touching the
         center post). If the cam is not a problem, a IC1000/suba-lite pad 
         will be tested. If this pad produces center slow polishing, there 
         is an other hardware issue. Please repeat alignments thoroughly.
  
B.1 - Running with ID/OD process with FULL IC1000/SUBA-IV pad.

A.1 - Running breakin with ID/OD process ID=8", OD=20". Erik B. will modify 
      the recipe and conditioing parameters. see below if a pad change is needed see below. 
      
A.2 - Running with ID/OD process, ID=8", OD=20". If a pad change is needed do 
      the following:
 
      1) Install a 22.5" SUBA-IV in the center of the platen
      2) Cut a 22.5" IC1000 down to 20" (use the pen/string in the room)
      3) Cut a 8" hole in the center of the IC1000
      4) Mount and center the IC1000 ontop of the SUBA-IV
      5) Break in and check MQCs.


          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.275                          CMP.B1                           275 of 292
SUBPAC::LANDRY                                       29 lines  17-OCT-1995 19:41
                     -< Hardwire them sensors........... >-
--------------------------------------------------------------------------------
A couple of things occurred today.........

*  Product did flow henceforth from the system most of the day.

*  Carrier timeouts happened again, this time head 5.  They wanted
   to finish of the product lot before they had us address it.  Once
   available, the A shift and Bob Gugerty sprang into action.  We found
   that the Molex connectors were indeed the problem.  Therefore, we 
   decided to totally eliminate the problem with ALL the heads by remov-
   ing the Molex connectors, a straight hardwire job.  After completing
   the task, we exercised the heads numerous times without failure.  We
   can't be assured, for a couple of days, that the problem is totally 
   fixed or not as this problem has been intermittent.  Time will tell.

*  The scrubber stopped rotating today.  The problem, like the carrier
   timeouts, has been intermittent.  The problem has been narrowed down
   to a small section of the electrical circuit.  No, not the normal 
   pivot spot within the flex-sheathing.  This time it is located in
   the general area of the connector at the scrubber head.  It's either
   in the wire or at the pins themselves.  While flexing the wire in that
   vacinity, the scrubber started to work.  Repeated attempts at flexing
   the suspected area to bring back the failure was futile.  Eng/Opts
   requested to take the system back into their control.  The problem 
   WILL be back.  If it happens on your shift, try to pull some wire
   slack back to the scrubber head, cut off 3-4 inches of wire, and
   put on new pins in the connector.  

      						"A"


          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.276                          CMP.B1                           276 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              13 lines  18-OCT-1995 06:25
                      -< carrier errors still with us... >-
--------------------------------------------------------------------------------
    
           The carrier error once again joined us.  It only happens
       when the carriers go out to load on #5.  It didn't error when 
       it came back from the polish table.  Bob and I looked at the 
       switch that was adjusted and tried to move it.  It got worse.  
       We put everything back and now it is just intermitting.  Rich
       is limping it though right now but Bob will have to take a peak
       at it in the morning.
    
           The other problem we were seeing was the conditioner arm
      timing out because the machine didn't know it was home.  Therefore,
      it wouldn't go down.  After (2) reboots, thje machine ran fine
      since.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.277                          CMP.B1                           277 of 292
SUBPAC::LANDRY                                       43 lines  18-OCT-1995 18:47
                 -< pad change, head alignment, slurry hoses >-
--------------------------------------------------------------------------------

Some things that transpired today with the system.........

*  The system received a new pad first thing this morning.  The old
   pad had a >800 wafer total on it.  It was smooth!!!!!!

*  While the pad change took place, we had the opportunity to perform
   the weekly PM.  All items were performed.  A good training spot for
   Brad, who was pro-active with the tasks!  See checksheet for details.

*  The system came up quick from the pad change with excellent results
   using an entire pad (hold the donut)!

*  During a short recess, JP.....Bob Guug's and I checked into the carrier
   timeout issue that intermittently exists.  We checked out the head to
   load cup alignment for head 5 (the current problem child).  It was
   discovered that the "out" direction travel to the load cup wasn't quite
   going far enough.  We performed head 5 calibration.  Ran the head in
   and out several times to assure repeatabillity.  Just like with the 
   hardwiring, only time will tell if the timeout still exists.

*  With the recess time ticking away, we moved on to head two performance
   problem.  Using DOS, we attempted to move the arm out.  No response!
   DCX board 2 lit the fault LED for head two travel.  Reset and tried
   again.  Same results.  No voltage command was evident between the DCX
   board and the motor controller.  This equates to the problem existing
   from the DCX board to the motor controller.  The wires in this loop
   remain stationary, they don not go through the flex-track.  I'm finding
   it hard to believe that the cabling has a break in it.  We'd like to
   look at the DCX board connector next and possibly a DCX module (or
   daughter board as some have called it) swap.  The DCX modules are the
   small boards mounted to the DCX main board, similar to the DCX modules
   we've changed out on the Strasbaugh's IBM computer.  

*  By the way, the system did not exhibit any carrier timeouts on C shift
   today.  Neither was there a problem with the conditioner up/down
   movement.

*  Late afternoon, one of the slurry hoses on the system sprung a small
   leak.  It was only leaking as the pump was used.  The bottom pump
   was the culprit.  Anyway, both hoses were changed out at this time.

      					Shift C and Me

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.278                          CMP.B1                           278 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"               9 lines  19-OCT-1995 06:28
                     -< carrier error, scrubber stopping >-
--------------------------------------------------------------------------------
    
    
             B1 ran great all night/morning without too many problems.
         The carrier error popped up again.  I gave head #5 a tug and
         started running again.  The scrubbers weren't turning during
         the scrub cycle.  After checking out the connector and belts,
         the scrubber started working again.
    
                                               D shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.279                          CMP.B1                           279 of 292
ASDG::POIRIER                                        16 lines  19-OCT-1995 19:39
                     -< Errors / New Invertor Panel Door. >-
--------------------------------------------------------------------------------
    
    * Errors:
    
      1 "Carrier Time Out" error occured during the shift. The system will
      be going down for a cable upgrade beginning tomorrow morning at 6:00 a.m.
    
    
    
      Electrical Panel Door:
    
      The door on the Invertor Panel has been changed to an upgraded type 
      with louvers for added cooling.
    
                                                               
                                                          John & Brad.  
      

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.280                          CMP.B1                           280 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"               8 lines  20-OCT-1995 08:46
                        -< Cable swapping has begun... >-
--------------------------------------------------------------------------------
    
        The carriers were removed from the tool and place in water
      on the table next to A1.  The brushes were placed in wet sink.
    
        All (4) reps showed up at 6am and have started the replacement
      of the cables.
    
                                                  D shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.281                          CMP.B1                           281 of 292
ASDG::POIRIER                                         8 lines  20-OCT-1995 19:37
                    -< Cable Replacement Activity - Update >-
--------------------------------------------------------------------------------
    
    * Cable Upgrade Activity:
    
      The replacement of the cables has been completed and the system has
      been powered-up. The Speedfam reps are currently checking their
      connections and troubleshooting a problem with head #3 not moving.
    
                                                                John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.282                          CMP.B1                           282 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              13 lines  21-OCT-1995 06:44
--------------------------------------------------------------------------------
          
           The REPS found they connected up (5) cables in the
       wrong connectors.  This was discovered after they had 
       all sorts of head problems.  Nothing blew up however...
       The cables were re-connected to their correct locations
       and they fired up the machine.  They are now getting a
       "table not responding error" at boot-up.  They didn't
       touch either table so they are troubleshooting this
       now.  Head #2 appears to fixed now.  They founbd a broken 
       wire at the terminal strip area in the multihead.  They
       will be back hear in the a.m. to continue the upgrade.
    
                                            BJ

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.283                          CMP.B1                           283 of 292
ASDG::POIRIER                                        13 lines  21-OCT-1995 19:52
                       -< Vendor Cable Upgrade - Update >-
--------------------------------------------------------------------------------
    
    * Cable Upgrade:
    
      The Speedfam reps verified proper functionality of the rotation and
      oscilation on all 5 heads. They ran into a problem with the Polishing 
      Table rotation this morning. They have identified the problem as a
      bad board in the Table Invertor. They have ordered the part and
      expect it to arrive at Logan on a 7:00 a.m. flight tomorrow. They
      are planning to pick-up the part at 7:00 and arrive back at HLO
      by 10:00 a.m to install the board and continue trouble shooting the
      system.
    
                                                          John & Brad.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.284                          CMP.B1                           284 of 292
FABSIX::B_FINN                                       17 lines  23-OCT-1995 19:21
                                 -< passdown >-
--------------------------------------------------------------------------------


	Reps replaced RS232 -  RS485 converter module. Table ran fine.
	Changed pad. (see entry in EST).  Installed carriers and attempted 
	breaking runs. Found particle (rust in conditioner bath)  disassembled 
	conditioner and bearing and bearing seal, cleaned rust out and 
        re lubed.  Reassembled.

	Started running the remaining breaking wafers and load cup 2 down sensor
        errored. Rep is moving sensor closer to the pedestal now.

	All 5 heads are on line

	Were Close.


					A shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.285                          CMP.B1                           285 of 292
PATE::GEE                                            71 lines  24-OCT-1995 07:16
                     -< Flipper Assemby / poor uniformity >-
--------------------------------------------------------------------------------
 	* After the 2nd run the multi-head went to polish and stopped.
	  There was no errors, we tried to re-start but the system
	  wouldn't respond.

		- Checked the IO's and found that #2 load cup was not
		  detected down, but infact it was down. We adjusted the 
		  pedestal and sensor to finish the run.

	* The unload flipper was missing #2 & #3 wafers everytime.


	Called field service, they arrived at 10:00pm.

		- Re-adjusted the #2 load cup pedestal and down position
	          sensor.

		- Put the unload cups back into the designated feed thru
		  holes.

		- Adjusted the unload cup pedestal up and moved the sensor.

		- Adjusted the unload flipper fingers.

	They left at 11:30.

	* We ran 25 wafers then mqc's. The mqc's range came out between 
	  7.7% and 9.0 %, 2 heads within spec. Eric B. made some 
	  adjustments and we ran another 25 wafers then another mqc, 
	  the range is now 5.5% to 12.0%, one carrier within spec.

	* The system is down for poor uniformity and load flipper problems.

-----------------------------------------------------------------------

	* Problems the Speedfam rep.'s should address this morning:

		1. After the first 25 wafers the load flipper began
		   mis-aligning the #2 wafer in the #2 load cup every- 
		   time. We had to manually adjust the wafer position 
		   in the cup to complete the break-in and mqc wafers.
		   This happened once in # 3 also.

		   It seem that we are constantly re-adjusting the load
		   and unload flippers. The positions seem to be drifting
		   or maybe the system vibration is moving the flipper 
		   assembly slightly.

		   The system should stay down until the positions are
		   re-adjusted.

		2. Unload cups not seating down all the way into the 
		   index table.

		3. We have the wrong unload cup pedestal. The one
		   we have in the unload cup is made for the load cups,
		   it's to thin. They need to order use a new one.

		4. The Scrub Station brushes are not rotating everytime. 
		   We had to start it by manually rotating the motor shaft 
		   the first time, then it ran fine for the night. 
		   The motor maybe sticking or there is a mechanical
		   bind in the scrub assembly somewhere. We checked the 
		   wiring and connector, they seem o.k. The scrub error 
		   didn't repeat so it's had to say what the actual 
		   problem is. 
		  
	                            		B shift
  




          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.286                          CMP.B1                           286 of 292
SUBPAC::LANDRY                                       38 lines  24-OCT-1995 19:46
                               -< todays events >-
--------------------------------------------------------------------------------
At the start of the shift, the system was experiencing load problems
with the flipper to load cup position.  Also, a pad change was taking
place along with the carriers being checked for proper ring extention.

*  A new pad was installed.  It was a modified version of the ID/OD
   style.  This pad had the inner portion (only) removed, with the 
   remainder of the pad left intact.

*  The carriers were all re-installed after ring extention verification
   was complete.

*  A cassette of wafers was loaded up and run using the "TRANSFER" 
   program.  No loading mishaps were seen.  In fact, it was loading
   very smoothly. ???

*  So, the Opts folks ran some break-ins and MQC's all was well, except
   carrier three's uniformity was +9%.  So, a couple of more break-ins
   were run.  The system came up with a "wafer not detected on carrier 4"
   error.  The wafer was checked for proper placement (which was found
   to be good) and off they went.  A few runs later and the problem 
   re-surfaced.  The pressed a button to continue, and to the polish table
   it went. CRASH!  Head 4 spit a wafer.  So.........

*  Head alignment was performed.  Unfortunetly, head 4 was dead nuts on.
   A couple of other heads were off, so alignment was performed.  The Index
   Table and Multihead positions were checked prior to any adjustments.
   During the adjustments, the limit sensor for head # 2 had to be adjusted
   for the proper head movement needed to complete this alignment.

*  Load flipper calibration is currently taking place.

*  The system will need a new pad (bare table now) and the carriers re-
   installed.  Three of the original carriers are still good.  Jim and
   Larry built the other two.  The two they built are on the carrier re-
   build table.  The original three are staying wet in the tray near the 
   system.

      						A shifters

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.287                          CMP.B1                           287 of 292
PATE::GEE                                            31 lines  25-OCT-1995 07:26
                              -< Head #4 vacuum >-
--------------------------------------------------------------------------------
	 Replaced the pad and carriers after the wafer break today.
	 Began running break-in wafers. We got an error "vacuum no
	 found on head #4".

	 We checked the vacuum gauge for head #4, there was no vacuum.
	 Removed the vacuum line going into solenoid valve #40, vacuum 
	 was present.

	 Checked the voltage to the valve connector we had 23.5v.
	
	 Checked the connector, it looked o.k. 

	 Toggled the vacuum in manual, it worked everytime. In auto
	 we would get the error every other run.

	 We thought the valve maybe sticking so we replaced it.
	  
	 The problem was still there.

	 -----------------------------------------------

	 Field service has been contacted.

	 * The vacuum works in manual everytime. In auto mode it
	   works about every other time.
	 
	                            		B & D shift
  




          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.288                          CMP.B1                           288 of 292
FABSIX::B_FINN                                       16 lines  25-OCT-1995 19:07
                                 -< passdown >-
--------------------------------------------------------------------------------


	Carrier 4's vacuum working intermittently.  Troubleshot and found
 open in common wire to solenoid 40. Found open in the wire. Replaced wire from 
 Multi Head terminal strip to solenoid molex. Ran 25 wafers aok.  Changed 
 to a new ID OD pad. Wrote and stored pad conditioner recipe # 3 which centers 
 conditioner on pad for equal overhang and removes all oscillation from 
 conditioner arm per Speedfam and Larry's recommendations.  
 Engineering has been on it, and it's running fine.

 Cut two extra speedfam ID OD pads and placed on vidmar cabinet.

 
				                 A shift



          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.289                          CMP.B1                           289 of 292
ASDG::POIRIER                                         8 lines  26-OCT-1995 19:37
                     -< System Performance / Rep Support >-
--------------------------------------------------------------------------------
    
    * System Performance:
    
      The tool ran well during the shift with no problems.
    
      Darren McDougal from Speedfam has returned to Vermont but
      will be on call until tomorrow at 5:00 p.m.. After tomorrow,
      call the 24hr Speedfan service line for support. 

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.290                          CMP.B1                           290 of 292
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"               7 lines  27-OCT-1995 06:52
                         -< Weekly PM almost complete >-
--------------------------------------------------------------------------------
    
    
         Most of the weekly PM was completed except for lubrications
      of the scrubber, paddle and conditioner gimbal.  The other items
      will be submitted in the checksheet topic.
    
                                                    D shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.291                          CMP.B1                           291 of 292
ASDG::POIRIER                                        16 lines  27-OCT-1995 19:10
                   -< Unload Air Cylinder Pedistal replaced >-
--------------------------------------------------------------------------------
    
    * Unload Air Cylinder problems:
    
      The system errored for "Load Cups Not Down" this morning. We checked
      the I/O's and found #97 (Unload Air Cylinder Sensor) not lit. We 
      decided to change the Air Cylinder Mushroom Pedistal from the thin
      type to the thick style per Speedfams request. The thick type
      increase the Cup travel by 1/8". After the Unit was reassembled,
      the Cup sensor was readjusted and the "Unload FLpper" pick-up
      position was taught a new location.
    
      Reminder: For Speedfam support tonight and thru the weekend, call
                their 24 hr support line. All reps have returned to Az.
    
                                                     John & Brad.
     

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 62.292                          CMP.B1                           292 of 292
SUBPAC::LANDRY                                       26 lines  30-OCT-1995 18:56
                   -< exhaust work, monthly PM, pad change >-
--------------------------------------------------------------------------------
Lots of activity on this baby today...........

*  The system ran up until about noon time.  At that time, many
   things happened with the system.

      -  There was a scheduled down situation to allow some contractor
         work to be done on the exhaust line.  The work centered around
         preparation for the much awaited Speedfam (B2), which will be
         brought into the fab on Wednesday.

      -  Some other contractor work performed was the laying of PVC pipe.
         The polisher remained running as this work didn't impact the
         system from running (no direct tie_in/break_in done on B1's
         facilities inputs)

      -  The Monthly PM was performed in this time frame.  And just in
         time, as the slurry lines were mighty thin.

      -  The pad was changed out.  The conditioner wheel was broken down,
         cleaned, and re-assembled.

*  Later on, the system errored for NO rotation/Inverter error.  The 
   problem was traced back to the "wafer detection sensors" being wet.
   They were dried and away we go.

      						A Shifters
17.105B1 wafer breakage due to loose screwsSUBPAC::BJUBINVILLENFD Fire - RescueThu Nov 30 1995 08:4711
    
        Many 3mm clamp ring screws were found loose after the
     tool had crashed.  An exercise wafer slid out of carrier #1,
     slid under #5 (scratching the product wafer) then crashing 
     into #4, scratching the product wafer there as well.  Carriers
     #1, #4 and #5 were replaced due to the crash.  The screws on
     carriers #2 and #3 were found loose.  They along with the three
     new carriers now have an exposure of -7 to -7.5.  John is 
     running to the hardware store to look for a different type of
     washer such as a star washer.
                                                  BJ
17.106(2) wafer crashesSUBPAC::BJUBINVILLENFD Fire - RescueFri Dec 01 1995 08:3517
    
     
           The tool crashed twice tonight, not "bad" crashes
      but a crash is a crash.  The first time, a product wafer
      broke in carrier #2 and sent "sand" sized particles to the
      other carriers.  (4) product wafers were all scratched.
      The carrier rings were scorn and all pulled off.  After
      looking into them later, some of the films were ok and
      some weren't.  I was able to rebuild some.  The second
      crash came from #4 and was cracked during polish.  The 
      wafer was unloaded in two pieces.  There were some 
      particles on the pad but the carriers were ok.  The
      only thing that was common was that it looks like both
      wafers broke at the notch.  ENG has the ball.
    
                                           BJ
    
17.107Load Cup height adjustmentASDG::POIRIERFri Dec 01 1995 19:3120
    
    
    Load Cup Height Adjustment:
    
    Samples of the broken wafers from last nights breakage incident
    indicated that a wafer may have fractured during the load sequence.
    Therefore, we decided to inspect the load cup height adjustment.
    
    Inspection was extreamly dificult due to the vague instructions in the
    manual (Speedfam was called for clarification), very limited access
    to the adjustment rods & rusted extention rods and lock nuts. (We had
    to run a 5/16 - 18 die down some of the extention rods and tap-out one
    of the tapped holes in an Air Cylinder Piston).
    
    All 5 load cups have been adjusted and tightened in place. D shift,
    please double check the load cup air cylinder sensor adjustment on all
    5 air cylinders as we are running out of time.
    
                                                                  
    
17.108Spindle Alingment OutSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 02 1995 07:3217
    
         The cups up/down was checked for sensor detecting and
      speed and the flipper handling.  #2 was a little fast going 
      up so it was slowed down to match the rest.  Everything else
      was ok...  The tool ran 45 wafers before #5 cracked on the 
      table.  The carreirs were ok.  we checked the index table
      to spindle alignment and found spindle #2 and #5 way out.
      #2 and #5 as you know were the ones that kept crashing.
      Upon completing this task, I noticed all of the load and
      unload cups were in the wrong places.  The MUST be put
      back in their own holes.......   Ops ran 50 wafers with 
      no problems.  MQC's ran and passed uniformity but had a 
      a small scratch on one wafer.  C shift OPS will be informed
      and they will either run it or do (5) more wafers to verify.
    
                                               D shift  
      
17.109Vacuum Errors / load Errors / Wafer Fly-outFABSIX::R_GEESun Dec 03 1995 06:5436

              <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]CMP.NOTE;1 >>>
                                    -< CMP >-
================================================================================
Note 19.19                   CMP.B2 EE/PE PASSDOWNS                     19 of 19
ASDG::POIRIER                                        26 lines   2-DEC-1995 19:35
                -< Vacuum Errors / Load Errors / Wafer Fly-out >-
--------------------------------------------------------------------------------
    
    Vacuum Failure:
    
    The system alarmed for "main vacuum too low" at approx 5:00p.m.
    I found the vacuum pump off, breaker #6 tripped & the check valve
    bound-up with dried slurry. I replaced the check valve and reset the 
    breaker. The vacuum rose to 25 in hg. ( The old check valve was
    cleaned. Its on the toolbox minus the spring which is under the floor).
    
    We then attempted to boot-up and experienced a "fatal error". We powered
    down the system and tried again. This time everything looked normal
    until the sytem errored for "carriers 3,4 & 5 not seated in cups
    properly". We recovered from this situation by powering down and
    rebooting.
    
    We then ran 15 wafers with the transfer program with no problems. After
    that, we changed to the break-in program. During the first run, a wafer
    slid out of carrier #1 during lift off. The wafer slid to the edge of
    the table as the table came to a stop. The wafer did not break. B
    shift, check the spindle to cup aligment on cup #1. D Shift found 2
    spindles off center with the load cups last night. Maybe they're  
    drifting out of locaton. The system is in BGN_REPAIR in workstream.
    We left it in Wet Idle Mode.                                     
    During load while the conditioning screen was up.
    
       

17.110Vacuum ProblemsFABSIX::R_GEESun Dec 03 1995 06:5536
	* Checked spindle #1 load cup alignment do to "A" shifts
	  passdown, it looked good.

	* Tried to run the Transfer program, but came up with a
	  Low vacuum to all heads error.

	  - Checked head #4 & #5 vacuum gauges, they both read 15hg.

	  - Powered down and checked the vacuum at the tank. It read
	    26hg, but when you start it up and pull vacuum on all 5 heads
	    it would read 15hg and never recover.
	
	  - Checked the wafer positions on all 5 carriers to verify
	    the wafers where seated correctly, all looked good. 
	  
	  - Replaced the check valve with the old cleaned up one.

	  - Cleaned the inside of the vacuum pump.

	* Began running again, vacuum increased to 26hg at the spindles.
	  
	  - Ran 25 wafers with no problems, started mqc's.

	* Came up with a "low main vacuum" error.

	  - Found the breaker #13 had tripped, also #6,(carrier #4 & #5 
	    rotation).

	  - Reset Breaker #6.

	  - Reset Breaker #13, this took many tries before it would stay on.

	* Ran 25 more wafers then MQC's with no errors, the system is up
	  in RTC.

						B shift
17.111vacuum errorFABSIX::B_FINNSun Dec 03 1995 19:2818

	Would not pick wafers up on table (carrier 5). Low vacuum alarm.
Pump was on. Vacuum tank was at 0 "hg.  Check for appropriate vacuum after 
check valve,  found NO vacuum, check directly at output of pump, found NO 
vacuum although pump was on.   Dismantled pump, diaphrams were fine, but 
filter had been caught in screw in NPT fitting at pump input. This, plus some 
dried slurry was choking off input side of the pump. Took pump, diaphrams, 
filters and fittings apart cleaned and reassembled. Replaced scored vacuum line 
and tested. Vacuum was at pump but not past check valve, installed original 
style check valve.  Tested aok. 

Adjusted cup locating pins for Load 5 and Unload 2. This will prevent the cups 
from twisting.

					A shift


17.112Bridge failure / broken waferFABSIX::R_GEEMon Dec 04 1995 07:2819
	* After stopping wet mode the bridge wouldn't move.

	  - The carriers went up then an error was displayed
	    "Timeout waiting for Bridge at polish". 

	  - Tried to re-home the bridge but no luck, had to reset the 
	    computer. The error never repeated.

	* Broke a monitor wafer on the pad around 6am. 

	  - A wafer slid out from under carrier #4, replaced the pad
	    and carriers #3 & #4. Checked vacuum to the head, it was 
	    25hg, good. Checked to see if the wafers are seating into 
	    the carrier properly, looked good. 

	  - Ran 50 breakin wafers, the system is ready for MQC's.


						B shift
17.113SUBPAC::LANDRYTue Dec 05 1995 20:2927
    The infamous low vacuum problem cropped up again today.  Here's the
    details of the problem/fix.
    
    *  The system errored for "no wafer on heads 2 & 5"
    *  The vacuum tank gauge was checked and we were pulling only 18" Hg
    *  Removed the check valve/needle valve from the loop and the system
       pumped down to @ 22" Hg, still not the optimal situation
    *  Broke down, cleaned up and re-installed the check valve and needle
       valve.  Slurry (still damp) was found in the check valve.  The
       needle valve was ok.
    *  After re-instalment, check the performance.  The pump down response
       was slow, but making it to 23" Hg.  Ran three cycles of the transfer 
       recipe without failure
    *  At this point, Opts began running again.
    *  The system went down very shortly thereafter for low vacuum.
    *  Again, found slurry in the check valve.  Rebuilt it.
    *  The pump was also changed at this time for slow pump down 
       performance.  The pump isn't new, rather, it's  
    *  Found some slurry in the vacuum line between the tank and the pump
    *  Flushed the hose to remove any "hidden" slurry blobs.  
    *  Retried pumping performance.  The tank pumps down to 25" Hg and
       never gets below 20".  Very acceptable at this time.  
    *  Ran a cassette of wafer using the transfer program and tank vacuum
       was steady, staying over 22" throughout the testing.
    *  Opts has the system back
    
    					Doc  /  Brendan  /  Guug's
17.114cup not down time out errorSTRATA::RIDLONEliminate the obviousWed Dec 06 1995 05:3322
    System alarmed for cups not down time out error twice during the shift.
    Once production was through this is what I found.
    
    * Both errors occured while unload cup 2 was over the load cup #3 
      solenoid (labeled "Send load cup 3 up" in pnuematics cabinet).  
    * Both times the solenoid was pushed down by hand. 
    * Unload cup 2's up and down movement was compared to unload cup 1's
      and was found to have more friction.
    * The shaft and hole for ul2 cup was cleaned which improved the up/down
      movement. However this didn't correct the problem. 
    * Exercised load cup 3's solenoid and found that down sensor wasn't
      being met (dig-in #94).
    * Loosened the setscrew on the down sensor and moved it out of the  
      way which gave the solenoid the clearence it needed to go all the
      way down. 
    * While the solenoid was down adjusted the sensor so that dig-in 94
      was satisfied.
    * Exercised load cup 3's solenoid several times to insure that the 
      plunger wouldn't hit the sensor and that dig-in 94 was being
      satisfied.  
    
					B shift
17.115passdownFABSIX::B_FINNWed Dec 06 1995 19:1420



	Elevator #1 failed to index. Rep added new connector to pins on power 
        cable.  System ran half a lot then faulted for the same error.  Check 
	in manual and elevator 1 would not index or home.  OHM'ed power cable 
        to motor. aok  Swapped cables (power and tach between elevator motors 
        and problem moved to elevator #2) isolated to tach cable. Ohm'ed out 
        and found two bad wires. One was open and one was intermittent (broke 
        while handling.)  Cut cable back and re solder all wires to 
        tach D connector.  Elevator ran fine.

	
	Water level at cassette  #1 was low and would not float wafer to 
        cassette.  Adjusted water level and sticking float sensor.



							Brendan/ Bob
17.116vacuum and rotating problemsSUBPAC::BJUBINVILLENFD Fire - RescueThu Dec 07 1995 09:0819
    
    
       We came into vacuum problems at the beginning of the shift.
    I inspected the check-valve and it was clogged.  I also found
    the needle valve clogged.  After they were replaced, the vacuum 
    went right to 25.  I ran a 25 wafer verify run and had no problems.
    About three hours later, there was a new problem.  The MHD picked
    up (5) wafers and went to the polish table and everything satarted 
    to rotate as it should.  Just before the carriers were ready to
    go down, all (5) carriers and table stopped rotating and errored
    on all saying they weren't rotating.  Now breakers were tripped 
    and now wafer detects were tripped.  There was a red LED on #5
    of the #2 MDX board but I wasn't able to find out what it was. 
    After I moved the MHD to LOAD, the tool errored saying the MHD
    didn't reach LOAD.  That problem went away.   I spoke with Bob 
    Guggerty on the phone and he was baffled.  He was suppose to
    be here around 7:45am so he should be in the fab.
    
                                        BJ
17.117Vacuum Tank Popet Valve rebuilt / Head #4 Wafer Loss SensorASDG::POIRIERThu Dec 07 1995 19:1745
    
    * Vacuum Tank Popet valve:
      -----------------------
      The reoccuring "Main Vacuum Too Low" error was traced back to the 
      vacuum tank popet valve. This valve was replaced approx 3 months ago.
      The internal o-rings looked worn and were replaced. The valve was
      reassembled and installed back into the tank. The tool has been
      running for approx 5 hours with no vacuum errors.
    
      Having experienced this situation twice now, it appears that the 
      following senerio takes place:
    
      1) The system errors for "Main Vacuum Too Low".
      2) We disassemble the check valve and find it bound-up 
         and full of dried slurry residue.
      3) We replace the check valve only to discover the exact same
         conditions two days later.
      4) Soon thereafter, the vacuum pump fails due to the ingestion of
         large quantities of slurry.
    
      What we learned:
      ---------------
      The vacuum tank popet valve must completly close and seal every time
      it's activated or slurry will leak by and contaminate the check
      valve, needle valve & vacuum pump. 
    
      Close inspection of the popet valve revealed a questionable design
      where the sealing surfaces for the internal o-rings were of unusual 
      design. There were sharp edges in bore of the valve that appeared
      to shave the edges of the main sealing o-ring. The condition of this
      valve was not very reassuring after only 3 months of service.
    
      Bob Gugerty, our Speedfam field service rep, is filing a report on
      this situation with the hope of recieving a better quality valve
      sometime soon.
     
    
    
    * Head #4 Wafer Loss Detection Sensor:
      -----------------------------------
     The Wafer Loss Sensor on head #4 was repositioned. It lookslike it was
     sensing the side of the carrier.
      
    
      
17.118Table and Carriers not rotatingSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 05:1811
    
    
          From C shift, the table and carriers weren't spinning
        in polish step.  The problem went away.  It popped up
        on us and I checked the program they were using which
        was I_o_LDA, and found in segment #3, which is the 
        polish step and found all RPM's set to 0.  They are
        suppose to be at 37.  Tool ran fine for the rest of 
        the morning.
    
                                            BJ
17.119GremlinsSTRATA::RIDLONEliminate the obviousMon Dec 11 1995 02:5810
    System alarmed because wafer wasn't making into elevator 1's cassette. 
    Wafer wasn't making it into cassette because elevator 1 wasn't full
    of water. The operators were able to "manipulate" the wafers into the
    cassette so the run was continued. Traced some signals and found that
    the elevator 1's quick dump valve was being commanded open by dig-out
    70. Once the run was completed the elevator was cycled out of and back
    into service which corrected the problem. Cycled it a couple of more
    times and problem didn't reoccur.
    
    						B shift
17.120Slurry pump linesFABSIX::R_GEEMon Dec 11 1995 06:525
	* Blew a slurry pump line. 

		- replaced both slurry pump lines and cleaned the system.

						B shift
17.121FABSIX::B_FINNMon Dec 11 1995 19:3329
	Weekly pm completed.  Assisted by Larry C. and Jim S.

	Upon startup, low vac error occurred found pump breaker tripped. 
Continually tripped, Tim found pump dead headed.  I adjusted bleedoff 
needle valve. aok

	In the late afternoon Jim was running a laminate split with program 
I/O ILD b.  The range was high he felt it was due to carriers 1+4 so  he took 
them off line.  Sometime into the run he notice that the heads were not 
oscillating.  I checked for DXC and infornor errors but all was fine. 
Program was fine I had Larry C. confirm. 

Put all five heads back on line without powering down.  Tied 5 
wafers on on programs I/O ILD A, I/O ILD B and I/O ILD C. System ran fine.

Took head 1+4 off line and ILD A past but B+C FAIL for NO OSCILLATION, and no 
system errors.

Jim found after 100% measure that heads 1 + 4 are not a problem, all 5 heads 
are back on line.   Suspect software problem.  But system just ran another lot 
aok.



Brendan



17.122Drain valve Elevator 2LUDWIG::RIDLONEliminate the obviousTue Dec 12 1995 06:4215
    Elevator cassette #2 did not fill with water during product run. 
    We checked the usual signals (dig-out 68-71) and found every thing as
    it should be, ie, the drain was closed and the quick fill was on
    (electronically). Looking at the drain valve it appeared that it was
    still open. Wacked it with a hammer a few times and it closed. Once
    production was through with the system we removed the lower portion of
    the drain valve and disassembled. The spring inside was painted and
    some of the paint had come off which may have contributed to the valve
    sticking close. Also found the oring on the plunger cracked, this was
    replaced. Cleaned it up and regreased with the teflon grease in the
    tool box. Reassembled and reinstalled and it tested fine. 
    
    
    						B shift
    
17.123passdownLUDWIG::FINNTue Dec 12 1995 19:269
	Wafers piled up at input of elevator #2, due to rinse not filling with 
water.  One wafer broke. Valve is sticking again.  Cleaned both elevators and 
scrub station.  Thanks for the help FRANK. Ordered a new valve through 
Speedfam.  Shut off Elevator two and we are running on elevator 1.  Warning 
placed on system until new valve is installed.   Valve should be in 
tomorrow fedx. 

						Brendan/Frank	
17.124passdownFABSIX::L_SHKOLNIKOVWed Dec 13 1995 07:009
     Upon startup, elevator #1 not responding error occurred went into hardware
enable screen everything looked fine but didn't work. Took elevator #2 back
in service - #1 started work again. Machine was running all night with on
elevator just did not use #2. Suspect software problem. But system just run 
another lot fine.
    Later got error message MOISTURE SENSOR ACTIVE. Checked bottom cabinet
on elevator side found about a gallon of water on the bottom and couple
loose fittings of elevator #2 drain pipe. Problem corrected. Machine running
fine since then.
17.125R Recieve Elev Dump Valve / Vac Tank Poppet ValveASDG::POIRIERSat Dec 16 1995 19:2019
    
    * Recieve Elevator Dump Valve:
      --------------------------
      The Recieve Elevator Dump Valve has arrived. It's on the Snap-on 
      tool box. Please install the valve at the first available opportunity. 
    
      The Right Recieve Elevator is currently off-line do to the dump valve
      sticking. 
    
    
    * Vacuum Tank Poppet Valve:
      ------------------------
      We're having problems again with slurry getting drawn thru the entire
      vacuum system. We believe the root cause is the Vacuum Tank Poppet
      valve not closing completely. A replacement valve arrived today. 
      Please replace the valve at the earliest possible opportunity. The
      check valve is clogging approx every 2 days. The old check valve was
      cleaned and stored in the pnumatics cabinet for a quick swap. The 
      system has been running non-stop for the last 3 shifts.
17.126new dumper inLUDWIG::RIDLONEliminate the obviousSun Dec 17 1995 03:333
    Replaced the dump valve on elevator 2 with the new one. Tested fine.
    
    					B shift
17.127Vacuum tank poppet valve / index table gearbox leakFABSIX::R_GEEMon Dec 18 1995 07:0713
	* "C" shift passdown indicated the vacuum tank poppet valve 
	   needed to be replaced.

	  Replaced the vacuum tank poppet valve, tested o.k. Dis-
	  assembled the old poppet valve, found no visible problems
	  with it. 

	* Noticed the index table gear box is leaking oil from
	  the rear shaft seal. Checked the oil level, it's still
	  above the bottom of the fill hole. We need to order
	  a seal and bottom case gasket to repair the leak.

						B shift
17.128pad/carrier changeSUBPAC::LANDRYTue Dec 19 1995 18:2310
    
    *  The pad and all 5 carriers were changed today.  
    
    *  The five used carriers were dismantled with 3 used plates being
       cased up for shipment.  The film was removed from the other 2
       plates.  DF "experimental" film was applied to them and will be
       used for experimental purposes in the near future.
    
    						A shift
    
17.129Main Vacuum to low.FABSIX::R_GEEWed Dec 20 1995 04:009
	* The system errored for "main vacuum to low".

	 Found no vacuum in the vacuum tank. Had to remove
	 and clean check valve, needle valve, all fittings.
	 Rebuilt and clean the top part of the vacuum pump.
	
	 Everything was clogged with dried slurry. 
					
					B shift
17.130ScratchesASDG::POIRIERFri Dec 22 1995 17:2611
    
    * Scratches:
      ----------
      The system has been scratching wafers since the begining of the
      shift. It was suggested that the inside edge of the Pad might be 
      responsible so we trimmed the inside edge of the Pad and rinsed
      everything down thoroughly. Still no improvement!  At this point,
      it doesn't make much sense to continue because the Pad will be
      removed very soon as part of the shutdown activity. We'll have to
      wait for results when the new Pad is MQC'd at start-up.
    
17.131SHUTDOWN COMPLETED, SCRATCHESSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 23 1995 01:1113
    

       Since the tool still has scratch problems.  This was
     logged down.  Tim D. said to keep it down and start the 
     shutdown procedure.  The checksheet has been completed
     with the exception of the slurry purge.  That will commence
     when the rest of the tools are shutdown.
    
       I left a PVC/SS setup with B2 for the N2/DI purge.  Disconnect
     the 3/4" elbow from the main shutoff valve above the floor and
     put the fitting on.  I used the existing N2 line for the 40psi
     purge.
                                               D shift
17.132Start-up for 12/29SUBPAC::BJUBINVILLENFD Fire - RescueFri Dec 29 1995 15:1012
    
    
        The DI came on but is still out of spec so we didn't want
      to keep it to the tool so we shut off the DI feed and opened the
      by-pass.  Facilities still have some by-passes opened, that's
      why we don't have very high pressure.
    
        The slurry lines are all back to normal and slurry is flowing.
      We put the filter on but but we didn't run the slurry through
      the machine because of the DI situation.
     
                                                       Mitch,John,Leon,BJ
17.133CMP.B1 START-UP COMPLETESUBPAC::BJUBINVILLENFD Fire - RescueSun Dec 31 1995 14:3518
    
                        CMP.B1 START UP CHECK SHEET
+++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++

1) RE-CONNECT DI WATER					       B/C/D-12/31/95	
2) POWER ON SYSTEM		                               B/C/D-12/31/95
3) CYCLE/RUN ALL DI FUNCTIONS FOR 15 MINUTE                    B/C/D-12/31/95
4) WIPE DOWN SYSTEM CLEAN ALL DRIED SLURRY	               B/C/D-12/31/95 
5) DRY TABLES, INSTALL PADS					
6) INSTALL SCRUBBER BRUSHES				       B/C/D-12/31/95	
7) INSTALL CONDITIONER					       B/C/D-12/31/95	
8) RUN WET IDLE PROGRAM FOR 15 MIN. (ENOUGH TO CLEAN CARRIERS) B/C/D-12/31/95
9) REPLACE SLURRY FILTERS				       B/C/D-12/31/95	
10) CONTACT ASHLAND CHEMICAL, REQUEST DI TO BE PURGE FROM      B/C/D-12/31/95
    SLURRY SYSTEM AND CHARGED WITH SLURRY			
11) CALIBRATE SLURRY FLOW				       B/C/D-12/31/95	
12) PUT IN WET IDLE AND TURN OVER TO MFG FOR MQCS	       	

17.134Head-4, Cond arm failureSTRATA::TDYERTue Jan 02 1996 15:3212
CMP.B1 - The system did not want to cooperate today durring MQCs. Head-4
         failed to rotate, Tim L. found the connector loose on DCX bd-2.
         The conditioner failed to go home. A broker wire was found on the
         connector for the "conditioner at home" sensors. Eventualy we
         ran MQC's. High particles were described as a "streak" by B shift. 
         We noticed that some of the wafers today after polish had a steak,
         a 1" by 3" stripe of THIN oxide. This may be a film quality
         issues from the CVD tools. The oxidemsy be less dense in the 
         striped area and polished faster. The thinner oxide causes problems
         when reading on the Tencor. All MQCs passed.

17.135broken wafersSUBPAC::LANDRYTue Jan 02 1996 19:4115
    See Tim D's note prior to this note.  I'll pick up where that one ends!
    
    The system passed MQC's and began to run product.  A wafer was lost
    from head 4, taking that and the next carrier out.  Cleaned up the 
    wafers, installed a new pad and it's currently awaiting a fourth
    carrier to come in from the rebuild room.  There are no plates left
    in house, so the system will run four heads for now.
    
    The reason for the loss of wafer may have been due to an interlock
    triggering the power to shut off on the system.  
    
    							Doc
    
    
    
17.136pressure plate tolerances out of specSUBPAC::LANDRYWed Jan 03 1996 19:356
    The system is down until new pressure plates arrive.  We have found
    that the last several plate changes have had a film curvature in excess
    of the 2 um that is spec'd.  Some plates have been found to be in the
    10-20 um range, resulting in extremely poor uniformity.
    
    						Doc
17.137Infranor's rewired / Scrubber RebuiltASDG::POIRIERThu Jan 04 1996 18:5418
    
    * Power Supplies & Motor Controllers:
      ----------------------------------
      The Infranor's were rewired with the Factory supplied wiring kits.
      The Connectors were partially pre-wired, it was still necessary to
      remove the Motor Controller Cable Wires from the old Connectors and
      fasten them into the new ones.          
    
    
    
    * Scrubber:
      --------
      The Scrubber was torn down and rebuilt today. The unit was adjusted
      and tested. It seems to be working fine.
    
    
    
      
17.139DOWNFORCE CHECKED/ADJ./RESULTSSUBPAC::BJUBINVILLENFD Fire - RescueThu Jan 04 1996 22:1022
    
    
          the downforce calibration was done on all (5) heads.
       Listed below is all of the data collected and original
       results of each head.
    
    ========================================================================
     CARRIER #  | INPUT COEFF.  | OUTPUT COEFF.  | INITIAL     | CURRENT
                |               |                | ACTUAL LBS. | ACTUAL LBS. 
                |               |                | @375        | @375
    ========================================================================
       #1       | .46182        | 6.19467        | 380 LBS     | 375 LBS
    ========================================================================
       #2       | .47710        | 5.99467        | 377 LBS     | 375 LBS
    ========================================================================
       #3       | .48765        | 5.82933        | 395 LBS     | 375 LBS
    ========================================================================
       #4       | .46972        | 5.95467        | 355 LBS     | 375 LBS
    ========================================================================
       #5       | .48201        | 5.99467        | 389 LBS     | 375 LBS
    ========================================================================
    
17.1381/3 B SHIFT P-DOWN - SCRATCHES, CRASHSUBPAC::BJUBINVILLENFD Fire - RescueFri Jan 05 1996 01:1627
From:	FABSIX::R_GEE "Bob Gee, Beeper #184  03-Jan-1996 0808"  3-JAN-1996 08:12:00.39
To:	@CMP_ALL
CC:	R_GEE
Subj:	Passdown, the notes file is down. Info for CMP.a1,CMP.B1,CMP.B2


	CMP.B1
        ------

	* The system was down until 3am due to scratches (?)
	  found on the mqc wafers. A second mqc was run with no 
	  scratches found.

	* At 6 am the system powered itself down in the middle 
	  of polishing for no apparent reason, crashing all 4 wafers.

		- we found that carrier #1 had been polishing
		  a broken wafer for some time. The wafer detection
		  didn't detect the crash due to the small size of 
		  the particles. 

		We thing the system powered itself down due to high
		amps. from  either the table or head #1. 
	
	      					

					B shift
17.141carrier 3 disabledSTRATA::RIDLONEliminate the obviousSun Jan 07 1996 05:3212
    Carrier 3 would not rotate with a wafer loaded and down on the pad.
    Rotated fine in wet mode. Checked the infranors and found the orange
    x5 motor control plug not seated properly on motor control #9. 
    Reseated and tried again but still the carrier would not turn. Pulled
    out mc#9 and looked at the orange connector. Found wires w9 and u7
    backwards. Put them in their proper slots and tested again and now the
    carrier wouldn't turn in wet mode. The w9 and u7 must be switched at
    the motor as well. Left the wires where they should be and removed the
    carrier and disabled it as well. 
    
    
    					B shift 
17.142rotation blues......SUBPAC::LANDRYSun Jan 07 1996 19:0213
    Found the prints to be wrong on the motor controllers.  Wires that were
    switched on the B shift HAVE been returned to their original places.
    The two wires of concern are U7 & W9.
    
    Checked all connections at the motor controller, all good.  Checked the
    connections at the DCX board.  They appeared to be fine as well.  Went
    into DOS and ran a rotation macro program to check the rotation.  All
    heads rotated.  At next MQC, the carrier should be reinstalled and 
    actual performance of rotation on the pad should be checked out.
    
    						Doc
    
    
17.143no rotation updateSUBPAC::LANDRYMon Jan 08 1996 17:2923
    The status of head three remains the same, rotation stops when the head
    is on the table under a load.  Eevn with the swap of the Infranor (B2,
    head 3), the problem still exists.  The next likely senario is a bad
    connection or broken wire.  The connections at the Infranor are good
    and the connection at the DCX board is good.  The other junction spots
    need to be looked at.  Ohming of the wires may need to be done if the
    connections prove to be fine.
    
    At the start of the shift, scratches were occuring on the wafers.  They
    could have been caused by silicon chips found in the index table or
    possibly from the rough cut ID of the pad.  The cut left several burrs
    on the top portion of the pad.  Also, Jim S cleaned other possibly
    areas as a plus.
    
    The system broke a wafer late morning.  The wafer came from head 5,
    taking head fours carrier/ring out.  The plate that was on carrier five
    is fine, but the ring will need to be replaced prior to re-using.  It's
    keeping wet in the tray near the sink.
    
    MQC's are being re-run as I type this memo.
    
    						Jim / Tim / Doc
    
17.144parts / help on the way......SUBPAC::LANDRYMon Jan 08 1996 17:337
    I forgot to mention in the prior note, Darren McDougal will be flying
    in with parts (Infranors, Wafer detection bd., etc).  Hopefully, he
    will arrive tomorrow, airports willing!
    
    					Doc
    
    							
17.145huge steps....backwards.STRATA::RIDLONEliminate the obviousTue Jan 09 1996 03:1825
    Got a call from Tim with a suggestion of checking the down force on
    head 3. With a setpoint of 300 the actual was around 290. We also ran
    some dummies through using bin1 breakin recipe which uses less down
    force. Still no head 3 rotation upon touch down. My next bright idea
    was to switch the cables from motor control 10 to motor control 9. Both
    the resolver (x1) and dcx board connectors (x2) were swapped and the
    five cables from the orange connector (that go to the motor) were
    swapped. W9 and U7 weren't swapped for head 4 so they were put onto
    mc#3 in there normal positions. All this proved for not as the carrier
    #4 wouldn't rotate. Put all the wires where they're supposed to go
    thinking everything would be fine but ran into some problems. Carrier
    #4 wouldn't rotate. Bob noticed that it looked like it was trying to
    rotate backwards so w9 and u7 were swapped on mc#4 and the head
    rotated. BUT it wouldn't rotate on the pad!!!! AGGGGHHHHHHHHHHH...shoot
    me. Ohmed out from x5 pins 6,7,8,9 on the motor controller to the motor 
    connector pins a,b,c,d on both motor #9 and #10. As it turns out both
    motors are pinned backwards, ie, W9 on x5-9 is supposed to go to pin C
    but the wire labeled w9 is soldered to pin A so you have to swap wires
    U7 and W9 on the orange plug x5. I know for a fact that the wires were
    NOT swapped before I removed them from mc#10 because I wrote their
    locations down. So I guess the question is how did it work before????
    Kind of long winded, hopefully its clear. Shitty week and its only
    3:20. 
    
    					B shift
17.146head 4 rotation problem re-appearsSUBPAC::LANDRYTue Jan 09 1996 19:3312
    *  See Tim Dyers general passdown note prior to this.
    
    Head four stopped rotating once again.  The head has been taken off
    line until tomorrow.  The reps are due in tomorrow to look at this
    re-occuring problem.
    
    The uniformity is good on all heads, put the MQC's show scratches on
    every wafer.  We ran the TRANSFER program, no scratches were found.
    So, the problem resides at touchdown / polishing steps.  Good luck.
    
    						Doc
    
17.147HEAD #4 NOT ROTATINGSUBPAC::BJUBINVILLENFD Fire - RescueWed Jan 10 1996 06:198
    
         Came in and had B1 down for scratches.  We replaced the pad  
      and brought it back up.  We used head #4 and it errored during
      the first initial (5) wafers for break-in.  I looked the wiring 
      on the infanor for the head and everything looked ok.  We started
      the tool back up and we didn't have a problem since.
    
                                                  BJ
17.148H3,4 NO ROT UNDER LOADLUDWIG::TDYERWed Jan 10 1996 10:3311
CMP.B1 - We replaced ALL the lugs at head-3's infanor connector and it worked
         under load. Head-4 also would not work under load today. At the
         terminal board on the left side of the multi-head, we swapped the
         3 motor phases. We swapped them at this location because access
         is best. Originally the phases from top to bottom on the terminal
         board was U, V, W (Sine, com, Cosine). They are now wired V, W, U 
         (com, Cosine, Sine). ALL heads tested OK, MQCs pass and system is
         up. I will schedule some time and we will re-label the wires in their
         current location.
 
17.149More scratchesSUBPAC::BJUBINVILLENFD Fire - RescueThu Jan 11 1996 07:0114
    

         The tool was MQC'ed twice and failed both times for scratches.  
      We installed a new pad and checked as much things as we could
      thing of as being the source of the scratches.  I know the brushes
      were changed so I swapped the lowers with the uppers to start.
      

              (5)  wafers -  no-touch, no scrub  -  no scratches
              (5)  wafers -  no-touch, scrub     -  no scratches
              (10) wafers -  I_O_MQC recipe, 120 second polish - scratches
              (5)  wafers -  I_O_MQC recipe, 120 second polish - scratches
                             no oscillation with positioning set at 129 on
                             both segments
17.150Particle Problems / Slurry Filter ChangedASDG::POIRIERThu Jan 11 1996 19:1410
    
    * Particle Problems:
      -----------------
      The system is still suffering from high particles. They appear as 
      scratches on the Surfscan due to the density of the particle
      distribution. However, they're really a group of densly spaced
      particles. After numerous trials to isolate the source, we're
      currently investigating the slurry supply. A new 10 micron 
      slurry fitler was installed on the tool at 7:00 p.m.
      Particle MQC"s will commence shortly. 
17.151Scratches - The saga continuesSUBPAC::BJUBINVILLENFD Fire - RescueFri Jan 12 1996 05:3031
    
         The conditioner was checked and this is the result...
    
                 @ 100%   -   53 lbs  -  book says 50 lbs
    
                 @ 67%    -   31.6 lbs
    
         The top panel was installed.  There was alot of metal shaving
      debris from aparent drilling.  Everything was wiped up.  The
      outside panels that access the spindles were removed and low and
      behold, the inner covers were all off and all of the slurry, metal
      chips and grease where evident and all ready to fall down below.
      There wewre even some broken chips of wafers in there.  Everything
      was vacuumed and cleaned out.  All of the inner and outer covers
      were put back on.  This should work right????     hahahahahahahahaha
    
                          NOT!     the scratches are still there and are
                                   actually worse.  They are #6-#10 wafers
                                   in the marked cassette next to the surf
                                   scan.  The I_O_MQC program was run.  The
                                   pre-particle test was done with PETOS-12K
                                   and post done with PETOS-6K.
                                            
    
               The pad that was on the tool when we cvame in last night
               was removed and I cut up three different sections and 
               placed them in seperate bags.  I left them on the back
               table marked for Tim and Frank to look at.
    
    
                                                   BJ
17.152SCRATCHES/VENDOR WORKFABSIX::B_WESTPHALFri Jan 12 1996 19:399
CMP.B1 - Down for scratches and work by reps.  The reps. upgraded the vacuum 
	 system with some new hose, a gauge that can be read from inside the 
	 vacuum cabinet and a moisture filter.  They also calibrated the down
	 force for each head to 377.  To fight scratches The multihead was 
	 blown out w/ N2 and the entire inside of the tool was wiped down.  The
	 Conditioner was taken apart and checked to if any particles were being
	 generated from inside.  It was also recommended by the reps that no 
	 grease be used to lube the conditioners gimble bearing.  The tool
	 has a new pad and presently sits at verify_u.
17.153SCRATCHES,SCRATCHES...FABSIX::L_SHKOLNIKOVSat Jan 13 1996 07:0510
    CMP.B1 - System still down for scratches. First run - no vacuum present on
         the carriers,found lose hose connector on multihead. Machine run
         fine (mechanically) after that. Failed MQC's for scratches.
         According to action plan BLOCKY conditioner has been installed and
         pad replaced.  


                                                    D-shift

                           
17.154wafer detection adjustmentFABSIX::B_FINNSun Jan 14 1996 18:026
	System alarmed  for "inverter error". Found wafer detect alarm on 
carrier 5.  Wafer detect was lit even when no wafer was present. We calibrated 
wafer detect sensor for carrier 5. Ran test wafers  aok

				A shift
17.155internal slurry line purgeSTRATA::RIDLONEliminate the obviousMon Jan 15 1996 01:5615
    Testing on A shift kind of pointed to the slurry lines being the source
    for particls. We only have two operators tonight so we thought it would
    be best if we rigged the slurry lines to purge di water for most of
    the night. We put in jumper lines from the slurry in to slurry out
    manual valves in the fluids box and then hooked up the slurry in lines
    to the two spare solenoid valves (sv-66, sv-67). We wired these valves
    directly to 24vdc and undid the slurry pumps so that the di could free
    flow. We also had to turn on dig-outs 74 and 75 (slurry flow 1 and 2
    solenoid valves) and whala, we had flow. We did run into one problem
    though, when we powered down the system to hook sv66&67 to 24vdc the
    vacuum pump wouldn't restart on power up. We installed the one that was
    in the room and it worked fine. We'll put everything back to normal
    before passdown so that testing can resume on A-shift. 
    
    					B shift
17.156major spindle work aheadSUBPAC::LANDRYMon Jan 15 1996 19:3818
    Slurry and pad expirements were run today, due to the scratching
    problem.  See Tim D's passdown for details.
    
    Rust is coming through at least three of the spindle assembly's.  They
    are currently being dis-assembled.  Parts are being ordered.  Much work
    ahead to recover from this situation.  Please organize parts when dis-
    assembling the heads.
    
    Head three is almost out, as Brendan is feverishly working away.
    
    The carriers need to be stripped down and thouroughly cleaned and
    re-assembled.  One has been completed, it's located on the cart with
    the pressure plate facing up.  Please wet it down occassionally. 
    Another carrier has been dis-assembled and cleaned, awaiting
    re-assembly.  It's on the bench next to the sink.
    
    					A Shift
    
17.157Spindle removalFABSIX::R_GEETue Jan 16 1996 07:0416
	Removed the four remaining spindle rotation motors, solenoid valves
	clear lines, fittings going to the rotary unions, gauges and rotary
	union guide rails. 
	Removed the quick release chucks and bellows from heads #1,3,4,5.
	The chuck from head #2 is rusted onto the rotary union shaft and
	will not come off, it has been lub'ed with oil and hopefully will
	free up for "A" shift.
	All the parts are located on a rack behind CMP.A2, each spindle
	has it's designated shelf top #5 bottom #1.
	One more carrier has been dismantled and is on the table next to 
	the sink. 
					B shift


	   

17.158new rotory unionsFABSIX::B_FINNTue Jan 16 1996 19:2011
	See Tim's pass down. Reps are finishing up the installation of rotary 
unions.  They have asked we continue the installation of the motors and 
solenoid manifold and hooked up all cables and sensors with the exception of 
head 2 which is awaiting a new spindle spline shaft from speedfam. Darren has 
also asked that the vacuum elbow/pressure gauge assembly not be hooked to 
the rotary union. All else should be connected, arms have been moved out to 
facilitatedinstallation of "that last bottom bolt". New vacuum tank installed.


					A shift
17.159Spindle workFABSIX::R_GEEWed Jan 17 1996 07:0759
	Action taken tonight:
	
	Head #1: 
		* Completed:

		  - installed, connected and tightened down the rotation
		    motor.
		  - installed the new middle and bottom solenoids with
		    new fittings and hose.
		  - installed the quick disconnect hub and bellows.

	        * left to do:

		  - install the gauge and stainless fittings going to
		    the rotary union.

	Head #2 : Nothing done.

	Head #3 : 
		* Completed:

		  - installed, connected and tightened down the rotation
		    motor.
		  - installed the new middle and bottom solenoids with
		    new fittings and hose.
		  - installed the quick disconnect hub and bellows.
		    NOT COMPLETELY ON OR BOLTED UP.

	        * left to do:
		  
		  - Finish installing the quick disconnect hub.

		  - install the gauge and stainless fittings going to
		    the rotary union.

	Head #4:				  
		* Completed:

		  - installed, connected and tightened down the rotation
		    motor.
		  - installed the new middle and bottom solenoids with
		    new fittings and hose.
		  - installed the quick disconnect hub and bellows.

	        * left to do:

		  - install the gauge and stainless fittings going to
		    the rotary union.

	Head #5: Nothing done.


	Carriers: Nothing done.
	
			B shift


	   

17.160FABSIX::B_FINNWed Jan 17 1996 19:3015
                                                 



			
	Cleaned/removed score on head 3 quick disconnect an reinstalled quick 
disconnect on head 3. Reinstalled motor and hardware solenoid manifold on 
head 5.  Installed quick disconnect on head 5.

Awaiting spindle spline shaft from speedfam on Head 2.

Darren says there is nothing to do on B1 tonight.

						A shift

17.161from the cats at speedfamSUBPAC::CMPThu Jan 18 1996 19:315
    brian, install left side terminal block plate , perform loadcell cals,
    change pad, wipe down inside of multihead,install top plate of
    multihead, and monitor vacuum tank for any possible leaks, do not worry
    about cycling wafers we will need to take a look at the robot in the
    am.
17.162grub work...SUBPAC::BJUBINVILLENFD Fire - RescueFri Jan 19 1996 06:577
    
         I started on some of the items in the previous reply but
      duty calls.  I got the top panel in place and was starting
      on the terminal block panel.  The load cell is on and is 
      warming up for the load cell calibrations.
                    
                                                    BJ
17.163down force and pmFABSIX::B_FINNFri Jan 19 1996 19:109
	Down force calibration successfully performed.  Robot would not home, 
traced to theta axis.  Speedfam Reps have hand carried the robot to ADE 
Newton,MA. for repair. Weekly pm was performed but we could not log time on 
RTC due to down status.  Check Sheet was submitted to notes file.

45 32" IC 1000 removed from fab 6 stock and placed in our cabinet
 
					C shift
17.164ROBOTSUBPAC::CMPFri Jan 19 1996 19:553
    HEY GUYS....THIS IS DARREN.....WE'VE GOT A ROBOT COMMING IN FROM
    PHOENIX TONIGHT...WE PICK IT UP AT LOGAN TOMMORROW AM....WE WILL BE IN
    TO INSTALL IT AND START CYCLING WAFERS.....
17.165FROM SPEEDFAM DUDESSUBPAC::CMPSat Jan 20 1996 17:085
    WE FINISHED THE ROBOT INSTALL AND CALIBRATION....CYCLED WAFERS FOR
    TRANSFER AND ON THE PAD...
    TO DO NEXT:  WIPE DOWN MHTA
    		 INSTALL COVERS
    		 START MQC'S OR ???????
17.166Carrier #1 stops rotating during polishFABSIX::R_GEESun Jan 21 1996 07:0716
	Replaced the rest of the covers and began running verification
	wafers. Every third run during the polish cycle the system would 
	error for "carrier #1 not rotating". 
	We noticed that on each run carrier #1 would slow down or sometimes
	stop.  The LED L1 on the Infranor for carrier #1 would light 
	indicating amps are to low.
	
	We checked and reseated the wire connectors going to the Infranor 
	and at the bulkhead connector panel on the multihead,  still not
	working.
.
					B shift


	   

17.167Head 1 rotation problem....continuedSUBPAC::LANDRYSun Jan 21 1996 19:3640
      Well, we looked into the Head 1 rotation problem tonight.  
      We unfortunetly have not found the problem as of yet.  The 
      system was returned to manufacturing to run on 4 heads for 
      the time being.

      A call was placed to the Speedfam Hotline.  No contact was
      made with anyone technical, though.  The answering service
      was to dispatch the call to the technical staff.

      So, here is what was checked and the outcome of each:

      *  Re-checked all connections at the Infranor and power supply.
         No problem observed.

      *  Ohm'd the wires out from the Infranor to the motor.  No
         problem found.  There were a couple of military connectors
         that we were able to tighten a bit, but don't suspect these
         to be of any consiquence.

      *  Checked the connectors under the flex track on the left side
         of the machine (table gear box side).  Again, a couple of
         connectors were not tight, but still don't think they were
         loose enough to cause the problem.

      *  Checked the connections at the DCX board (#2) and found the 
         connector associated with head 1 to be cocked a bit.  We
         thought we might have had it when we exercised and the first
         try at rotation was successful.  Tried to repeat the success
         but the next attempt, the Infranor L1 light came on dimly BUT
         the head continued to rotate.  A third try was a disaster as
         LED L1 light lit up good and the head promptly stopped.

      *  Some slop in the motor backlash was then looked at.  It was 
         adjusted, but it wasn't of any help.

      *  We then discided to halt at this time and turn 4 heads over to
         manufacturing for the shift.

      					A shift
17.168Carrier Arms not oscillatingFABSIX::R_GEEMon Jan 22 1996 06:5612
	The polish arms failed to oscillate in. They would move 
	to the outside OD position and stay there for the remainder
	of the polish cycle. No alarms or errors were displayed.

	Reset the computer and the problem went away.

	
					B shift


	   

17.169current reference adjustmentFABSIX::B_FINNMon Jan 22 1996 19:2311

	System had been running 4 heads due to intermittent current error. 
	Reps adjusted current reference for spindle rotation on head #1.
	Ran 3 runs to test system. Looks Fine. Turned back to manufacturing
        in "verify". Tim has asked that they run new MQC and OLD MQC ( lifting 
        off in the center of the pad and slightly off the pad, respectively), 
        and then inspect for Scratches.


						A shift
17.170ADE Robot failureSUBPAC::LANDRYTue Jan 23 1996 19:2516
    The system is down for an ADE robot failure.  Seems to work fine until
    it pauses while the multihead makes a wafer pickup at the index table.
    When the robot tries to resume working, it just sit there, ten times
    out.  We've rebooted a couple of time today to get there wafers through
    the system, but it's currently logged as DOWN.  The power supply was
    pulled out of it's casing for trouble shooting purposes.  It cycled 50
    wafers just fine.  As the power supply is put back into the casing, the
    problem re-occurs.  May be a connection/wire problem as these flex when
    the power supply is pushed back in.  Daniel and Rich are looking into 
    some error codes associated with the problem.  They have hooking up the 
    software (in our PC) for the ADE robot as I type.  They may need to
    find some specialty connectors/wires that we don't have in the room. 
    So, they may be asking for help finding it here.
    
    					A shift  /  S/F reps
    
17.171ADE robot still not functioningSUBPAC::BJUBINVILLENFD Fire - RescueWed Jan 24 1996 04:5027
    
         Took over running wafers through the tool after the Rep.
      went home.  We were using the TRANSFER program and bumped
      Segment #3 process time to 90 seconds.  We ran 75 wafers
      with this program.  Twice the robot went in and failed to 
      retrieve a wafer.  Both times the wafer count was around
      slot 11-20.  The error message indicated that the robot
      had missed the wafer and to just put the wafer on the pre-
      aligner.  For the last 25 wafers, we ran I_O_ILDA.  The 
      robot failed when it was ready to load wafers 6-10.  The
      robot didn't even try to go for the wafers.  The 3rd LED 
      is lit at this time.  It will only clear if the power is
      cycled to the controller.  I hooked up the PC as directed
      by the Rep. and checked the Robot Status screen.  I found
      bytes with a value of zero at the time of the error.
      
                               Byte #2  - Arm Ready
                                "   #11 - Brake Not Active
                                "   #15 - Servo Motor Inactive
    
         We tried several different profgrams and the tool will
      either it will fail right away or load 5 then fail.
                                                  
         The tool is in WET mode with the Robot still errrored.
    
                                                 BJ
    
17.172Robot problemsFABSIX::B_FINNWed Jan 24 1996 19:2216
	Robot was removed from the system and hand carried to ADE. ADE checked 
the robot out and found no problems.  Robot was brought back and cycled tested
300 times via pc and external software, without error.  Robot was then
reinstalled. Reps then found broken wires in the rs232 connect. Repaired.
They feel this happened today since this connection was checked last night 
by Rich and he saw no broken wires. ADE did suggest that it may be a noise 
problem since it occurs on the fifth wafer after multi head moves out. Reps 
just tried to disconnect ground , but the multihead is not liking this.  
Their still working on it.

					C shift


 

17.173VERIFY COMPLETED, NO ROBOT ERRORS, UP AND IDLE!!!SUBPAC::BJUBINVILLENFD Fire - RescueThu Jan 25 1996 05:5211

        After 25 more wafers were run with the TRANSFER program,
     25 more were ran with a 10 sec. polish.  A total of 100
     wafers were run with no robot timeouts.  Rich ran the MQC's
     and got scratches.  The program was coming off the edge, 129.  
     B2 was running at 120.  I had him run the MQC's with 120 and 
     the tool passed MQC's with no scratches.  The tool is finally
     UP and IDLE.  

                                         BJ
17.174SPEEDFAM REPS.SUBPAC::CMPThu Jan 25 1996 14:473
    INVESTIGATING PROBLEM WITH ROBOT, FOUND THAT IT IS INTERNAL TO ASSY.
    WE ARE LOOKING INTO A FLAKY CONNECTION TO THE BRAKE CAUSING THE
    Z AXIS TO FAIL.
17.175old robot out...new one going in!SUBPAC::LANDRYThu Jan 25 1996 19:4241
      The system is back down for a robot problem.  Many things were
      checked/rechecked today.  It's currently pointing towards the 
      robot, itself,  malfunctioning.  Speedfam reps are currently in 
      transit to the manufacturer (ADE in Newton, Ma), where they will 
      pick up another robot.  Ah, alas, here they are now with the
      new robot.  That was quick!  It will be installed now.
      Quick days happenings..........

      *  The robot failed early in the shift.

      *  It was suspected that the wire connector for the brake was the
         problem.  The wire connectors were soldered and testing was
         attempted.  

      *  Things were looking fairly good at this stage.  Therefore,
         position calibrations were performed and more testing commenced.

      *  The system ran into trouble, not picking up the wafer from the
         cassette.

      Other things as an FYI

      *  The robot has been removed from the system

      *  The squeak still exists in one of the carriers when rotating
         CLOCKWISE in the rinse station.  No squeak is exhibited when
         normal rotation (CCW) occurs.

      *  Couldn't get the system to pass initialization after the robot
         was removed.  Tried editing the HARDWARE.CMP file to remove
         the robot, pre-aligner, and mapper to see if we could get
         past this problem, no luck.  Upon startup tonight/tomorrow,
         the above three line items need to be verified as "on" in
         the above file.  1 = on     0 = off

      *  The system is currently powered down.  Thw carriers/pad are 
         being kept wet by the spray hoses.

      						Brad  /  Tim  /  Doc

17.176SMASHED WAFERSFABSIX::B_WESTPHALThu Jan 25 1996 19:454
CMP.B1 - Down for slipping wafers on the polish table.  Wafers from carriers 
	 4 & 5 were smashed while running product.  Two carriers were damaged,
	 the same were replaced along w/ pad. Tool is back to production for
	 MQCs.
17.177robot installationSUBPAC::BJUBINVILLENFD Fire - RescueFri Jan 26 1996 06:2614
    
         I tried... I tried...  I tried...
    
             Worked on the robot alignment all night and
         came real close but not enough.   If the (3) 
         robot bolts next to the leveling screws were a
         1/2" longer, I would have it.  I also found the
         metal strip that channels the vacuum on the end
         effector had broken its silicone or epoxy seal.
         This was without a doubt a direct problem with 
         the robot pick-up.  Wether it was like this all
         week or just from yesterday I don't know.  
    
              It's all your guys....         BJ
17.178ROBOT TROUBLEFABSIX::B_WESTPHALFri Jan 26 1996 19:0911
CMP.B1 - Down for problems with the robot.  Received errors for the Z axis 
	 while trying to level and cal the robot this morning.  The robot
	 was pulled and tested against the one pulled yesterday using different
	 cables and controllers.  The best combination was reinstalled.  Had  
	 problems with the wafers not being picked right because the wafers and
	 end effector were not parallel with each other.  A new end effector 
	 was installed and #2 cassette platform was shimmed accordingly to 
	 correct this situation.  The cal was completed and 50 wafers were ran 
	 to check for any further handling problems.  The system is back to 
	 production for MQCs.
	 	 
17.179power glitch wafer detect systemSUBPAC::BJUBINVILLENFD Fire - RescueSat Jan 27 1996 16:4513
              from this morning, ASDG was down...
    
                The tool passed MQC's and OPS ran production
             lots on it all morning.  There was a power clitch
             in CMP around 7:15am and the tool started having
             problems as well.  The tool was indicating that
             that the wafer detect system was picking up a 
             wafer slide-out and stopping the carrier rotation.
             If this is still happening during the day, have the
             Speedfam REPS look at the detection sensors to make
             sure they are ok as far as alignements are concerned.
    
                               D shift
17.180Robot ProblemsFABSIX::R_POIRIERSun Jan 28 1996 07:5412
    CMP.B1 - ROBOT PROBLEMS
    
    	Robot would not move when it was supposed to pick up wafers.
        We tried a number of combinations (new robot/new controller,new
        robot/old controller, etc.) and eventually ended up with the old
        controller and new robot.  There is still a problem with alignments
        for the robot. These alignments are necessary because the robot
        is coming in to high in the cassette.
    
    
    					B Shift
    
17.181power glitchSTRATA::FINNSun Jan 28 1996 19:2815
	Checked robot in both manual and auto, cassette 1 is fine but cassette 
two need to be leveled, (we say cassette not robot because robot is fine on 
cassette one). Changed pad and let them run cassette one to move the six lots 
in the room.

	Had power glitch and evac. Soaked carriers and conditioned pad. MQC's 
passed fine.

System is up and running 5 heads with warning on load cassette # 2.

 CMP.b1   1/28/96  Q CR QJMR0313a,P/II/S-IV   Pad was visibly worn
		   Lot #LR5-JP9                
    
					a shift
17.182robot and cassettes adjustedSTRATA::RIDLONEliminate the obviousMon Jan 29 1996 07:029
    Adjusted cassette 1 home and adjusted robot level so that when wafers
    were removed they wouldn't scrap in the cassette slots. Checked the
    alignment on cassette 2 and it was off. Removed the shims that C shift
    had installed and installed some in another area so that the wafers
    were level with the robot. Recalibrated cassette 2's home and then
    checked wafer removal and it was great. Placement on prealigner for
    both cassettes was also fine. 
    
    					 B shift
17.183Vacuum Pump fan brokeFABSIX::R_POIRIERTue Jan 30 1996 07:028
    A whining noise was noticed coming from the vacuum pump. The 
    fan on the left side of the Vacuum pump broke. It was up against 
    the fan cover and wasn't spinning. We replaced it with a fan from 
    the old Vacuum pump.
    
    
    					B Shift
    
17.186Carrier #1 not rotating error.FABSIX::B_WESTPHALSat Feb 03 1996 19:295
    CMP.B1 - Tool errored for carrier #1 not rotating while product was being run.
	 The operator reported that the wafer from carrier #5 was left on the
	 pad unbroken and the screen was locked up on him so the tool was 
	 rebooted.  A 15 wafer transfer was also ran by production to verify 
	 the tool.  The incident was then logged as an assist.
17.184System powered down / Carrier rotationFABSIX::R_GEETue Feb 06 1996 06:2515
	The system powered itself off in the middle of an
	Auto polish cycle, no reason found.

	Powered the system back up and began running dummy wafers.
	On the first run the system errored for heads #1,#2,#3 not 
	rotating. No error lights on the Infranor or DCX boards,
	everything was normal. The voltages at the Infranor were also
	correct. Powered the system up and ran 50 dummy wafers with
	no errors, released it back to production.

					B shift


	   

17.185new connectors for DCX/Cable connectSUBPAC::LANDRYTue Feb 06 1996 19:1910
      Today, the reps installed new connectors to the ribbon cables
      that go to the DCX boards.  The connectors were of the "clip"
      style, which locked the connector to the board, hence prohibiting
      the cables to become dislodged when cables are moved.  They also
      verified proper installation prior to returning to manufacturing
      hands.  Opts personnel ran 25 exercise wafers and MQC's to also
      assure functionality.

      					A Shift

17.187Carrier #1 rotation, pad changeSUBPAC::BJUBINVILLENFD Fire - RescueThu Feb 08 1996 06:2123


       The tool failed for MQC's once again.  The pad had almost
   700 wafers on it and the carriers all had near 1000 wafers each.
   We pulled off the carrier and pad.  We noticed a build up of
   some sort of a grime/grease on top of some of the carriers.
   These were the same one on since the particle/multihead cleaning.
   After everything was cleaned up and re-installed.  We had two
   wafer fly-outs on different occasions as well as several Carrier #1
   not rotating errors.  We checked the alignments and #5 and #1 were
   out but not enough to cause a problem and weren't the ones that
   had the fly-outs.  We checked the newly installed carriers with all
   of the stainless washers and found an exposure of -.011 to -.013 on
   all (5) carriers.  The initials on the carrier "tags" were "SG."  I
   don't know who that is but could someone notify her/him of the
   exposure issue?  We then looked at the 2nd problem which was with the
   carrier #1 rotation.  We cranked up the voltage to the max. (15.17V)
   as directed and gave it a shot.  The error was still there.  We pulled
   the carrier and ran with four heads.  MQC's came in fine.  I'm
   placing a WARNING for the HEAD #1 issue.

                                                   D shift
17.188High ParticlesASDG::POIRIERThu Feb 08 1996 20:3613
    
    * High Particles:
      --------------
    
      The system was thoroughly cleaned and MQC's were run. It passed
      for particles but failed for uniformity on one spindle. (10.5%)
      
      This may be due to only 30 break-in wafers being run. 20 more
      break-in wafers were run and a uniformity measurment is in
      progress.  We expect the uniformity to come in shortly.
    
      NOTE: We're currently running critically low on oxide monitors!
    
17.189Carrier #1 Rotation Motor will be replaced.ASDG::POIRIERThu Feb 08 1996 20:4613
    
    * Carrier #1 Rotation Problems:
      ----------------------------
     
      Carrier #1 Rotation errors have been continuing to plauge the tool
      all day. The Speedfam reps changed the Infranor motor controller
      and no change was noticed. They then measured the voltage on the
      Infranor IDC tap (pins 1 & 6) under load on spindle #1 & spindle
      #2. The voltage on spindle #1 was twice the voltage on spindle #2.
      Speedfam has concluded that Carrier #1 Rotation Motor is BAD. 
    
      The tool will go back "UP" tonight with 4 spindles active. The 
      motor will be replaced when a new one is recieved on-site. 
17.190Head rotation problemSUBPAC::BJUBINVILLENFD Fire - RescueFri Feb 09 1996 05:1811

       Went through some more things one head #1.  Dan and I
    checked the downforce (which was in) as well as some DCX
    and offsets to make sure nothing was changed.  A new motor
    is coming hopefully today.  Head #1 is still off line and 
    has a WARNING on it.  All of the panels except for the two
    for head #1 were put back on the tool.  A new pad was installed
    and was brought back up for Production.

                                                 BJ
17.191head-1 failureSTRATA::TDYERFri Feb 09 1996 18:3823
    
CMP.B1 - Down for head-1 troubleshooting. The system will be returned in the
         next hour with FOUR heads. Here's what we found today:

           Swapped the motor and resolver cables for head 1 and 5 on the
           RIGHT side of tb-13.
         * The problem moved to head-5 (head-5 stalled)
           We put the wires back to normal on TB13
           Swapped mtr-3 cable (head-1 rot) with mtr-12 cable (head-5 rot) and 
           their respective resolver cables on the LEFT side of the bulk-
           head connections under the polish area.
         * The problem moved to head-5.
         * So, we have a bad cable between the bulk head connector and the
           INFRANOR for head-1 rotation. (the infranor was already replaced)
           The connector itself could be bad,please inspect. If nothing 
           obvious is found, put system back together with heads 2-4 ONLINE.
           The reps will replace the cable next week.

         - At one point both head-5 and head-1 stalled, for this reason I have
           also request a flex track motor cable.
    
           TD/JP/Rich
17.192Head #1 rotation failures / new cables on the way.ASDG::POIRIERFri Feb 09 1996 19:328
    
    * Head #1 rotation problems:
      -------------------------
      After further troubleshooting, it was determined that the PAC-SCI
      Rotation motor for head # 1 was O.K. We now feel that a cable is
      causing the intermittent rotation failures on head #1. See Tim
      Dyer's passdown for a detailed explanation of the troubleshooting
      activities that took place today. 
17.193left send mapper controlSUBPAC::BJUBINVILLENFD Fire - RescueSat Feb 10 1996 07:3812
      The tool mapper was intermittingly erroring out and 
   was not mapping the cassetttes.  Swapped cable outputs
   and the problem went to the right send unit.  After than
   I swapped the cables all together and the there was no
   change.  The problem is within the Kennsington controller
   itself.  Until a board is ordered, only the right send
   elevator can be used.  On top of that, it has to be manually
   mapped (ie: tell the system where the wafers are).


                                      BJ
17.194scrubber upgradeSUBPAC::LANDRYTue Feb 13 1996 19:2612
    The scrubber motor failed today.  First thought it to be the wire,
    but a quick check indicated said wire was fine.  The motor torque
    was examined.  I could get it to spin with a little twist of the
    coupling.  Well, it was decided on that the scrubber up-grade would
    now take place.  At the time of this typing, all components have
    been installed.  BUT, while calibration was taking place, the
    adjustment pot on the board was dislodged.  Reps are attempting
    to re-attach the pot.  May need to improvise with another pot
    should this one not survive.
    
    						A Shift
    
17.195Oscillation positioning of carriers completedSUBPAC::BJUBINVILLENFD Fire - RescueFri Feb 16 1996 07:2931
 
       The arm oscillation calibration was performed and found
    all heads out of spec.  Below is the original results along 
    with the new results.

       The measurement is measured from the center of the table
    to the center of the head post which should be 14".

        HEAD         OLD         NEW
        ============================
        #1           12.64"   13.99"
        ============================
        #2           12.72"   14.04"
        ============================
        #3           12.85"   14.04"
        ============================
        #4           12.82"   14.04"
        ============================
        #5           12.76"   14.04"
        ============================
        

         the blocky style conditioner wheel was installed on the
    tool and broken in with ~1 hour at 67% downforce.  The tool
    was cleaned and the carriers were re-installed.  A new pad was
    installed but has not been broken in or conditioned.  It is still
    new and hasn't seen water on it yet.  I'm going to leave the 
    carriers rotating and rinsing.  I'll let EE decide about the
    new pad conditioning.
                                               BJ
17.196B1 now matches B2FABSIX::B_WESTPHALFri Feb 16 1996 19:405
CMP.B1 - The tool was given to process engineering and all of the process
	 parameters were adjusted to match B2's.  A door that protects the
	 track that the multihead rides in was damaged.  The reps. are ordering 
	 the new parts.

17.197scratchesSUBPAC::BJUBINVILLENFD Fire - RescueSat Feb 17 1996 07:127
      System was down at the beginning of the shift for scratches.
   The pad was replaced and broken-in.  There were no MQC wafers 
   available to test the machine.  Tool passed MQC's for uniformity
   and no scratches.....

                                          D shift
17.198Head #1 not rotatingFABSIX::R_GEESun Feb 18 1996 06:4710
	* The system kept erroring for head #1 not rotating.

		Tony (the field service Rep.) adjusted the 
		voltage at the Infranor from 12.5 to 15.0.

				B shift


	   

17.199Slurry filter changedFABSIX::R_GEEMon Feb 19 1996 02:034
    	Replaced the slurry filter do to unusual slurry flow. It
    	was last changed on 1-30-96.
    
    					B shift
17.200Scratches... still scratchingSUBPAC::BJUBINVILLENFD Fire - RescueWed Feb 21 1996 07:3315


      Still down for scratches.  After is failed once again
   for scratches, the pad was replaced.  It failed again for
   scratches.  I didn't come across any passdowns that said the
   old conditioner was put back on so we tried that and we
   still had the scratches.  We tried running the MQC's at
   115 on the pad with no oscillation and there were no scratches.
   We now would like to see what edge is causing the problem but
   we have no MQC wafers as of 7:30am and we won't know until
   then.  I would "hope" to see two wafers run from 115 to 95 
   and two run 115 to 135 to see where the scratches are coming
   from.             
                                       D shift 
17.201slurry pumps linesFABSIX::B_WESTPHALFri Feb 23 1996 11:125
    CMP.B1 - Down for a leak in one of the slurry pump lines.  Both line
    	     were replaced but one was replaced with medical grade
             masterflex tubing as recommended by Speedfam.  This was done
             so a comparison could be made between the 2 types of tubing. 
    	     
17.202Inverter errorSUBPAC::BJUBINVILLENFD Fire - RescueSat Feb 24 1996 05:1510
 
       Countless time the tool errored for "Inverter Error, record and
    hit button to continue."  We've yet to find the source of the problem.
    The error comes up over the index table during carrier clean/wafer 
    rinse after polish.  Kev from SpeedFam call in to check on things 
    and Barry told him about.  He said is was a software issue.  Instead
    of OPS logging a million ASSISTS, I threw a WARNING on it.

                                        BJ
17.203Inverter Timeout Error-Table RPM / Head #4 Rotation Errors.ASDG::POIRIERSat Feb 24 1996 19:2416
    
    INVERTER COMMUNICATION ERRORS:
    ------------------------------
    The system errored numerous times Today for "Inverter Timeout".
    The Speedfam rep (Tony) feels this is due to a recent change in
    the Table RPM. According to Tony, the system is trying to start
    the Table Rotation command but the Table hasn't stopped rotating 
    from the previous Segment. This causes a "Wait" error & eventually
    an Inverter Timeout condition. He is adding an extra Segment
    to the recipies to give the Table extra time to slow down.
    
    
    CARRIER #4 ROTATION ERROR:
    -------------------------
    Spindle #4 errored for "Rotation Failure" six times Today. 
          
17.204Broken waferFABSIX::R_POIRIERTue Feb 27 1996 06:5116
	Wafer slipped out of carrier # 4 and was crushed by carrier # 3. 
    	It appears the wafer detection didn't catch it in time.  We 
        noticed that the polish table appeared to be dry. After re-booting 
        the system we checked the slurry flow and saw no problem. Just 
    	before the crash we heard the table vibrating. Possible reason 
        for crash could have been no slurry flow. Nobody was watching the 
        tool when crash happened. The tool had been running all night. 
        We replaced three carriers. # 3 and # 4 due to the crash and # 5 
        because of bad uniformity. After replacing pad we ran 30 wafers 
        without any problems. 
	
	Carrier 4 was a 21 hole carrier as was carrier 5. # 1,2,3 were 
	all 29 hole carriers.
    
    					B shift
17.205mapper problemsSTRATA::FINNTue Feb 27 1996 19:3823
	Wafers dropping off the carrier at liftoff.  Process changed 
oscillation so liftoff occurs further out off the pad. 

	System errored for "Waiting for mapper response out".  Cassette would 
not see home switch as it tilted back.  We have no schematics. Swapped cables 
at mapper and problem moved to cassette two. Swapped both ends of both cables 
and problem returned to cassette one. Swapped tilt boards within the 
controller and problem stayed with cassette one.  Swapped cables at input to 
the controller and problem moved to cassette 2.  Problem is definitely in the
controller.  On installation the system magically began working. It remained 
working for one hour then failed for the original error. We disabled the 
controller. Operator must manually tell the system via the map page which 
slots contain wafers.  Cassette 2 is not working it wont tilt, cassette 1 is 
working. System is on warning. Tried to put cassette 2 on warning but rtc is 
down.  

	We notified reps that we took the controller off line, they may be 
able to get one here from Speedfam by tomorrow morning.  This is still a maybe,
Kevin or tony will call the room and confirm this availability in one hour.


		                        	A shift

17.206mapper info / vacuum tank dumpingSUBPAC::LANDRYWed Feb 28 1996 19:3718
    The Mapper situation was looked at today.  New parts installed, but
    unfortunetly the mapper is in the same shape.  It still is DOWN for
    automatic running.  Manual mapping is still required.  See Tim's
    daily passdown for game plan.
    
    Due to manual mapping, the vacuum tank needs to manually dumped every
    50 wafers.  Again, see Tim's instructions on procedure.  Was last
    dumped at 7:20 PM.
    
    Elevator #1 messed up at the end of the shift.  It would not home for
    the Opts folks.  On my first look at homing it, I noticed that the
    motor was driving, but extremely slow.  Once home position was finally
    made, the elevator motor drove at normal speed.  Retried homing
    sequence four times after that, with normal homing speed/performance
    every time.
    
    					A shift
    
17.208Crash!, No slurry flow!SUBPAC::BJUBINVILLENFD Fire - RescueThu Feb 29 1996 08:5222
        
    
          The system crashed early.  It appears that there was very
      little slurry on the table.  A wafer was dragged out of #4 and
      took out #3.  The slurry flow was checked and both pumps were
      ok.  We replaced the filter to make sure there was no blockage.
      We replaced the slurry lines just to be safe.  The medical line
      on pump #1 had a slit in it so we might of saved a blow out.
      After the pad was replaced and break-ins started, it was noted
      that there was no slurry coming out of pump #2.  After trouble
      shooting #2, it had juice but wasn't turning.  We checked all
      of the wiring and everything appeared to be ok.  We took the
      pump out of the machine and it worked???  We pulled the brushes
      out of the motor and they were pretty well spent.  Since this
      pump/motor is intermittent, Tim is having SpeedFam pull the pump 
      for the un-used #3 pump for the time being.  The (3) remaining
      carriers are on the table soaking next to the sink.  The pad 
      has seen lots of debris so it will have to be pulled.  Lots of
      hardware so have lots of fun...  call if me if you guys get
      snagged.
     
                                                BJ
17.209Slurry motor/kennsington controllerFABSIX::B_WESTPHALThu Feb 29 1996 16:257
CMP.B1 - Down from previous shift for a wafer crash.  Low slurry flow due to an
	 intermittent problem w/ slurry pump #2 was believed to be the cause.  
         Pump #2 was exchanged w/ #3 and both #1 & the now #2 pumps were 
	 calibrated.  The problem with the wafer mapping was also addressed by
	 Speedfam. The kennsington controller was swapped out w/ an old part 
	 that was tested good.  System seems to be mapping alright.  The tool
         was cleaned up and, a new prestamped pad was installed.  
17.210Spindle Rotation Error on Heads #4 & #5ASDG::POIRIERFri Mar 01 1996 19:277
    
    * Spindle Rotation Error:
      ----------------------
      The system errored for "Spindles 4 & 5 Not Rotating" Today.
      Circuit Breaker #6 was reset and the tool ran for the remainder
      of the shift with no further errors.
    
17.211mapper error is backSTRATA::FINNTue Mar 05 1996 19:4018
	Five breakin wafers were being run prior to product. Carrier 1 spit 
the wafer on the pad. Carrier 5 survived minus it's white ring, but carrier 
one needed to be replaced due to small tear in the carrier film. Pad was 
changed. 

Breakin wafers were started

	System errored for mapper at 7:05 PM as we were watching the screen.
Both cassettes A and B failed.  Systems exhibits same symptoms as last week. 
Cassette is also continously rocking back and forth looking for it's limit 
switch, as it did last week. Tony (SF) has been notified and will be onsite 
in approximately 30 minutes.  System was manually told which slots contained 
wafers so manufacturing could continue.


					A shift
                              
17.212Mapper issueFABSIX::B_JUBINVILLEWed Mar 06 1996 19:269
    
    
        Tony came in looked at the mapper issue.  He found wafer debris
     under the left send elevator.  The wafer debris seemed to of have
     be causing the elevator not to tilt properly and therefore causing
     a fault in the Kensington.  Ran 25 wafers using the transfer program
     without any problems.  System was turned over to production and the
     WARNING was lifted.
                                             BJ 
17.213head rotation problemsSUBPAC::BJUBINVILLENFD Fire - RescueSat Mar 09 1996 04:527

       System errored for head #4 not rotating after lift off and
    wouldn't clear.  The breaker the the head had tripped.  Reset
    breaker and rotation checked out ok.  Tool ran fine after that.

                                           BJ
17.214Pad change/Carrier #4FABSIX::B_WESTPHALSat Mar 09 1996 19:426
CMP.B1 - Tool went down for low removal rates.  The pad was replaced and 
	 process area was cleaned up by production.  While getting ready 
	 to run the breakins we received a carrier #4 rotation error.  No
	 breakers were tripped but the infranor was reading motor over temp
	 error.  Reset the system and the carrier ran without any further
	 errors.  Currently running MQCs.
17.215carrier 4 rotation errors and then someSTRATA::RIDLONEliminate the obviousSun Mar 10 1996 05:3514
    Had two instances tonight where carrier 4 failed to rotate during
    polish. First time we ended up emoing the tool to correct the problem.
    The second time I had the bright idea of just powering down carrier 4's
    infranor to clear the errors. The good news is that it worked the bad
    news is that the circuit breaker that powers mc10 also powers the
    mc13 (multi-head transport). Even though the multi-head was locked in
    at the home position it still lost its way and after polish it stopped
    halfway on the way back to the index table. Carriers 1 and 5 dropped
    onto the edge of the index table and pulverized the wafers. 2,3, and 4
    were ok. 1 product and 1 dummie were wasted. Cleaned up the mess,
    changed the carriers 1 and 5 (inspected the rest), and changed the pad. 
    
    
    						B-shift
17.216Carrier #4 No RotationFABSIX::B_FINNSun Mar 10 1996 19:1024

	Carrier four no rotation.  Passdown shows this as reoccurring 
problem.  Infranor for carrier 4 rotation showed all three lights 
illuminated indicating a motor temp error. We swapped carrier rotation 
#4 Infranor with that of carrier rotation #3.  After 20 wafers problem 
returned and did not follow the Infranor, it stayed on Infranor attached to 
motor/carrier # 4 indicating a real temp problem or intermittent 
connection problem. Checked rotation # 4's PTC  on terminal strip # 12, pins 
57 and 58. Found voltage to be + 57 millivolts.  Head 5 was 
+ 67 millivolts the rest were lower lower including a - 12 millivolts on 
head 3. Ohmed Cable back to X1 pins 1 + 2 of Infranor looked good. Tried to 
stress cables looking for intermittent opens found none. Tried system. 
Failed at wafer clean prior to wafer pickup for #4 rotation.  Found a 
REAL loose military connector on # 4 motor. Tried running in manual mode at 
both polish table and at index table to simulate failure, while monitoring 
tc with meter, but system would not fail. Changed program to a no 
touch, system ran fine. Tried breakin program system ran fine. Tried to give 
system back with all five head, but Jesus wanted to run four heads and 
test five heads later. System is up with head four off line. Carrier has been 
removed and is in water. 


						A shift
17.217carrier 4 workSTRATA::RIDLONEliminate the obviousMon Mar 11 1996 06:5613
From the work that A-shift did we thought it would be a good idea to 
install carrier 4 and re-mqc the machine. The mqc's passed but when 
they went to run product it faulted for carrier 4 rotation. Every time
we've seen it fault it's been the first or second set of wafers and always 
before touch down. This time we turned of the breaker to reset the 
infranor and, learning from our mistakes, we homed the multi-head. 
It finished the product lot without error. AFter they were through with 
the machine we decided to install a resistor on TB-12 pins 58-57 to bypass the 
motor tc. 30 wafers were ran with out errors. The resistor (100ohm) was 
removed and original tc wires reinstalled as well as all panels for shift 
change.

                                        B-shift 
17.218Mapper / Field Rep. beeper #FABSIX::R_GEEMon Mar 11 1996 07:0314
	At the end of A shift the mapper failed. After sitting for
	an hour or so it began working again. Field service was contacted 
	about the problem. It worked for the remainder of our shift.
	
	We were informed by Speedfam field service that Tony is
	nolonger in the area and to contact Kevin with any equipment
	issues. His beeper # is

			1-800-796-7363 pin # 1020584
	
				B shift

	   

17.219mapper failed/ offlineSTRATA::FINNMon Mar 11 1996 19:246
    
	Mapper failed. Shut mapper off at 5:30pm. Shut off controller and 
        reconfigured hardware file. Operators are mapping in manual.
	Reps are aware of the failure.
    
					A shift
17.220Mapper FABSIX::R_GEETue Mar 12 1996 07:3913
      * Send cassette #2 (B) wouldn't level for the robot to pickup
	the wafers. Since the mapper had 4 hours to cool down and
	is working again, we put it back on-line. 

      * Errored for spindle #1,#3, & #5 not rotating. No problems
	found pressed start and continued. The error never re-occured.
    
    
 	
				B shift

	   

17.221Mapper againFABSIX::R_GEEWed Mar 13 1996 06:5411
	The mapper failed again, we powered the mapper off and disabled
	it. An Assist and a Warning was logged in workstream to manually
 	map the cassettes. 
	After 3 hours we powered the mapper back up and enabled it,
	all is well for now. The warning was removed. If anybody is 
	interested, the mapper lasted 24 hours before failing.
	
				B shift

	   

17.222ground addedFABSIX::B_FINNWed Mar 13 1996 19:045
Reps added grounds to cassette tilts to help eliminate mapper errors.
Mapper is on , no problems today.

						A shift
17.223crashFABSIX::B_WESTPHALFri Mar 15 1996 19:4211
CMP.B1 - Down for breaking a wafer from carrier #3 in segment 5.  Suspect that
	 the problem was caused by a dip in slurry flow due to the operator 
	 reports of a loud vibration previous to the crash.  The slurry filter 
	 was checked and looked good but was replaced as requested by engineer-
	 ing.  The slurry flow for both pumps was out by about 100ml/min.  The
	 same was corrected.  The process area was cleaned up, pad replaced w/
	 a prestamped.  The index table was drained and wiped out.  Only
	 carrier #3 needed to be changed out.  Gave the tool back to
	 production. 

17.224Wafer breakageSUBPAC::BJUBINVILLENFD Fire - RescueSat Mar 16 1996 07:0016

    
           System slurry flow dropped to almost nothing and halted
        the table and caused a crash on a hot lot to boot.  I took
        a look at the slurry filter and it was caked.  I took the
        filter off of A1 to get B1 running.  Slurry flow was back
        to normal after that.   The was cleaned up and I put new
        carriers on #3 and #4.  MQC's were run and #3 was they only
        one that failed MQC's for Unif. twice, ~13%.  Due to the
        condition of the area, #3 was removed from the tool and
        the tool is running with (4) heads.  A WARNING is on the
        tool for #3.


                                               BJ
17.225carrier 4 not rotSTRATA::RIDLONEliminate the obviousSun Mar 17 1996 03:388
    Carrier 4 errored for no rotation during process. Same old error where
    the infranor faults for overtemp before the carrier touches down.
    Killed the breaker (cb#6) for that carrier, homed the multi-head, homed
    the index table and restarted. Finished the lot with no other problems. 
    
    
    					B-shift
    
17.226Carrier 4 rotation errorFABSIX::R_POIRIERMon Mar 18 1996 07:3511

Early in the shift we had a carrier rotation problem with 
Head #4. It was just before Touchdown. Infranor #4 alarmed
(L2-Sensor). Reset the circuit breaker CB6, re-homed the
Multihead, and continued running the tool.

Also, the right load cassette didn't map during a run. The
left one mapped ok, and the system didn't alarm out but the
right one had to mapped manually.

17.227carrier 4 rotation errorsLUDWIG::RIDLONEliminate the obviousTue Mar 19 1996 02:4214
    We've had a total of 5 carrier #4 rotation tonight. Took the tool when
    they were through and removed the covers for carrier 4. Found the
    PSI/VAC gauge destroyed from rubbing on the cover. The screws on the
    back of the gauge were found down below the spindle gear. Alot of
    plastic from the gauge was also found in this area. Looks like some of
    the plastic got into the spindle bearing. Rotation by hand seems to be
    smooth and is fairly easy. Cleaned up as best we could, replaced the
    gauge and adjusted the elbow so that the gauge wouldn't rub any more.
    Tried to run some breakins but it alarmed for carrier 4 rotation on the
    first set. Eight lots in the room other wise I'd tear the spindle out.
    The powers that be need the machine so carrier 4 has been disabled. 
    
    					
    					B-shift
17.228head 4 / mapperSUBPAC::LANDRYTue Mar 19 1996 18:3216
    The system went down for a pad change today.  While this was to happen,
    the Reps (Kevin and Ian) were allowed to take a look at the Head 4 no
    rotation issue.  I don't have the full story as I wasn't able to work
    with them on this.  This what I do know:
    
    	-  Head four will remain off line until some parts come in.  A
           cable is suspected as the problem.  I think it's a cable within
           the MHTA, but check with them for exact details.
    
    	-  The mapper also messed up again.  Same old same old.  They have
           some items coming in to use with the PC which will help in the
    	   trouble shooting scheme.  The mapper was shut-off for awhile
           and then turned back on.  It's currently functioning for now.
    
    							Doc
    
17.229No Head OscillationSUBPAC::BJUBINVILLENFD Fire - RescueWed Mar 20 1996 03:2915

  
         System kept failing MQC's for bad uniformity.  Too
    much edge removal.  During the running of the MQC's, I
    noticed the heads weren't oscillating at all when they
    were suppose to.  Aprrox. 60 wafers were were run with the
    heads half way off the inside of the pad, thus, too much
    edge removal.  The heads wouldn't oscillate in MANUAL mode
    either.  Hard booted entire tool and the heads functioned 
    fine after that.  OPS tried running MQC's again but the
    inside of the pad had taken to much of a beating.  The
    pad was replaced and MQC's were started over.

                                             BJ
17.230Head #4 Cable Upgrade / Load Cassette Mapper ControllerASDG::POIRIERThu Mar 21 1996 20:2615
    
    * Head #4 Rotation Cable:              
      ----------------------
      The Speedfam reps have completed the cable upgrade. A 50 wafer
      verification run using the Handling recipe has been sucessfully
      completed. The system is now running MQC's.
    
    
    * Load Cassette - Kensington Controller:
      -------------------------------------
      The Reps didn't have time to replace the Kensington Controller 
      this afternoon. Therefore, please continue to manually map
      the cassette and manually drain the vacuum tank tonight.
      The reps are planning to complete the Controller replacement
      tomorrow.
17.231Load Cassette Mapper Controller Replaced.ASDG::POIRIERFri Mar 22 1996 19:269
    
    * Load Cassette Mapper Controller:
      --------------------------------
      The Speedfam reps finished replacing the Kensington Controller
      at 7:15 p.m.. The system is now performing a 25 wafer verification
      run. After this run, full MQC's will be needed.
    
    
      They also repaired a door switch.
17.232transducer installedLUDWIG::FINNTue Mar 26 1996 19:4514

	Installed (temporarily) one transducer and one omega display to capture
pressure and vacuum on carrier 5. Installation is not only to to check for 
early mortality and judge the effectiveness of the transducer. If all 
works out we will in stall all five and mount their displays in a panel.

Index table would not home. It would move 2 cups then seize. DCX error 1 on 
board one. Power failed and check infranors aok reset DCX board cable 
and module, still had error. Check motor in DCX mode all ran fine, once table 
saw the home switch system ran fine, could not duplicate. Up and running


					A shift
17.233CMP.B1 shutdown for facilities tie-in completedSUBPAC::BJUBINVILLENFD Fire - RescueThu Mar 28 1996 06:1254
   

  SHUTDOWN:  5:00 AM ON 3/28/96

  STARTUP:   9:00 PM ON 2/28/96


                    CMP.B1 SHUTDOWN CHECKSHEET
                    --------------------------


                                                        Tech.       Date
                                                       -------     -------


  1.  Change RTC status to SCHEDULED DOWN.               BJ        3/28/96

  2.  Remove all carriers.  Any carrier that
      has a wafer count higher than 500,
      break the carrier down.  If any carriers
      have less than 500, place in DI water and
      and mark a indicated.                              BJ        3/28/96

  3.  Remove all wafer receive cassetts.                 BJ        3/28/96

  4.  Remove and discard all polishing pads.             BJ        3/28/96

  5.  Remove all PVA scrub brushes and submerge         
      in DI water.                                       BJ        3/28/96

  6.  Perform a complete clean of each system
      to remove dried slurry.  Wipe down and dry
      all wet surfaces.                                  BJ        3/28/96

  7.  Remove and discard all slurry filters.             BJ        3/28/96
                                                         
  8.  Disconnect inlet and outlet lines on each
      filter housing and remove housing.  Connect
      inlet and outlet lines together with a 3/8"
      union.  This will completely bypass the 
      filter housings.                                   BJ        3/28/96

  9.  Close main incoming DI Water valve to each
      system and open the DI Water return valve
      (UDR).  This will allow all incoming DI Water
      to flow to UDR.                                    BJ        3/28/96

 10.  Power down each system.                            BJ        3/28/96


 11.  Power down each Neslab TCU.                        BJ        3/28/96
    
    
                                                                       
17.234motor #5 problem, DI flow, slurry stuff..SUBPAC::BJUBINVILLENFD Fire - RescueFri Mar 29 1996 06:3331

           First off.. after the startup, carrier #3
      failed real bad for uniformity (18%).  Replaced with
      a new carrier.

           Found slurry dripping all over the conditioner
      arm and coming out of the multihead.  Blown slurry
      line, replaced them both.

           Multiple head #5 rotation errors, soft and hard
      boots fixed at first then it died all together.  I took 
      it off-line and paged Ian to let him know what's up.

           Once MQC's were ran, some small scratches were
      seen.  This is most likely caused from the slurry
      line that blew.  A new pad was installed and MQC's
      have run and we are waiting for the results.
   
           While admiring the polishing, excessive water
      was seen under the entire multihead.  While watching
      the carrier/wafer cleaning, gallons of water was spraying
      everywhere.  The DI supply inside the tool was set to 
      40 psi.  This could of been the cause for #5 to flake out.
      The DI was turned down to spec, 20 psi and there was no
      more excessive spraying.  Rechecked the rotation on #5
      and its still dead.  There was water found up inside the
      motor #5 cavity and around the newly mounted read-out
      gauge.

                                       BJ
17.235ARM 5 ROT ERROR AFTER GORE CABLE UPGRADEFABSIX::R_TRANFri Mar 29 1996 21:2621
    SPEEDFAM CONTINUED TO TROUBLESHOOT ARM 5 ROTATION ERRORS. FOUND THAT
    ARM 5 WOULD SERVO OUT WHEN POLISHING AT DOWNFORCE OF 377LBS, OK AT
    250LBS(ERROR WAS SHORT CIRCUIT!!).
    CHECKED DOWNFORCE CALIBRATION ON ARMS 1 & 5 ALL O.K. 
    CHECKED SLURRY FLOWS - ALL O.K
    REPLACED PAD AS BRAD BROKE CHIP OF GLASS FROM BEAKER WHEN CHECKING
    SLURRY FLOWS.
    RERAN MACHINE AND ERROR ON ARM 5 ROTATION HAPPENED AGAIN.
    SUSPECTED THAT AS  THE ONLY THING THAT HAS CHANGED IN THE SYSTEM IS THE
    NEW GORE CABLES IN THE MHTA THAT MAYBE THE IMPEDENCE OF THE CABLES WERE
    DIFFERENT CAUSING THE ERRORS.
    
    SWAPPED INFRANORS 3 & 5 AND RAN MACHINE WITH 20 WAFERS 10 ON I_O_BKIN
    AND 10 ON IAN PROCESS (I_O_MQC WITH REDUCED SEG 4 TIME OF 20 SECS).
    
    ROTATION ERROR HAS DISSAPPEARED. THE ONLY EXPLANATION WE CAN DETERMINE
    IS THAT THE ORIGINAL INFRANOR ON ARM 5 WAS MORE SENSITIVE TO IMPEDENCE
    CHANGES THAN THE INFRANOR ON ARM 3.
    
    IAN L. / KEVIN C. (SPEEDFAM)
    
17.236rotation and MHTA errorsSUBPAC::BJUBINVILLENFD Fire - RescueSat Mar 30 1996 06:3617

         Before touchdown, heads #4 and #5 errored for no rotation.
      Found breaker #6 was tripped.  Had to EMO tool and reset
      breaker.

         Three times during the night, head #3 failed for not rotating
      before touchdown but I didn't find any visible evidence as to why.
      After the third time, I went through all of the wiring and voltages
      on #3 and found everything ok.  I ran manual polish with rotation
      and oscillation for 15 minutes and there were no error.  OPS started
      back up again and as of 6:30am, no problems were reported....

         Tool timed out for multihead not arriving at load.  Tool
      showed no signs of problems or restrictions.  Soft booted
      once again....
                                         BJ
17.237ARM 3 INFRANOR ROT ERRORSSUBPAC::CMPSat Mar 30 1996 16:297
    IT NOW APPEARS THAT DUE TO THE MANY ARM 3 ROTATION ERRORS THAT THE
    INFRANOR WE REPLACED WAS ALSO BAD. I REPLACED ARM 3 ROTATION INFRANOR
    (WHICH WAS ORIGINALLY ARM 5'S) WITH A KNOWN BRAND NEW ONE.
    
    RAN 25 WAFERS ALL O.K. NOW EARL IS RUNNING MQC'S
    
    IAN LOWBRIDGE(SPEEDFAM)
17.238slurry problemFABSIX::R_GEESun Mar 31 1996 08:1812
	A wafer slid out from carrier #4 and crashed carrier #3.
	There was a loud vibration sound heard just before the crash,
	and the pad was dry. We suspect that the slurry somehow  
	stopped flowing. We checked the slurry pumps and hoses,
	also check the slurry flow, all looked good. We replaced
	the 2 carriers and the pad. We need to keep an eye on the 
	slurry flow during polish.

				B shift

	   

17.239493c/index oil leak/monitor/MHTA coverSUBPAC::CMPTue Apr 02 1996 16:4018
    during the PM this morning we fitted a replacement MHTA flap cover
    (back right) and investigated index table gearbox leak.
    The leak appears to be in multiple places. We recommend
    removing/install new seals or replace the gearbox when the index table 
    repairs are being done.  
    
    Ran speed493c with NO PROBLEMS WHATSOEVER!!. We recommend changing to
    493c on B1 ASAP.This will enable process engineers to develop process 
    changes required due to the faster transition time between segments in
    493c.
    
    Couldn't complete monitor check due to lack of communication cable. we
    will obtain one in order to complete task.
    
    looked at calibration of polish pad temp sensor , have calibrated it in
    a coarse manner. Will fine tune when 493c is in use.
    
    Kevin Colman / Ian lowbridge
17.240Multihead TravelFABSIX::R_GEEWed Apr 10 1996 08:122
    The multihead wouldn't move after coming out of wet mode. Had 
    to reboot the system to get it to home.
17.241boot up errorsFABSIX::B_FINNTue Apr 23 1996 20:3545
	System crashed head 4 during polish segment 7. System and 
        silicon debris cleaned up. 

	Cup to carrier alignment checked. Heads 1 2 and 5 were out of 
        alignment. but head 4 which broke the wafer was in.
        We aligned all cups including arm and index table axis,  when 
        we verified the new settings it was discovered that the index table's 
        new home position had not updated. We tried over compensating in the 
        CCW direction and realigning in the CW direction, (due to known 
        software bug). This failed 3 times. We went into dos and 
        verified current home setting, on the 5 alignment it stored 
        the new index value.

	Upon final bootup index table would only home approximately 1' 
        then stop and errored for "time out". We entered DOS DCX mode 
        and manually positioned the index table for "home" ( aprox. 
        one inch from the home sensor). It then worked automatically from 
        there.
	
        Pads and carriers removed. Facilities work completed. Pads and 
        carriers replaced. System booted fine, we were attempting to 
        run 5 wafers to verify flipper to cup alignment, but Robot failed 
        to go get the wafer from the cassette. NO ERRORS until 
        flipper could not find wafer at the prealigner. We tried to 
        use manual mode to move robot without success. When told in manual to 
        " get wafer from slot 1" error was " Time out waiting for Robot 
        response. Rebooted and tried again, but Index table failed to home,
        had to use DCX mode again to home the table. 
        ALSO had dcx failure board 1 led #1. Rebooted and table homed fine, 
        but now arm four would not home. Rebooted and Arm 3 would not home. 
        No DCX or infranor errors on any or the arm problems. Entered DCX and 
        moved both arms manually fine. Rebooted and System errored for "c:\ 
        bad sector found" "IGNORE _ ABORT_ RETRY. At FS recommendation we 
        reseated CPU and VME still errored for Bad sector found. We chose 
        ignore and continued through init until it errored on at "homing arm 
        3". We shut Arm 3 off in the config file and the problem moved 
        to head 4. FS is going to get a new DCX card for us from their 
        office. (all the oscillations are on card 2). But does not explain
        other errors! 

    	Also intermittent problems with init on elevator 1 and pad
        conditioner up/down, had to be homed in manual
	
	
17.242corruption aboundsSTRATA::RIDLONEliminate the obviousWed Apr 24 1996 02:1016
    Continuing from last note....FS replaced the DCX#2 board which didn't
    help and then they replaced the infranor for carrier 3 oscillation
    which also didn't help. ROTGO and OSCGO commands worked fine. FS then
    booted system up with speed493 software. Initialized fine. Ran some
    breakin wafers through and everything worked fine. We also aren't
    seeing the handler delay that was seen on B2 last week. We did have a
    minor problem with slurry flow but that was isolated to the filter down
    stairs being clogged. Ashland promptly rectified that situation. So to
    summarize, 491 software is corrupted. Upon initialization it won't
    perform various functions such as homing arm #3. If 3 is disabled the
    problem moves to 4. If 4 and 5 are disabled the problem moves to the
    pad conditioner. 
    
    
    					B-shift w/FS
    
17.243aarrrrggggghhhh.STRATA::RIDLONEliminate the obviousWed Apr 24 1996 07:2213
    The tool failed to polish. It would alarm consistently for carriers not
    rotating. The first time was due to the poor slurry flow from the
    filter being clogged. However, subsequent times the carrier down force
    was seen to go as high as 500 lbs upon touch down. Whilest
    troubleshooting broke a wafer in manual polish (what a waste of a
    program) so machine is being cleaned up and pad replaced. Carrier 4
    will also be (or need to be) replace as well. Not sure if the down
    force issue is due to using the 493 version startup so it wasn't
    pursued that intensely. The infranor for carrier 5 rotation died during
    the manual polish (again, what a waste of a program) so it will need to
    replaced. 
    
    					B-shift
17.244continuedSTRATA::RIDLONEliminate the obviousWed Apr 24 1996 07:335
    News flash......the infranor for carrier 5 rotation is now working????
    The power led was off but now its on and it does rotate...amazing.
    Carrier 4 has been replaced and Roger is now working on the pad. It
    could be installed by passdown but then again it might not be as well.
    Its been one of those nights.
17.245hard disk / NUSTRATA::FINNWed Apr 24 1996 20:1424
	System had bootup problems. Norton utility was used and found 
        cross linked files and corrupted files.  These files were then deleted.
	Norton utilities was used to defragment and cleanup disk.

	CPU and Hard drive was borrowed from B3. Swapped B3's CPU, but BIOS was 
        configured for b3's hard drive. Swapped both as set and system booted 
        fine, with the exception of head 5 rotation. Changed infranor. System 
        booted aok.

	Put B3's hard drive and CPU back in B3. 

	Reps ordered New CPU and Hard drive

	Reps worked with Software Engineers in AZ. Kevin and Rich undeleted
        the previously deleted files at Az.'s request. System booted fine.

	Carriers were reinstalled. New pad was installed and conditioned twice.

	Jesus is running breakin's.	

	New disk and CPU coming in tomorrow as a back up if this goes down 
        again. 
					A shift
17.246Software problems, downforce checkedSUBPAC::BJUBINVILLECMP Equipment EngineeringThu Apr 25 1996 07:2840
           At the beginning of the shift, a wafer was left
       on the table at lift-off.

           We had problems the entire shift.  After multiple
       rotation errors after every run of (5) wafers, it got
       out of hand.  The multihead failed to move at one point
       and the DCX board #1 was reset to clear up the multihead
       transport problem.  After a while, the multihead failed
       again and Kevin suggested switching from 493 to 491.  
       At start of the process, it failed once again for rotation 
       errors on all (5) heads.  Kevin suggested trying to do the 
       downforce calibration on all (5) heads.  The data is below.
       The tool has started a 50 wafer verify with touchdown.  
       Time will tell...

       CARRIER       PRE                       POST
       ==============================================
         1       I-.43234                 I-.46429
                 O-6.83289                O-6.29708
                  413lbs                  377 lbs
       ==============================================
         2       I-.46543                 I-.45312
                 O-6.25199                O-6.54377
                   365lbs                  377lbs
       ==============================================
         3       I-.43735                 I-.44988
                 O-6.63926                O-6.31586
                   390lbs                   377lbs
       ==============================================
         4       I-.46371                 I-.49868
                 O-6.12997                O-5.62069
                   403lbs                  377lbs
       ==============================================
         5       I-.46486                 I-.44615 
                 O-6.23607                O-6.55172
                  361lbs                   377lbs 


                                                   D shift
17.247offsets and software kicksSUBPAC::BJUBINVILLECMP Equipment EngineeringFri Apr 26 1996 05:5068
           
            We came into problems with B1 again.  After seeing
        the offset results from C shift, it did correlate to the
        adjustments we did on the downforce Wednesday night.  
        We went through the downforce procedure again and found
        the downforce out.  The only way this could of happened 
        is if the zeroing on the offsets didn't go through.  We
        got all (5) heads back into spec.  Below are the results.
        We checked the spindle to center of polish sweep calibration
        and (4) out (5) heads were in spec.  Head #1 was out around
        .3 inches.  This will have to be manually adjusted, not 
        through the software.  The home switch will have to brought 
        out to bring it in.  The results are below as well.
                    
            During the running of the 1st (5) verify no-touch wafers,
        the tool came up with a LOW VACUUM error when the wafers were
        on the carriers.  START continued the run without touchdown
        and it unloaded fine. The tool failed to keep running the
        rest of the wafers after the 1st (5) had unloaded so START 
        needed to be hit to finish the other 45 wafers.

            After coming out of WET MODE, the multihead timed-out
        going to home.  A softboot didn't work.  The SHORT CIRCUIT
        LED was lit on the Infranor.  The tool was powered down and 
        the wiring was checked at the Infranor without signs of a
        problem.  The tool powered back up and the multihead was
        fine.  

       CARRIER       PRE                       POST
       ==============================================
         1       I-.46429                 I-.45312
                 O-6.29708                O-6.23873
                  390lbs                  377 lbs
       ==============================================
         2       I-.45312                 I-.44988
                 O-6.54377                O-6.31565
                   400lbs                  377lbs
       ==============================================
         3       I-.44988                 I-.45531
                 O-6.31565                O-6.2069
                   390lbs                   378lbs
       ==============================================
         4       I-.49868                 I-.46658
                 O-5.62069                O-5.99204
                   353lbs                  377lbs
       ==============================================
         5       I-.46615                 I-.44668 
                 O-6.55172                O-6.43501
                  389lbs                   376lbs 


         HEAD-TO-POLISH-CENTER

           HEAD        PRE-TEST       
           ====================
           1             13.777
           ====================
           2             14.032
           ====================
           3             14.101
           ====================
           4             14.094
           ====================
           5             14.053
           ====================
       

                                 D shift
17.248Broken WaferFABSIX::R_GEESun Apr 28 1996 07:5212
 	Spit a dummy wafer from head #3, it was the last run of 
	20 wafers. The wafer came out during segment #4 of the
	TOUCHUPB recipe. We checked all the alignments and everything 
	looked good. Checked slurry flow and that also looked o.k.
	Ran 25 wafers using the TRANSFER recipe with no vacuum lose 
	or mis-alignments. The cause of the fly-out is unknown
	at this time, released it back to production.
	
				B shift

	   

17.249Cat Scratch FeverSUBPAC::BJUBINVILLECMP Equipment EngineeringThu May 02 1996 08:2913
    
    
             Tool still down since yesterday with scratches.
         We did two pad changes and a lot of cleaning.  Nothing
         is evident as to what it is at this time but it related
         to the pad.  We had a no-touch run done and nothing
         showed up.  The last run showed a 30mm scratches on #4.
       
             Two pads later and we're still dead in the water. 
         Rich is running more break-ins and is going to check
         particles once more.  They should be done by 8am.
    
                                         D shift 
17.250software problemsSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat May 11 1996 06:1810

       Tool kept having rotation errors in segment #3.  The 
    software for SPEED491 was brought back on line.  No problems
    with tool since.

                                D shift


      
17.251STRATA::SWENSONThu May 16 1996 20:349
                          B1 events for today.


1.   CB 4 TRIPPED
    
2.   HAD ROTATION ERRORS ON HEAD 4 AND 5 
    
3.   ALSO MAIN VACUUM ERROR TODAY.
17.252New filter traps for vac pumpSTRATA::RIDLONEliminate the obviousTue May 21 1996 05:148
    Found some replacement filter traps for the vacuum pump in 94 stock.
    Also found some replacement filter elements. The new traps were
    installed on B1 and B2. Two more traps are in the top of the tool box
    along with the replacement elements. A note will be sent to Annie
    Hargraves to add these items to CMP.B's parts lists.
    
    
    					B-shift
17.253Head Rotation Errors / Infranor MC Adjustments Made.ASDG::POIRIERFri May 24 1996 20:1013
    
    Rotation Errors:
    ---------------
    
    The system experienced "Head Rotation" errors this afternoon. Kevin
    from Speedfam came in and made the following adjustments to the 
    INFRANOR Motor Controllers:
    
    Motor Controller #1 - Voltage was increased from 12 vdc to 13 vdc.
    
    Motor Controller #3 - Voltage was increased from 10.5 vdc to 12 vdc.
    
    A 25 wafer verification run was sucessfully completed at 6:00 p.m. 
17.254Carrier #1 not rotatingFABSIX::R_GEESun May 26 1996 07:4515
	The system was erroring for "carrier #1 not rotating",
	we would restart and it would complete the polishing
	cycle. This was happening on every run. We found that 
	carrier #1 & #2 would stop rotating during step #3 of the 
	polish cycle, no visual errors found at the Infranor or 
    	DCX boards. Took carrier #1 off line and ran 24 monitors 
	with no errors. Put carrier #1 back on line and the error 
	was gone. The error happened one other time later in the night
	on the first run of a metal 3 lot, again no errors found.
	
		
				B shift

	   

17.255ROBOT CONTROLLER ERRORSUBPAC::CMPTue May 28 1996 18:043
    THE ROBOT CONTROLLER FOR THE TOOL IS SCHEDULED TO ARRIVE ON WED.
    MORNING, THE PROBLEM WITH IT'S OPERATION WAS TRACED TO THE CONTROLLER
    NOT GENERATING 28 VOLTS FOR THE PROM SAFETY BOARD.
17.256Robot Controller Replaced / Head #2 Rotation ErrorsASDG::POIRIERWed May 29 1996 20:3413
    
    Robot Controller:
    ----------------
    The Robot Controller was replaced this afternoon. New points were
    taught and a verification run was completed.
    
    
    
    Head #2 Rotation Errors:
    -----------------------
    The tool is going DOWN at shift chage for repeated Rotaion Errors on
    Head #2.
    
17.257Head rotation errorsSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu May 30 1996 07:4030

         Came in with tool down for head #2 rotation problem.
    After tweaking Infranor to 13V, we checked all (5) heads
    with 400lbs downforce.  Head #1 stopped rotating every time
    (4) seconds into polish in segment #3.  No errors on the
    infranors, nothing!  All of the voltages were fine.  We 
    tweaked rotation #1 and #2 Infranor to 13.5V but #1 kept 
    failing at 400lbs and every run at 377 lbs.   We took #1 
    off line and the tool ran flawless.  Due to the workload 
    and B2 struggling through MQC's we turned B1 back over to 
    production with (4) heads to lighten up the workload.  The 
    tool failed MQC due to a recent spec change which only 
    allows for 300 removal rate range across the (5) carriers.  
    OPS could not bring tool into spec because off-set program
    needs (5) inputs.  OPS tried putting in n/a and 0000.0
    but it would take n/a and 0000.0 produced bogus readings.
    (ie. 132 for an off-set)  Well, the tool can't run with (4)
    heads.  I took the tool back and swapped out the infranor
    for #1 rotation with a new one from SF office.  The voltage
    was first set to 12.5 and #1 stalled out on the pad again
    with errors on the infranor.  The voltage was then boosted up
    to 13.5 and it still failed.  We tried running the gradual
    downforce increase using the I_O_BREAKIN recipes but after
    the third run, we left a wafer on the table at lift-off.
    Kevin now feels its a pad issue.  We've changed the pad and 
    currently giving it a long pad condition.  After that, we 
    are going to try break-in wafers  again.

                                                   D shift
17.258BAD HEAD ON B1SUBPAC::CMPThu May 30 1996 17:4516
    WE FOUND A PROBLEM WITH HEAD 1 STOPPING ON SEGMENT 3, AFTER
    TROUBLESHOOTING, WE FOUND A BAD CABLE OR CABLES IN THE FLEXTRACK.
    PARTS ARE NOW ON ORDER AND WE WILL UPDATE AS TO THE ARRIVAL DATE AND
    TIME.
    
    THE SEQUENCE OF EVENTS FOR TROUBLESHOOTING ARE AS FOLLOWS:
    
    1)  MOVE CABLES AT BULKHEAD CONNECTOR BETWEEN HEAD 1 AND 4 TO DIVIDE 
    PROBLEM TO MULTIHEAD OR INFRANORS BACK TO THE CPU CABINET.
    2)  PROBLEM STILL WITH HEAD 1, SO WE MOVED CABLES AT BULKHEAD BACK TO 
    NORMAL CONFIGURARTION, THEN MOVED CABLES AT TB13 BETWEEN HEADS 1 AND
    5. THE PROBLEM THEN CHANGED TO HEAD 5 AND THE FAULT NOW ISOLATED TO
    THE CABLE IN THE FLEXTRACK.
    3)  DURING TROUBLESHOOTING WE MONITORED THE VOLTAGE AT THE INFRANOR
    AND MONITORED THE CURRENT (AMPS) AT TB13.
    
17.259Head #1 Rotation Cable Is BadASDG::POIRIERThu May 30 1996 20:377
    
    Bad Cable:
    ---------
    Head #1 Rotation Cable is bad. Head #1 is off line and we're running
    with 4 heads until the cable arrives tomorrow.
    
    
17.260B1 FIXSUBPAC::CMPFri May 31 1996 19:332
    THE TOOL IS BACK WITH ALL 5 HEADS, THERE IS A POSSIBLE SOFTWARE ERROR
    THAT CAUSED THE PROBLEMS.
17.261ROTATION AND DROPPED WAFERFABSIX::J_SWENSONSat Jun 01 1996 20:524
    1. WAFER SPIT AT LIFT OFF ON CARRIER 2.
       SWAPPED CARRIER 2 AND 3 ERROR WOULD NOT REPEAT.
    2. ROTATION ERROR ON CARRIER 1. TRIPPED BREAKER (RESET)
       REBOOTED AND RAN ERROR, DID NOT COME BACK.
17.262Carrier #1 rotation errorFABSIX::R_GEESun Jun 02 1996 07:5111
	The system was erroring for "carrier #1 not rotating",
	we would restart and it would complete the polishing
	cycle, this was happening on almost every run. The system
	was erroring in segment #3 of program #31 ILD1. We tightened 
	the connections on TB13 that go to head #1 Resolver. The error 
	never reappeared. 
		
				B shift

	   

17.263HEAD 1 UPDATESUBPAC::CMPMon Jun 03 1996 09:117
    AMMENDMENT TO FRIDAYS PASSDOWN:
    
    REPLACE BOTH MOTOR AND RESOLVER CABLES IN THE FLEXTRACK, STILL HAD
    ERROR ON HEAD 1, ADJUST VOLTAGE ON INFRANOR TO 14.5V AND REBOOTED
    SYSTEM WITH 491 SOFTWARE. RAN WAFERS WITH OLD SOFTWARE THEN BROUGHT
    BACK 493C SOFTWARE AND RAN WITH THAT VERSION. 493C IS NOW THE CURRENT 
    SOFTWARE RUNNING.
17.264carrier 1 not rotatingSTRATA::RIDLONEliminate the obviousTue Jun 04 1996 08:276
    Tool errored for carrier 1 not rotating again. Recipe this time was 
    touchupc. Only happened on first 5 wafers. Backed out and restarted 
    with no other mishaps. 
    
    
    					B-shift
17.265Carrier #1 rotation FABSIX::R_GEETue Jun 04 1996 08:407
	We had carrier #1 rotation errors about 6 times tonight. No
	specific pattern or problems noticed.
 		
				B shift

	   

17.266carrier 1STRATA::RIDLONEliminate the obviousWed Jun 05 1996 03:182
    Carrier 1 rotation errors on first 5 wafers of an ILD2 recipe. 
    The rest ran fine.
17.267rotation errorsFABSIX::L_SHKOLNIKOVThu Jun 06 1996 08:598
    
	The system was erroring for "carrier #1 not rotating",
        this morning. The system was erroring on 4th second in segment #3
		
				D shift

	   

17.268It's back online.SUBPAC::CMPThu Jun 06 1996 19:191
    The tool is back online for your processing pleasure
17.269Rotation Errors / CDATA.CMP File ChangedASDG::POIRIERThu Jun 06 1996 20:1923
    
    Carrier Rotation Errors:
    -----------------------
    This morning we observed that the actual Down Force readings on the
    status screen were approx 25% higher than the 377 lb. setpoints 
    during a Rotation Error event. The error occured with 4 sec remaining
    in Segment #3.
    
    After observing B2, we noticed that the initial Down Force (spike)
    was only around 5%. We decided to look in the DOS files for any
    differences between B1 & B2. We found a difference in the CDATA.CMP
    file. The "CLOSED LOOP INTERVAL" in B1 was 18 and was 36 in B2.
    
        (CLOSED LOOP INTERVAL was changed from 18 to 36 in CMP.B1)
    
    
    85 wafers were run at 410 lbs. Down Force with no Rotation Errors. 
    The initial Down Force (spike) at touch down in segment 3 is now
    in the 5% - 10% range. The system was returned to Manufacturing 
    at 3:00 p.m.
    
                                                          C Shift.  
                                                                  
17.270Wafer track #1FABSIX::R_GEETue Jun 11 1996 07:518
	The system was alarming for "wafer in cassette #1 wafer 
	track", after it was put into wet mode. We tightened a loose
	connection on the wafer present sensor and adjusted it, (#98).

				B shift

	   

17.271Process doorsFABSIX::R_GEETue Jun 11 1996 07:5110
	Put the 4 side doors back onto the alignment tracks. 
	We need to get new rollers for these door to ensure
	they stay on the tracks. The rollers are in very bad 
	shape, but are working for the time being. Please
	be gentle when opening and closing the side doors.

			       		B shift 



17.272Head #1 Rotation Error - Experimental Pad On TableASDG::POIRIERWed Jun 12 1996 20:2911
    
    Head #1 Rotation Error:
    ----------------------
    We experienced 1 Rotation Error on C shift today. Process Engineering
    was running an experimental pad at the time the error took place.
    Larry C descibed the pad as "harder and more aggresive than the normal
    IC1000 pad". Also, we observed that the ramp overshoot was in the
    normal 5% range in segment #3. We're of the opinion that the Rotation
    error was caused by the additional friction of the experimental Pad. 
    
                                                               C Shift. 
17.273B1 slurry calibratedSUBPAC::KRUPAFri Jun 14 1996 17:168
    Slurry was calibrated after the slurry delivery problem.  
    Results:
    		set-point	Pre-adjust	Post-adjust
    		-----------	-----------	------------
    Pump #1	500mL/min	~350 mL/min	~480 mL/min
    Pump #2	500mL/min	~400mL/min	~500 mL/min
    
    -Earl/Rich/Frank
17.274moisture sense module offSTRATA::RIDLONEliminate the obviousMon Jun 17 1996 01:225
    After B2's leak detection module was found to be off B1's was checked
    and also found to be off. It was turned back on. 
    
    				
    						B-shift
17.275Lost a wafer at liftoffFABSIX::R_GEETue Jun 18 1996 08:279
	Lost a dummie wafer from carrier #5 at liftoff, it 
	didn't break. Recovered the wafer and restarted the 
	system. Checked the vacuum system, everything looks
    	normal.
 
				B shift

	   

17.276slurry detect/mapper problemsLUDWIG::RIDLONEliminate the obviousWed Jun 26 1996 17:147
    Installed the slurry detect sensor, line, and amplifier. Emo's the tool
    to hook the amplifier up to 24vdc and when I brought it back up the
    mapper wasn't working properly. The reps reseated the ic's on the
    control cards inside the mapper which apperently fixed the problem. 
    Reinitialized and everything started working properly. The slurry
    detect isn't operational as of yet. Keep the amplifier on A for the 
    time being.
17.277Particles are back...SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Jun 28 1996 08:1810
    
          Problems with particles all night.  
    
          * cleaned entire tool - failed for particles 
          * no touch - particles passed
          * cleaned condtioner and carriers - failed for particles on 
            #1 and #2.  the rest barely passed.
          * swapped out carriers #1 and #2 - passed particles barely
            all others failed bad.
          * replacing pad one more time
17.278wafer breakage...SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Jun 29 1996 09:2013
    

          found source of particles... :^)

          A wafer slid out of carrier #1 and nailed #5.  Small
       pieces of #1 managed to get around the table cause 
       scratches on the rest of the wafers.  Tool was stripped
       down and cleaning was started but not finished.  
       
          total for the crash - (2) carriers - #1 & #5
                                (5) product wafers

                                                BJ
17.279Carrier to CUp Alignment Problems/Table Drains/Rust on TableASDG::POIRIERSat Jun 29 1996 20:2930
    
    Polish Table Drains:
    -------------------
    The rectangular shaped Teflon blocks that the table rings drain into 
    have been drilled out. The slurry flowing off the table now drains
    through the bottom of the blocks instead of flowing into tubing and
    draining through the base of the tool. All 3 tools now drain slurry
    from the table in the same way. This change will eliminate the repeated
    clogging we have experienced with the old draining set up. 


    Rust Found On Table Under Pulley Ring:
    -----------------------------------      
    In order to drill out the Teflon drain blocks, we had to remove the
    pulley ring that is mounted at the TOP of the table. We discovered a 
    large quantity of rust sandwiched between the ring and the table. 
    Both rings and the table were thoroughly cleaned. 



    Carrier To Cup Alignment Problems:
    ---------------------------------
    The tool was down from this morning for a wafer crash.  We continued to
    clean where D shift left off.  We had considerable problems with the 
    carrier to cup alignments. We tried aligning the head to the cups using
    the standard procedure. We're still haveing trouble getting head #2 to
    load properly. B Shift will pick it up from here.  
    
                                                        C Shift.
             
17.280cup to carrier alignedLUDWIG::RIDLONEliminate the obviousSat Jun 29 1996 23:5418
    Picking up where C-shift left off we came in and checked the cup to
    carrier alignment. All 5 were off by a considerable amount. Before
    moving anything we checked the locate.cmp file and found that the
    "offset from home" number for the index table was at -497. Went back to
    calibrate it and we were able to get all five to align by moving the
    index table a couple thousand steps. Final "offest from home" number
    was -2403. We then tried screw up the alignment by first making the cpu
    reinitialize the DCX boards (opening the process doors when they
    shouldn't be takes care of that), rotating the index table, resetting
    the cpu board, and then finally emoing the tool. After each step we
    checked the locate.cmp file and the -2403 number never changed. Ran 5
    wafers through using the transfer recipe and had no problems. Checked
    the locate.cmp file one more time and "offset from home" was still at
    -2403. Told production they could have it back but to let us know
    before they ran so we could watch it. 
    
    
    						B-shift
17.281MC11 infranor changedeLUDWIG::RIDLONEliminate the obviousSun Jun 30 1996 06:5314
    Production was short handed tonight so they couldn't get right to
    MQCing B1 so took the oportunity to continue with the slurry detect
    installation. Powered down the tool to remove the amplifier (to
    relocate) and upon power up the infranor for the index table didn't
    power up. This happened while calibrating the index table earlier in
    the night but it powered up when we emo'd the machine again then. This
    time it wouldn't power  up no matter what we did. We found one in the
    cabinet so it was changed. Other problems in the room prevented the
    completion of the slurry detect installation (Bob had vaca) so this
    will be finished tonight. Tool powered up fine after the infranor was
    swapped. 
    
    
    						B-shift
17.282SLURRY DETECT ACTIVATEDLUDWIG::RIDLONEliminate the obviousMon Jul 01 1996 03:285
    The slurry detect amplifier and probe are installed and adjusted.
    Slurry detect in hardware.cmp has also been turned on.
    
    
    						B-shift
17.283Fly-out from head #5FABSIX::R_GEEWed Jul 03 1996 08:2910
	A product wafer slid out from under carrier #5 during a 1 sec.
	polish (step #5). It was the first run off a PMD touch-up using
	the "LIGHT" recipe. Three product wafers had to be scrapped.
	We cleaned up the mess, checked the alignments and loading,
	replaced the pad and carriers #3,#4, & #5.
 
				B shift

	   

17.284Start up for B1LUDWIG::SWENSONFri Jul 05 1996 20:2412
    Jumpered di pressure switch
    powered system up. System hung up waiting for vacuum level at 
    carrier # 4. powered down and rebooted with no problems.
    installed 5 new carriers
    performed conditioner arm cal, found to be approximately @93 lbs @ 67%
    set to 76. 3 +/- 1 lb
    cut primary pad.
    retrieved new brushes from stockroom for scrubber.
    
                                  C Shift
    
    
17.285mapper problem FABSIX::L_SHKOLNIKOVSat Jul 06 1996 08:2115
           
	System errored for "Waiting for mapper response out".  Cassette would 
not see home switch as it tilted back. Cassette B (#2) is continuously rocking
back and forth looking for it's limit switch and the mapper moving up and down
in the same time.  
Swapped cables at mapper and problem moved to cassette one. Swapped both ends
of both cables and problem returned to cassette two. Swapped tilt boards within
the controller and problem stayed with cassette one.  Swapped cables at input
to the controller and problem moved to cassette 1. Problem is definitely
within the Kennsington itself. I disabled the controller. Cassette 2 is not 
working it wont tilt and scan, cassette 1 is working. Rich (SF) has been called
but did not call back.  

                                                     Leon
17.286vacuum pump movedSTRATA::RIDLONEliminate the obviousSun Jul 07 1996 06:227
    Relocated vacuum pump to the black box below the floor.
    Test fine.
    
    
    					B-shift
    
    
17.287Carrier #1 not rotatingFABSIX::R_GEESun Jul 07 1996 08:0614
	This morning the system began erroring for carrier #1 not 
	rotating. The problem was that during step #3 carrier #1 
	would stop rotation, no errors on the dcx boards or Infranor.
	The actual downforce display looked good, only overshooting by
	2 to 3 lbs. 
	Checked all the connections at the infranor and everything
	looked ok. Checked the voltage at the infranor and it is at
	12.10, no change was made. Conditioned the pad for 300 seconds
	and ran 25 wafers with no more errors.

				B shift

	   

17.288Index table out of adjustment / air cylindersFABSIX::R_GEETue Jul 09 1996 06:3316
	Repaired water leaks at #2 and #3 load cup air cylinders, plus
	fix other leaks.
	Installed all 10 cups and adjusted the pedestal to cup
	clearance. Checked the index table horizontal runout, it 
	is within spec but will need to be brought in tighter. When 
	the table rotates there is excessive vibration. The vibration 
	is due to the inside of the index table rubbing against the 
	carrier scrub station. The index table will need to be centered 
	better to eliminate the excessive vibration.
	When the table is in the home position it is rubbing between
	load two and load four. 

				B shift

	   

17.289index table and such......SUBPAC::LANDRYTue Jul 09 1996 20:3525
      We continued on with the Index Table install today.  The following
      info contains work performed so far.

      *  The new Index Table has been installed.  Runout was done by
         B shift.

      *  10 new cups were installed.  Cups were specifically matched to
         the respective feed-throughs.

      *  All six cup lift cylinders were replaced with a new style 
         cylinder.  Upon checking for leaks, they were several found.
         It appears they all have been repaired over the last couple of
         shifts, but continual monitoring for a few days may be of a
         benefit.  

      *  The new mapper was installed and is functional.

      *  PLEASE refrain from using this system tonight as the reps still
         have several items to do.  The system is down to them for these
         upgrades!
      				
      						A shift  /  SF reps
      

17.290upgrade progressSUBPAC::CMPTue Jul 09 1996 20:374
    
    
    Completed adjustments of the cups and multihead to table, installed
    the new mapper system and will finish up in the A.M.
17.291upgrades and repairsSUBPAC::CMPWed Jul 10 1996 17:294
    The index table work has been completed and the tool tested with
    wafers. An upgrade for the conditioner ring has also been installed
    at the same time. The mapper has been replaced and calibrated due to
    a failure.
17.292LOAD CUP #! sensorSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Jul 11 1996 08:0312

       The sensor or connection for LOAD CUP #1 DOWN sensor 
     is junk.  Due to the fact that I can't get into the bottom
     of the machine because of my belly (stop laughing!), I 
     can't seem to find where the sensor connection is.   
     Since the sensor is not detecting CUP DOWN, the multihead
     can't move.  I went into SERVICE and hit SERVICE CARRIERS
     which brought the carriers onto the table and I have two
     DI guns spraying on everything to keep things wet for now.

                             BJ
17.293upgrade updateFABSIX::J_SWENSONThu Jul 11 1996 20:272
    Reps continued to wrok on upgrades all day and system will remain down
    for eng. Reps will perform further upgrade verifications tomorrow.
17.294Invertor Communication Errors / IO_ILD3 recipe corruptedASDG::POIRIERSat Jul 13 1996 20:4438
    
    Invertor communication error while running IO_ILD3 recipe:
    ----------------------------------------------------------
    The system errored for "invertor communication" on every attempt
    to run the IO_ILD3 recipe at the begining of C shift today.
    Repeated reboots did not remmedy the situation. The Multihead
    would move to polish position, pause & timeout for invertor 
    communication. There was no spindle rotation or table rotation
    and Infranors & DCX boards were completely lit up with error lights.
    
    However, the BREAKIN, MQC, IO_ILD1 & IO_ILD2 recipies ran with no
    problems. It appears that the IO_ILD3 recipie was corrupted by
    EMOing the tool during a run. The person operating the tool at the
    time was experiencing a "Main Vac Too Low" error and hit the EMO to
    recover.
    
    Speedfam recommends exiting back thru the menus on the touchscreen
    to DOS before EMOing the tool. This will facilitate a CLEAN shutdown
    in the software and will avoid the corruption of files.
    
    We recovered from thgis situation by performing the following:
    
    1) Choose the PROCESS edit/view/copy button from the CMP-V Main Menu.
    
    2) Choose the PROCESS copy/erase/rename button from the Process Menu.
    
    3) Erase the corrupt IO_ILD3 recipe.
    
    4) Copy a "known good" IO_ILD1 or IO_ILD2 recipe and insert in place
       of the file just erased.
    
    5) Rename the known good file to IO_ILD3 and change the run time to 400
       sec. (IO_ILD1, IO_ILD2 & IO_ILD3 recipes are identical except for
             the run time).
    
                                                        C Shift.
    
       
17.295mapper errorSTRATA::FINNSun Jul 14 1996 20:2310
	System originally failed for mapper error. Ops techs EMO'ed the system. 
	Upon bootup, system would not init passed "homing arm 1". We were paged 
        and system was placed down at that point. Found screen locked up and 
        DCX error 2 on board 2. Reset DCX board and error cleared but screen 
        frozen. Reset CPU bd for a soft reboot. And all I/O's failed except
        com board. Reset cpu again, same thing. EMO'ed system and brought it 
        back up, tested mapper aok.  

						A shift
17.296High particles and scratches / conditioner tubFABSIX::R_GEEMon Jul 15 1996 08:1015
	The system failed mqc's for high particles and scratches.

	When we were cleaning the system we found what is believed to
	be dry sticky Contact Cement on the conditioner disk and floating 
	in the conditioner disk rinse tub. Also, some found on of the 
	load/unload cups. We cleaned the entire system and replaced the
	pad. We found the conditioner overflow tube had been broke and 
	repaired with the Contact Cement, the Contact Cement was flaking 
	off and contaminating the system. We replaced the overflow tube
	with one found in the Speedfam cabinet.
	                                 
			                       B shift

	   

17.297high particles.....SUBPAC::LANDRYWed Jul 17 1996 20:3520
    The system is currently down for high particles.......
    
    Check and/or eliminated:
    
    Throughly cleaned the polish area, index table, all cups, conditioner
    wheel (including the tank).
    
    Scrubber brushes disabled from program and particles were still high.
    
    Flushed spindles to assure that no contamination/rust was present.  It
    was clean.  
    
    Run the "TRANSFER" program and particles were good.  This run included
    the use of the scrubber program.  Also have to feel that the carriers
    are fine as they obviously hold the wafers!
    
    So, it looks like the pad needs to be changed, at least.  May want to
    dissemble the conditioner gimbal should a new pad fail.
    
                                               A shift
17.298Particles....SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Jul 18 1996 07:3929

       Tool had high particles when we came in.  We took the
    conditioner and assembly apart.  There was dried up slurry
    inside.  The gimble assembly was taken apart and dried out.
    The drain tube was sitting too high.  It was unscrewed almost
    3/4" from the base.  The brushes we swapped from top to bottom.
    The tool was cleaned up and a new pad was installed.  The first
    round of MQC's failed for (1) head with high particles but the
    other (4) were just barely in spec.  MQC's were run one more 
    time and then all (5) failed for high particles.  It would make
    sense that as more wafers were run on the tool, the particle
    count would get higher then it would make sense it had something
    to do with the pad.  The pad were taken from different lots.
    The options left would be:
 
           change the conditioner 
           change the brushes 
           check slurry flow

        The slurry flow came up 150 ml/min too high on each pump.
     Do the math and that 300 ml/min extra during the polishing.
     The flow was brought back into spec at 551 ml/min each pump.
     still failed.....

         Leon is starting another pad change now.


                                              D shift
17.299particlesSTRATA::WESTPHALWed Jul 24 1996 20:389
CMP.B1 - Down for high particles.  After cleaning, checking scrub recipes, and 
	 numerous MQCs it was determined by process and process engineering 
	 that we may have a problem with the wafers.  This conclusion was
	 reached because Gwen rechecked her MQCs, using some old MQC wafers 
	 from CMP.B3's process acceptance.  The MQCs were good.  EEG was
	 notified and accepted  the particle check done with CMP.B3's the old 
	 MQC wafers.  The tool was returned to production.  EEG will get with 
	 the Engineer for the DEP tools and work out the issue.

17.300DCX.CMP corruptedFABSIX::R_GEETue Jul 30 1996 07:5012
	A monitor wafer from carrier #3 stuck to the pad at liftoff.

	Had to reset the computer to recover due to the touch screen 
	freezing up. When we tried to initialize the system nothing
	would home. Found the DCX.CMP file had been corrupt when we
	reset the computer. Re-entered all the variables and intialized
	the system with no problems. This is the second time today
	this happened.
			                       B shift

	   

17.302B1 upgradeSUBPAC::CMPThu Aug 01 1996 15:433
    busbar upgrade done for infanors on b1, made adjustment dcx. cmp file
    as per upgrade, installed new style rollers to doors, checked and
    adjusted scrubber rpm.
17.303carrier time-out errorsSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Aug 02 1996 06:4618

         Tool faulted three times for time-out waiting for
    carriers to go to home, load and clear respectively.
    According to workstream, it timed out three times on
    C shift for waiting for carriers at load.  The tool 
    would let us get out of the program every time however.
    Soft and hard boots didn't work.  After the tool had been 
    off for a while it was powered up and we lost the Infranor
    for the multihead.  I personally don't think it was related
    to the carrier issue because it failed again after Leon
    replaced the Infranor.  After speaking with Rich from SF.,
    he had me look at some files and eventually delete the 
    current DCX.CMP and replaced it with the back-up file,
    DCX.OLD.  After 50 wafers were run for VERIFY, there were
    not problems at this time and the tool was turned back over 
    to OPS.
                                            D shift
17.304PARTICLESSTRATA::WESTPHALSat Aug 03 1996 20:2814
CMP.B1 - Down for high particles and install of new ball bearings in the
	 conditioner arm.  The tool never came up from the PM.  Pulled all the 
	 cups and cleaned the index table. Cleaned the conditioner and process
	 area.  Swapped out the pad with a prestamped one from stock.  The ball
	 bearings for the conditioner arm came in this morning and were
	 installed but we also need the seal which protects the bearings.  This
	 was never ordered.  Speedfam was notified about the seal and will 
	 order it ASAP.  Until the old seal was cleaned and reinstalled until
	 the new part arrives.  MQCs were run on the toolbut they failed for
	 high particles.  Process Engineering was contacted and they requested 
	 that a filter be placed inline with the tool and the pad changed again
	 and the MQCs repeated.  

17.305High Particles Continue / Pads and/or Slurry suspectedASDG::POIRIERSat Aug 03 1996 20:3552
    
    High Particles:
    --------------
    B1 still has not come up from the Monthly P.M. due to high particles.
    Repeated cleanings and pad changes have not yielded acceptable 
    particle levels. (2000+)
    
    John Swenson reviewed the SPC charts for particles on all 5 polishers
    from July 2nd thru Aug 2nd. He found the total mean particle level
    at or above the UCL on all 5 tools. In addition to TOTAL particles,
    we have been consistantly failing for > than 250 LARGE partilces.
    
     Total mean - particles:  7/2/96 thru 8/2/96 on all 5 polishers.
    
     Strasbaugh:
    
     Carrier:  Left     Right
     ------------------------
     CMP.A1    732      1135     UCL = 1000
    
     CMP.A2    925      1112     UCL = 1000 
    
    
     Speedfam:
    
     Carrier:     1         2         3        4        5
     --------------------------------------------------------
     CMP.B1   1626        1502     1569      1712     1710      UCL = 1500
    
     CMP.B2   1300        1609     1412      1325     1226      UCL = 1000
    
     CMP.B3   955         976      950       1080     979       UCL = 1000
    
    
     Frank Krupa was called at home this afternoon. He would like B & A
     shifts to try a few things in an attempt to isolate the source of
     the particles. He suspects PADS or SLURRY at this point.
    
     1) Temporarally install a point of use filter on the incomming slurry
        line on CMP.B1. Change the pad and run an MQC.
        NOTE: The last 3 pad changes yielded 2000+ adders and were over the
              250 limit for LARGE particles.
    
     2) If the slurry filter experiment does not help, continue changing
        pads making sure to use a different lot number each time.
    
        Frank has asked that all particle data be posted in notes and sent
        to him for evaluation. Please remember to collect particle data
        for every change of condition during the evaluations.
        
    
                                                              C Shift.
17.306filter installedSTRATA::RIDLONEliminate the obviousSun Aug 04 1996 06:5523
    Following the game plan given to C-shift by Frank we installed a point
    of use filter. When we opened the slurry valves in the vmb, however,
    the slurry flowed backwards through the filter (yes, we installed it on
    the "slurry in" line). We followed the "slurry-in" line back to the vmb
    and found it connected to the slurry return main line. Swapped the
    lines in the vmb so that the slurry would be filtered going into the
    machine and not out of it. A monitor run was performed and the 
    particles were still high (large >350 to <700 and total >1500 on 
    4 carriers and ~1300 on carrier 1). The pad was pulled 
    and a hand cut one installed (how do you check the pad_change data in
    workstream?). Several black specs were seen on the pad after
    installation. They came off with water but a thorough inspection
    was done prior to installation and no black  specs were seen. It is
    believed that they were coming from the red handled roller as alot of
    black crap was around the Teflon ends. It was taken apart and
    thoroughly cleaned. Another test run still came out with hi particles,
    large hovered around the 500 range and totals around 2000. Another pad was 
    installed from another lot. Particle test data was not available as of 
    posting.


                                        B-shift
     
17.307PARTICLESB1.TXT 8/4/96FABSIX::M_PROVENCHERSun Aug 04 1996 07:4636
	MQC B1 08/04/96
	PAD LOT# (6EN-JP4(NOT SURE))
	PRE PARTICLE DATA	POST PARTICLE DATA
      #1-5/32/0			 #1-388/1396/0		
      #2-5/39/0                  #2-580/1910/0
      #3-3/39/0                  #3-615/1955/0
      #4-3/29/0                  #4-398/1278/0
      #5-1/16/0                  #5-405/1395/0

RESULT : PARTICLES FAILED MQC CRITERIA, PAD WAS PULLED FOR NEW ONE.

	MQC B1 08/04/96
	PAD LOT# 6BR-JP7 (PAD #4)
	PRE PARTICLE DATA       POST PARTICLE DATA
      #1-3/22/0			 #1-587/2366/0
      #2-5/27/0                  #2-487/1849/0
      #3-2/06/0                  #3-478/2060/0
      #4-4/17/0                  #4-455/1786/0
      #5-4/15/0                  #5-531/2290/0

RESULT : PARTICLES FAILED MQC CRITERIA, PAD WAS PULLED FOR NEW ONE

	MQC B1 08/04/96
	PAD LOT# 6BR-JP9 (PAD #10)
	PRE PARTICLE DATA	POST PARTICLE DATA
      #1-8/22/0 	         #1-526/2160/0
      #2-2/14/0                  #2-408/1620/0
      #3-7/16/0                  #3-398/1618/0
      #4-1/18/0                  #4-458/1707/0
      #5-3/36/0                  #5-472/1921/0

RESULT : PARTICLES FAILED MQC CRITERIA, PAD WILL BE PULLED AND RE-
	 -PLACED WITH A NEW PAD FOR THE NEXT SHIFT. MQC'S FOR THESE 
         THREE RUNS, WERE TAKEN FROM LOT #MA 2614.MQC'S TO BE MEASURED
	 FOR THE NEXT RUN, WILL BE TAKEN FROM LOT #MA2611.

17.308FABSIX::B_FINNSun Aug 04 1996 20:4629
	Check pad data history via workstream. Confirmed that lot jp5 
had not been used on b1 since pm/particles. Cut out and installed 
6BR-JP5, cleaned pad conditioner and tub (both were in good shape).



	Failed mqc's for high particles. 

	0-16-0	556-1939-0
	4-16-0  659-2524-0
	0-06-0	700-2503-0
        4-05-0	584-2091-0
	4-27-0	642-2175-1.23


	According to cmt data jp4 was the last pad to pass for Earl. 
Installed 6en-jp4 pad 4 on system. 
	During breakin's robot failed to go get the first wafer. 
System does not realize this until flipper misses wafer at 
pre-aligner. placed wafer on prealigner but system hung up.

	Exited to DOS and re init system. re mapped and started 
again. Robot did not go get first wafer again. Exited to DOS and 
powered down. Brought system up and it ran fine.  Now running 
breakin and MQC wafers.


						 A shift
17.309new carriers/downforce calLUDWIG::RIDLONEliminate the obviousMon Aug 05 1996 07:1031
    After 6 different pads and a point of use filter installation it looks
    like it might not be a pad or slurry problem. Brendan mentioned that
    spc shows that particles took a "dive" right after the monthly was
    performed on Friday. We checked spc and even though particles were up
    and down before Friday (8/2), since then they have been  out of spec
    high. We had process read the last monitors for removal rate and
    uniformity. The rate was low at ~2800/min and uniformity was below 
    9% on all carriers. We rechecked the down force calibrations for the 
    conditioner and carriers since, in our estimation, these would be 
    the only things on the monthly that could effect particles. The 
    conditioner was found to be at 58lbs with 77 requested. Using note 
    23 as a guideline (its also in the pm spec) this was adjusted so 
    that the actual was 77lbs. The carriers were also found to be low 
    by roughly 20 lbs on all carriers. These were also adjusted per 
    the speedfam tech manual so that the actual on the screen and the
    sensotec read 377 +/-1%. Not quite  what we expected to see to
    attribute to the particle problem. The last pad put on by A-shift had
    the lowest particle counts compared to the other 5 pads used in the
    past two days so the same lot number was installed. New carriers were
    installed to eliminate them as a possible source of particles. 
    During the first set of MQC's two mqc wafers slipped out during lift
    off. The second set of mqc's made it through fine. The wafers may
    have slipped out due to inproper procedure. According to the process
    spec 5 wafers are to be run using bin1, 5 using bin2, and 10 using
    Bkin. The first mqc's ran 5 each using bin1, bin2, and bin3 and then 
    only 5 using bkin. Second set of mqc's passed particles. Large were
    316 to 231 and totals were 780 to 1074. 

               
                                              B-long-winded-shift

17.310PARTICLESSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Aug 08 1996 07:5114

           Tool failed particles....  no kidding...

           A new pad installed at shift change and failed.
        a new pre-cut pad from STK-6 was installed.  Break-ins
        just started.

        ps... B2 passed with old spec values for particles.
              It has a pre-cut pad on it....


                                 D shift
 
17.311ParticlesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Aug 09 1996 07:2822


           The wafers that used for evaluation for the
       EXT_80lbs run on C-shift were, none-usable before
       D shift started.  We ran wafers on D shift and
       the failed for particles
   
        1 - pre-3-0-18  post-571-0-1959
        2 - pre-6-1-20  post-637-0-2006
        1 - pre-3-0-14  post-562-0-1925
        1 - pre-8-0-18  post-1437-0-2143
        1 - pre-4-0-14  post-1430-0-2185

            We brought over the 55-gallon drum of slurry but
        we couldn't find the 5u filter that EE instructed.  We
        tried a 1u filter but we could only get a maximum flow
        of 300 ml/min.   NEXT..........  


                                 D shift
 
17.317B1 Particle testFABSIX::J_SWENSONSat Aug 10 1996 20:3554
                              CMP PASSDOWN 
                              ------------

CMP.B1
------

At start of tests had to do several reboots for robot errors, then 
ran the tests below prior to lunch, discontinued testing due to wafer crash
on standard io_mqc recipe; wafer came out of carrier 5 after 8 seconds into the
polish segment; changed out all carriers due to cracks on the pps rings at the
helicoils.  Had difficulty restarting tool due to low vacumm; discovered 
nitrogen pressure was too low. 




                    Particle Test Results for 08/10/96

     Tool B1                             Carriers

                         1          2           3          4          5  


1. Fab6 Slurry        628-2432-0    D        603-2222-0      D       522-1904-
   Normal
   (MQC_RECIPE)       



2. Fab4 Filtered      328-1124-0    D        728-2584-39.6   D       847-3417-0
                    
   Slurry                                     


3. Fab6 Slurry        316-1018-0    D        240-1153-0      D       191-709-0
   water flush         


CMP.B3
------

System crashed 30 seconds into the 50 second segment 6.

Removed carriers 1,2 and 5 and replaced.
cleaned tool replaced pad,reset counters.
system returned to production for MQC's.


CMP.B2
------

Weekly p.m. completed
    
    
17.312RobotFABSIX::R_GEEMon Aug 12 1996 08:258
	The robot wouldn't respond, found one of the cables going 
	to the robot was loose. Re-seated, repositioning the cable
 	and tightened the connectors. Ran 100 wafers using the transfer 
	program with no errors.	    
			                       B shift

	   

17.313Load flipper / unload cup / broken waferFABSIX::R_GEETue Aug 13 1996 08:1821
	The load flipper was missing wafers at the aligner. 
	Adjusted the load flipper to aligner position back 
	10 steps.
	
	The unload cup down sensor wasn't seeing the cup down,
	adjusted the sensor.
	
	Left breakin wafers on the table at lift off many times
	tonight.

	Many computer and screen problems tonight, believed to be
	created by excessive vibration.

	Broke 2 breakin wafers on the pad in step #6, the wafer slid 
	from carrier # 1. We think maybe the new water polish step 
	may not be a good idea with a new pad. 
		
    	                       B shift

	   

17.314B1 back and proud of it.SUBPAC::CMPTue Aug 13 1996 21:5512
    First off,, What the heck happened???
    
    Found the alignments off on the robot, load and unload flippers, and
    the elevators. Realigned the robot, load and unload flippers, and
    the elevators, also made adjustments to the load and unload cups
    up and down speed. Checked the multihead arms and they looked good.
    I reloaded 493C software and saved 491 for a computer museum.
    
    
    
    
                                              Rich(Mr. Goodwrench)
17.315loosing its mindSTRATA::RIDLONEliminate the obviousWed Aug 14 1996 07:5912
    Production tried to mqc the tool due to b2 and b3 being down. However
    several faults prevented this. First off the carriers down at carrier
    clean wasn't being seen. Then carriers out to load cups wasn't being
    seen. Checked the dcx files and nothing seemed to be out of the
    ordinary. Tried recalibrating the carriers to load cup positions which
    worked. Both the index and carriers 4 and 5 positions had to be
    changed. Tried to run again but tool still wasn't seeing the carriers
    down at the clean station. The DIGin for the reed switches were being
    satisfied (41-45). Running out of time for this one as well. Will have
    to leave down for A-shift eval.
    
    						B-shift
17.316b1 is backSUBPAC::CMPWed Aug 14 1996 20:145
    Continued work on the problem of software, first made some changes to
    file (hardware.cmp) then made calibration changes to correct problem
    handling wafers. The tool is back to operations, but they will be doing
    some recipe changes for a while. Also replaced all the cmp files due to
    corruption.
17.319b1 alignmentsSUBPAC::CMPThu Aug 15 1996 14:534
    Found the calibrations were off for the robot,load and unload flippers
    and theelevators, as well as the index table alignments. Realigned the
    table to the carriers, spent most of the day with table alignment.
    Reloaded .cmp software files due to possible corruption.
17.320adjustments and a public service message.SUBPAC::CMPThu Aug 15 1996 19:5014
    Made more adjustments to the load flipper fingers for better loading
    into the cups, also made minor adjustments to the aligner and robot to
    cassettes.    Just a word about the flipper alignment, when you have
    aligned the fingers to drop the wafer into the cup and you see the
    fingers are NOT aligned to the prealigner when they go to pick up a
    wafer, adjust the prealigner at that time (if you adjust the fingers
    to the prealigner then they will not be aligned to the cups) The 
    prealigner is the last thing to adjust for wafer pickup. After the
    prealigner is adjusted then it may be necesary the calibrate the
    robot to prealigner.........
    
    
    
                                            Rich (Mr. Goodwrench)
17.321Load flipper and elevator #1 errorsSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Aug 16 1996 08:1523

        A fifty wafer verify run was done after Marco Inc. meeting
     got over.  There wasn't any handling issues except for an
     (2) "pinging"  noises when the flipper was picking up the 
     wafer at the aligner.  So, the tool was turned over to OPS 
     for pad break-in and MQC's.  It was loading the finally round 
     of wafers and the load flipper failed to pick up a wafer at 
     the aligner because the two outside fingers got under the wafer.
     The flipper at load cup alignment was perfect so I nudged the 
     aligner to the left (looking at the front of the machine).  The
     robot at aligner positioning was still ok.  The tool handled
     great after that.  Elevator #1 dumped the DI water but failed to 
     fill back up.  There was air going to the quick fill valve and
     DI was getting by.  We were looking into the 3-way dump valve 
     when it started working again.  If it had failed again, we
     we going to pull the valve assembly out but it worked fine
     all morning.  It failed for the first round of MQC's barely due
     to uniformity and we are waiting for the next next results now.

                                                    BJ and Dave
 
 
17.322update for thurs.SUBPAC::CMPFri Aug 16 1996 09:115
    This is an update for Thursday.
    
    
    On thursday we also installed a new lift cylinder on load cup 1
    due to the fact that it was not lifting smoothly.
17.323water leakSUBPAC::CMPFri Aug 16 1996 11:332
    There was a report of a leak, possibly coming from lift assy. from load
    cup 1, found a screw that had a bad seal and replaced the seal.
17.324sticky cyl.SUBPAC::CMPFri Aug 16 1996 17:225
    Replaced the keyboard, found load flippers not fully opening
    intermitently, the air cylinder will be replaced. Found on thurs.
    that the air pressure to the cyl. was at 30lbs. when reduced to 
    correct limit of 20lbs. the problem started to happen. ? was it
    turned up to compensate for stickiness???
17.325Flipper Gripper for the "Gipper"SUBPAC::CMPFri Aug 16 1996 19:223
    Replaced the air cylinder for the load flipper, 2 more will arrive
    tommorow at package plus, John or John will pick up and replace unload
    side due to age.
17.326Load Flipper Problems ASDG::POIRIERFri Aug 16 1996 20:4852
    
    Load Flipper Performance Problems:
    ----------------------------------
    
    Load Flipper - The load flipper malfunctioned intermittently today.
                   We repeatedly found the flipper at the pre-aligner
                   with the left fingers against the side of the wafer 
                   and the right fingers on TOP of the wafer. We think
                   the right fingers are slipping off the side of the 
                   wafer and riding on top. This of course causes the
                   flipper box sensor to leave the "wafer present" position
                   and make the "flipper closed " position, thus the error.
    
    
                   We opened-up the flipper box and replaced the Bimba Air
                   cylinder. 25 wafers were run without incident running
                   the No Touch recipe. We decided to run an additional 
                   25 wafers using the Touch-Up -A recipe since most of 
                   failures occured during an actual polish recipe.
                   The system broke a wafer on the first run!
    
                   Details: A wafer slid out of head #5 and caused
                            head #4 to break a wafer five seconds into
                            segment #7. The Touch-Up-A recipe runs water
                            only during the 7th segment. It is supposed to 
                            have 313 lbs of down force. B1's down force
                            was 377 lbs. We suspect this as the cause.
                            Fresh 50K oxide monitors we being run at the
                            time of the breakage.
                   
     Where do we go from here? : We feel the following items should be
                                 checked:
    
                            a) Flipper box alignment. (The Index Table
                                                       position was changed
                                                       earlier in the week)   
                            
                            b) Flipper Box mechanisim. (We inspected it and
                                                        changed the air
                                                        cylinder)
    
    
      Current Status:  Carriers 3 & 4 need replacement.
    
                       Pad needs replacement.
    
                       Tool needs to be cleaned.
    
    
      Workstream:     Changed from VERIFY-U to BGN_REPAIR.
    
                                                              C Shift.
17.318Flipper issuesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Aug 16 1996 22:5623

          All of the break-in programs were successfully updated by OPS
      per PE.  PE to change over polish recipes. 

          The load flipper was mis-loading at the start of MQC's.
      the mis-loading was causing scratching.  The load flipper
      fingers were loading the wafer off center.  They were 
      adjusted to compensate.

          After lunch, the unload flipper was starting to have
      problems.   It was intermittingly going under the wafer
      on the inside flipper arm fingers.  It was physically
      bent lower????  Rather than bend it back, it was shimmed
      to make both sides level.  If someone want to bend it...
      I didn't hear about it.

          The load flipper box makes a clunk now and then when
      going from the load cup to the aligner.  It doesn't appear
      to be causing a problem.  Rich (Speedfam) was notified and 
      will look at look at it first thing in the morning.

                                                  BJ
17.327Load flipper and float'in wafers...SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Aug 17 1996 08:0915

       The tool was cleaned up after the crash on C shift.
    Manually verifying load flipper.  The flipper failed to 
    pick up a wafer after 15 wafers.  Found flipper just under
    wafer at align.  Re-cal'ed flipper back 20 steps.  Tool ran
    25 wafers without any problems.  A new pad and (2) new carriers
    were installed.  Tool ran 25 wafers with MQC's and passed
    after second round of MQC's. 

       Later in the morning, several wafers were floating out of 
    slots after the wafer was sent into the cassette and the 
    elevator indexed.  Adjusted water level to compensate.

                                            D shift  
17.328Elevator Dump Valve Sticking / DI Water Leak In Fluids CabinetASDG::POIRIERSat Aug 17 1996 20:4019
    
    Elevator #1 dump valve:
    ----------------------
    The dump valve got stuck in the open position twice on C shift today.
    It happened on D shift once last night. We couldn't find one in the 
    Stockroom or the Speedfam Reps office. We should probably replace this
    valve as soon as new one can be obtained.
    
    
    DI Water Leak In Fluids Cabinet:
    ------------------------------- 
    A small puddle of water set off the "Leak Detection Sensor" alarm this
    afternoon. We wipped it up and observed the cabinet for 10 min. No
    leaks were found. We looked in the cabinet an hour later and found a
    puddle of the same size. It appears the leak is origionating in the
    manifold area. It may be leaking only during an auto cycle when certian
    solinoid valve are energized? 
    
                                                                 C Shift. 
17.329load flipper looseLUDWIG::RIDLONEliminate the obviousSun Aug 18 1996 04:275
    Load flipper missing wafers every so often. Found the right thumb screw
    loose. It didn't have any problems once it was tightened.
    
    
    					B-shift
17.330Broken oxide monitor from head #5FABSIX::R_GEESun Aug 18 1996 07:1710
	A dummy wafer slid out from under carrier #5 and broke
	on carrier #4. 6 seconds into step # 7 of the TUP-f 
	recipe. No reason for the crash was found, we replaced
	the pad and carriers #4 & #5. The dummy wafer that slid
	out still had oxide on it, the program also looked correct.

			                       B shift

	   

17.331Carrier 5 changedSTRATA::FINNSun Aug 18 1996 20:296
    
    MQC'S failed uniformity, particles passed, Kevin changed carrier #5.
    Mqc's passed. Running well all day.
    
    					a shift
    
17.332Broken waferFABSIX::R_GEETue Aug 20 1996 08:269
	A product wafer slid out from under carrier #1 and broke
	on carrier #5. 6 seconds into step # 7 of the TUP-f 
	recipe, same thing  as saturday night. No reason for the 
	crash was found, we replaced the pad and carriers #1 & #5.

			                       B shift

	   

17.333head 1 rotation error.SUBPAC::CMPTue Aug 20 1996 18:5410
    Head 1 rotation errors, found that head 1 would stop during seg. 3 of
    recipe due to rotation stopping, swapped cables for head 1 rot. with
    head 2 rot. and still had problems with head 1. Ordering cables for
    bulkhead to term. block and from term. block to motor.
    Cables will arrive in A.M. 
    
    
    
    
                                                 Rich (Mr. Goodwrench)
17.334head 1 rotation errors.SUBPAC::CMPTue Aug 20 1996 19:045
    This is an addendum to the last message.
    
    
    Earlier in the day we had the same problem and replaced the infranor
    and reseated the cable at the bulkhead.
17.335bad motor, bad,bad motorSUBPAC::CMPWed Aug 21 1996 20:018
    Well here's the poop,  After testing the cables from the bulkhead up to
    the motor for rotation 1 and swapping the infranor from rot.2 with
    rot.1 and swapping the dcx cables for rot.1 with rot.2, it finally
    seems to be the motor. One thing that was noticed when the motor is
    running is that there is a thumping coming from the motor then it just
    stops and errors. So the order has been placed for a new one as well as
    a quick dump valve due to sticking open. Parts will arrive tomorrow in
    the A.M.
17.336Head #1 Rotation Motor On Order / Elevator #1 Dump Valve StickingASDG::POIRIERWed Aug 21 1996 20:4019
    
    Head #1 Rotation Errors:
    -----------------------
    Head #1 continued to error for "Rotation" on every recipe containing
    a value of 377 lbs of down force. It ran fine on recipes with a value
    of 313. All cables relating to head #1 were verified good. The Infranor
    motor controller for head #1 was verified good as well. 
    
    A replacement PAC-SCI motor for Head #1 Rotation was ordered from
    speedfam, it will be here at 10.00 a.m. tomorrow. The system was 
    returned to manufac with head #1 off line. (Warning status) 
    
    
    Elevator #1 dump valve:
    ----------------------
    This dump valve was stuck in the open position once again today. The
    Speedfam rep has ordered a replacement.
    
     
17.337rotation to down condition.SUBPAC::CMPThu Aug 22 1996 19:525
    The problem with head1 rotation is put on hold due to the fact that the
    CPU went south, a new cpu and disk setup was tried but the disk had
    no software installed, so the order went out for a fully loaded disk
    to be delivered, also will contact Xycom in A.M. for info on old disk
    so that we might get going with old disk and new CPU.
17.338Hard Drive Problems / Head #1 Rotation FailureASDG::POIRIERThu Aug 22 1996 20:3915
    
    CPU & Hard Drive Issues:
    -----------------------
    The CPU & Hard Drive were replaced this afternoon. We're having
    problems with the set-up. Speedfam is shipping another hard drive.
    This one will be pre-loaded with software. Delivery at 10:00 a.m.
    
    Head #1 Rotation:
    ----------------
    Still searching for cause. Rotation motor verified good. Will resume
    troubleshooting when hard drive comes up. 
    
                                                             C Shift
    
    
17.339B1 is up again.SUBPAC::CMPFri Aug 23 1996 19:2411
    Well we finally got back up, the CPU and disk drive was replaced. The
    nice thing about the disk is that it's twice the capacity as the old
    one. Still continuing to work on the stalling problem, J.P. and B.W.
    are looking at a few more items and I will check in to see how they
    made out. 
    
    
    
    
                                                      Rich (Mr. Goodwrench)
    
17.340Hard Drive Replaced / Elev #1 Dump Valve Replaced / Head #1 Rotation ProbsASDG::POIRIERFri Aug 23 1996 20:1420
    
    Hard Drive Replaced & Software Loaded:
    -------------------------------------
    The replacement hard drive, pre loaded with dos, arrived at 11:30 am
    today. All necessary files were then transfered to the hard drive from 
    CMP.B3
    
    We are now back where we left off 24 hrs ago, troubleshooting the head
    #1 rotation problem. The effort is now concentrating on the cabling &
    connections in the down force control loop. 
    
    
    
    Elevator #1 Dump Valve:
    ----------------------
    The sticky dump valve in elevator #1 has been replaced. 
    
    
    
                                                               C Shift.
17.341Continuation on carrier #1FABSIX::O_WHITFORDSat Aug 24 1996 08:0527
     Continued with troubleshooting the carrier #1 & polish table stalling. 

     Just for grins, we measured the waveform and voltage at TB 16, pins 3 & 6.
These are the signal lines leading back to the DCX/Board #8 thru the ACRO-MAG
modules for carriers 1 & 2 respectively.

     We took some baseline readings at 377PSI (Same proceedure as downforce 
checks) and found carrier 1 at -1.74VDC and carrier 2 at -1.72 VDC. Good so far.

     Loaded up 5 wafers and made a test run using a break in recipe. 

     Got a lot of moaning and serious vibrations during max downforce/
oscillation but no stalls of either carrier #1 or the polish table. Measuring 
the load cell output at TB16/pin3, the output ranged from -1.66 VDC to -1.88 VDC
while carrier #2 remained fairly stable ranging from -1.66VDC to -1.72 VDC. 

     We disconnected the cabling from the load cell, and both input & output
ends of the load cell amp to inspect the pins & sockets. Found nothing obvious
to indicate a connection problem......

     Had to drop it at that point, starting Weekly PM on CMP.B3

              Dave & Leon




17.342Head #1 Rotation Error Problem - Update.ASDG::POIRIERSat Aug 24 1996 16:4840
    
    Head #1 rotation failure problem:
    --------------------------------
    It appears that the problem is with the analog control loop wiring
    for head #1's down force control. We have sucessfully run a full
    cassette of wafers at 377 lbs of down force. (IO_BKIN recipe)
    
    D shift removed and retightened every connection on terminal boards
    14,15, & 16 last night. They also removed and reconnected cable #2
    on the right side lower bulkhead . They sucessfully completed 1/2
    cassette of wafers and passed down to us this morning. We continued
    with a verification run and experienced a head #1 rotation error on
    the fourth run.  
    
    Richie from Speedfam arrived this morning with a new Analog I/O cable
    for cable #2 ( head #1's feedback cable). The cable was run outside
    the tool from the right bulkhead connector to terminal board #16 on
    the Multihead. We have just finished running a full cassette of 
    wafers at 377 lbs. of down force with NO head #1 rotation errors.
    However, head #5 slipped out a wafer near the end of segment 5 and
    it broke under carrier #4. We are now in the process of cleaning up
    and resuming the verification run. We will need to run at least 4
    cassettes of wafers before we're satisfied that we've got resolution
    on this problem.  
    
    This passdown is being sent out early because I'm leaving at 4:00 p.m.
    today. Check out Bradley's passdown at 8:00 p.m. for the latest info
    on the head #1 rotation error saga.
    
                                                       C Shift.
    
                            
                                  NOTICE
                                  ------
     The IO_BKIN recipe on CMP.B1 has been modified. The time in segment #5 
     has been reduced to 20 seconds from 120 seconds. We did this to reduce
     the amount of oxide we were removing from the wafers. We weren't able
     to obtain fresh 50K wafers until 2:30 this afternoon. Please remeber
     to change the time back to 120 seconds before returning the tool to
     manufacturing.
17.343unload solenoid plunger bentLUDWIG::RIDLONEliminate the obviousSun Aug 25 1996 06:279
    Removed the spacers in the unload flipper. Checked the arms on the
    marble stone and both were perfectly flat. Checked the arms for level
    with out spacers and both were level. Checked unload cup and it was not
    level to the flipper. Found the unload solenoid plunger(mushroom) bent.
    Replaced with spare from cabinet. Checked alignment and was much
    better. 
    
    
    					B-shift
17.344Carrier #1 rotation Adjusted Infranor voltageFABSIX::R_GEESun Aug 25 1996 07:3916
	The system was erroring for carrier #1 rotation stopping.

	We turned the voltage on the Infranor for carrier #1 rotation
	from 10.36 volts to 10.50 to try and stop the carrier rotation
	problem. Found that the LED on L1 would still flicker during
	step three, so we raised the voltage to 10.60, it no longer 
	flickered.
	Ran 200 wafers using a modified BREAKIN recipe with 390 lbs
	of downforce in steps #3, #4 & #5 and only 20 second of polish
	time. No errored during the 200 wafers run, released the tool
	back to production with all 5 heads on line. Removed the Warning
	in Workstream.
			                       B shift

	   

17.345all carriers not rotatingFABSIX::R_GEETue Aug 27 1996 07:5219
	The system errored twice for carriers not rotation in step #2 
	of the IO_ILD1 recipe. The first time we simply restarted and
	the system ran fine. The seconds time we tried to restart but 
	the system acted like it was lost, it began polishing with no
	carrier rotation and the carriers all the way in. The system
	had to be EMO'ed to recover.
	After the system was re-initialized we tried to duplicate what 
	had happpened by running 1 breakin run with the second run being
	IO_ILD1. 8 seconds into step #5 of the IO_ILD1 recipe a wafer
	slid out from under carrier #4 taking out carriers #3 & #2.
 	We cleaned up the mess, replaced the pad and 3 carriers, ran
 	the pad breakin procedure then MQC's. We then ran the breakin
	recipe followed be the IO_ILD1 recipe with no problems, ran
	good the rest of the night.

			                       B shift

	   

17.346Head rotation errors / head to index table alignmentFABSIX::R_GEEWed Aug 28 1996 08:0036
 Head rotation error

	The system errored for carriers #1 & #2 not rotating in step
	#4 of the 1P_ILD1 recipe. This is the 3rd time we have had this 
	problem running this particular recipe. Found no lit LEDs anywhere,
	everything looked o.k. We exited out and changed the recipe to 
	1P_ILD2 and edited the time to match recipe #1, then restarted with 
	no errors.
	We renamed 1P_ILD1 to ILD1_BAD now #42 and copied 1P_ILD2 to 1P_ILD1.
	Ran the recipe, the first run had the error again then ran fine after
	that, 50 wafer. Looks like the intermittent rotation problem is back, 
	but seems to be only happening while running 1P_ILD1 recipe immed-
	iately after BREAKIN.

Broken wafers / head to index table alignment

	Later in the morning a wafer slid out from under carrier #3 taking
	out #2 and both wafers. The fly-out happened at the beginning of
	step #3 of the 1P_ILD1 recipe (touchdown). We cleaned the mess
	and checked index table to head alignments, found all heads out
	with #3 out drastically.
	Changed the offset from home position in the locate.cmp file from
	-2799 to +194 and also head in out positions. We now have heads
	#2 & #4 looking good, #3 maybe and #1 & #5 out in different index
	table directions. 
	
	How we stand now.
	There is a new pad on the system, the 3 good carriers are wet and 
	labeled on the table, 2 new carriers next to the vise. The align-
	ment fixtures are in. Field service needs to evaluate the alignment
	problem.   

			                       B shift

	   

17.347head#2 rotation error.FABSIX::L_SHKOLNIKOVThu Aug 29 1996 08:088
    	The tool was alarming  for carrier #2 not rotating twice
        this night.(both times IO_ILD1 at segment#4) First time
        system was rebooted and in the second time we turned the 
        voltage on the infranor for carrier #2 rotation from 10.3 v 
        to 10.6. Run 50 wafers using IO_ILD1 with 30 sec of polish 
        time with no errors.

                                                 D shift
17.348Head #2 Rotation Errors / Dump Valve LeakASDG::POIRIERThu Aug 29 1996 20:1215
    
    Head #2 Rotation Errors:
    ------------------------
    Four verification runs were sucessfully completed running the
    TUP_A & IO_ILD1 recipies. (2 runs using TUP_A & 2 using IO_ILD1) 
    The tool wasn't run by Ops this afternoon. They will begin running
    MQC's on D Shift tonight. 
    
    
    Elevator #1 Dump Valve:
    ----------------------
    A small DI leak was discovered where the dump valve connects to the
    PVC drain. A little teflon tape did the job.
    
                                                
17.349Head #1 Rotation Error During Segment #3.ASDG::POIRIERSat Aug 31 1996 20:2414
    
    Head #1 Rotation Error:
    ----------------------
    The system errored for "Head #1 Not Rotating" during segment #3 of the 
    IP_ILD2 recipe. Since the carriers are not touching the pad during
    seg #3, and are seeing no frictional forces, this event appears to be
    unrelated to the head #1 rotation errors we struggled with last week.
    
    The DCX board showed an error for Head #1 rotation. A soft boot took
    care of the problem. 
    
    
                                                               C Shift.
     
17.350CARRIER TO CUP ALIGNMENTSSUBPAC::CMPTue Sep 03 1996 11:4112
    HEAD TO CUP ALIGNMENT ON B1
    
    
    ON WED. 8/28 TOOL1 HAD A CRASH AND THE HEAD TO CUP ALIGNMENT WAS
    CHECKED, AND WAS FOUND TO BE OUT OF ALIGNMENT. THE CAUSE OF THE
    MISALIGNMENT COULD BE A) WHEN THE DISK WAS REBUILT THE SOFTWARE INFO
    MIGHT HAVE BEEN SLIGHTLY DIFFERENT THAN THE TRUE ALIGNMENTS NEEDED.
    B) THE TOOL MIGHT HAVE BEEN POWERED OFF DURING SOFTWARE THAT WAS
    RUNNING. IN ANY CASE THE TOOL WAS ALIGNED AND TESTED AND TURNED
    BACK TO PRODUCTION WITH NO FURTHER PROBLEMS. THE ADJUSTMENTS THAT
    WERE NEEDED WERE RESET INDEX HOME POSITION, AND CARRIER IN/OUT
    POSITIONS.
17.351New Conditioner ring planSTRATA::KRUPAWed Sep 04 1996 17:4324
CMP.B1 Plan for tonight:

Summary: To install a new conditioner ring that has higher cut-rate in an
         effort to increase B1's removal rates to the level of B2&B3.

Plan:	- Do not remove pad
	- Check pad step height to ensure there is >60 mils if it isn't 
	  install new pad
	- Remove old ring and save it for Frank labeled "CMP.B1's old ring"
	- Install New conditioner ring w/serial #216
	- Run conditioner manually at 80% downforce for 30 minutes on old pad
	- Change pad
	- Record total number of wafers on the tool(Note this is the total
	  running counter that is not re-settable). This is being done to 
	  track the total # of wafers on the conditioner ring.  Record this
	  number and serial# in notes with a comment that a new ring was 
	  installed.
	- Perform std 10 minute condition followed by normal breakins and
	  MQCs.
	- If removal rate has increased be sure to pass it down to next shift.
	- Call me if there are any questions, concerns, or problems.

thanks,
-Frank
17.352Tank #2 Dump valveASDG::POIRIERWed Sep 04 1996 20:3610
    
    Tank #2 Dump valve:
    ------------------
    The valve stuck open causing the tank to remain dry during an unload
    sequence. A product wafer broke in the unload track. The valve was 
    replaced and the tool was cleaned. The system should have alarmed for
    the low water level. We're investigating why it didn't alarm.
    
                                                     
                                                          C Shift.
17.353Robot Deservoed Itself TwiceASDG::POIRIERFri Sep 06 1996 19:508
    
    ROBOT:
    -----
    The Robot de-servoed itself twice today. The Robot controller was 
    powered down both times to recover. Please monitor, root cause unknown.
    
                                                            C Shift.
    
17.354Robot problemsFABSIX::O_WHITFORDSat Sep 07 1996 08:053
    Had one robot error with not loading wafers. Again, powered down the
    robot controller for a short time (@ 5 min.) restored power to
    controller and soft boot the tool. Working fine since then.
17.355wafer breakage, recipe re-doingSUBPAC::LANDRYSat Sep 07 1996 19:3624
    The system broke a wafer today.  The details (as best to our knowledge)
    are this:  The I_O_BKIN program was being utilized.  The damage occured
    during segment 3.  There were exercise wafers (plenty of oxide), and
    the 3rd consecutive run of that program.  From the Opt's people, I was
    told that this was a problem the day before.  So, I deleted the
    I_O_BKIN recipe, transferred I_O_BIN3 to the vacated spot, and edited
    the necessary data points.
    
    The pad was changed along with all five carriers.  Due to a lack of man
    power, the system could not be taken care of at the time o the mishap.
    This caused the carriers to dry off, hence why they were all changed.
    
    Upon system startup, breakin wafers were run.  Bin 1, 2, 3 all ran
    fine.  I_O_BKIN was then run and the system faulted for no Head 1
    rotation.  We checked it out and found nothing unusual.  The Infranor
    was reading 10.08 volts.  We kicked this up to 11 volts and re-tried.
    Well, as luck would have it, the system shimmied pretty good on the 
    next attempt.  This was caused by a lack of slurry flow to the table.
    The room pressure was low.  Called Ashland, they swapped over to a new
    filter and pressure returned.  We then ran 5 consecutive runs using the
    newly edited I_O_BKIN recipe.  It ran all wafers fine.  The system was
    given back to Opt's for MQC's.
    
    						Brad  /  Doc
17.356dump 2STRATA::RIDLONEliminate the obviousSun Sep 08 1996 03:338
    Tub 2 wasn't filling in wet idle. Found dump valve not closed
    completely. Removed lower portion and relubed all orings. The oring for
    the shaft was dry and the shaft had oring residue on it because it was
    dry which increased the friction and didn't allow the valve to close
    all the way. Relubed and tested X10 with no problems. 
    
    	
    					B-shift
17.357rotation problems on head 1SUBPAC::CMPMon Sep 09 1996 19:1211
    the head rotation problem.....
    
    
    today we swapped the acromags for head 1&2 with backfill 3&4, it did
    not correct the problem. swapped an analog board in stock and still
    no result. checked voltage on infranor on pins 1&5 and saw a dramatic
    increase compaired to head 3's infranor (4.2v compaired to 2.2v) and
    made an adjustment to infranor to 14.5v to compesate until the problem
    is resolved. At this time we have run 25 wafers and will run 25 more
    to test for operation for the night shift. Also being fixed is the pins
    for the wafer loss plug due to wear and tear.
17.358add-on to previous noteSUBPAC::LANDRYMon Sep 09 1996 20:149
    Just a quick add on to the previously entered note......
    
    	The DCX daughter board was swapped out as well.  Problem remained.
    
    Also, the second cassette of 25 wafers ran successfully.  The system is
    being returned to manufacturing for now.  
    
    	            Doc (for Dave who is faithfully working in the fab)
    
17.359crash, head 5SUBPAC::LANDRYMon Sep 09 1996 20:329
    Another quick add-on.............
    
            The system crashed during mid-afternoon.  This while the DF 
    settings were being played with.  The wafer slipped out of head 5,
    which took out head 4's carrier.  This carrier was replaced along
    with the pad.
    
                                     Dave
    
17.360pad conditioner problemSUBPAC::CMPWed Sep 11 1996 09:176
    On Tuesday afternoon when the Tool was rebooting we noticed that the 
    pad conditioner arm went beyong it's normal distance for travel and 
    did not home properly. Upon further investigation it was discovered
    that the target on the gear that senses home and pad position had come
    loose and destroyed the limit switches. A parts order was made and
    they will be here in the morning (Wed.).
17.361condtioner repairedLUDWIG::FINNWed Sep 11 1996 20:4020

	Field reps installed switches for B1's conditioner. Ran test wafers 
        had trouble with upper cabinet door interlock.
	
	Rich asked that system be placed up and idle.

	I installed the kick plate and swing out meter rack for the pressure
        transducers. One meter is broken. Need to order a new one. Hookup and
        calibration will begin next week.

	Pad was changed by Jim S. due to bubble. (installed a pre cut)

	System broke wafers during breakin seq 3. Carriers 2+3 replaced as well 
        as pad.

	MQC wafers were run and are now being read.

    
    					A shift
17.362End of shift crashFABSIX::O_WHITFORDFri Sep 13 1996 08:5910
    Flyout from carrier #2, crashed into carrier #1 before tool got stopped
    completely.
    
    Cleaned up fragments from polish table basin, removed pad and cleaned
    polish table surface & drain lines. Installed two fresh carriers in
    positions 1 & 2.
    
    Needs new pad, MQC's. Tool still down.
    
    Leon & Dave
17.363dump valve stickingSUBPAC::CMPFri Sep 13 1996 09:173
    Last evening b1 had a sticking dump valve on elevator 1, removed the 
    dump valve and cleaned and lubed the mechanism, also checked and
    readjusted the spring.
17.364B1 checks and adjustmentsSUBPAC::CMPFri Sep 13 1996 12:175
    B1 had a crash this morning and we check the following items for proper
    operation. Slurry flow, made an adjustment to pump1 400ml min to 500ml-
    min, pump 2 about 450ml min to 500 ml min.
    Checked the carrier to cup alignment and found it to be within spec.
    
17.365Carrier #2 crash againFABSIX::O_WHITFORDSat Sep 14 1996 06:4618
    While running new break in wafers on recipe PMD, segment #7, wafer slid
    out of carrier #2 & "chain reactioned" into carriers 1 & 5. Wiped out 3
    wafers and carriers 2, 1, & 5. This was the first run after tool had
    been in "wet idle" mode for a little less than an hour.
    
    Cleaned up fragments, dust, mud, blood & beer.....
    
    Replaced pad.
    
    Checked carrier to load cup alignment for load cups 1, 2, & 5. Found to be
    within spec.
    
    This is the 3rd crash this week on D shift, the second which involved a
    wafer sliding out of carrier #2.
    
    Dave & Leon
    
    
17.366solenoid #2 water leakSTRATA::RIDLONEliminate the obviousSun Sep 15 1996 06:378
    Moisture sensor alarm during process. Sensor on right side under index
    table in puddle with a continuous drip coming down from index table.
    Isolated to solenoid for load cup 2. Removed and found one of the
    sealing washers for the three allen bolts was ripped. Stocked item.
    Replaced and so far leak has been stopped. 
    
    
    						B-shift
17.367Dumb vavles.STRATA::RIDLONEliminate the obviousMon Sep 16 1996 05:3112
    Dump valves for elevators failing to close completely again. Looks like
    relubing with the teflon super lube isn't working. Di water apparently
    breaks this stuff down. In fab4 we had to get a special grease for 
    teflon valves to rebuild, I have some feelers out trying to find the
    stock number (in any) for this grease. I would suggest calling 
    TEQCOM industries (I tried but they apparently don't work a compressed
    schedule) at 714 979 7024 to see if they can ship us some grease. 
    Removed both valves and relubed with nsk grease, seems to be working
    so far. 
    
     					B-shift
    							 
17.368carrier 2 no roSTRATA::RIDLONEliminate the obviousMon Sep 16 1996 05:326
    Carrier 2 not rotating alarm during segment 4 of an ILD1 recipe.
    Exited out and restarted with no other errors. Did not witness actual
    fault so its not known if carrier 1 stopped as well. 
    
    
    					B-shift
17.369rotation problems, Infranor adjustmentsFABSIX::R_GEEMon Sep 16 1996 20:357
    Several rotation stalls today, Head 1 and 2. Checked the voltage on
    head 1 Infranor, it was totally max'd out at 15.3 volts.  Dropped it
    back to 14.5 volts.  Raised the voltage on head 2 to 12.5 volts per
    Tim D's request.  The system is currently in it's 24 hour MQC.
    
    					A shift
    
17.370Vacuum transducers / slurry soleniod / broken waferFABSIX::R_GEETue Sep 17 1996 20:4114
	Flushed out all the slurry lines and fittings with DI water 
	for 30 minutes. Replaced slurry solenoids #70 & #71 and the 
	two check valves.
	Later in the day, at the end of a 400 second polish a wafer was 
	found in #3 unload cup broken, with wafer particles on the pad. 
	Found the new solenoid #70 had failed, replaced the solenoid and
	tested o.k.

	Installed transducers on all 5 heads to monitor vacuum, still need
	to be wired.

					A shift
	

17.371Loud vibrationsFABSIX::O_WHITFORDWed Sep 18 1996 07:1413
    Operator EMO'd tool during period of high vibration during segment #7
    of M1 recipe. This was the same lot which crashed earlier in the day.
    
    3 monitor wafers were involved. All monitor wafers were inspected and found
    to have a good amount of oxide remaining on them.
    
    Checked slurry flow from pumps 1 & 2. Both were in spec at 550ml/min.
    Checked water flow at 1800ml/min, this too was in spec at a little over
    900ml/30 seconds.
    
    We found no hard fail this time.
    
    Leon & Dave
17.372CMP.B1 unload cup#2 sticking problemFABSIX::J_SADINFreedom isn&#039;t free.Wed Sep 18 1996 10:198
    
    
    	Unload cup #2 sticking. found guide pin had walked out slightly.
    Tapped guide pin back in and smoothed out guide pin channel with
    scotchbrite. Reinstalled and cycled 30times. No failures.
    
    
    	jim
17.373dump valve fixSUBPAC::CMPWed Sep 18 1996 19:353
    The sticky dump valve problems on the tool should now be history. Today
    we installed the positive close feature, ran an air line from the
    regulator to the dump valve.
17.374vibration problemSTRATA::TDYERThu Sep 19 1996 18:5148
CMP.B1 - Vibration problem:
 Created 3 recipes, vib_1 (step7=300), vib-2 (step7=313) and vib-3
(step-7=330). Ran 5 wafers on each recipe results were:
	vib-1 = vibrated in step 7 head4
	vib-2 = no vibration
        vib-3 = no vibration

 Notice all offset were the same except head-4, decide DF on 4 may be out.
 Calibrated DF on all 5 heads, head4 WAS out by ~20lbs.

 Ran the following:
	vib-1=vibration head2
	vib-2=vibration head1
	vib-1=vibration head3
 Created an MQC recipe "270_mqc" with 270lbs of  DF in the clean step, step-7.
 NO vibrations, particle were:   tot	lrg	~scrL
				425	79	.2
				220	70	.24
				249	79	.9
				274	58	.6
				292	60	.4
Create a MQC recipe with 250lbs of DF in clean step, particles were:
                                tot	lrg	~scrL
				425	79	.2
				220	70	.24
				249	79	.9
				274	58	.6
				292	60	.4
Repeated the 250lb test results were:
				413	112	0
				290	90	0
				341	109	0
				204	70	0
				406	97	0
Repeated the 250lb test again results were:
				362	91	0
				294	66	0
				330	82	0
				441	71	0
				291	84	0

No hardware discrepancies were found. The table and or carrier shaft bearings
on this older tool may be wearing down and make the tool more susceptible to
vibration. DF in step-7 will be lowered to 250lbs in In ALL recipes with a 
313lb STEP7. We will continue to monitor and research the problem in conjunction
with the head-1 rotation problem.

17.375Robot ErrorFABSIX::J_SWENSONFri Sep 20 1996 19:335
     Robot became de'servo'ed.(ILD2 recipe) Cause unknown.
     Powered controller on and off still no go.
     Exited to DOS, softbooted system, robot came up.
     Cycled wafers through the no touch program, returned system back to 
     manufacturing.
17.376Robot Squeaking/Unload cup 2 not loweringFABSIX::J_SWENSONFri Sep 20 1996 19:578
    An add on from last reply:
    noticed robot would squeak during the first five
    slots of a cassette in the bottom position.
    Unload cup #2 did not lower once tonight. will
    pass on to B-shift.
    
                  C-shift
                       
17.377Robot not respondingLUDWIG::WESTPHALSat Sep 21 1996 20:437
    
	 
CMP.B1 - Down for robot errors.  The robot kept deservoing.  Tried move robot 
	 on the cal screen.  Tried reseting the robot controller w/ no luck.  
	 Powered the tool down for a few minutes and checked the cables, 
	 Everything seemed alright.  Powered it back up and had no more 
	 troubles.
17.378robot/ul cup 2LUDWIG::RIDLONEliminate the obviousSun Sep 22 1996 06:0010
    1. Robot deservo'd. Soft boot corrected.
    2. Unload cup 2 not going down all the way. Had to push the pin in more
    to get it to start going down but then it hung up with 1/4 left to
    travel. After swapping cups around to eliminate some things we ended up
    simply switching unload cup 2's gasket (on B1 it's the black gasket on
    the index table) with unload cup 3's gasket. Up and down movement was
    smooth after that. Tested several times with no errors. 
    
    
    						B-shift
17.379flextrack cabling updateSUBPAC::LANDRYSun Sep 22 1996 20:086
    Here's an update on the cabling issue.  I talked with Kevin Coleman
    today.  He has ordered the necessary cables that reside in the flex-
    track.  Should see them here by Tuesday morning.  The system has been
    running fine since the "B" shifts/Kevin's repair.  
    
    						A Shift
17.380exSTRATA::RIDLONEliminate the obviousTue Sep 24 1996 06:0210
    Unload flipper missing wafer every now and then. Found that wafer was
    not level with flipper grippers. Had this problem a couple of weeks
    ago. Same fix here. The plunger on top of the unload solenoid was bent.
    Removed and installed the one that we took off a couple of weeks ago
    (someone did a good job in straightening in out, thanks). Checked
    alignment and it was flawless.
    
    
    						B-shift
    
17.381correctionSUBPAC::CMPTue Sep 24 1996 13:201
    the last update refers to tool b2
17.382Pad ProfileSTRATA::KRUPAWed Sep 25 1996 18:1313
    Performed pad profile on last pad.  Was changed on 9/24 after only 965
    wafers.  Results were:
    	location	thickness
    	--------	----------
    	ID		65  &  65
    	1/4 RADIUS	68  &  69
    	1/2 RADIUS	72  &  73
    	3/4 RADIUS	77  &  78
    	OD		82  &  83
    
    This indicates that the conditioner is placed too far toward the ID.
    Changed the ID/OD of the conditioner recipe from 9.3" to 9.6" for this
    tool.
17.383Head #1 Rotation ErrorsASDG::POIRIERThu Sep 26 1996 20:219
    
    Head #1 rotation errors:
    -----------------------
    The system experienced head #1 rotation errors twice in the last hour.
    A soft reboot cleared-up the problem both times. Due to time constraints,
    we haven't been able to persue this problem. Passing info on to D
    shift.
    
                                                         C Shift.
17.384groundingSUBPAC::CMPFri Sep 27 1996 17:071
    The grounding of the robot and the mapper are now completed.
17.385exLUDWIG::RIDLONEliminate the obliviousSun Sep 29 1996 22:0419
    This is what happened last night (Saturday night). Carriers 4 and 5
    stopped rotation during mqc's (with mqc's on the carriers). Down forces
    displayed on the screen were in the 460 range with head 5 at 550 lbs. 
    Carrier 1 offset was -18, carrier 5 offset was +19. Before the heads
    stopped rotation there was excessive vibration as well. When we exited
    and restarted there was also excessive vibration. Since the monthly pm
    was due and given the above data we decided to perform the monthly. 
    We expected to see some major differences in the down force cal but
    as will see we did not. Carrier 1  was 5 lbs high, Carrier 2 and 3
    were 10 lbs low, carrier 4 was 12 lbs high, and carrier 5 was right on.
    BUT on carrier 5 the actual df on the screen read ~400lbs while the
    sensotec read ~377. The other 4 carriers were the same, that is, the
    sensotec reading equaled the actual reading on the screen. The only
    other thing to note is we made some adjustments to the robot. It was
    way off at cassette 1. I'm surprised it picked up the wafer at all.
    
    
    				B-a-day-late-and-a-dollar-short-shift
    
17.386rotation problemsLUDWIG::RIDLONEliminate the obliviousMon Sep 30 1996 07:327
    Had three instances when carrier 1 and 2 stopped rotating. All three
    times no one was watching the machine. Twice the recipe was ild3 in
    segment 3. The third time it was tup_e in segment 3. All three times
    the recipe was exited and restarted with no problems. 
    
    				
    						B-shift
17.387Inverter / Wafer Detection issueSUBPAC::LANDRYMon Sep 30 1996 21:3641
Rough day.......

      The system was erroring out for an "Inverter Error" this morning.
Upon arriving in the fab, found the system had powered off when the 
"stop" button on the run screen was hit.  So, we rebooted the system
to check out the problem.  After initialization, we attempted to unload
5 wafers from the Index Table.  All was well until the 4th wafer was
tried.  At this time, the system halted.  The system didn't alarm or
timeout.  We found the screen to be locked up.  Soft-booted and re-
initialized the system.  We didn't get far.  The system timed-out for
"one or more cups not down".  Checked the I/O's for the cups, sure
enough, cup 4 sensor wasn't being made (dispite it's actual proper
down position).  Checked the sensor functionality and found it to be
operational.  Made the necessary physical adjustment of the sensor.
Re-initialization went fine.  

      Returned to check on the original problem of the Inverter.  We run
some runs and could not duplicate the problem!...???  Gave the system back
to opts.  They ran fine for awhile, then the Inverter error came back.
The opts people logged an assist, exited from the run screen, and restarted
the run.  It ran again for several hours.

      Soon after, opts had a lag time in running the system.  We loaded up
a cassette of wafers and ran the NO_TOUCH recipe.  Several runs into the 
running, the wafer detection board lit up for wafer 3.  NOTE: this will
cause the table to stop and eventually cause an inverter error.  Dave
and Rich looked at the sensor for this head, re-setup the sensor and tried
again.  Soon after, the board lit again, this time just the wafer detection
sensor LED was illuminated.  They feel the board is defective.  We tried
to draw one from stock, but that turned into a frustrating endeavor.  The
part has an assigned "digital" number, but the part they showed me was
not the board.  The stockroom people showed us several bins of un-assigned
Speedfam parts not yet assigned a number.  No luck there.  Kevin C / Dave
have contacted Speedfam, who say the boards were shipped in August and
should be here.  Thats where we're leaving the passdown at.

      So, we have some issues to work and decisions to make.  B shift is
going to attempt to locate the board first.  If no luck, they will contact 
Tim D to have the options presented to him for dissitation.

      					A shift  /  Reps
17.388wafer non-senseSTRATA::RIDLONEliminate the obliviousTue Oct 01 1996 05:2013
    Continuing from A shift. We were able to find the wafer detect board in
    the stock room and installed it, HOWEVER, it was defective. It always
    showed a wafer loss alarm. Troubleshot the board and found that u2 pin
    8 was low. The inputs were also low and this was a nand gate. Got the
    other board from stock and it worked fine. Returned the defective board
    to stock. Ran some test wafers and we still got an inverter alarm in
    segment 1 every now and then. Cal'd the sensors and found that sensor 3
    was flaky. Replaced and ran a boat of wafers using Bkin recipe (had to
    keep the carriers up as they're running out of BREAKIN wafers) and
    didn't get any alarms. Running another set through now.
    
    
    						B-shift
17.389three boats thruSTRATA::RIDLONEliminate the obliviousTue Oct 01 1996 06:435
    Ran three boats with the modified (carriers staying up) bkin recipe. 
    Nothing more will be done on this shift due to the upcoming fire drill.
    
    
    						B-shift
17.390b1 rotation error/fixSUBPAC::CMPTue Oct 01 1996 20:4633
    Had problems with the head 1 rotation again. 
    Started the fix we have been waiting for, by checking the fairchild
    regulators. We found the adjustments to be slightly low as follows.
    
    Head1-55.0
    Head2- 58.2
    Head3-55.9
    Head4-56.7
    Head5-57.3
    Made the neccesary adjustment to 60.0 and testing with 25 wafers.
    We found the voltage on the infranor to spike to 4.7 - 5.0 so the
    next step was to put in the new gearbox. When the gearbox was removed
    we found the gear at the bottom to have grease on it. Upon putting
    things back together and running again monitoring the voltage at the
    infranor we saw that the voltage would only spike to 2.7 - 3.0 A
    noticeable differance, but we had a crash that we are now cleaning up,
    possibly due to some debris or damage to the pad????
    We also did the downforce after the reclibration to the fairchilds
    and found these results.
    
    Head 1- 413.2   Adjusted to 376.1
    Head 2- 442.4   adjusted to 376.5
    Head 3- 413.4   adjusted to 376.7
    Head 4- 407.6   adjusted to 378.4
    Head 5- 423.7   adjusted to 377.4
    
    After the crash is cleaned there is 45 wafers to run and monitor the
    infranor.
    
    
    
    
                                                   Rich (Mr. Goodwrench)
17.391<finish up on carriewr stallsFABSIX::O_WHITFORDTue Oct 01 1996 23:5610
    Working with Speed Fam rep (Dave) after chrash from previous shift, we 
    replaced 3 carriers and a new pad was installed. (looks like hell, 1st time
    either of us has had to custom cut one) and ran 45 test wafers on the
    BIN3 recipe. We do get some slight vibrations from time to time but not 
    consistant. Visual on the wafers show some look to be on the light side. 
    (close to showing bare silicon) Tool completed the remaining 45 wafers
    with no carrier stalls. Turned over for MQC's.
    
              Dave & Brian
    
17.392HEAD #1 STALL OUTSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringWed Oct 02 1996 07:4215


            Came as REP was leaving.  The tool had 10 more wafers to 
       go as part of the 50 wafer verify.  Everything looked good and
       after the 50 were done running, the tool was turned over to
       OPS for MQC's.  During the second set of (5) wafers, the MQC
       wafers, head #1 stalled out on the table just after touchdown. 
       The tool had to be EMO'ed.  Per SF rep, the rotation Infranor's
       for heads #1 and #2 were brought from 10.7 volts up to 14 volts.
       OPS ran another 25 wafers for verify and it went fine.  MQC's
       were run twice and the removal rate on three heads were below
       spec (2400).  A new pad was installed and conditioned.  The
       tool is now ready for MQC's.
                                            D shift  
17.393vibration plan of attackSUBPAC::CMPWed Oct 02 1996 13:015
    Had a wafer crash early this morning, so this is what has been done so
    far as part of the plan for the vibrations. Checked the down force
    calibrations and found they were very close (6-7 lbs. low), installed
    a new slurry 1 pump, cleaned the tool fully, made adjustments to slurry
    flow on all recipes to bring them up to 900 ml.
17.394vibrationsSTRATA::SWENSONWed Oct 02 1996 20:1413
  Finished cleaning the tool. Put 3 new carriers on, and proceedded to  run
  the vibration test.  All recipes  had there slurry supply doubled, and 
  slurry pump 1 has been replaced.   
  In segment five 17 sec. of 30, of vib_.test recipe the tool crashed.
  Down force on all five heads have been performed. 
  Tool has been cleaned  and pad broken in.
  The vibration test recipe has been modified and is presently running
  a vibration test of 100 wafers.
  25 wafers has been successfully ran as of now. 



                  C-SHIFT
17.395WAFER CRASHSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Oct 03 1996 02:5716
    
    
    rechecked downforce using the point method , then rechecked with
    carrier on tool. At about 65 wafers into the 100 wafer run we had a
    slide out--- carrier 2 went into 5. Cleaned the tool again replaced Pad
    and two carriers this happened at step 7( water ). Checked all recipes
    at this point and found that step 5 had a ramp in it that it did not
    need, reason being we were allready at set point and also at step 7
    needed a ramp being that we are now applying less down force.
    Adjusted all wafer loss sensors for reason being that they were not
    catching wafers on slide out. Restarted 100 hundred wafer run again and
    caught a slide out on head 4 . Were at 85 wafers into it . what was
    noticed that head 4 load cup was sticking causing miss load. We will
    run 25 more wafers at this point and turn back to production to run
    their MQC . It is now 2 A,M, will answer all question tomorrow.
    KEVIN--- Speedfam REP
17.396removal rates issuesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Oct 03 1996 07:2011

        The finished up its VERIFY run and MQC's were run.  The 
     tool failed for removal rate and range.  OPS re-ran MQC's 
     and it failed for removal rate again.  The removal rate 
     chart shows a decline in MQC results since the slurry
     flushing.  That with the fact that so much has been done 
     to the downforces, it is in ENG hands for where to go from
     here...   Awaiting ENG by phone via page.

                                            D shift  
17.397VIBRATION FIXSUBPAC::CMPThu Oct 03 1996 19:567
    CONTINUED WORK WITH THE PROBLEM OF VIBRATIONS. A SUSPECTED PROBLEM WITH
    POLISHING WITH WATER STARTING WITH SEGMENT 6 WAS INVESTIGATED AND FOUND
    TO BE THE CAUSE OF THE VIBRATIONS. A CHANGED TO THE RECIPES BY THE
    PROCESS ENGINEER WAS MADE BY CHANGING THE AMOUNT OF TIME THE WATER WAS
    RUN DURING SEGMENT 6 THUS FIXING THE PROBLEM. A PARTICLE CHECK WAS MADE
    AND THEY WERE WELL WITHIN LIMITS. AT THIS TIME WAFERS ARE BEING RUN AND
    THE TOOL WILL BE TURNED OVER FOR PRODUCTION.
17.398Wafer breakage due to stall on #3SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Oct 05 1996 07:0411


         Tool crashed in BIN3 after head #3 stalled out on the table
    5 seconds into segment #3.  Tool cleaned up and carriers #2 and
    #3 replaced.  The tool has not passed MQC's for the uniformity 
    issue so it is still in VERIFY-U.  Crash form entered in note #34.

                                               D shift

     
17.399MQC failure/Down to ENG.FABSIX::J_SWENSONSat Oct 05 1996 20:163
       Tool was put down to eng. failure to mqc 3x h to h variation  and
       std for three carriers. 
    
17.400Poor uniformity & removal rate / downforceFABSIX::R_GEESun Oct 06 1996 20:4213
	Failed MQC's for uniformity and low removal rates, we
	tried a new pad and carriers with no improvement. Checked
	the downforce using the plate and found heads #1, #3, & #4
	out around 25 lbs. Head #5 was just slightly out and #2
	was within spec. Checked the conditioner downforce and it was 
	also out high by 11 lbs. Brought all downforces in within spec.	

	Ran another set of MQC's, they failed for uniformity, 10 to 12%
	but the removal rate was good. 

					Doc & Bob
	

17.4011/4 turn incoming slurry valveSUBPAC::LANDRYMon Oct 07 1996 20:3911
The system was still experiencing "low removal rate" this morning.  We
went through the slurry "package" looking for any slurry related causes.
Checked the DI flush solenoid to assure that the water wasn't leaking
by, hence diluting the slurry content.  The solenoids were shutting just
fine, no leakage at all.  We then checked the fluids cabinet for any
possible problems.  Found the incoming valve for slurry pump 2 was in
the shut-off position.  Opened the valve, flowed slurry to assure no
crystilization had occured.  All was fine.  Re-rans MQC's, everything
is now in spec!  

      					A shift
17.402wafer loss detectedSTRATA::RIDLONEliminate the obliviousTue Oct 08 1996 06:009
    While running 5 breakins before product, one slipped out of carrier 4
    went under carrier 3 and tripped the wafer loss detect system. Checked
    carrier's 3 and 4 for damage, replaced carrier 4. Both wafers (the one
    from 4 and the one on 3) were still intact. Checked pad for possible
    damage with none found. Conditioned the pad and returned system to 
    production. The step this happened in was step 5.
    
    
    						B-shift
17.403494d software testedSTRATA::FINNWed Oct 09 1996 19:2739

       	System was taken at 6am test of new software (494 d). Backed up old 
       	493 software and recipes. 

       	Hardware portion of modification ie break out boards and analog boards 
       	upgraded. New software installed. Slurry/water line length on pumps 
       	one and two to match the other Speedfam systems.

	System was started up and appropriated software/calibration files 
        altered for new software.
 
	Several software bugs were noticed and documented by Speedfam F.S.

	Down force was plotted on chart recorder, and as promised no DF over 
	shoots were recorded with 494 d software. However carrier one stalled 
        several times with out a significant increase in DF on the chart 
        recorder or on CRT screen!
	
        A New DCX mother board was installed in DCX1,  by SF reps.

	System was reverted back to 493 software and associated hardware. 
	IO_MQC recipe seqment six was changed to 10 sec's and breakin and MQC 
	wafers run. Particles passed and all recipes with the exception of 1 
        sec - 2 sec touchup recipes were modified to reflect 10 secs in seq six.

	While these breakin/MQC wafers were being run the Chart recorder 
        monitored 493 Downforce. No spiking or overshoots were recorded. 

	System is being cleaned and will have a pad change. Full MQC will be 
        needed before returning to manufacturing.

	Original DCX mother board will be reinstalled tomorrow or whenever 
        manufacturing will give it up for two hours so SF field service 
        can monitor DF, due to the mysterious disappearance of the DF spiking 
        problem.

					A shift
                                                                 
17.404Scratches and (1) stallSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Oct 10 1996 07:3712
    
    
             Tool failed for scratches and got logged down.
          Looked at pad and found some debris on pad.  Rinsed
          pad then scrubbed it.  MQC's were re-run and all
          MQC's variables passed.
    
             Tool had one "carrier #1 not rotating" and we were
          able to recover and keep running without any problems
          for the rest of the morning.
                                             D shift
          
17.405ERP replacedFABSIX::B_FINNThu Oct 10 1996 19:0414
		
	System was taken down for troubleshooting of head rotation 
        error and df overshoot. Connector X2 on carrier #1 infranor was 
        examined aok. Pressure regulator for carrier 1 was replaced. 
        
	Brad observed a spike to 490 on carrier one followed by a 
        rotation stall.
	
        Plotted downforce again across several run. Traces were smooth 
        with no signs of overshoot

        Downforce will be checked before returning system to 
        manufacturing. 
                        
17.406carrier rotation errorFABSIX::J_SWENSONFri Oct 11 1996 20:266
    Infranor for carrier one adjusted to 14.02, per Brendan, We had carrier
     rotation problems for head one, no down force spike segment #4,
    4 seconds into segment of a 5 second segment. Down force at the time
    was 399.
    
                            C Shift 
17.407Incoming slurry problemSUBPAC::LANDRYSun Oct 13 1996 20:169
    Head one stalled out this morning.  At the time of the occurrance, opts
    personnel noticed that the slurry distributions was at a trickle pace. 
    Checked the incoming slurry to the system.  Many, many bubbles were
    present.  Quickly checked the other systems in the room and all were of
    the same state.  Halted processing across the board and notified
    Ashland.  Ashland found the filter clogged.  Repaired problem.  No
    other head rotation problems the rest of the shift.
    
                                                   A Shift
17.408494 software installSTRATA::FINNMon Oct 14 1996 17:3710
	System is down for install/debug/eval of 494 software and repair of 
carrier one head rotation problem. Reps have installed new hardware 
(ie breakout boards). Software has been installed and critical parameters set.
Rep is currently working to debug for our tool, (sys #1).

Please leave tool in remote screen in wet mode.  System is not available for 
manufacturing.

						A shift
17.409b1 software updateSUBPAC::CMPMon Oct 14 1996 20:158
    Software is updated to 494D, there is a little more debugging to be
    done and hope to be completed this evening. The tool will be in wet
    mode waiting for process verification tomorrow.
    
    
    
    
                                                  Rich (Mr. Goodwrench)
17.410hang, stall, and crashSUBPAC::CMPTue Oct 15 1996 21:356
    work was wrapped up for the day and the tool was turned over for
    production. When MQC's were run the screen locked up, when that was
    corrected there was a warning of polish table not rotating, when that
    was corrected then there was a crash. The crash I believe was caused by
    a very old pad (over 1000 wafers). We cleaned the mess and is now
    going through the breakin series.
17.411Wrap up after crashFABSIX::O_WHITFORDTue Oct 15 1996 21:586
    Worked with SF rep, replaced pad (1170 wafer count) and heads 4 & 5.
    (Slide out occured from carrier #5 and crashed into #4...) carriers had
    aprrox 340 wafers each... Turned over to ops for verification and MQC
    runs.
    
                   Dave
17.412Nasty deviation resultsSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringWed Oct 16 1996 07:5612


         Tool failed real bad all night for deviation.  Heads 
      #1, #3 and #5 were ~16% and the OP couldn't bring it in.
      All three carriers were replaced, after 3 rounds of MQC's,
      the average deviation was 16% with #1 being as high as 24.5%.
      ENG taking the tool in an hour for Rep. work.

                                               D shift

     
17.413rotation #1 trouble shooting STRATA::FINNWed Oct 16 1996 18:3718
	Continued troubleshooting rotation errors. 
	dxc mother board vme analog i/o, acromag, breakoutboards, epr,
        infranors and infranor power supplies, have all been eliminated todate.
        Chased phase wiring to motor, Swapped motor wires at bulkhead between 
	carriers one and carrier two. Carrier 1 stalled but errored for 
        "carrier 2 rotation fail". This narrowed our search from bulkhead to 
        motor assembly. Ran new cables and tested, still failed.  Swapped 
        output of pressure transducer from carrier one to carrier two. 
        Still failed on carrier one. 
	Motor and gear box have already been swapped. 
	Checking parts supply to mechanical prints for bearings.

	System will be down through the night for bearing change.

						A shift

						
17.414Wrap up for the dayFABSIX::O_WHITFORDThu Oct 17 1996 00:035
    Assisted SFreps & observed work in progress. The workday concluded with
    the decision that oscillation arm #1 will have to be removed from the
    MHTA to facilitate bearing replacement. Due to design and construction,
    this cannot be accomplished in the tool. Work will resume on this in
    the AM.
17.415bearings changed FABSIX::B_FINNThu Oct 17 1996 19:0718
	Continued disassembly of carrier one motor arm assemble. 
Oscillation arm was removed. Disassembled shaft and bearings, 
removed lower bearing and took to machine shop to be pressed out. Two new
bearings were pressed in.  Reassembled Oscillation arm. Reps are 
currently reassembling carrier rotation assembly. New carrier quick 
connect could not be found in fab six SR. We looked twice. Kevin will 
order. Old quick connects  non contact surface is galled, but release
is functional, and is being temporarily reinstalled. 

Down force and vacuum switch will need to be adjusted. System should 
be tested at 400lbs of df for reoccurrence of carrier stall.  

At this time F.S. feels they will not be done until tomorrow, but if 
conditions change, J.P. John or Brad will update this passdown.


					
17.416passdownSTRATA::FINNFri Oct 18 1996 18:0516
	System had vacuum switch adjusted for head one. Pad changed.
	Carriers installed. Tried to run but had low slurry flow. Needed 
        purge of lines. Down force could not be performed on carrier one due to 
        input coefficient not updating. Swapped outputs of load cell amp 
        to carrier 2 with similar results.  Checked stock room for loadcells 
        and amps. (note: these are a matched set). But the items in stock are 
	not matched . Confirmed this by calling Speedfam with serial numbers of 
	components in stock. Amp was installed but problem was the same. 
	Loadcell was installed with new am and input coefficient now changed 
        when recalibration took place but the output was none linear. Reps 
        adjusted span on Loadcell amplifier.  Input coefficient looks good. 
        Downforce is being checked. System will need 100 wafers @ 400 lbs of 
        downforce to test rotation stalling error fix.  Please call Tim with 
        result.

17.417Head #1 Load Cell Calibration Activity.ASDG::POIRIERFri Oct 18 1996 20:1510
    
    * Head #1 Rotation Bearing Replacement Activity: - Update
      ---------------------------------------------
      As of 7:10 p.m. the Speedfam Reps are working 
      on the head #1 load cell amplifier calibration.
    
      Refer to Tim Dyer's passdown for appropriate 
      plan of action.
    
                                                            C Shift.
17.418Parts on the wayFABSIX::J_SWENSONSat Oct 19 1996 20:376
    
        Tool is awaiting parts. Part's will be Monday, there seemed to be a
        confusion with delivery, over night versus 2 day. Carriers were
        removed and are presently being soaked and tool powered down.
    
                                   C-Shift
17.419passdownFABSIX::B_FINNMon Oct 21 1996 19:1615
	Analog board was changed, io cables were swapped. Pad Conditioner
would not descend. Fixed cables, pad conditioner works properly but 
high (max outputs) on manual io screen and scrubber. Tried new io boards (replaced 
100k pot on original board. Az office wanted card cage and supply changed. 
It was done, same problem. Io outputs from break outboard 8 were 
lifted per Tim, problem stopped. Lifted one at a time and isolated to 
Chiller output.  Lifted output for chiller at back of system still had 
problem. Found Broken leg to a pot on "advanced motion control 
assembly". Repair continues.

	Quick connect and loadcell/amp arrived fedx. Carrier ones 
quick connect has been changed.

		
17.420further updateSUBPAC::CMPTue Oct 22 1996 12:445
    Monday Night.......
    Found that the wires to the new BB for the 494 upgrade were in the
    incorrect positions, why??????????? rewired correctly and problem
    disappeared. Continued with the load cell problem and will finish
    up tomorrow.
17.421passdownSTRATA::FINNTue Oct 22 1996 19:1123
	Monitored loadcells to insure proper reading dead weight and under 
pressure, right at output of amp. Found erroneous input coefficient to be 
caused by load cell. Rich noticed a spacing difference and discovered a 
washer/spacer missing, causing load cell to press on outer edge of S block and 
not on load cell. Replaced missing spacer, tested aok. Performed Down force to 
all head. Tried to had several false starts due to slurry flow, manual/ auto 
software bug, and bad pin at i/o for wafer blow off carrier one.
That was all straightened out by late afternoon.

Changed pad ran condition and breakins up to 377. aok. Next ran 400lbs 
downforce and received CARRIER ONE ROTATION STALL. Bearing did not fix the 
problem. 

FS rep are splitting up into two groups, Daniel and Cesar will work into the 
night (considering they have been here since 8am). Rich and Kevin will go home 
and be back in the am. 

Relooking at the cables


				 

17.422Tues night updateSUBPAC::CMPWed Oct 23 1996 10:219
    Late entry for Monday night........
    
    Continued nto work the head stall problem with 2 Rep's from Phx.
    Completed the rebuild of the polish arm and troubleshooting of
    the load cell. During the test the head stalled and made checks
    of what had been done to date. Installed cables from dcx to infranor,
    and bulkhead to TB13, also from TB13 to motor(all installed outside
    of tool for testing), will continue tomorrow with shift coverage
    and more information from Phx.
17.423correction of last messageSUBPAC::CMPWed Oct 23 1996 10:232
    Please make a note that the previous message was labeled Monday night
    but is for Tuesday night. Sorry very tired lost track of what day.
17.424exFABSIX::B_FINNWed Oct 23 1996 18:5930
	
	Continued troubleshooting carrier one rotation problem.

	Acromat mother board replaced. No change.

	Motor from carrier one was installed on carrier three, and 
        carrier three motor placed on carrier 1. Carrier stall occurred 
        on carrier #3 now. No stall occurred on carrier #1. Run was 
        repeated with similar results of carrier #3 rotation fail combined 
        with a spike in infranor current.

	In other words the problem has been successfully been move to 
        carrier #3 by merely swapping motors.

	A new (well from the SF office) motor been installed on 
        carrier #3. It failed for rotation as well. No wires were 
        touched just motor swapped. Phasing was verified and found to 
        be correct.

	Bad motor from #3 was swapped back into carrier one to 
        verify the problem. Carrier #1 failed for rotation.

	Speedfam is sending another motor tonight reps will pick 
        up at the airport(11pm) .
        I asked that a "know good motor" be taken off a system for
        shipment.
                          
	A 200 wafer cycle test at 400lbs will need to be run at 
        completion of fix.

17.425passdownFABSIX::B_FINNThu Oct 24 1996 19:0325


	Two motors are being shipped in from Speedfam. 

	Please install the first motor in carrier one. If it fails 
	swap motors between head 3 and 1. If the problem moves to head 3 put 
	next (second) motor in carrier 3 and try it. If it fails Call it a 
	head 3 problem turn head 3 off line. Begin cleaning for return to 
	manufacturing in the AM.


	If one of these motors passes fine! If not please clean the 
        system through top to bottom inside out. Perform any 
        outstanding pm's with the exception of DF which has just been 
        done, repeatedly. Leave system down. Engineering will bring system 
        up tomorrow with new software on four carriers. 

	The problem is the motor but more info is required from 
        speedfam and motor OEM.

	



17.426motor repairSUBPAC::CMPThu Oct 24 1996 21:485
    Braved the Boston traffic to pick and bring back the 2 motors this
    afternoon, installed the motor and ran 20 wafers before the room was
    closed down to fix slurry/particle problems. The test will continue
    when the slurry comes back online. By morning hope to have some good
    news.
17.427passdownFABSIX::B_FINNFri Oct 25 1996 17:3228
	Another motor was shipped in from speedfam and was installed 
on carrier 3. This motor was said to be a "know good motor taken off 
their training system". It's year of manufacture was 1995. It failed 
for rotation stall carrier #3.

	System was wiped down. Infranor #3 set to 14.5 amps and loop 
interval set to 75. (200 test wafers were successfully run at an 
exaggerated df (400lbs) last night. Rotation alarm tolerance has been 
set to 15% from 10%.

	MQC were run and failed particles. No surprise given what the 
        system has been through. Also wafers were center fast.

	large		total		scratches

	355		2619		14.8  
	364		1974		0
	498		3232		0
	619		4201		0
	382		2045		7.4

Note: Scratches look like slurry
      Uniformity was center fast, pad was changed last night due to 
      breakage during test runs. 
 
System will need to be cleaned cleaned cleaned.

17.428Rotation Motor Changed / High ParticlesASDG::POIRIERFri Oct 25 1996 20:0926
    
    Motor replacement update:
    ------------------------
    A replacement rotation motor was installed into head #3 today.
    The ID tag on the motor indicated that it was manufactured in 1995.
    With the Infranor set to the recommended 10v, the head stalled on 
    the first run. The voltage was increased to 14v and the tool was 
    returned to manufacturing. Although 14v is higher than recommended,
    Speedfam is confident that the motor will not stall with the higher
    voltage. This conclusion was based on the results of last nights 
    200 wafer run with the Infranor set to 14v.
    
    
    Particle Problem:
    ----------------
    The system failed MQC's for high particles and uniformity. 
    
    Action taken thus far:
    1) Replaced Pad, Carriers & Brushes.
    2) Cleaned polish area, multihead & index table.
    3) Ran slurry purge for 10 min.
    
    D Shift, please calibrate slurry flow (flow doesn't look right)
    and run MQC's again.
    
                                                               C Shift.
17.429ParticlesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Oct 26 1996 05:019

        Tool down at shift change for particles and slurry pump
      calibration.  Both pumps were low and adjusted to proper
      flow cal.  The tool probably failed particles because of the
      slurry issue because after a brief cleaning and after the
      slurry system flushing, the tool passed particles.
   
                                        D shift
17.430tolerance adjustmentSUBPAC::CMPSat Oct 26 1996 13:525
    During some analytical checks for the head stall problem, we
    experienced a out of tolerance warning for rotation, this appeared to
    be similar to past warnings showing a slower speed for rotation.
    Made an adjustment to the tolerance factor by raising the percentage
    from 15 to 30. 
17.431"ROTATION OUT OF TOLERANCE" ERRORSLUDWIG::WESTPHALSat Oct 26 1996 20:2510

CMP.B1 - Down for "ROTATION OUT OF TOLERANCE" errors and some voltage readings 
	 that needed to be taken by SFAM.  Rep in formed me that the error
	 response time needed to be adjusted temporarily to correct the error
	 problem.  John H. & Rich B. hooked up an O-scope and a Dranetz 
	 analyzer to the power in on various carrier rotation motors.  Tool was
	 put to END_REPAIR.


17.432tolerance errors....SUBPAC::LANDRYSun Oct 27 1996 19:2125
Today on this system, we experienced some strange behavior.  Let me
eloborate for you.

1)  We occassionally had "tolerance errors for pressure" on head 5.
    Not knowing the new software yet, this alarm is foreign to me.  I
    can't offer anymore into it. 

2)  Another strange occurance was the head up movement of head 5.  It 
    was slow in raising on several occassions.  It also alarmed for 
    "all heads not up on MHTA" one time.  The head was creeping up on
    this occurance.  In between measuring wafers, I quickly looked into
    the air lines for this.  All hoses were connected, and no leaks were 
    observed.  There may be a pinching situation occuring as it only
    happens at the polish table.  It raises fine at the Index Table.
    We'll need to keep an eye out on this.  

3)  We occassionally had "rotation out of tolerance" errors.  No surprise
    here.  This is in the bundle of rotation concerns.

4)  During one of the runs, a pressure tolerance alarm sounded.  It appeared
    to happen at the end of the cassette of wafers.  When silencing the 
    system the MHTA stayed out over the Polish Table, and didn't unload
    the last 5 wafers???? Looks like the computer burped??

                                              Doctor
17.433ep regulatorLUDWIG::RIDLONEliminate the obliviousMon Oct 28 1996 04:1013
    Head 5 failed to come up after polishing. It also removed ~2000
    angstroms less than the other 4 heads. Isolated problem to EP regulator
    not releasing the air to the head. Replaced the regulator and performed
    down force cal. Ended up having to do all 5 heads due to a mix up. We
    used the point method to calibrate and then rechecked with the carriers
    on. The regulator for head 2 will need to be replaced soon. It drifts
    while the others don't. Would have done it tonight but there aren't
    anymore. Breakins and mqc's running now. All heads went up after
    polishing. 
    
    			
    					B-shift
    
17.436Wafers left on PadFABSIX::R_POIRIERTue Oct 29 1996 06:5116
	We had a problem with wafers coming out at lift-off. The 1st
	time was on carrier 2. We took the carrier off and noticed
	the white ring had a couple of deep nicks in it. We replaced
	the ring and tried again. Again, carrier 2 left the wafer
	on the pad at lift-off. We replaced carrier 2 and ran more
	break-in wafers. This time the wafer came out of carrier 1.
	We found water in the vacuum pump filter, cleaned it out,
	replaced the filter and started running wafers again.
	No problems since then.




					B Shift

17.437EP's checked, condition ass. badSUBPAC::LANDRYTue Oct 29 1996 19:3340
A rough ride with this beast today.....

The system failed early this morning for high particles and poor 
uniformity.  As an add-on, the slurry filter down stairs had been
changed about an hour prior to the above MQC results.

The particles were failing on 3 heads badly with the other 2 heads
just barely in.  The major problem was the uniformity, which was
pretty bad.  Head 4 was exceptional out.  ALL the heads were polishing
center fast.  The ID/OD was adjusted to it's limit without any suc-
cessfull results.  Therefore, we decided to look into the DF on the
heads.  We first checked the EP regulators.  Heads 1, 2, and 5 needed 
adjustment at the top end (60 PSI).  Now that these were adjusted, it
was absolutely necessary to perform DF calibration.

Upon performing head one DF cal, we could not get the system to repeat.
Found the wiring hook up to the EP regulators unsatisfactory.  Two of the
three wires were only attached by two strands.  Stripped back and tried 
again, same poor results.  We decided to remove the wiring from the 
stand-alone chart recorder.  Checked DF again.  All was stable and 
repeatable.  Moved on to head two, same fluctuating senario.  Again,
removed the chart recorder wiring from the loop and all stabilized.  Need
I say we removed all the chart recorder wiring at this point?  Got all
heads to be at 377 lbs DF.

Put a new pad on the system, and began the pad conditioning cycle.  Whoa!
What is the condition wheel doing?  It wasn't rotating uniformly.  It 
would speed up / slow down.  Definately a negative affect to the condi-
tioning of the pad.  Gimbal was good though.  Tore down the feed-through
assembly.  Found much rust and a less than desirable lower bearing.  We
attempted to get the bearing out of the assembly without damaging it.  
It was virtually an impossible task.  Therefore, a decission was made to
get the bearing to work as best as possible, and get a new assembly 
shipped in.  We were able to free up the bad bearing enough to make the
wheel rotate as should.  A new assembly IS NEEDED.

At currently writing, the system has passed particle checks.  It needs
the uniformity check done next.

                                    A shift
17.438Pad change, tolorence errorsSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringWed Oct 30 1996 06:1410


       Tool failed MQC's for bad uniformity and OPS couldn't bring
    it in.  A pad change was completed.  We are awaiting results.
   
       Just an FYI, the "tolerance" errors were seen throughout our
    shift....

                                        D shift
17.439Recipes,carriers etc.FABSIX::L_SHKOLNIKOVWed Oct 30 1996 07:4216
	The system went down for wafers coming out at lift-off. The 1st
	time was on carrier 2. I checked the vacuum and it appears to
        be fine. When we checked the recipes we found that the segment#9
        was different compare with B2. Frank was called and recipe was
        changed to match B2. No problem since then.
        Later tonight the tool was alarming for load miss the wafer.
        So load flipper has been adjusted - works fine now.
        
        Continued with high uniformity. The tool failed MQC's four times
        tonight. Looks like the problem related to carriers issue. Did not 
        have a chance to check the carrier curvature ( it should be closer 
        to -2u). 


                                                         Leon
17.440UNIFORMITY OUT OF SPEC.LUDWIG::WESTPHALWed Oct 30 1996 19:2317
CMP.B1 - Down for poor uniformity.  Ran the MQCs.  The uniformities for 1 & 4
	 were high and the rest of the carriers were barely in (around 9.4). 
         swapped heads 1 & 2 to see if the problem followed the carrier 1.   It
	 seemed to follow the carrier so it was decided to swap them all.  The 
	 MQC was ran again.  The removals and uniformities are as follows.

                            REMOVAL      STD

                             4736        10.9
                             5076         9.1
                             5120         4.8
                             4758        12.7
                             4773         8.4

        Adjusted the offsets to see if that will make any difference.  The MQCs 
	have been ran and are waiting to be read by the following shift.
17.441Still out of Spec.FABSIX::L_SHKOLNIKOVThu Oct 31 1996 07:3314
    Still down. Uniformity is way out of spec. Three times run the MQCs.
    After first MQC'S has been run the uniformity for carr.#4 was high
    12.7% and ops couldn't bring it in by running the offset program, so
    carrier #4 was replaced and we attempted to run the MQC's again.
    Well now all 5 heads came in for deviation right in spec. and failed
    for removal rate between 2000�  and 2495�  (spec lower control limit
    for the r/r is 2400� ). We checked the conditioner downforce calibration
    and it was looking good - 74 lb. So pad was changed(we put the same pad 
    which sitting on B3 now) after that and removal rate looks good and
    uniformity is way out of limit between 9.6% and 19.7% on carr#5.


                                                      Leon

17.442Carrier rotation out of tolerance error.FABSIX::J_SWENSONThu Oct 31 1996 10:4220
    System errored several times for carrier rotation out of tolerance.
    
    
                     carrier
                        1         23RPM
                        2         24RPM
                        3         25RPM
                        4         24RPM
                        5         24RPM
    
    
    Carrier's 1 and 3 were strobed and the deviation was +/- 2 RPM during
    process.
    
    
                              Screen also said 66 sec left in process.
    
    
                                      C-Shift
                                         
17.443POOR UNIFORMITYSTRATA::WESTPHALThu Oct 31 1996 19:366
CMP.B1 - Down for poor uniformity issues.  Frank brought the STD in by
	 adjusting the position of the conditioner and the IDOD for carriers.
	 The tool passed MQCs and was returned to production.


17.444Wafer dusting...SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Nov 01 1996 06:0012

    
 
              The tool crashed after a loud short vibration followed
          by a ROTATION TOLERANCE ERROR and a monitor wafer sliding
          out of #2 and "dusting" against #1, taking #1 product wafer
          as well.   The tool was cleaned up and we are completing 
          several items on the MONTHLY PM.  A crash report was posted
          in NOTES.   

                                        D shift
17.445please record for SFFABSIX::B_FINNFri Nov 01 1996 15:3410
CMP.B1 - Kevin Colman is looking into the carrier rotation 
         tolerance alarms. He has asked that the following info be 
         record for any further incidents.

	Process name
	Segment number
	Total time remaining

Thanks
17.446B1 issuesFABSIX::J_SWENSONFri Nov 01 1996 19:2624
    
       * a couple of things were forund during p.m.
    
    
      - broken wire for load cup five sensor (repaired original, could not
        locate duplicate in the stockroom.
     
      - pad conditioning % of down force was set to 99% (set back to 67%)
    
      - not able to perform rpm check for polish table due to software.
        did not want to take the chance of scratching table.
    
    
      - slurry valves for both pumps are turned off.
    
      - the valves will need to be opened and slurry flushed for 20 - 30
        minutes.
    
      - the pad needs to be conditioned and broken in.
    
      - the tool has the carriers rotating and spray on.
    
    
                            c-shift
17.447Software Locking UP During Wafer Loading SequenceASDG::POIRIERSat Nov 02 1996 19:3621
    
    Software Repeatedly Locking Up:
    -------------------------------
    The Software locked up on 6 out of 7 consecutive runs. Each time we had
    the same condition. The lock up occured while the wafers were being
    loaded into the index table. Two wafers were in the load cups, one
    wafer was in the load flipper (home position), one wafer was in the 
    pre-aligner and one wafer was on the robot. We ran our test runs on 
    the ILD1, ILD2 and ILD3 recipies with the same results. We suspect that
    that a corrupted file is to blame. The Speedfam 1 800 number was called
    (per John Hoekstras aprroval) and phone support was requested. 
    Kevin from SF informed us that the Software on B1 has been locking up
    whenever the MANUAL POLISH mode is used. D Shift informed us that they did
    indeed use the manual polish mode last night to perform the table rpm
    check on the monthly P.M. Kevin is calling us back at 7:30 p.m. for
    info on our last test run (it locked up). It looks like the software
    will need to be reloaded. B shift is picking up the phone support with
    Speedfam at this point.
    
                                                            C Shift.
    
17.448???LUDWIG::RIDLONEliminate the obliviousSun Nov 03 1996 05:076
    Performed a soft boot. Ran 25 wafers using ild1 with no problems. Ran
    25 using notouch with no problems.  Would have liked to have run ild2
    and ild3 but the room was short of breakins. Tool is up waiting mqc's.
    
    
    						B-shift
17.449Mapper B / Rotation toleranceSUBPAC::LANDRYSun Nov 03 1996 19:2719
Today's happenings.......

*  We experienced NO problem with the system halting while loading up
   the Index Table.  Looks like the B shift reboot took care of that.

*  We did have a couple of rotation tolerance errors.  Nothing unusual 
   was found.  Acknowledged the alarms and resumed running.  Oh ya, then
   the friendly "reminder" error followed to tell you it just errored.

*  The cassette B mapper had a problem.  The tilt would activate as
   should, the wafer scanner would scan as should, but it would stay
   in this loop (tilt, scan, tilt, scan) until it had a "timeout for
   mapper B" error.  We checked all the connections, all were fine.
   Felt for any overheating at the transistors on the controller, none
   were hot, just warm.  Went back to retry mapping, but the screen was
   locked up.  Soft-booted the system, all mapping functions returned 
   to normal.

      					A shift
17.450reload of 493C softwareLUDWIG::FINNMon Nov 04 1996 19:1419

A group decision was made between Equipment, Process and Field Service 
Engineering, to revert to 493 c software. This was in response to the "carrier 
rotation errors and manual/reboot bugs. 

Hardware was changed at the breakout board and 493c reloaded. Autoexec 
file edited. To prevent or lessen the chance of a stall due to pressure 
overshoots the infranors on carriers 1 and 3 have been increased to 14.5 amp. 
The interval loop and delay in cdata.cmp file were increase to 75.

3 breakins recipes were increase in time in segment 6 from 3 secs to 
10 secs, all other recipes checked in segment 6 to insure 10 sec's.

Ran breakin 2-3, Io_breakin, and runs in io_mqc without error.

Gave system to manufacturing for real MQC.

					A shift
17.451alignmentsLUDWIG::RIDLONEliminate the obliviousTue Nov 05 1996 04:299
    Alarmed for main vac low while running breakins for mqc's. Carrier 3's
    wafer wasn't seated in the carrier. Performed cup to carrier alignment 
    and ended up having to recalibrate the mhta load position, carrier 5's
    load position, carrier 3's load position, and index home position to
    get all the heads in spec. Ran some wafers through to test with no
    problems. 
    
    
    					B-shift
17.452Software burp?SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Nov 08 1996 06:5127

         How do you explain this one....
    
            The OPS tech running the tool set up for two different
        product lots with two different recipes for the two boat.
        This is the "out-lined" form:

        - The tool somehow cooked (5) wafers.  (5) Metal-1 wafers were
          run at ILD1 recipe.
        - The tool left (5) other wafers in the unload cups and double 
          slotted (10) total wafers.
        - The OPS tech first noticed the (5) "cooked" wafers
        - The tool then started alarming for multiple unload errors
          due to the tool not being able to handle the double stacked
          wafers.

           The process ended and I went into the recipe/mapping section
        to double check the OPS mapping/recipe line-up.  Everything was
        in order and we couldn't see what could of have gone wrong.

           The tool was powered down for 10 minutes then powered back
        up.  The tool ran 25 break-in wafers without any problems.
        The tool was handed back over to OPS.

                                                
                                        D shift
17.453conditioner DF verifiedSUBPAC::LANDRYMon Nov 11 1996 18:587
    Per request......
    
         We checked the Conditioning Down Force.  Found it to be just fine.
         No adjustment was made.
    
                                         A shift
    
17.454cmp.b1 downforce calibrationFABSIX::J_SADINFreedom isn&#039;t free.Wed Nov 13 1996 15:1112
    
    
    	re-calibrated down force on all heads. 
    
    	pre-adjust	post
    1-  377		377 (no change)
    2-  360		376.9
    3-  376.5		376.5 (no change)
    4-  367.5		377.5
    5-  367		376.8
    
    
17.455High Particles / Broken Scrubber Drive BeltASDG::POIRIERThu Nov 14 1996 19:2322
    
    High Particles:
    --------------
    The system went down at 6:00 p.m. for high particles. (approx 2000)
    The Pad and Scrubber Brushes are being changed. MQC's to commence on
    "D" shift.
    
    Note: The other Speedfam tools have particles under 1000.
          Both Strasbaughs have around 400 adders.
    
    
                                                          
    
    Broken Scrubber Brush Drive Belt:
    --------------------------------
    While replacing the Scrubber Brushes, we noticed that the Scrubber
    Drive Belt was broken. The belt has been replaced and the Scrubber 
    is being reinstalled at this time.
    
    
                                                            C Shift.
    
17.456DRIED CARRIERS, LOAD FLIPPER, LEAKSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Nov 15 1996 06:1218

          The tool wasn't in WET MODE when we came in at 10pm from
      the Comm. meeting.  the carriers were inspected and they
      looked ok but wouldn't cut.  After numerous MQC runs, the
      OP couldn't bring in the range and dev.  All (5) carriers
      needed to be changed. 

          While loading the 1st wafer for break-in, the load flipper
      went 2" past the wafer and hard stopped.  The flipper wouldn't
      recover so we have power down the tool to free it up.

          The 1/4" DI fitting on the scrubber was stripped and
      DI was spraying on the scrubber motor connector.  The
      entire bulkhead fitting needed to be replaced.  It was
      metric to boot... 
                                                
                                        D shift
17.457High Particles / Tool Cleaned & Pad ReplacedASDG::POIRIERSat Nov 16 1996 19:2713
    
    High Particles:
    --------------
    The system failed MQC's this afternoon for high particles.(3000+)
    All other CMP Polishers had low particles. The tool was thoroughly
    cleaned and the pad was replaced. Mfg is now running MQC's.
    
                                                     
    
    
                                                       C Shift.
    
    
17.458broken wafer, leaking valveSUBPAC::LANDRYSun Nov 17 1996 20:1431
Oh boy, what an afternoon with this baby........

The system somehow broke a wafer in the Unload Module, more specifically,
in the water track of elevator #1.  It's not conclusive of how it actually
occured, but a computer malfunction is suspected.  The system didn't get
any alarms when the breakage happened.  The Opts tech stopped the system
when he noticed the fragments.  Something else strange about this problem
was that there were several wafers stacked on top of each other on the 
pedestals of the scrub station.  The Unload flipper just kept stacking them
up.  Also, one of the scrubber brushes ended up in the water track.  A total
mess to say the least.

After cleaning up the debris and installing new brushes, we attempted to
run a cassette using the NO_TOUCH recipe. The system immediately came up 
with a Main Vacuum Low alarm.  We couldn't clear it.  Checked the gauge in
the pneumatics cabinet, whoa, there was water in the cabinet.  The water 
was passing through the electronic/pneumatic solenoids.  We were able to
trace the source to an internally leaking PVC valve in the fluids cabinet.
The water is backing up through the incoming air supply to the valve, which
is terminated in the pneumatics cabinet.  The valve is for the Polish Tub
Rinse.  We don't have one in stock.  Rich Barrett was to call back to see
if we located one.  If not, he will have one sent out ASAP.

Also, that the water has made it to the solenoid block, water has been sent
to various parts of the system.  Those lines may need to be blown out prior
to moving on.

We did run wafers to check on the unload sequence.  All looks fine.  

                                             A shift
17.459runningSTRATA::RIDLONEliminate the obliviousMon Nov 18 1996 06:1410
    We removed the air line quick disconnect from the process tub valve and
    installed a barbed fitting and then ran a hose from it to the drain. We
    then manually opened the valve so that the polish tub is constantly
    rinsing. We then disconnected all the air lines at specific functions to
    bleed out any water that might still be in the manifold block. ONce we
    gave it back to production it alarmed for no cassette present in elev
    1. Found the right sensor needed to be adjusted. Its been running fine
    now for about 8 hours. 
    
    					B-shift
17.460Carrier 2 Rotation errorsFABSIX::R_POIRIERTue Nov 19 1996 06:2810

	We had problems with Carrier 2 rotation errors. It happened
	twice. We checked the Infranor voltage level and adjusted it 
	from 10.05 volts to 13.0 volts. Problem hasn't happened since.



					B shift

17.461rebuilding valveSUBPAC::CMPTue Nov 19 1996 10:066
    Over the weekend there was a problem with the polish tub rinse valve
    leaking back into the air line. Rebuilt the valve with new O rings
    and there are no more leaks.
    
    A recommendation is to rebuild this type of valve at the first sign of
    wear as O rings are really really cheap.
17.462pad arm fixSUBPAC::CMPTue Nov 19 1996 19:003
    Today the pad arm would not oscilate. Found the infranor was bad,
    replaced and also readjusted the infranor for rotation on head 2
    back to 10V.
17.463test motor installed/ deinstalledLUDWIG::FINNWed Nov 20 1996 19:3611
	New motor was obtained by speedfam which had had its resolver adjusted 
by 30' to help our wafer stalling problem. It was installed but made 
significant noise. It was then deinstalled and the original motor was 
installed.

Test wafers were successfully run and system returned to manuf.


						a shift

17.464Head #2 stallSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Nov 21 1996 05:4318
             Tool errored for head #2 not rotating in segment #5.
         Tool would not recover so we had to reboot.  




                                                
                                        D shift

                           \/  \/
                            (__)
                            (@@)
                     /_______\/
                    * |     ||  \ Brian's deer that got away....
                      ||----||
                      ~~    ~~            
                           
17.465System Paused in Seg #1 / Broken Wafers.ASDG::POIRIERThu Nov 21 1996 19:289
    
    Broken Wafers on Heads 3 &4:
    ---------------------------
    The system broke 2 wafers during seg #5 today. The system was paused by
    the operator in seg #1 and then restarted. The tool was cleaned and
    the pad & carriers were replaced.  We've had no further problems. 
    
    
                                                        C Shift.
17.466ParticlesFABSIX::J_SWENSONSat Nov 23 1996 18:2211
   -  Tool failed for particles
    
   -  Cleaned tub,carriers and conditioning ring.
    
   -  Mqc's ran again and failed for particles.
    
   -  PAL filter changed over to Osmonic's particles passed and system
      returned for manufacturing.
    
    
                                      C-Shift      
17.467crash, head 3 the culpritSUBPAC::LANDRYMon Nov 25 1996 22:0416
    The system crashed a wafer (actually, 2) today.  The system was running
    the IO_MQC recipe.  Right before it crashed, the system alarmed for "No
    rotation on head 3".  Cleaned up the small mess, changed the pad and
    carrier 2 only.  Checked the carrier the wafer spit out of for damage,
    proper extention ring height, etc.  Everything looked fine.  The
    extention ring measured in at 8.0, with zero damage to the ring.  So we
    put the carrier back on the system (same head, #3), ran breakins and
    MQC's.  All is well.  The crash form holds all the vitals, should you
    need additional info.
    
    If the system faults for carrier 3 rotation, please take the head off
    line and enter a warning status for said problem.  Re-qualification of
    the system will be necessary as uniformity will surely change.  MQC's
    will be a must.  This request comes from Tim D.
    
                                                   A shift
17.468index infranor diedFABSIX::M_RIDLONEliminate the obliviousTue Nov 26 1996 04:4014
    While unloading the last 5 wafers from a touch up lot the scrubber
    didn't time out and go back home (or in other words, it kept on
    scrubbing and scrubbing and scrubbing and scrubbing and...you get the
    picture). I was able to stop the process and enter manual mode so the
    cpu wasn't locked up. I was also able to stop the scrubber in manual
    mode, hmmmmm. Tried to rotate the index table and it didn't move, ah
    haaa. Found that the infranor had died. Replaced and everything worked
    fine. Theorize that the cpu was waiting for the index table to rotate
    so as to continue unloading. When it didn't the scrubber kept going and
    going and going and.....you get the picture. Like I said...it's only a
    theory.
    
    
    						B-shift 
17.469another test motorFABSIX::B_FINNTue Nov 26 1996 19:2316
FS installed modified rotation motor on carrier #3. It failed 
(sounded like the last test motor, ie golf balls in a garbage disposal).  
Original motor was reinstall aok.

Contingency plan per Tim:

"
     IF HEAD-3 HAS ANY INSTANCES OF ROTATION STALLING, TAKE 
     HEAD-3 OFF LINE AND RE-MQC WITH 4 HEADS AND RELEASE TO MFG. 
     WE DO NOT WANT TO RISK BREAKING PRODUCT WAFERS.
							"


							A shift

17.470ParticlesFABSIX::J_SWENSONFri Nov 29 1996 19:1027
      High particles -
    
      system failed twice for high particles.
    
      Tool was cleaned after the first failure, carriers also.
    
      Cmp.B2 and cmp.B3 are running well below spec
    
    
                   B1              B2                 B3
    head 
    
    
     1          341-1873-0     142-352-0          143-641-0
     2          336-1768-0     111-369-0          172-725-0
     3          342 1345-0     86-268-3.0         212-785-0
     4          330-1788-0     57-192-0           185-881-0
     5          392-1683-0     42-179-3.0         240-618-0
    
    
             Only 6 Speedfam carriers left and no plates in-house
    
             Will place call to Tim for action plan
    
    
    
                               C-Shift
17.471exFABSIX::J_SWENSONFri Nov 29 1996 19:3920
          Action plan for B1 after talking with Tim.
    
    
     - while running wafers, time how long the water is on in segment 6
       (should be 2-3 seconds at the end of the segment, no more)
    
     - clean the index table and the unload tracks with peroxide 
    
     - Do not use fresh carriers from the carrier rebuild room. (6 left)
                                                             (no plates)
     - Perform clean with peroxide on unload elevators also.
           
     - as a last resort, put tool on barrel to isolate house system.
        
           
            This is a long shot since B2 and B3 are so clean.
           
                          
    
                                C-shift
17.472Particles... begone!SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Nov 30 1996 03:1719
        The tool was down at the beginning of the shift for 
      particles.  Before the action plan was implemented, it
      was decided that another boat of 25 wafers were run with
      (5) more particles wafers as well.  The tool passed 
      particles after they were run.


                                                
                                        D shift

                           \/  \/
                            (__)
                            (@@)
                     /_______\/
                    * |     ||  \ Brian's deer that got away....
                      ||----||
                      ~~    ~~            
                           
17.473tool lockupSUBPAC::LANDRYMon Dec 02 1996 19:3718
    Couple of problems back to back today.........
    
    Whilst the tool was running, the system stopped processing due to a
    Process Door open fault.  The doors were closed and the system running
    resumed.  Unfortunetly, the tool timing sequence messed up and the run
    kept on polishing.  Needless to say, the five wafers were toasted.  The
    "stop run" button would not actuate to halt processing.  Finally, the
    system decided to stop.  We then tried to re-sart the system, but a
    multitude of things were not going right.  The heads never dropped,
    although they oscillated.  The carriers didn't rotate as well.  The
    flippers halted operation.  We couldn't end or stop the run. 
    Therefore, a soft-boot was performed.  After initialization, the tool
    was put to the DOS prompt and a a hard power down was performed.  The
    system was then repowered up and initialization was performed.  All
    functions were working properly.  Ran a quick test run to verify tool
    performance, all is well.
    
    					A shift
17.474Carrier # 3 dropping wafersFABSIX::R_POIRIERTue Dec 03 1996 07:235
	Replaced Carrier # 3. It was dropping wafers in the 
	clean tub after polishing. Ran boat of dummies and
	had no more occurances.

17.475wafer loss/inverter errorFABSIX::B_FINNTue Dec 03 1996 19:1612
B1 failed in 1st_pmd run with 5 monitors for carrier rotation 
(2+3) and inverter error. Wafer lost board showed a fault for wafer 
loss, but no wafers were out of the carriers. Carrier one wafer loss 
sensor had samll droplets of water splashed around it on the bottom of 
multihead, and was very dim. Wiped up water and  adjusted carrier one's 
wafer detect sensor per the procedure and tested the other four sensors. 

Ran test wafers. GOOD TO GO!

					A Shift

17.476sensory perceptionFABSIX::M_RIDLONEliminate the obliviousWed Dec 04 1996 03:3710
    Alarmed for process door open during process. Operator checked doors
    and found them closed and restarted to continue processing. The tool
    then alarmed for heads not rotating in segment 5. Thats when we saw the
    Bat signal. We checked all the doors and found that the right side
    index table process door sensor wasn't seeing the door. Turned out to
    be just a loose crimp on one of the pins. Corrected and resumed
    polishing with out a hitch. 
    
    
    						B-shift
17.477Wafer detect sensors adjustedFABSIX::R_GEEWed Dec 04 1996 19:057
	Wafer detection tripped stopping the table and carrier 
	rotation, no wafers had slipped out. Calibrated all 5 wafer 
	detection sensors, ran 50 wafers, all set.

					A shift
		

17.478Head #3 taken off line until further noticeSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Dec 07 1996 02:0414
          The tool alarmed for Head #2 and #3 not rotating in
       segment #4.  I assuming it was caused by #3 but no
       one saw it actually happen.  I booted back-up fine 
       with no rotation errors.  When OPS went to run the next
       (5), head #3 stalled out in seg. #4 again.  Per E.E.,
       (T. Dyer),  head #3 has been taken off line and will need
       MQC's checked before they can run product.  I'm leaving 
       the carrier on the tool to protect debris from getting
       up inside the spindle or stuff coming out...

         
                            D shift, oh, it's just me tonight!
                    
17.479motor fix for stallingSUBPAC::CMPThu Dec 12 1996 08:296
    Yesterday we replaced the motor on head 3, the motor originaly stalled
    and it was suggested from Phx. that we adjust the resolver to match the
    Infranor. After the adjustment the motor was checked and it worked fine
    ( no current spikes or stalls) a small vibration was noticed and a
    software correction was made that stopped the noise(the software
    was adjusted to stop an over driving situation to the motor).
17.480RESET FILE PARAMETERSSUBPAC::CMPThu Dec 12 1996 19:176
    NOTICED THAT THE HEAD STALLING WAS HAPPENING MORE FREQUENTLY TODAY,
    THIS WAS DUE TO THE FACT THAT YESTERDAY WE SET THE CLOSED LOOP AND 
    INTERVALE TIMER BACK TO 36 AFTER THE MOTOR WAS REPAIRED. THE REAL
    REASON THE NUMBERS WERE CHANGED IN THE FIRST PLACE WAS THE SPIKING
    THAT WE SAW WITH 493C SOFTWARE AS WELL AS THE MOTOR STALLING.
    NOW THE NUMBERS ARE BACK TO 75.
17.481Bad uniformitySUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Dec 13 1996 06:566
               A pad change was needed due to failing MQC's
          for bad uniformity.  The pad replaced and the tool 
          passed MQC's. 

                                          D shift
17.482Inverter Comm Errors - Suspect Wafer Loss BoardASDG::POIRIERSat Dec 14 1996 19:4722
    
    Inverter Communication Errors:
    -----------------------------
    The system errored for Inverter Comm this afternoon. The LED was lit
    for carrier #1 on the wafer loss detection board. The wafer loss sensor
    for carrier #1 was adjusted and a sucessful 25 wafer verification run
    was completed. The tool was then put Up and Idle.
    
    On the next run, the inverter communication error returned. Again, the
    LED for carrier #1 on the wafer loss detection board was lit.  The 
    detection board was jumped out and the system did not fail for the 
    inverter. Every time the detection board was activated, the system
    would shorty thereafter fail for inverter comm.
    
    Although there is no physical evidence of failure on the wafer loss
    detection board, we're of the opinion that something is wrong with it. 
    (The last time this board failed on B2,there were scorched components)
    
    We're passing on to B Shift for replacement of the Wafer Loss Detection
    Board.
    
                                                        C Shift.
17.483wafer loss pcbFABSIX::M_RIDLONEliminate the obliviousSun Dec 15 1996 20:255
    Replaced the wafer loss board. Recalibrated head 1's wafer loss
    sensor. Ran 15 wafer verification (room a little low on wafers).
    
    
    					B-shift 
17.484wafer loss infoFABSIX::M_RIDLONEliminate the obliviousSun Dec 15 1996 21:079
    Troubleshot wafer loss board replaced last night. Found that the board
    is faulty. U4 pin 3 should in a hi state with pins 1 and 2 at a low
    state (nand gate) in order for this board to output the correct
    signals. U4 pins 3 is at -0.46 volts while 1 and 2 are at .166 volts. 
    Obviously this gate is malfunctioning. These boards are failing at a
    higher than acceptable rate, in my opinion.
    
    
    					B-shift
17.485new proximity sensorFABSIX::M_RIDLONEliminate the obliviousMon Dec 16 1996 23:596
    Replaced the proximity sensor for the right front process door. The
    led was burnt out and this is the one we had problems with last week
    as well. Tested fine. 
    
    
    					B-shift
17.486readjust voltageSUBPAC::CMPWed Dec 18 1996 15:014
    Lowered the voltage on the rotation Infranors to 10.5V and ran 25
    wafers. Set the down force to 400 pounds and ran a 30 second process
    on no crash. No stalls noticed and the down force stayed within a good
    limit.
17.487Heads stall - not machine issueSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Dec 19 1996 05:0921
          Head #2 stopped rotating at touchdown.  Tool alarmed
       and halted.  Head #2 voltage was sitting at 10.5V.  Re-set
       tool and started over.  It failed two more times on #4 and
       then a third time on head #5.  The voltages were tweaked
       back up to 12V.  The tool failed again on head #2.  Spoke
       with Rich on the phone.  Checked ramp time...  it was ok.
       On thing noted, it was only product wafers it was screwing
       up on.  On the third run, 5 send-aheads were run under the
       same recipe and they ran fine.  Ran the product again and it 
       stalled on head #2.  We then ran 25 wafers under the same 
       recipe without any problems.  We then ran (5) wafers from
       an ILD3 lot using the same recipe and it worked.  The 
       original lot was put back on the tool and it failed again.
       The ILD3 lot continued to run fine on the tool with the
       voltages back to 10.5V and the ILD1 lot was put on CMP.B3 
       and that ran fine.  I know it is unlikely but this happened 
       to a lot about a year ago.... 
         

                                      D shift
17.488Head stalls... carriers guilty as chargedSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Dec 20 1996 06:1835
          Head #2 stopped rotating 10 seconds into seg. #4.  Tool alarmed
       and halted.  Head #2 voltage was sitting at 12.5V.  Re-set
       tool and started over.  It failed one more time on #2.  
       A ILD1 run went first then failed on an ILD2 lot.  I can't
       explain..   yesterday seemed so understandable..  The downforce
       was checked on all (5) heads and they were in spec.  A pad change
       was done after that then we turned it back to OPS.  Tool went
       through break-in without any stalls...  Tool failed MQC's and
       a pad and carrier change was prompted.  The exposure was checked
       and (4) out of the (5) carriers were out of spec -  (2) were at
       -5, (2) at -6 and one at 7.5.  It makes perfect sense now why 
       certain lots were able to run on the tool and some weren't.
       The ILD1 lot that was on the tool failing yesterday was obviously
       one the thinner side and the white ring was touching the pad.
       The ILD3 lot was thicker and the rings weren't touching.  On a 
       side note, the slurry was checked and one came in a 450ml and one
       a 500 which were both suppose to run at 550ml/mn.  The pad was
       replaced with a expiration date of 1/97 as compared to the one on 
       the tool which was 9/96.  The carriers were replaced and is now
       conditioning and is ready for break-ins again.

             HEAD - PRE - POST
             =================
             1      375    377
             =================
             2      384    377
             =================
             3      374    377
             =================
             4      382    377
             =================
             5      372    377
             =================
                                             D shift
17.489FABSIX::B_WESTPHALSat Dec 21 1996 19:4514
    
CMP.B1 - Down for wafer crash.  A wafer came out of head #2 and crashed head
	 #1 6 seconds into segment 7.  The requested down force was 250lbs but
	 the heads were ranging from 357 - 388.  The tools uniformity was out
	 and removal rates were barely in.  Cleaned up the process area and
	 conditioner.  Changed the pad (135 wafers) with a lot# not used before
	 and carrier #1.  As requested by EEG the "closed loop interval" and 
	 "closed loop delay" variables in the CDATA.CMP file were changed from
	 100 for each to 75 for each.  This was tested and then the variables
	 were changed 75 and 36 respectively and the test was repeated.  The 75
	 and 75 arrangement seemed to work best so the variables were changed 
	 back (75&75) and a 50 wafer marathon was run.  Pad was broke in and
	 the tool sets in wet idle needing only MQCs.

17.490head-2 stallLUDWIG::TDYERMon Dec 23 1996 15:556
    
    Head-2 stalled at touch down on recipe tup_a (377lbs of DF). Pleanty of
    slurry, no vibrations, screen displayed 398lbs on head-2. Re-started
    with out error. Long term fix is being worked with Phoenix.
    
    TD
17.491multiple problemsSUBPAC::LANDRYTue Dec 24 1996 13:3338
Lots of fun with this baby today.......

Right after the morning passdown, we were beeped for a major crisis.  
Seems the previous shift started the "rocket" lot on this tool.  Just
as A shift was coming in, the tool came up with a "process door open"
fault.  Opts tried stopping the processing, but the touchscreen was 
locked up.  Therefore, the EMO switch was used to halt processing. NO
wafers were over polished and the intergrity of the rocket lot was in
deed intact.  Nice save Lads!

Well, the doors were checked.  One of the process doors was on the edge
of making/breaking the sensor.  This was noticed after rebooting the 
system.  We checked the Bimba air cylinders to see if they were locking
the doors in-place.  Low and behold, nary a one was working.  All four
door Bimba's were jammed closed.  Found all to have bent shafts.  There
were only two in stock, so thats all that was changed out.  We changed 
out the two bimba's on the left side of the system (Inverter cabinet side).
A note has been sen5t to the Speedfam reps to have several sent to us.

After resuming production, the tool ran fine for a few runs, then the 
system halted running.  It was in the middle of unloading the Index Table,
the fourth wafer would have been next.  No alarms were heard, nor were any
error codes visible on the screen.  So why did it stop???  Pressed the stop
button, then pressed the start button to resume.  Still no action.  Hit the
stop button again, exited the processing screen, re-entered the processing
screen, then tried resumming the run.  The alarm went off for "one or more
cups not down".  Checked all cup I/O's and found the Unload cup I/O not
illuminated.  Checked the lift cylinder/pedistal/sensor to find that the
pedistal was not retracting enough to make the sensor.  The retaining nut
on the pedistal shaft had loosened up, the pedistal backed out a few threads
with the retaining nut rotating down each time.  The nut was actually pre-
venting the the cylinder to retract properly.  We re-set-up the height/sensor
so that it would function properly.  We then secured the nut (I've been told
it's best to keep your nuts secured).  Of course, the unload cup height needed
adjusting AND the unload pick up spot re-calibrated.  We performed said func-
tions.  Tested the results........smooth, slick.

                               A Shift
17.492Christmas crash.FABSIX::L_SHKOLNIKOVThu Dec 26 1996 07:3011
   
     The system crashed a wafer (actually, 2) tonight.  The system was running
    the IO_MQC recipe.  The wafer came out of carr.#5 and took #4. Wafer detect
    system didn't see the wafer however right after it crashed, the system
    alarmed for "No rotation on head 3".  Cleaned up the mess, changed the pad
    and carriers 4&5.Checked the carrier #5 to load cup alignment- slightly
    off (actually 256 steps out). All is well.  The crash form holds all the 
    vitals, should you need additional info. Just completed weekly pm. and
    released tool for manufacturing. 
                                          
                                          Leon.
17.493Mapper #2 failureSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Dec 27 1996 04:4115

         Cassette #2 mapper failed to map cassette.  Hard booting
     the Kinsington didn't correct it.  A softboot wouldn't work
     either so a hard boot was needed to bring it back to life.


         According to all the shutdown/start-up in the life and
     days of CMP.B1, the mapped/robot association has had problems
     within 24 hrs or a startup.  I don't want to sound superstitious
     or anything but should we make sure a controller is in stock 
     for next week?


                                              D shift
17.494Head to Head Removal Rate & Uniformity ProblemsASDG::POIRIERFri Dec 27 1996 14:5948
    
    Head to Head Removal Rate & Uniformity Out Of Spec:
    ---------------------------------------------------
    The Electo Pnumatic regulator calibration was checked, results were as 
    follows: (@4095)
    
            Before      After
            ------     ------
     #1)    59.8       60.0
     
     #2)    59.9       60.0
    
     #3)    60.4       60.0
    
     #4)    60.3       60.0
    
     #5)    60.2       60.0 
    
     These results were close to nominal. The Spindle Down Force was then
     checked and the results were as follows:
    
               Before        After
               ------        -----
    
    Head #1)   376.4         n/a
    
         #2)   378.7         n/a
    
         #3)   373.5         377.9
    
         #4)   372.5         376.8
    
         #5)   378.2         n/a
    
    
       Again, these results were close to nominal. Our attention was then
       directed towards the pad. The pad peeled off the table with very
       little force. It had only 295 wafers run on it. The pad was replaced
       and MQC's are now being run.
    
       FYI: These head to head removal rate & uniformity problems occured
            yesterday on B3. In each case the pad had run approx 300
            wafers.  (Possible Pad Problem?)
            
    
                                                             C Shift.
    
            
17.495Head to Head Removal Rates Still Out Of SpecASDG::POIRIERFri Dec 27 1996 19:4229
    
    Head to Head Removal Rate Still Out Of Spec:
    -------------------------------------------
    Action Thus Far:
    
    1) EP Regulators Calibrated
    
    2) Down Force Calibrated
    
    3) Pad Changed
    
       Status changing from VERIFY to DOWN.
     ---------------------------------------------------------------------
    Suggestions:
    -----------
    a) Check the Pad Conditioner down force calibration.
    
    b) Try a different lot # Pad.
    
    
    FYI: The removal rates on all 3 Speedfams have dropped slightly from
         the normal 3000 to 3200 A range.  We've been seeing 2700 to 2900
         A today. 
    
    FYI: The house slurry sytem pressure was up to this morning. The filter
         was removed and replaced with an osmonics type filter.
    
    
                                                            C Shift.
17.496Pad change - uniformity/R.R.SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Dec 28 1996 07:0910

           Pad change due to removal rate issues.  A pad
       change was done but head #4 failed for uniformity
       which was ~18% and the others were ok.  After the
       carrier was replaced, the tool came in after a couple
       MQC runs.


                                              D shift
17.497Neumatic adjustment, Semi-annual PMSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Jan 02 1997 08:2313
    
         B1 is still down.  It needs slurry flow cal. and (4) lubrication
      items on the semi-annual.
    
         On the new PM spec, there were two items that were different
      between the tools.  One was Carrier Head balance which was 3.2
      which is now 6psi and Scrubber down was 80 psi and it is 20 psi
      now.  The downforce was checked and was out as a result of the 
      change and was brought back into spec.  The scrubber was tested
      ok a well.
    
                                              D shift
    
17.498High Particles / Cleaned up okASDG::POIRIERThu Jan 02 1997 19:3613
    
    High Particles:
    --------------
    The tool failed MQC's for high particles this morning. The index table
    & Cups were wiped with alcohol and rinsed with DI. The carriers were
    also scrubbed and rinsed. The next MQC was within spec (on the high
    side).
    
    FYI: B2 & B3 MQC'ed fine for particles. The House Slurry system appears
         to be OK.
    
    
                                                                 C Shift.
17.499Index table INFRANOR changedFABSIX::B_WESTPHALThu Jan 02 1997 19:5310
CMP.B1 - Down for wafers not being picked up from the index table.  The carrier-
	 to-cup alignments were checked and all were found to be out in the 
	 same manner.  They were out CCW so far that the carrier would not even
	 fit into the inner part of the cup.  It was also noticed that the 
	 index table was vibrating excessivly. The Index table INFRANOR was 
    	 swapped out and the voltage adjusted to 13.6v.  The alignments were 
    	 done again and the shafts for load cup air cylinders were lightly 
    	 greased.  Operations ran a boat of wafers (BREAKIN/MQCs) with no 
    	 further problems.

17.500oem vac filtFABSIX::M_RIDLONEliminate the obliviousSun Jan 05 1997 03:134
    Installed OEM vacuum filter and another gauge.
    
    
    					B-shift
17.501pad condFABSIX::M_RIDLONEliminate the obliviousMon Jan 06 1997 06:395
    Observed that the pad conditioner is not centered on the pad while
    conditioning. Could not address as the tool was running and would be
    until shift change. 
    
    						B-shift
17.502Conditioner recipe #1 ID/OD changedFABSIX::R_GEEMon Jan 06 1997 18:436
	Changed the ID/OD for pad conditioner recipe #1 from 9.3 to 9.5
	so the conditioning ring is centered on the pad during the 
	conditioning cycle. 	
					A shift
		

17.503pump lines replacedFABSIX::M_RIDLONEliminate the obliviousTue Jan 07 1997 06:236
    Tool alarmed for carriers not rotating in segement 10 of a TUP375 run.
    Found that the line in slurry pump1 had failed. Replaced both pump
    lines and continued the run.
    
    
    						B-shift
17.504index table vibrationsSUBPAC::CMPTue Jan 07 1997 16:174
    Looked into trouble with index table vibrations.
    Checked for a binding and found nothing really bad, but tried another
    infranor and found that it cured the problem. It appears the infranor
    that was replaced the other day was marginal in it's operation.
17.505particles...SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Jan 09 1997 06:5714


         Tool failed for particles after pad change and
      weekly PM.  The tool was cleaned... again... and
      was re-run and failed for particles.. again...
      The tool was cleaned once again along with another
      pad change.  It was noticed that the conditioner wasn't
      catching all of the pad so the O.D. position was changed 
      from 9.2" to 9.5".
      


                                              D shift
17.506Particle Problem / Slurry System Flushed & Filter ChangedASDG::POIRIERThu Jan 09 1997 19:1019
    
    Particle Problem:
    ----------------
    MQC's failed for particles shortly after shift change this morning.
    Particles on B2 & B3 were running close to the UCL. The house slurry
    system was flushed with D.I. water and the house filter was changed.
    Slurry flowed thru the new filter for one hour before the system was
    returned to manufacturing.
    
    The system passed MQC's for particles after the flush and filter change
    but they were higher than expected.
    
      Total                   Large
      -----                   -----
    700 -1300               200 - 250 
    
    
                                                             C Shift.     
                                                                 
17.507prealignerFABSIX::M_RIDLONEliminate the obliviousTue Jan 14 1997 06:227
    Load flipper was occasionally hitting the wafer in the prealigner. We
    ovserved that the wafer, when aligned, wasn't centered between the
    forks of the flipper. We moved the base of the aligner to center the
    wafer. Tool loaded fine after. 
    
    
    						B-shift
17.508Conditioner Wheel ReplacedASDG::POIRIERThu Jan 16 1997 19:0715
    
    Conditioning Wheel Changed:
    --------------------------
    The Conditioning Wheel was changed on the tool Today. The new diamond
    wheel I.D.# is 196. The total wafer count on the tool at change over
    was 109,501.
    
    FYI: Tim D. edited the maint counter screen. There is now a condition-
         ing cycle counter. The counter adds 5 counts for each cycle. Since
         the conditioner sweeps once per cycle, remember to divide by 5
         to get the correct count.
    
        
                                                                C Shift.
     
17.509wafer loss errorsSUBPAC::CMPMon Jan 20 1997 12:307
    Arrived to find the interlock intermitently tripping causing errors.
    Checked the wafer loss circuit and  found that # 3 was tripping in
    segment 5, upon further checking we found the pad was worn to the
    point that the dark of the metal table was bleeding through causing the
    sensor to trip the wafer loss detect. When measured the pad was only
    at 52 which is well below the point of replacement. The pad was
    replaced and the tool brought back up.
17.510Broken Wafers In Unload Elevator TankASDG::POIRIERWed Jan 22 1997 19:3811
    
    Broken Wafers In Unload Elevator Tank:
    -------------------------------------
    A slight mishap occured at shift change tonight. Two wafers slid out of
    a cassette as it was being removed from the unload elevator. The wafers
    slid partially into the water track as the elevator was homing. There
    are broken pieces of wafer in the bottom of the tank. Passing on to 
    "D" shift for clean up. 
    
                                                       C Shift.
    
17.511Elevator crashFABSIX::O_WHITFORDLead, follow or get out of the way....Wed Jan 22 1997 21:257
    Drained elevator #1 tank and cleaned up wafer residue. Flushed tank
    (fill/drain) several times and then filled and observed water level to
    make sure the dump valve was not leaking due to chunks getting past  
    the seive and down the drainline. Looks to be holding o.k., returned
    control back to ops.
    
                            Dave W.
17.512homed load moduleFABSIX::M_RIDLONEliminate the obliviousTue Jan 28 1997 07:005
    Tool would not map cassette b. It was obvious that tilt B had lost hits
    home. Homed the load module and mapper worked fine on both A and B.
    			
    						
    					B-SHIFT
17.513Unload Lift Cylinder Sensor Replaced / Cup #2 Sleeve HonedASDG::POIRIERThu Jan 30 1997 19:1124
    
    Unload Lift Cylinder Sensor:
    ---------------------------
    System errored for "one or more cups not down".  It appears that the
    table rotated with a cup not completely down. There was a mechanical
    collision between load cup #2 and the unload lift cylinder.
    The sensor on the unload lift cylinder was replaced.
    
                                                                 
                                                        C Shift.
    
    
    
    
    Load Cup #2:
    -----------
    Cup #2 failed to reach its down position tonight. Speedfam honed the
    sleeve for cup #2 in the index table. Verification run in progress.
    
    
                                                             C Shift.
    
    
    
17.514sticking cupSUBPAC::CMPThu Jan 30 1997 19:123
    The tool had a sticking unload cup (2), found the shaft was a little
    tight in the table. Honed the opening to allow the cup to move freely.
    
17.515Load station problemsFABSIX::O_WHITFORDLead, follow or step aside.Thu Jan 30 1997 22:3615
    Load station problems. 
    
    1)   Operator reported that the load robot wasn't placing wafer in
    pre-aligner properly. Described as it wasn't actually setting the wafer
    into the pre-aligner, it was being "raked" off the robot arm. 
    
    2)   When the load flipper came to the pre-aligner, the fingers closed
    over top of the wafer.
    
         We exercised the load sequence after powering down the tool,
    leaving it set and restoring power. The exercise sequence went fine
    with no evidence of either problem descibed above.
    
                               
    
17.516unload pedistal bend_ageSUBPAC::LANDRYSun Feb 02 1997 20:0033
The system experienced a crash today: "Load Cup 4 vs Unload lift pedistal"

Whilst unloading wafers, somehow, the lift pedistal didn't retract enough
to allow the next cup to safetly pass by it.  This caused the bottom of
Load Cup 4 to severly bend the lift pedistal of the unload cup.  No alarms
were experienced with this mishap.  The Opt Tech noticed that the system 
had ceased operating and was sitting stagnant.  Upon investigating the
reason why, he found that the cup currently positioned at the unload 
position was angled badly.  Our initial checks found that the unload 
pedistal was bent over, wedging Load cup 4 and halting the Index Table.  
The respective Infranor was fine and NO breakers were tripped.  No alarms
or timeouts were prevalent.

We had to exit out to DOS and use the DCX commands to drive the table back
off of the cup.  The cup was then removed and checked out.  No apparent 
damage was found.  The respective hole for that cup was checked out as well.
Again, no apparent damage.  The locating pin was present, so that rules out 
the cup rotating as the cause.  We then checked the lift cylinder/pedistal.
The pedistal portion of the assembly was bent over about 60 degrees.  It
was removed.  The lift cylinder didn't appear to be damaged, as it's up /
down movement was smooth and consistant.  We checked the sensor gapping;
it was found to be at .015".  This was regapped .010" and a new pedistal
was installed.  We verified that the sensor on/off LED performance was
working as should.  Height alignment was performed at this time.  Also, 
unload flipper at unload cup calibration was performed.

As for unload cup 4's performance after the crash.....we checked it out at
the load cup station and the Index Table home position.  The cup was lifting
smoothly and consistantly.  It appears to be fine as wafers are loading
quite well at both cup 1 position and head 4 position.  We watched it's 
performance for a couple of cassette loads of wafers.  All is well.

                                                A shift
17.517Load cup 5 sensor adjustFABSIX::B_FINNSun Feb 09 1997 20:089
	Errored for one or more cups not down. Found Load Cup 5 I/O 96 
not made. Found sensor touching mushroom. Adjusted to an 10 mils and 
sensor did not see mushroom. Adjusted gap to 8 mils, aok. Ran up and down 
in manual mode and cup sensor performed aok. Seemed more 
sensitive that usual during the position adjustment. 
					
							A Shift

17.518robot adjustmentSUBPAC::CMPTue Feb 11 1997 07:302
    Made a slight adjustment to the robot positioning to the aligner. It
    was off slightly which could have caused a problem.
17.519lift assy repair for 2-3-97SUBPAC::CMPMon Feb 17 1997 11:578
    Found the tool down for cups not down. Found load cup 4 had crashed
    into the lift assy. causing the mushroom to bend over and damage the
    sensor assy. Replaced the air cylinder, sensor and mushroom then set
    the sensor gap. During the repair we discovered the springs on the
    bottom of the cup would not allow the center to fully reseat and the
    lift plate on the bottom hit the mushroom causing the damage. We also
    replaced the springs and plate due to the damage. Returned tool to
    operation.
17.520low slurry flowFABSIX::R_GEEMon Feb 17 1997 19:338
	Errored for low slurry flow. Found air in the slurry lines.
	Called and paged Ashland chemical they said they did change filters
	without notifying use, they were told that we must be notified.
	Manually pumped slurry and restarted the tool.

					A shift
		

17.521crash, ? whySUBPAC::CMPWed Feb 19 1997 17:024
    Cleaned up the crash that occured during a product run. The wafer in
    head 5 turned to very very small pieces and ruined the rest of the
    wafers. The cause is unknown and was discovered in step 11. Cleaned
    the tool and returned to production.
17.522Wafer breakage - no productFABSIX::B_JUBINVILLEFri Feb 21 1997 08:3313
    


          Tool crashed...  It looks like it came out of head #4 and
      took out #5.  It was discovered that the wafer detect system
      was physically jumped out so that's why the tool kept running.
      Without the detection system on line, the tool will keep
      running until it lifts off the table from which it will check
      for vacuum.  It didn't have good vacuum, thus it stopped
      at seg.#11 as well as it did two days ago.  Initial clean-up 
      was started and the tool has been turned over to C shift for 
      alignments and possible downforce.
                                                 D shift and Brad
17.523Replaced Slurry Pump #2 Motor/ Calibrated Down ForceASDG::POIRIERFri Feb 21 1997 19:4815

CMP.B1 - Down for wafer crash. Continued clean up where D-shift left off. 	
	 Cleaned index table and cups.  Checked the down force and carrier to
	 cup alignments.  Made corrections where ever needed.  Conditioned pad 3
	 times and returned it to production.

                          Down force using point to point method

                               found at      adjusted to
                                 347           354.2
                                 370           353.4
                                 357           354.7
                                 349           353.8
                                 368           354.1
17.524Inverter errorsFABSIX::R_POIRIERWed Feb 26 1997 07:1816
	During the 3rd run of a product lot the tool stopped in
	segment # 1 for an "Inverter Error". The tool also shut
	itself off. We removed all wafers and re-homed the tool.
	We then ran a boat of dummy wafers and during the 4th
	run the same error occured. Again in segment # 1.
	I thought the pad looked a little thin, it had 1300 
	wafers 	on it, and I thought it might be the same problem 
	we had on B2 a few months ago where the wafer detect was
	causing the errors, so I changed the pad and ran some
	break-in wafers. Mitch also adjusted the wafer detects.
	Tool is running ok.


					B Shift

17.525computer lock up X 2SUBPAC::LANDRYSun Mar 02 1997 19:3316
    The tool locked up twice today for different problems......
    
    1)  The first lockup happened mid morning.  Seems the Multihead was
    stalling during it's travel.  It required a softboot to recover.
    
    2)  The second lockup happened late in the shift.  The condition arm
    froze above the polish table.  No "red lined" alarms appeared on the
    screen, but the audible alarm was sounding.  Tried a quick softboot,
    but no luck, as the polish table would not stop rotating.  So, a
    hardboot was then performed, but we came upon a bit of trouble when
    head # 3 would not home.  Found that the overtravel switch had been
    made.  So, rebooted, skipped over initialization, went to DOS, executed
    the DCX command and drove the motor off of the switch.  Rebooted again
    and accomplished a successfull initialization of the system.  
    					
    						A shift 
17.526LOADING PROBLEMSFABSIX::B_WESTPHALThu Mar 06 1997 20:5214
CMP.B1 - Down for loading problems.  The tool did not pick up the wafers
	 properly from cups #2 & #3 after it was returned to production after
         the monthly pm.  Checked the carrier to cup alignments (all carriers
	 out) and the indexer needed to be moved clockwise 3500 counts.  The 
	 proper adjustments were made and saved.  When the tool was homed and 
	 the alignments verified the index table needed to be moved @3500
 	 counts counterclockwise again.  We assumed that the INFRANOR was bad. 
	 So the same was swapped and the alignments done again with the same 
	 result as before.  Speedfam figured out that when we were rehoming w/o
	 exiting the screen it goes back to the previous home position (as a 
	 safety feature).  The Tool was returned to production for Breakins &
	 MQCs and is now having particle problems.

17.527elevator #1STRATA::FINNMon Mar 10 1997 17:5925

	System halted during process with arms "in" at secondary clean station.
	Conditioner disk was over polish table.  VIA the opto 22 i/o's - the 
        system knew that the conditioner disk was up, but did not have 
        indication of its lateral location. All arms showed they were at 
        home position. Table was running. Error read "carriers not arriving at 
        load".


	Aborted run. Upon bootup, Carrier 3 did not home. Needed to reset DCX 
        board. Retrieved product wafers. Ran arms and pad conditioner and 
        watched sensor. No errors.
     
        Operation said Elevator #1 had been cross slotting. Check alignment and 
        found elevator low. Verified with several carriers and it was 
        consistently low. Recalibrated Elevator #1  and checked with jig and 
        wafers @ top, middle, and bottom slots. 

	Ran 25 no_touch wafers. AOk


 						A shift / Kevin

 
17.528Load problem Carrier #2FABSIX::R_POIRIERWed Mar 12 1997 05:2017

	During a touch-up run, tool errored for low vacuum.
	Carrier #2 didn't load a wafer in the 3rd run. Couldn't 
	open doors without stopping process. Stopped process,
	and checked for vacuum on Carrier #2 by placing the
	wafer on the carrier. Vacuum was present. Unloaded all 
	wafers, and exercised load cups. Cup #2 showed no problems. 
	Carrier #2 looked ok. Ran some dummy wafers and Carrier #2 
	load each time with no problems. Returned tool to Operations. 
	Tool ran fine the rest of the nite.




						B Shift

17.529Wafer left in load cup 2FABSIX::J_SWENSONThu Mar 13 1997 19:3914
    Wafer left in load cup 2
    ------------------------
    
    - Found load cup 2 not functioning properly.
      replaced springs.
    
    - Vacuum low replaced vacuum pump and filter.
    
    - Ran wafers to verify, NPF
    
    - tool ran rest of shift without incident.
    
    
                                 C-shift/Dave,Kevin,Rich
17.530wafer slide on CMP.B1FABSIX::J_SADINFreedom isn&#039;t free.Sun Mar 16 1997 19:5610
    
    
    	System slid a wafer out of carrier five and it hit carrier four. No
    smashed wafers, but we did lose a piece of the retainer ring on carrier
    five, and the ring on number four had some damage also. Replaced both
    carriers and put system to verify for mqcs. 
    
    a-shift
    
    p.s. - dummy wafer slid during segment five of a PMD touchup. 
17.531Wafer detect/infranor issueSUBPAC::BJUBINVILLECMP Equipment EngineeringWed Mar 19 1997 06:3216
        While completing the weekly PM, it was suggested
   last week that the wafer detect system be checked in a 
   different fashion.  Actually go into MANUAL POLISH and
   trip a detector to see if the table and heads would
   actually stop.  Well, I put a wafer under the detector
   for head #3 and everything stopped...  great... It works!
   The tool then froze up.  It eventually errored for an
   Inverter Communication error.  We a hit a button to continue
   and it started homing the arms.  Guess what...  #3 wouldn't
   home.  The infranor was out...  Tried resetting the tool
   and it was dead...   how 'bout that, two in one night!
      


                                              D shift
17.532Pad and carriers replacedSUBPAC::BJUBINVILLECMP Equipment EngineeringFri Mar 21 1997 06:0712



           Pad needed to be replaced due to the step height
      and carrier #1, #2 and #3 were replaced for wafer count.
      OPS had a tough time trying to bring in the uniformity.
      We just replaced carrier #4 and we are waiting for the results.



                                              D shift
17.533Carrier Rotation FailureASDG::POIRIERSat Mar 22 1997 19:3710
    
    Carrier rotation failure:
    ------------------------
    The system errored for carriers not rotating in segment 3 at 7:10 p.m.
    All 5 carriers stopped rotating in seg 3. Infranors were had power and
    no errors were observed. Unloaded wafers from hot lot and placed tool in
    wet idle. Passing on to B Shift for repair.
    
    
                                                          C Shift.
17.534Carrier #2 rotation infanorFABSIX::R_GEEMon Mar 24 1997 14:117
	Errored for carrier #2 not rotating in segment #3 twice. Adjusted
	the Infranor voltage from 12.54 to 13.0


					A shift
		

17.535carriers not rotatingFABSIX::J_SADINFreedom isn&#039;t free.Tue Mar 25 1997 14:5011
    
    
    	Multiple carrier not rotating errors (head 2 and head 5). Tried
    swapping the head 2 infranor earlier, but received bad infranor from
    stock. Replaced old infranor and adjusted voltage to 14.70. Booted up
    ok but began erroring again shortly after install. Error can be cleared
    and the recipe restarted, so production is currently running like this
    until the problem can be properly diagnosed. Error only seems to occur
    on the first run of five wafers (ok after that). 
    
    a-shift
17.536Carrier #1 rotation errorFABSIX::R_POIRIERWed Mar 26 1997 06:2415

		B1 errored for Carrier 1 not rotating during
		segment 4. Carriers had not yet touched the
		pad. Unable to recover using the Touchscreen.
		Had to do a soft-boot. Unloaded all wafers
		and increased Infranor for carrier #1 rotation
		from 12.3v to 13.0v. Checked rotation and it
		worked fine. Gave tool back at 1:30 am and it 
		has been running fine ever since.



					B shift

17.537uniformity problem head 3FABSIX::J_SWENSONThu Mar 27 1997 19:4311
    Carrier 3 uniformity problem
    ----------------------------
    
    - Carrier 3 failed twice @17% for uniformity
    
    - Replaced carrier and uniformity came in @9.5%
    
    - This happened last week also and the carrier was swapped out
      and uniformity came in.
    
                                   C-Shift
17.538Carrier #5 not being detected at polish tableSUBPAC::BJUBINVILLECMP Equipment EngineeringFri Mar 28 1997 05:1921
         Tool was running product when the OPS tech noticed
     there was an error on the screen for "timeout for one
     or more heads not down."  The screen and process timer
     were froze.  The tool was running TUP_375 and was in 
     segment #3.  For the fact that the tool was in touch-up,
     all heads were down and rotating as well as the table
     and most importantly, product...  the tool was EMO'ed.
     Upon restart, the tool came up fine and all of the 
     Carriers At Polish REED switches appeared to be reading
     ok.  We ran 15 VERIFY wafers and it was fine.  The OPS
     tech went to run a lot and it timeout again, waiting
     for one or more heads not down but they were.  The 
     inputs were quickly checked and head #5 INPUT wasn't 
     on.  The tool was again EMO'ed.  We ended up having to 
     adjust the positioning of the REED switch for CARRIER #5
     AT POLISH TABLE.  VERIFY wafers were run and the system
     was running fine.  Good to go......


                                              D shift
17.539Rotation ErrorsASDG::POIRIERFri Mar 28 1997 19:3218
    
    Head Rotation Failures:
    -----------------------
    While running TUP_A this morning, carrier #2 stopped rotating 4 seconds
    into segment 4. The TUP_A & 1P_ILD3 recipes were deleted. New recipes
    were copied and edited to make new versions. 55 wafers were then run on
    the NEW TUP_A with no further problems.
    
    
    After lunch, 20 wafers were run on IO_ILD1 & 20 Wafers on IO_ILD2 with
    no problems. At 3:30 the system errored for for " all 5 carriers not
    rotating" in segment #4 while running TUP_125. The product wafers were
    removed from the tool and 50 test wafers were run using TUP_125.
    There have been no rotation errors so far. Passing on to D Shift to
    continue the evaluation.
    
                                                                C Shift.
    
17.540Tool verifiedSUBPAC::BJUBINVILLECMP Equipment EngineeringSat Mar 29 1997 03:507
          Tool ran an additional 25 on top of C shifts 50 wafers
       with no rotation errors.  Tool was turned over to OPS and
       passed MQC's.


                                              D shift
17.541Carrier to Cup ALign/ Down Force Cal / Infranor ReplacementASDG::POIRIERThu Apr 03 1997 20:2246
    
    Wafers Sliding Out of Head #4:
    -----------------------------
    A wafer slid out of head #4 this morning and collided with the wafer in
    head #3. Both product wafers were broken. The tool was in seg 3 of the 
    TUP_B recipe when the slide out occured. 
    
    This afternoon, a monitor wafer slid out of head #4 at touch down but
    did not break. The system was running the IO_MQC recipe at the time.   
    
    
    The manufac tech running the tool told us that the head to head removal
    rate was out of spec. 
    
    
    The following action was taken:
    ------------------------------
    * The "carrier to cup" alignment was performed on all 5 heads.
      Head #4 needed adjustment.
    
    
    * The down force calibration was performed on all 5 heads.
      
    
     We decided to run a 25 wafer verification run after the adjustments 
     were made. The tool vibrated on the 3rd set of 5 while running the 
     TUP_B recipe. The wafers were a little thin for a 50 sec polish.
     The cycle stop button was pressed to prevent the last 10 wafers from
     running. This action caused the tool to lock up.
    
    
    The current situation is:
    
    * Head #4 won't home.
    * LED #5 on DCX board #2 is lit. (Head #4 oscillation)
    * The Infranor for Oscillation Arm #4 error LED's are lit.
    * The Infranor for Rotation Arm #3 error LED's are lit.
    
    Action:
    
    * It's 7:15 p.m.
    * We're ordering replacement Infranors from the stock room.
    * Ritchie from Speedfam is comming in to replace the Infranors
      with "D" Shift. 
    
                                                                 C Shift.
17.542repairs to InfranorsSUBPAC::CMPThu Apr 03 1997 21:192
    Replaced a Infranors, head 3 rotation, head 4 osc. returned tool to 
    wet mode to continue the verify.
17.543Wafer/water track #2 errorSUBPAC::BJUBINVILLECMP Equipment EngineeringSat Apr 05 1997 06:127
    
        Twice wafers got stuck trying to go into elevator #2.
     slot #1 position was moved down 100 steps.  No problems
     since.
    
                                   D shift
    
17.544WAFER CRASHFABSIX::B_WESTPHALSat Apr 05 1997 20:2310
CMP.B1 - Down for a wafer crash.  A wafer came out of head #3 & crashed both
	 carriers #3 & #4.  Checked the carrier to cup alignments and all
	 looked good.  The downforces were checked.  Head #4 was at 346.1 lbs. 
	 and head #5 was at 461.2 lbs.  Both heads were brought in to 354 lbs.
	 Some water/slurry was found in the vacuum's moisture trap.  The filter
	 was changed.  No definiate cause for the crash was found.  The tool 
	 was cleaned and the apropriate consumables were changed out.  Returned 
	 the tool to production. 

17.545extension ring culprit to crashSUBPAC::LANDRYSun Apr 06 1997 20:2119
    
Crash!!!!
    
    Started the shift off today with a crash on this system.  A wafer had
slid out of carrier #3.  Said wafer then preceeded to take carrier #2
out to lunch.  Wafer detection kicked in, halted the table as should, 
and prevented any further damage.  Upon further review of the intricate
details, it was discovered that carrier #3 had four (4) screws missing
from the extention ring.  The screws were not found in rinse tub area,
it appears that they were inadvertantly missing upon installation on
the system.  

Regardless of the carrier problem, we quickly checked the head to cup
alignment on that head.  It was fine!  Changed the pad, installed new
carriers on heads 2 & 3, and ran a successful breakin procedure.  MQC's
were run, but a few adjustments were made due to uniformity and removal
rates.  The 2nd set of MQC's ran fine.

      					A shift
17.546Wafer slipped out of carrier # 1FABSIX::R_POIRIERMon Apr 07 1997 08:3113

		During a process run (TUP200) a wafer came out
		of carrire # 1 at touch-down 8 secs. in Seg.3.
		Rebooted tool and unloaded all wafers. WE took
		off carrier # 1 and found nothing wrong. 
		Ran a boat of dummy wafers, no problems. Ran
		product and again saw no problems. Gave tool
		back to production.


					B Shift
 
17.547dropped wafer carrier 3FABSIX::B_FINNTue Apr 08 1997 20:4610
Dropped dummy wafer from carrier 3, four secs into seg 3 of touch up d recipe.
Wafer did not break due top waf detect. Checked all alignments, found cup to
 carrier out on all but cup four with cup 3 being the worst. adjusted checked 
cup to flipper aok. 

returned to manuf.

							A Shift

17.548LOST A WAFER/SCHEDULED PAD CHANGEFABSIX::B_WESTPHALSat Apr 12 1997 20:2915
    
CMP.B1- Down for an error for "carriers not rotating".  The tool had errors for
	"carriers not rotating".  It was at step 3 with zero seconds of process 
	time expired.  The heads had not even touched down.  Noticed there was 
	a wafer in the bottom of the polish basin and all wafers to be unloaded
	from the index table were present except for head #1's.  All five 
	carriers had wafers on them ready to be processed.  Deduced that 
	carrier #1's wafer did not blow off during the unload cycle for some 
	reason.  Surface tension must have held the 2 wafers in carrier #1's
	pocket.  When the carriers began to turn in steps 1 & 2 the surface
	tension broke between the two wafers triping the wafer loss sensor.  
	Ran a full boat of wafers w/ no errors.  Returned the tool to
	production... The tool later went down for a sheduled pad change. 
	tool is presently running MQCs.
    
17.549Elev 2 belt brokeLUDWIG::STANDINGWed Apr 16 1997 20:534
    CMP.B1  Elevator 2 belt broken, use elevator one only.  And run 25
    wafers max.  Belt ordered, will be in tomorrow.
    
    
17.550Elevator, unif., pad changeSUBPAC::BJUBINVILLECMP Equipment EngineeringThu Apr 17 1997 01:0821
    
    
        Elevator #2 kept alarming for not being at home.  Since
     belt not on tool, the elevator was disabled in the hardware
     enable menu.
    
        Tool failed MQC's - pad change and weekly PM completed 
    
        On a side note, when we checked the wafer detect system,
    we tool heads #1, #2, #4 and #5 off line to do the test.  When
    we ran the WAF_DET recipe, heads #3 and #1 oscillated out even
    though only head #3 was on line?!?!?!?!?!?!?  Impossible?? I did
    it.  Of course when we tripped the detection board with a wafer
    and stopped the tool, Head #1 eventually timed out for not being 
    at home and the screen froze.  The tool had to be EMO'ed and we
    didn't lose an infranor.  That's a first...  Tool ran fine after
    that.  We are just waiting for MQC results.
                              
                                                  D shift
                              
    
17.551DECONTAMINATION STARTEDFABSIX::B_WESTPHALFri Apr 18 1997 21:458
CMP.B1 - Down to external for decontamination.  The slurry lines were purged w/
	DI.  The DI & slurry lines are all currently being purged w/ N2. 
	Only a few solenoid valves can be purge at one time or the power supply
	will load down and cause all solenoids to chatter.  wiped down the
	cups and index table.  Carriers were removed.  Slurry and DI were shut
	off.

17.552Decon continuesFABSIX::O_WHITFORDLead, follow or step aside.Sat Apr 19 1997 07:159
    Continued with decon. 
    
    Purged city water line with N2 for @ 3 hrs and then blocked off all
    drains and purged elevator, scrubber, index table and polish table GCW
    drains for @ 6 hrs. Nocied some water build up in slurry lines,
    reconnected N2 to slurry inlet and it is still purging.
    
    
                                             D shift
17.553Decontamination Punch List - UpdateASDG::POIRIERSat Apr 19 1997 20:3022
    
    DECONTAMINATION PUNCH LIST:
    --------------------------
    All items on Bruces' list have been completed with the exception
    of "locking down the Multihead". Speedfam plans on doing this on
    Monday. Please double check the Vacuum, Plumbing & electrical
    connections in the Automation Frame before separating it from
    the Main Frame. We're pretty sure we got everything but its
    possible we missed something. 
    
    The vacuum trap along with the gages and vac pump fittings have been
    stored in the vacuum pump cabinet under the floor. The electrical
    connections for the vac trap solinoid & pump have been labled and
    stored in th vac pump cabinet as well. 
        
    The chiller has been drained and unplugged.
    
    The platen chillers lines have been purged.                  
    
    
                                                               C Shift.
    
17.554removed pressure transducersFABSIX::B_FINNSun Apr 20 1997 20:107

	Removed pressure transducers from each carrier arm. Replaced 
slurry lines from valves to carrier arm. Removed pressure transducer wiring. 

					A shift

17.555b1 removed from roomFABSIX::B_FINNWed Apr 23 1997 20:1410

	B1 was removed from the room today. Temporary wall was reinstalled. We
reinstalled the guards around the exposed B1 facilities hookups. Hooked up
the VAC wands. Coughlin's removed the mega slurry alarm boxes and conduit, 
only on one box has been reinstalled and no conduit.


					A shift